A Complete Bibliography of Publications in the Journal of Cell Biology: 2015–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

02 June 2023 Version 1.07

Title word cross-reference

1 [AANLL^{+}20, RCA^{+}21]. 14 [KMW20]. 2 [MNvdS^{+}20, SCK^{+}20a, SCK^{+}20b]. 3 [HHGR21]. ^{2+} [BS20b, LLK^{+}21, PMSO^{+}23, SIP^{+}23, YCC^{+}21]. _{4} [SMC^{+}20]. α [BJSOS^{+}20, BJSOS^{+}21, EMEZ^{+}20, FOR^{+}20, GCS^{+}20, GLM^{+}22, KST^{+}22, LHL^{+}23, LGS22, LSOM23, MMDK^{+}22, RGP^{+}22, SMS^{+}20, SGL^{+}23, WM23]. α5β1 [HAL^{+}23]. β [ACPR21, BP22, Bog21, EMEZ^{+}20, GL20, GCS^{+}20, HMT^{+}21, JLS^{+}22, KST^{+}22, LCB^{+}23, MOS^{+}20, MSX^{+}21, NKS^{+}21, SIP^{+}23, SMM^{+}21, SCN^{+}23, SPKP22, WXM22, WM23, XGD^{+}23, ZGR^{+}22]. β_{2} [SMC^{+}20]. Δ [HVPM20, LGB^{+}21]. F [MMDK^{+}22]. γ [LTL^{+}20, WTU^{+}21, ZPŠS21]. κ [HKK^{+}20]. N [RVNS21, SYW^{+}20].

-actin [MMDK^{+}22]. -barrel [SMM^{+}21]. -catenin [BP22, HMT^{+}21, NKS^{+}21, MMDK^{+}22, SGL^{+}23, GL20]. -cell [SIP^{+}23]. -coronavirus [JLS^{+}22]. -dependent [LLK^{+}21]. -heavy-spectrin [SCN^{+}23]. -integrin [SMC^{+}20]. -OFF [BSC22]. -phosphate [HHGR21, RCA^{+}21].

-selection [ACPR21]. -TAT1 [RGP^{+}22]. -terminal [SYW^{+}20]. -terminus [RVNS21]. -Tubulin [FOR^{+}20, LSOM23, WM23]. -TuRC [WTU^{+}21].

/calmodulin [YCC^{+}21]. /CK1 [LTL^{+}20].

1 [ANRS^{+}20, AMG^{+}20, CLL^{+}21b, ESH^{+}23, HLB^{+}22, HZN^{+}21, HSU^{+}20, JDKK^{+}22, KST^{+}23, LGL^{+}23, MWF^{+}23, PFPB^{+}20, QZX23, SHBF^{+}20, SSO^{+}20, STY^{+}20, TRJ^{+}20, XHF^{+}20, ZPŠS21, ZAR^{+}21]. 1.2-mediated [IvCD^{+}21]. 1/NBAS [WLW^{+}22]. 1/Rhotekin [YLH^{+}21]. 10/bZIP [LJJ^{+}21]. 11 [HVPM20]. 13 [PPB^{+}21]. 14 [MLQ^{+}21]. 146 [TMG^{+}21]. 170 [HBDC^{+}20]. 19 [CS21b, CS21c, CS21d]. 1C [SV22].

2 [CSG22, CLC^{+}21, HAW^{+}22, KGVK^{+}23, LZZ^{+}21, LSG^{+}22, PCZ^{+}23, SJL^{+}22, WCC^{+}23]. 2/3 [LZZ^{+}21]. 2G [GL20].

3 [BLZ^{+}21, FIK^{+}20, HGG^{+}23, LYL^{+}23, LZZ^{+}21, MVM20, RH23, YLC^{+}21]. 3-dependent [BD20]. 3/4 [MLS^{+}22]. 3/Lrp1 [ZTL^{+}23]. 30 [LMJ^{+}20]. 33 [CM21]. 3D [LQS23, vLEM^{+}20, DES^{+}23, LQS23, MSX^{+}21, SMFC^{+}22]. 3D-Speckler [LQS23].

4 [HRB^{+}21, MLS^{+}22]. 4-kinases [ZLJ^{+}22]. 40S [KPA^{+}20, HGK20, KPA^{+}16]. 43 [DSY^{+}22, GWR^{+}21, HCL^{+}21].

5 [FIK^{+}05, PBPBS22]. 5-phosphatase [DWA^{+}22]. 5/SPG11/SPG15 [HHGR21].

6 [YJX^{+}20, ZLJ^{+}23].

7 [VRSN23, WYL21, WZK^{+}23].

8 [WLW^{+}22].

A1 [YCC^{+}21]. AAA [JBV^{+}20]. Aberrant [FFZ^{+}22, DRW^{+}23, HCWX^{+}22, ICMM20, KSS^{+}20b, KSS^{+}20c]. Abi1 [QLC^{+}20]. ABI2 [JCL^{+}23]. abscission [PZ21]. absence [HESH^{+}22, MPKB^{+}20, Tev20, TG21]. Acb1 [CGBMC20]. ACBD5 [KHB^{+}22]. ACBD5-VAPB [KHB^{+}22]. accelerates [RDW^{+}20, YM21]. accessory [LMM^{+}23]. accomplished [WME22]. according [eSG23]. accretion [RMA21]. Accumulated [LWZ^{+}23]. accumulates [KAS^{+}22]. accumulation [AVC^{+}22, OYJJ23, WZK^{+}23, ZPG^{+}23]. accurate [LZT^{+}23]. acentrosomal [CVT^{+}21]. acetylated [RDL^{+}20]. acetylating [AZR^{+}22]. acetylation [ALPH20, RGP^{+}22]. AChRs [ORCT^{+}20]. acid [ATTF20, TB20a, TTM^{+}21]. acidification [LRM^{+}20]. across [LSX^{+}22]. act [Dor20, Zar20, ZAR^{+}21]. Actin [WH22, ALPH20, BCC^{+}21, BPF^{+}21, BB20, BG22, CFK^{+}22, CJS^{+}21, DJI^{+}21, EYC^{+}20, FLJ^{+}22, GSC^{+}20, Gui21, JCL^{+}23, KBH^{+}22, LAH^{+}21, LDH^{+}21, MTCL^{+}23, MRWK^{+}22, MYM^{+}21, MMDK^{+}22, NVPP20, PMB^{+}22, PKC^{+}22, PLG^{+}23, PMB^{+}20, RBL22, SHBF^{+}20, SHGG21, SJL^{+}22, Sir23, SV22, VFL20, WB20, WPS22, WRG23, YLH^{+}21, YMAS20, MLS^{+}22]. actin-based [PKC^{+}22]. actin-bundling [CJS^{+}21]. actin-independent [WPS22]. actin-membrane [MTCL^{+}23]. actin/mitochondria [APL^{+}21]. action [MNC20]. activate [FAMQW22, FDG^{+}21]. activated [HTL^{+}21, Tai22, ZCD^{+}21]. activates [FCHM20, PKY^{+}20, PZ21, WLM^{+}21, ZRO^{+}23]. activating [GSL^{+}23, PCGB20]. activation [BMS^{+}22, BLU21, CFK^{+}22, DHB^{+}21, HGN^{+}21, HGG^{+}23, IvCD^{+}21, LPMA^{+}22, LGL^{+}23, LGS22, MRH^{+}23, RLK^{+}20, SKX^{+}23, STY^{+}20, TSP21, TRHS23, VCS^{+}22, VGK^{+}21, ZLS^{+}21, ZLJ^{+}23]. activators [SdCS^{+}22]. Active [CLR^{+}20, KYR^{+}22, Tev20, HAL^{+}23]. activity [BED^{+}21, CFD^{+}20, CH22, CSS20, CMN^{+}22, DLZ^{+}20, DHTP22, FHM^{+}22, FOR^{+}20, FLW^{+}23, GCNL21, GDB^{+}20, HDW^{+}21, KLCM^{+}23, KAH^{+}21, LCB^{+}23, LSOM23, MSH^{+}20, MPKB^{+}20, MC21, PZWW21, PGW^{+}21, RRBW^{+}21, SSR^{+}22, TRJ^{+}20, WB20, WZK^{+}23, WCL^{+}23, ZLS^{+}21]. activity-dependent [CH22, HDW^{+}21, MC21, PGW^{+}21]. activity-induced [TRJ^{+}20]. Actomyosin [CH22, BED^{+}21, CHS^{+}22, EJBB^{+}20, FRO^{+}20, KSM^{+}21b, MHGM22, SMS^{+}20, SMC^{+}20, WLM^{+}20, ZGR^{+}22]. acts [KMD20, NR22, WLW^{+}22]. acute [CLL^{+}21b, YCC^{+}21]. acyl [BBP^{+}20, RCF^{+}22]. ADAD2 [XYG^{+}23]. ADAM23 [KGVK^{+}23]. Adaptability [WB21]. adaptation [HCB^{+}23]. adaptations [TWT20, WM20]. adaptor [AHY^{+}21, GSL^{+}23, HH21, KB22, RLS^{+}20, ZSJE20, dAC^{+}22]. adaptors [CJK^{+}22, FC21, SBV^{+}20, WPCB^{+}21]. adapts [BPF^{+}21]. Adding [LC20]. adenomatous [EYC^{+}20]. Adenoviral [DRC^{+}20]. Adherens [OHY^{+}20, PVYJ^{+}21, SMS^{+}20, YKSC^{+}22]. adhesion [BW23, BNV^{+}23, CFV^{+}21, CLH^{+}20, GMB^{+}20, GGFBR^{+}22, HI21, LXJ^{+}23, SGL^{+}23]. adhesions [AKN^{+}22, HAL^{+}23, JKL^{+}22, MMDK^{+}22, RRBW^{+}21, Tan23, WZtM^{+}20]. adhesive [VOR^{+}21]. adipocyte [SHD^{+}21]. adipogenesis [APL^{+}21, EM22, SRUdC^{+}22]. AdoMet [BVYW20]. ADP [CGK^{+}22, KSP^{+}21]. ADP-ribose [KSP^{+}21]. ADP-ribosylation [CGK^{+}22]. adrenal [MND^{+}20]. adult [LTL^{+}20]. Advocating [MP21c]. Afadin [SMS^{+}20]. affinity [CT20]. after [CLL^{+}21b, RCH^{+}20, WMS^{+}21]. against [KKZ^{+}22]. aggregation [VGO^{+}23]. aging [LJJ^{+}21, LTL^{+}20, PHAM^{+}20, RG23, SHGG21]. Agudo [MP23b]. AIF’s [MRG^{+}20]. airway [SCK^{+}19, SCK^{+}23]. AIS [Let20]. AIS-located [Let20]. AKT [MRL^{+}21, CDLZ^{+}22, Smy22]. al [AR20]. ALAL [AMG^{+}20]. ALAL-1 [AMG^{+}20]. Albert [WMA^{+}23]. Algorithms [LZT^{+}23]. ALIX [LMRG20]. ALIX- [LMRG20]. ALK3 [GGFBR^{+}22]. ALK4 [GKM^{+}20]. alleviates [SLL^{+}21, SLL^{+}23, SPT^{+}09, SPT^{+}21]. Allosteric [MRH^{+}23, BJR^{+}21]. allosterically [FCHM20]. allostery [RGP^{+}22]. allow [DES^{+}23, MHN20, SWS^{+}21a]. allows [LRL^{+}20, WRG23]. along [DNVP23, WKX^{+}21]. ALS2 [KKN^{+}21]. alterations [AMG^{+}20, SDD^{+}22]. altering [WXM22]. alternate [CYU^{+}21]. Alternative [MLL^{+}20]. alters [BGM^{+}21, GPEC^{+}23, RGK^{+}22]. amino [ATTF20]. amoeboid [KRH^{+}20]. Amon [VM21]. among [JJ23]. AMPA [CFD^{+}20, GLGL^{+}21]. amphipathic [CT20]. amphisome [KB22]. amphisomes [ZBM^{+}22]. amplification [DSMB20, KVG^{+}20]. amplified [MTD20]. amplifying [WTS^{+}21]. Amyloid [ESH^{+}23]. Ana2 [MSR^{+}20, SWN^{+}22]. Ana2/STIL [SWN^{+}22]. anabolic [ZWJ22]. analyses [KST^{+}21]. Analysis [AMG^{+}20, CLH21, VLdRADJ22, ABM^{+}23, BBPS23, DES^{+}23, LZT^{+}23]. analyzer [LQS23]. Ananthanarayanan [MP21c]. anaphase [DPM^{+}20, SBEB20, SWS^{+}21a, ZVL^{+}23]. anchor [LD21, Mar21, TWY^{+}22]. anchored [AHY^{+}21, CLC^{+}21, CM21, LWZ^{+}23, MOS^{+}22]. anchoring [GCL^{+}21, MTCL^{+}23]. anchors [ARM23, BPF^{+}21, LM21, OMK^{+}22, WAOS^{+}21]. ancient [BD20]. aneuploidy [SRW^{+}21, Ver21]. Angelika [VM21]. angiogenesis [CKM^{+}20, EM20]. Angulin [SFO^{+}21]. Angulin-1 [SFO^{+}21]. anillin [MSC^{+}20]. anillin-like [MSC^{+}20]. animal [KMJ^{+}23]. Anisotropic [BRD^{+}21]. ANKRD24 [KLB^{+}22]. ankyrin [CYL^{+}20]. ankyrin-B [CYL^{+}20]. annexins [FCCH21]. ANO5 [FCCH21]. Antagonism [HPO^{+}23, JMC^{+}20]. anterograde [BS20a]. anti [LGS22]. anti-tumor [LGS22]. antibody [MRG^{+}20, SSHC21, TSL^{+}20]. Antigen [GLM^{+}22, BEM^{+}23, LWG^{+}22]. Antigen-derived [GLM^{+}22]. antioxidant [LRL^{+}20]. antioxidants [CGBMC20]. AP [BLZ^{+}21, HHGR21]. AP-5 [HHGR21]. AP-5/SPG11/SPG15 [HHGR21]. apart [MYM^{+}21]. APC [SGW^{+}20, ZVL^{+}23]. APC-Cdh1 [SGW^{+}20]. APC/C [ZVL^{+}23]. APEX [NGG^{+}20]. APEX2 [TPM^{+}21]. apical [AHvR^{+}20, BBM^{+}23, BRD^{+}21, CH22, HSSK20, LRM^{+}20, MHGM22, OHHR23, RBL22, SLP^{+}22, SCK^{+}19, SCK^{+}23, ZLS^{+}21]. apico [HMT^{+}21]. apico-basal [HMT^{+}21]. apicomplexan [BDT^{+}22]. aPKC [DLZ^{+}20]. APLNR [TJAG^{+}21]. apoptosis [FCT^{+}20, SdCS^{+}22, YSC^{+}02, YSC^{+}21]. apparatus [Bur21, GVD^{+}20a, GVD^{+}20b]. appendages [KRHP^{+}21, VHPP^{+}20]. archetypal [RKLJ22]. architecturally [WRG23]. architecture [MP22g, TWH^{+}21, WJW^{+}22]. area [WB20]. ARF [WDRRF^{+}23, KCP^{+}21]. Arf/Rab [KCP^{+}21]. ARF3 [Cas23a, SFC^{+}23]. Arf6 [OMK^{+}22]. ArfGAP [XGD^{+}23]. Arfs [PBPBS22]. arginine [CYU^{+}21]. Argonaute [ANRS^{+}20]. Argonaute-1 [ANRS^{+}20]. ARHGAP17 [KLCM^{+}23]. ARL13 [DZA^{+}22]. ARL3 [LSX^{+}22]. Arl8b [RCM^{+}23b]. Arp2 [BD20, LYL^{+}23, MVM20]. Arp2/3 [BD20, LYL^{+}23, MVM20]. Arp2/3-dependent [BD20]. arrays [YKSC^{+}22]. arrival [CBC^{+}20]. arsenic [JTM^{+}23]. arsenic-induced [JTM^{+}23]. arsenicals [LL22]. arsenite [LL22]. artifacts [SSHC21]. ARV1 [LWZ^{+}23]. assemble [BTF^{+}20, NR22, SdRVH^{+}21, TNLPF20]. assembled [EYC^{+}20, HAL^{+}23]. assemblies [CLL^{+}21a]. assembling [CS20]. Assembly [WMS^{+}21, AH20a, BZD^{+}21, BVYW20, BP20, BOW^{+}22, CWX^{+}21, CYH^{+}21, CAS23b, DCK^{+}20, FDA21, GKRL^{+}23, Goo20, HESH^{+}22, HGK20, KSWC22, KMW20, LSD^{+}20a, LYL^{+}23, LW20b, MHGM22, MRG^{+}20, PK23, RSB^{+}23, RVNS21, SHLS22, SPS^{+}20, SHGG21, SWT^{+}22, SLH^{+}20b, TOL^{+}20, WMS^{+}20, WTU^{+}21, WLM^{+}21, YKSC^{+}22, ZXW^{+}20, ZFZ^{+}23]. Assessing [CPS^{+}22]. assessment [FBVD^{+}22]. assigning [SHA20]. assisted [FHM^{+}20]. associate [CBC^{+}20, KPA^{+}16, KPA^{+}20]. associated [AKN^{+}22, BZD20, BOW^{+}22, CWKP23, CKM^{+}20, GPEC^{+}23, HJL^{+}22, PHAM^{+}20, RBL22, RLK^{+}20, SBV^{+}20, SWS21b, TG21, YLH^{+}21, ZXY^{+}23]. association [ALC^{+}20, KVG^{+}20, RRBW^{+}21, ZAK^{+}22]. Astral [DdCVT22, ZVL^{+}23]. astrocyte [BC23, CPS^{+}22, LWL^{+}23]. astrocytes [Bez22, IMR^{+}23]. Asymmetric [DCK^{+}20, MDV^{+}21, WM23]. asymmetries [MMKM21]. ATFS [LGL^{+}23]. ATFS-1 [LGL^{+}23]. Atg13 [FAMQW22]. ATG16L1 [FWP^{+}20]. ATG16L1-WD40 [FWP^{+}20]. ATG2 [BBPS23, DTG23]. Atg39 [CMT^{+}21, MOK^{+}22]. Atg7 [LWL^{+}23]. ATG8 [JMKS^{+}23]. ATG8-dependent [JMKS^{+}23]. atg8ylation [JWB^{+}22, CNL^{+}21]. ATG9 [BBPS23, OWY^{+}23, OTOF21]. ATG9A [CDD^{+}22, YKK^{+}20]. atlastin [BSC^{+}23, CMN^{+}22, KBN^{+}21, LZZ^{+}21]. atlastin-1 [KBN^{+}21]. atlastin-3 [BSC^{+}23]. atlastins [JMY^{+}23]. ATM [HCB^{+}23, PZ21]. ATPase [FWP^{+}20, HJL^{+}22, MAW^{+}22, RLV^{+}20]. ATPase/TORC1 [LGL^{+}23]. ATR [VZQ^{+}21]. ATR-mediated [VZQ^{+}21]. atrophy [HGG^{+}23, RH23]. attachment [DKCT21, SSZL21]. attachments [ARCM20, GOR^{+}20, SKN^{+}21]. attack [MP23b]. attention [Tar21]. augmented [WBR^{+}20]. Aurora [BDD20, CRZ^{+}21, DKCT21, HHT^{+}20, INM^{+}21, LZC^{+}20, PKY^{+}20, PRB^{+}20, PCGB20, SBEB20, SKS^{+}23, TSP21, ZBY^{+}21]. autocrine [KIV^{+}20]. autoinhibition [ALC^{+}20, CMN^{+}22, QZX23]. Autologous [JFM^{+}22]. Autolysosomal [RCM^{+}23a]. automated [GMD^{+}23, LSS^{+}23, LQS23]. autonomous [CSS20, NPdC^{+}21, SIP^{+}23]. autophagic [AAF^{+}20, KAH^{+}21, ZPG^{+}23, ZXY^{+}23]. Autophagosome [MLS20, BBPS23, CCV^{+}21, CMT^{+}21, DTG23, OTOF21]. autophagosomes [OWY^{+}23, hYKO^{+}20a, hYKO^{+}20b, hYKO^{+}21, ZBM^{+}22]. Autophagy [CLL^{+}21b, GG20, TWT20, AT21, AAF^{+}20, Alm21, BZC^{+}21, BBPS23, CDD^{+}22, DSY^{+}22, EZB^{+}20, FAMQW22, FCHM20, HJL^{+}22, JKZ^{+}22, JMKS^{+}23, KJ23, LZZ^{+}21, MOK^{+}22, NWZ20, NSB^{+}21, RKLJ22, RZN^{+}22, SYW^{+}20, SNL^{+}22, TKK^{+}20, WCG^{+}22, XZJ^{+}21, ZLW23]. autophagy-lysosome [WCG^{+}22, XZJ^{+}21]. Auxilin [HSU^{+}20]. averaging [TML22]. Avinoam [MP22e]. Axin [BP22, NKS^{+}21]. axis [BMS^{+}22, CW23, KKPH^{+}21, MRL^{+}21, MdCT23, SLS^{+}23, SPKP22, SMC^{+}20, YKK^{+}20, ZTL^{+}23]. axon [AH20a, BMM^{+}20, CYL^{+}20, FPMS^{+}21, KMD20, LPMA^{+}22, MPKB^{+}20, NBI^{+}22, TOL^{+}20, WKX^{+}21, YMAS20]. axonal [AVC^{+}22, CCV^{+}21, Hök22, LPMA^{+}22, SHBF^{+}20, WLM^{+}20]. axoneme [GVA20]. axonostasis [RCS22]. axons [BS20a, FSC22, KGVK^{+}23, Pro20].

B [BDD20, CYL^{+}20, CRZ^{+}21, DKCT21, DWA^{+}22, HHT^{+}20, HKK^{+}20, LZC^{+}20, PKY^{+}20, SBEB20, WH22]. B1 [DOA^{+}22, HLGD20, JMB^{+}20]. B1-Cdk1 [JMB^{+}20]. B56 [BZD^{+}21]. back [Hic22]. back-up [Hic22]. bacterial [PMB^{+}22]. BAF [KAS^{+}22]. balance [Cas23a, LGB^{+}21]. Balancing [BCdS22]. band [SGN^{+}20]. BAP1 [YLH^{+}22]. barbed [Gui21, SHGG21, WRG23]. barrel [SMM^{+}21]. barrier [CPC^{+}20, CHS^{+}22, GNL^{+}20, HSF^{+}23, LWL^{+}23, PL22, SCK^{+}19, SCK^{+}23, VRSN23]. basal [CVT^{+}21, GVA20, HMT^{+}21, McW23, RSWP20, SvDSW^{+}20]. base [FDG^{+}21]. based [ABM^{+}23, GGFBR^{+}22, KSM^{+}21a, LWG^{+}22, LYL^{+}22, MTCL^{+}23, MLS^{+}22, PKC^{+}22, RCA^{+}21, SHA20, YSR^{+}21]. basement [GKRL^{+}23]. basis [AGH^{+}22, SBEB20]. basket [KWV^{+}23, VV23]. BBSome [DZA^{+}22, LSX^{+}22, SNN20]. BBSome-dependent [DZA^{+}22]. BBSome-mediated [SNN20]. BDNF [RH23]. BDNF-TrkB [RH23]. beating [NYN^{+}21]. before [SS22]. behavior [EM20]. being [MP23a]. bend [MTW^{+}23]. bent [BC23]. best [KRC^{+}22]. between [BG21, CL21, DSG21, HRB^{+}21, HPO^{+}23, JMC^{+}20, KST^{+}21, KMK21, LD21, LLLR20, LRL^{+}20, LLW^{+}21, SNYA^{+}21, VOR^{+}21, XGD^{+}23, ZXW^{+}20]. Beware [MP23b]. beyond [BLU21, MP21b, ZWJ22]. bi [KMW20]. bi-orientation [KMW20]. BICD [SBV^{+}20]. Bidirectional [ZCD^{+}21, CGCR^{+}22, CBC^{+}20, YM21]. bilayer [HYX^{+}20, SMM^{+}21, ZY21]. bile [BBM^{+}23, CG21]. Bin1 [LLX^{+}21]. Binding [FPMS^{+}21, BJSOS^{+}20, BJSOS^{+}21, BWA^{+}23, GOR^{+}20, JKZ^{+}22, KSS^{+}20a, KKP^{+}21, KSP^{+}21, PGD^{+}20, TTM^{+}21, WAK^{+}20]. binds [FER^{+}23, LKMM^{+}23, RCM^{+}23b]. Biochemical [WTU^{+}21, AHLR22]. bioengineered [CPS^{+}22]. biogenesis [AANLL^{+}20, BWK^{+}21, CCFN^{+}20, CYR^{+}21, CWX^{+}21, CEM^{+}20, CM21, DTG23, EEW^{+}22, ESX^{+}20, GPL^{+}21, GMCO^{+}22, JGN^{+}20, KB22, LC20, LM23, MLS20, NPdC^{+}21, PTS^{+}22, RLV^{+}20, SJL^{+}22, WHN^{+}21, XYG^{+}23, YW21, YJX^{+}20, ZJDR22]. biology [Dri20, LVMFL20, O’D20a, PGDD21, SSZL21, SH20, WM20]. biomedical [GPES21]. biomolecular [GMC^{+}20, SCB^{+}20]. biorientation [SWS^{+}21a]. biosensor [MVM20]. biosynthesis [HSSK20, LWZ^{+}23]. biosynthetic [KSS^{+}20a]. biotinylation [CLH^{+}20]. BiP [AAR^{+}21]. BiP-mediated [AAR^{+}21]. bipolar [CYH^{+}21]. bipolarity [GNL^{+}20]. birth [MRA20]. bistable [YPM^{+}21]. bleb [RCA^{+}21]. bleb-based [RCA^{+}21]. BLOC [BLZ^{+}21, JDKK^{+}22]. BLOC-1 [JDKK^{+}22]. BLOC-1-AP-3 [BLZ^{+}21]. blood [LWL^{+}23]. BLT1 [SMC^{+}20]. BMPRII [GGFBR^{+}22]. BNIP3 [GCW^{+}23]. BNIP3/BNIP3L [GCW^{+}23]. BNIP3/BNIP3L-mediated [GCW^{+}23]. BNIP3L-mediated [GCW^{+}23]. Bo [MP22a]. bodies [RSWP20, XYG^{+}23]. body [GVA20, MdCT23, MP22e, RVNS21, SvDSW^{+}20]. Böke [MP22d]. bone [BCS^{+}21, ZTL^{+}23]. boost [CW23]. boosts [LLK^{+}22]. Border [MGM22, BCC^{+}21, Köh21]. Borealin [WDJ^{+}21]. BORG3 [FRO^{+}20]. bound [HZN^{+}21, KMK21, LHS^{+}22, PZWW21, SKN^{+}21]. boundary [SNYA^{+}21]. bovine [SdRVH^{+}21]. brain [AR20, BWEHS21, GWR^{+}21, KNiY^{+}21, LLC^{+}20, LWL^{+}23]. branch [NBI^{+}22]. Branched [EYC^{+}20, KBH^{+}22, SV22]. branches [WKX^{+}21]. branching [CYL^{+}20, ZHHJ22]. BRCA1 [JFM^{+}22]. BRCA1-A [JFM^{+}22]. BRCA2 [DMR^{+}20]. BRCA2-deficient [DMR^{+}20]. Brd4 [DHB^{+}21]. breakage [RDW^{+}20]. breaks [KMJ^{+}23]. Breakthrough [VM21]. breast [ASK^{+}22, FFZ^{+}22, SPS^{+}20, TMG^{+}21]. bride [Ver21]. bridge [CL21]. bridges [GSLH^{+}21]. bridging [KB22, TRHS23]. brings [Dri20]. Brinkley [GPES21]. broad [WPCB^{+}21]. broad-spectrum [WPCB^{+}21]. broken [CBJ^{+}21]. BTLA [XHF^{+}20]. Bub1 [CML20, HHT^{+}20]. Bub3 [CML20]. BubR1 [HGN^{+}21]. Bud1 [WPS22]. budding [CWN^{+}23, DNVP23]. buffering [MP22c]. Building [Goo20]. builds [KRHP^{+}21]. bulk [MHN20, MC21, PGW^{+}21]. bulkheads [BBM^{+}23, BRD^{+}21]. bundling [CJS^{+}21, FLJ^{+}22, NYN^{+}21]. bypasses [CSD22, SHGG21]. bZIP-mediated [LJJ^{+}21].

C [ABB^{+}22, RFL20, YSC^{+}21, BZC^{+}21, CMN^{+}22, KAS^{+}22, SYW^{+}20, ZVL^{+}23, YSC^{+}02]. C-ferroptosis [ABB^{+}22]. c-Myc-induced [YSC^{+}21]. C-shaped [SYW^{+}20]. C-terminal [BZC^{+}21]. C. [CSG22, DPM^{+}20, HČK^{+}20, JBV^{+}20, LGL^{+}23, LMJ^{+}20, RCH^{+}20, TP20]. C17iso [ZHW^{+}21]. C9orf72 [ATTF20, CYU^{+}21]. Ca [BS20b, IvCD^{+}21, LLK^{+}21, LPMA^{+}22, PMSO^{+}23, SIP^{+}23, YCC^{+}21]. Cab45 [HBS^{+}20]. Cactin [MGM22]. cadherin [EM20, GMIC^{+}20, HVPM20, HMT^{+}21, SFC^{+}23]. cADPR [Hök22, LPMA^{+}22]. Calcineurin [ZSJE20, PGW^{+}21]. Calcineurin-dependent [ZSJE20]. calcium [BRB^{+}20, CW23, GKFR20, Hök22, LYS^{+}20, MBG^{+}23, VCS^{+}22, VOR^{+}21, ZCD^{+}21]. calcium-independent [BRB^{+}20]. Calcoco1 [Yam21, NSB^{+}21]. CALCOCO1-mediated [NSB^{+}21]. calibration [LQS23]. calmodulin [YCC^{+}21]. calreticulin [VGK^{+}21]. calreticulin-mediated [VGK^{+}21]. cAMP [SHLS22]. CAMSAPs [CVT^{+}21]. canalicular [CG21]. Cancer [CKR^{+}20, ASK^{+}22, AMG^{+}20, BDR20, BDS^{+}21, BW20, Cas23a, FFZ^{+}22, GPEC^{+}23, HPO^{+}23, SDD^{+}22, MTD20, MP22f, O’D20a, RCDMM20, SMFC^{+}22, SKF^{+}23, SPS^{+}20, TMG^{+}21, TWT20, TG21]. cancer-associated [GPEC^{+}23]. cancers [MOS^{+}20, VGK^{+}21]. canonical [AT21, Cas22, HJL^{+}22, JKZ^{+}22, LSD^{+}20a]. Can’t [MMC20]. capacity [LFD^{+}21, OZW^{+}21]. Cappin [Sir23]. capping [BWA^{+}23, Sir23]. Caprin1 [KPA^{+}16, KPA^{+}20]. Captive [MP22a]. captures [CMT^{+}21]. carbon [RCDMM20]. carboxyl [RCF^{+}22]. Cargo [GNML^{+}20, AANLL^{+}20, CWKP23, DZA^{+}22, DF22, EEW^{+}22, HH21, QZX23, RCM^{+}23b, SNYA^{+}21, SBV^{+}20, TEH^{+}20, WAK^{+}20, XGD^{+}23, ZLJ^{+}22]. cargo-selective [CWKP23]. cargoes [ARCM20]. carrier [LML^{+}21]. carriers [BLZ^{+}21]. CARTS [WHN^{+}21]. cartwheel [CWX^{+}21]. Cas12a [FHM^{+}20]. cascade [KCP^{+}21]. cascades [ZMMM^{+}20]. Caspase [HTL^{+}21, EE22]. caspase-3 [EE22]. cassette [GPEC^{+}23]. catalyst [BSC^{+}23]. catalytic [RGP^{+}22]. Catalytically [CBS^{+}21]. catastrophe [FAHZ21, VZQ^{+}21]. catenin [BP22, BJPH^{+}20, HMT^{+}21, MOS^{+}20, NKS^{+}21, SMS^{+}20, MMDK^{+}22, SGL^{+}23, GL20, vdGM22]. cathepsin [ZLJ^{+}23]. cationic [ATTF20]. caught [Dor20]. cause [ITB^{+}23]. caused [Hök22]. causes [KYR^{+}22, KNiY^{+}21]. causing [FFZ^{+}22]. caveola [LMM^{+}23]. Caveolae [PKA20, ZAR^{+}21]. caveolin [AANLL^{+}20, ZAR^{+}21]. caveolin- [AANLL^{+}20]. Caveolin-1 [ZAR^{+}21]. cavin1 [ZAR^{+}21]. Cavin4 [LLX^{+}21]. CBP [WJL^{+}23]. CBX2 [BZD20]. CD2AP [WB20]. CD4 [MWF^{+}23]. CD47 [SLS^{+}23]. CD47-QPCT [SLS^{+}23]. CD47-QPCT/L [SLS^{+}23]. CD8 [BMS^{+}22]. Cdc14 [FAMQW22]. Cdc20 [HGN^{+}21, ZVL^{+}23]. Cdc20-mediated [ZVL^{+}23]. Cdc31 [RVNS21]. Cdc42 [KLCM^{+}23, ZMMM^{+}20, FBR^{+}21, GCNL21, GC22, LD20, RLK^{+}20, WZZ^{+}23]. CDC42EP5 [FRO^{+}20]. CDC42EP5/BORG3 [FRO^{+}20]. Cdh1 [SGW^{+}20]. Cdk [SWN^{+}22, YPM^{+}21]. Cdk/Cyclin [SWN^{+}22]. Cdk/Cyclin-dependent [SWN^{+}22]. CDK1 [HSL^{+}20, JMB^{+}20]. CDK4 [YJX^{+}20]. CDK4/6 [YJX^{+}20]. CDK5RAP2 [WMS^{+}20]. CDKA [STY^{+}20]. CDKD [STY^{+}20]. CDKD-dependent [STY^{+}20]. Cdr2 [OMK^{+}22]. Cdt1 [RCA^{+}23]. Cell [AMMK^{+}22, FSZ^{+}22, HW22, MBG^{+}23, Ove21, SPRWB20, SH20, ASK^{+}22, ABB^{+}22, AMFW^{+}21, AR20, AHvR^{+}20, AO20, BCC^{+}21, BCdS22, BDH^{+}21, BS20b, BEM^{+}23, BWA^{+}23, Bog21, BMS^{+}22, BNV^{+}23, BKR^{+}22, CFV^{+}21, CDD^{+}22, CNL^{+}21, Cas22, CKR^{+}20, CLL^{+}21a, CKM^{+}20, DCK^{+}20, DYW^{+}20, DWA^{+}22, DHTP22, DRC^{+}20, EE22, EM20, FBR^{+}21, FFZ^{+}22, FIK^{+}05, FIK^{+}20, GMC^{+}20, GCNL21, GGFBR^{+}22, GLM^{+}22, HI21, HDG22, HGG^{+}23, HRS^{+}20, JIBK23, KLC^{+}20, KSM^{+}21b, KBH^{+}22, KNiY^{+}21, LCM22, LMS^{+}21, LDE^{+}22, LAH^{+}21, LXJ^{+}23, LDH^{+}21, LVMFL20, MS20, MHS^{+}20, MSB^{+}21, SDD^{+}22, McC21, MGM22, MdCT23, ME21, MP21a, MP22i, MMC20, MMKM21, NYN^{+}21, NTA^{+}21, NMO^{+}22, O’D20a, OMI22, OMK^{+}22, OHY^{+}20, PGDD21, PWW^{+}20, Ped22, POL^{+}20, PAS^{+}22, RWSZ^{+}20, RG23, RS22, RSWP20, SIP^{+}23, Sea21, SMFC^{+}22, SKF^{+}23, SLES20, SGL^{+}23]. cell [SLP^{+}22, SSZL21, SBBJ21, TNC^{+}20, Tev20, TMG^{+}21, TWT20, WB20, WDL^{+}20, WXM22, WH22, WM20, WHE^{+}22, XHF^{+}20, XVW^{+}23, YMH^{+}20, YLC^{+}21, ZLS^{+}21, ZPŠS21, ZWJ22, ZAK^{+}22, vLEM^{+}20]. cell-to-cell [BS20b]. cells [ACPR21, BDR20, BHS^{+}21, BG21, BCS^{+}21, BDD20, CKR^{+}20, CZTL21, CVT^{+}21, Dri20, DMR^{+}20, DLK^{+}21, Dus21, FLW^{+}23, GNL^{+}20, HZN^{+}21, HYL^{+}20, JRGH21, KPM^{+}22, Kin21, KRH^{+}20, LHL^{+}23, LD21, LW20a, LYL^{+}23, LWG^{+}22, MHN20, MTCL^{+}23, MWF^{+}23, MND^{+}20, MA20, MP22f, MP22g, MMC20, MSX^{+}21, PDW^{+}20, SLS^{+}23, SPKP22, SCB^{+}20, STS21, Tai22, TG21, UZS^{+}23, VZQ^{+}21, WHE^{+}22, WBH^{+}21, YSR^{+}21, ZVM^{+}20, dCS^{+}21]. Cellular [MP22h, BGM^{+}21, FMY^{+}21, GMD^{+}23, GH20, KSM^{+}21b, KRH^{+}20, LJT^{+}22, MLQ^{+}21, NBI^{+}22, PK23, PKA20, TRHS23, VTL^{+}20, WBR^{+}20, VGO^{+}23]. Cellulose [WCL^{+}23]. CENP [ARCM20, MSJ20, SRW^{+}21]. CENP-A [MSJ20]. CENP-F [ARCM20]. center [O’D20a]. centers [CLR^{+}20]. central [HESH^{+}22, KNA^{+}22, RCA^{+}23, SBEB20, ZBM^{+}22]. Centralspindlin [DNVP23]. centrin [RVNS21]. Centriole [CYH^{+}21, SWN^{+}22, WMS^{+}20, CWX^{+}21, GGA21, HLB^{+}22, IWI^{+}21, KNA^{+}22, LNY^{+}22, NPdC^{+}21, PKD^{+}20, PSC^{+}20, SYQ^{+}22, TWH^{+}21, VHPP^{+}20, VDC^{+}20]. Centriole-independent [WMS^{+}20]. centrioles [KSS^{+}20c, KSS^{+}20b]. Centromere [LZC^{+}20, AGH^{+}22, BDD20, CD21]. Centromere-localized [LZC^{+}20]. centromeres [DCK^{+}20]. Centromeric [CZTL21, FOR^{+}20, WLM^{+}21]. Centrosome [Zar20, AHQ20, MTR^{+}20, OZW^{+}21]. centrosome-linker [AHQ20]. Centrosome-localized [Zar20]. centrosomes [RFL20, SdRVH^{+}21, VDC^{+}20, WHE^{+}22]. CEP164C [ATS^{+}21]. CEP192 [CYH^{+}21]. CEP350 [KNA^{+}22]. CEP55 [ZBY^{+}21]. Cep57 [IWI^{+}21]. Cep57L1 [IWI^{+}21]. CEP97 [LNY^{+}22]. ceramides [LKMM^{+}23]. cerevisiae [FDA21]. CFTR [HVPM20]. cGAS [KAS^{+}22]. CGRP [LYS^{+}20, GKFR20]. chain [MLL^{+}20, RCF^{+}22, TSL^{+}20]. chains [Ike20, SNN20]. Chan [MP23a]. chance [O’D22]. change [BSC22]. changes [KYR^{+}22, KHV^{+}22, MTR^{+}20, RLK^{+}20]. channel [LLLR20, VOR^{+}21, ZCD^{+}21]. channeling [PKH^{+}20]. channels [KGVK^{+}23, WLBS20]. chaperone [AAR^{+}21, EZB^{+}20]. chaperone-mediated [EZB^{+}20]. Chaperoning [ZY21]. characterization [TWH^{+}21]. characterizes [STvT23]. CHC22 [CCFN^{+}20]. checking [MS23]. checkpoint [ACPR21, BP20, CSS20, HL21, JMB^{+}20, PKY^{+}20, PZ21, WLM^{+}21]. chemogenetic [FHM^{+}22]. Chii [MP23a]. chimeric [BEM^{+}23]. Chk2 [PZ21]. Chlamydomonas [DZA^{+}22, LLW^{+}20]. chloride [WZK^{+}23, ZLJ^{+}23]. Chm7 [TTM^{+}21]. CHMP2B [DSY^{+}22]. CHMP7 [PSS^{+}20]. cholerae [JKZ^{+}22]. Cholesterol [LSG^{+}22, WZG22, JKZ^{+}22, LWD^{+}21, LHS^{+}22, WHN^{+}21]. cholesterol-binding [JKZ^{+}22]. cholinergic [ZVC^{+}21]. chorein [HSW^{+}22]. chromatid [RDL^{+}20]. chromatin [BCWM21, MSJ20, ME21, PSP^{+}21, SBEB20]. chromosomal [BZD20, RDW^{+}20]. Chromosome [INM^{+}21, SDD^{+}22, TP20, BDT^{+}22, CML20, CBJ^{+}21, CSOG^{+}20, FDSR22, KTT^{+}22, LZC^{+}20, MS23, PCGB20, SPL^{+}20, SWT^{+}22, WLBS20]. Chromosomes [GNL^{+}20, DG22, FDSR22, MP22h, MYM^{+}21, SPRWB20, Tev20, WDJ^{+}21]. CI [RCM^{+}23b]. CI-M6PR [RCM^{+}23b]. Cilia [BC23, DCRDC^{+}22, DSG21, FY20, GSC^{+}20, GVA20, LLW^{+}20, MND^{+}20, SNN20, SCL^{+}21, ZBY^{+}21, DZA^{+}22]. Ciliary [SvDSW^{+}20, FDG^{+}21, IMR^{+}23, LSX^{+}22, LSD^{+}21, MKD^{+}21, NYN^{+}21]. ciliogenesis [AT21, KRHP^{+}21, PKD^{+}20, PRB^{+}20, SYQ^{+}22]. ciliopathy [KRHP^{+}21]. cilium [SKX^{+}23, SIP^{+}23]. circadian [KKPH^{+}21]. circuit [MLS^{+}22, WJL^{+}23]. cis [BLZ^{+}21, LKMM^{+}23]. cis-Golgi [LKMM^{+}23]. cis-SNARE [BLZ^{+}21]. Citrullination [GSB^{+}20]. CIZ1 [SWT^{+}22]. CK1 [LTL^{+}20, DSY^{+}22]. clarifies [LJT^{+}22]. CLASP2 [GOR^{+}20]. Class [BJSOS^{+}20, BJSOS^{+}21, EZB^{+}20, LLBC^{+}20]. Clathrin [CJK^{+}22, MTW^{+}23, Smy22, CDLZ^{+}22, CSD22, CCFN^{+}20, CS20, CMM^{+}20, HAL^{+}23, HSU^{+}20, KBB^{+}23, MLL^{+}20, PHMD20, RLS^{+}20]. clathrin-coated [CDLZ^{+}22, CS20]. clathrin-dependent [KBB^{+}23]. clathrin-mediated [CMM^{+}20, HSU^{+}20, PHMD20]. claudin [HSF^{+}23, VRSN23]. claudin-7 [HSF^{+}23, VRSN23]. claudins [SFO^{+}21]. Clb4 [ZVL^{+}23]. ClC [WZK^{+}23]. ClC-7 [WZK^{+}23]. Cleaning [SMK20]. clearance [ICMM20]. cleavage [SRK22]. CLEM [LSS^{+}23]. CLH [ZLJ^{+}23]. CLH-6 [ZLJ^{+}23]. CLIP [HBDC^{+}20]. CLIP-170 [HBDC^{+}20]. Clipping [VRSN23]. clocks [GH20]. close [DG22, HL21]. Closer [GY20]. Closing [vdGM22]. closure [MLS20, PSS^{+}20]. cluster [KGVK^{+}23]. clusterin [SMK20]. clustering [DWA^{+}22, LSG^{+}22, ORCT^{+}20, ZMW^{+}22]. clusters [BTF^{+}20]. Cnm1 [EBZC^{+}21, CL21]. coactivator [ANRS^{+}20]. coat [SNYA^{+}21, SLH^{+}20b]. coated [CDLZ^{+}22, CS20, MLL^{+}20, Smy22]. coats [MTW^{+}23]. code [ALPH20]. coenzyme [BBP^{+}20]. cofactor [KJ23]. coherence [KRH^{+}20]. cohesin [RDL^{+}20, SPL^{+}20]. cohesion [CZTL21, PAS^{+}22, RDL^{+}20]. COL17A1 [NTA^{+}21]. coli [EYC^{+}20]. collaboration [LRL^{+}20]. collagen [GKRL^{+}23]. collapse [OCLB21]. collar [SNYA^{+}21]. collateral [CYL^{+}20, CW23]. collecting [BED^{+}21]. Collective [KIV^{+}20, BCC^{+}21, EM20, MGM22]. colocalization [VLdRADJ22]. colon [TG21]. color [VLdRADJ22, VVW^{+}23]. colorectal [BDS^{+}21, MOS^{+}20, SKF^{+}23]. combinations [DJI^{+}21]. Combinatorial [SLH^{+}20b]. combine [AH20b]. comes [MP22c]. comfort [DG22]. common [WDL^{+}20, WM20]. Communicating [LVMFL20]. compaction [BCWM21, ME21]. compartment [AFB^{+}20, ESX^{+}20, SIP^{+}23]. compartmentalization [SDD^{+}22, SRK22]. compartmentalizes [GGFBR^{+}22]. compartments [CCFN^{+}20, FSC22, KMK21, ZFZ^{+}23]. compensation [KKZ^{+}22]. compete [BMM^{+}20]. competence [HGK20]. Competition [LLW^{+}21, BGM^{+}21, Sir23, WRG23]. complements [RWSZ^{+}20]. complete [GSP^{+}20]. Completion [WKX^{+}21]. complex [ATTF20, BZD^{+}21, BLZ^{+}21, GSL^{+}23, GBBT^{+}22, Goo20, HHGR21, HLGD20, HTL^{+}21, HHD^{+}20, HČK^{+}20, JFM^{+}22, KKZ^{+}22, KSWC22, KWV^{+}23, KMW20, KRHP^{+}21, KST^{+}22, LM23, LDH^{+}21, LZZ^{+}21, LW20b, MP23c, MRG^{+}20, NKS^{+}21, OYJJ23, OHHR23, PSA^{+}23, PHT^{+}23, RCA^{+}23, SKX^{+}23, SLL^{+}21, SLL^{+}23, SYW^{+}20, SLP^{+}22, Tar21, WKC^{+}22, WBH^{+}21, YLC^{+}21, ZXW^{+}20]. complexed [HSF^{+}23]. complexes [CPC^{+}20, CLH21, CLH^{+}20, CWAT20, FPZ^{+}22, GG20, KPA^{+}16, KPA^{+}20, KST^{+}21, MHGM22, PCZ^{+}23, SGN^{+}20, TKK^{+}20, YM21]. complexity [RCDMM20]. component [BCM^{+}22, GSL^{+}23, MGM22, PSA^{+}23]. components [BS20a, HČK^{+}20, ZPŠS21]. Composition [BDT^{+}22, BW20, GM23, JMY^{+}23, WBH^{+}21, YZW^{+}20]. comprise [OWY^{+}23]. comprising [WJL^{+}23]. compromises [IMR^{+}23]. Computational [KST^{+}21]. concentrates [RCF^{+}22]. condensate [CLL^{+}21a, MKO^{+}21]. condensate-organized [MKO^{+}21]. condensates [GMC^{+}20, SCB^{+}20, WCG^{+}22, ZPG^{+}23]. condensation [KPA^{+}16, KPA^{+}20]. condensin [KTT^{+}22, PCGB20]. condensin-dependent [PCGB20]. condensing [SPRWB20]. conditional [VLdRADJ22]. conditions [SHGG21]. cone [SHBF^{+}20]. confined [SWS21b]. Conformational [dAC^{+}22, BSC22, SMM^{+}21]. Confounding [WHA20]. congression [PCGB20]. connections [SvDSW^{+}20]. connects [TMG^{+}21]. connexin [KIV^{+}20]. connexin-43 [KIV^{+}20]. connexins [LRL^{+}20]. Consensus [BOW^{+}22, Bur21]. Conserved [BVYW20, VFL20, AAF^{+}20, FER^{+}23, GDB^{+}20, SSR^{+}22]. constitute [AHvR^{+}20]. constitutive [BSC^{+}23, PSA^{+}23]. constrain [WB20]. constricting [SCN^{+}23]. constriction [BJAR^{+}21, CH22]. construction [WRG23]. contact [AGW^{+}20, AO20, BCS^{+}21, BCM^{+}22, CCH^{+}21, DCG^{+}23, EBZC^{+}21, FC21, FIK^{+}05, FIK^{+}20, GMCO^{+}22, KSN^{+}22, KWdB^{+}20, KST^{+}22, LAH^{+}21, LYL^{+}22, MND^{+}20, McC21, MdCT23, PWW^{+}20, SV22, TNLPF20, VBG^{+}22, WHN^{+}21, dDFGP^{+}21]. contact-induced [FIK^{+}05, FIK^{+}20]. contacts [BEM^{+}23, DZA^{+}20, ESX^{+}20, KSM^{+}21b, KHB^{+}22, KMK21, LW20b, SvVV^{+}23, SFO^{+}21, WR22, WYL21]. contain [PDW^{+}20]. containers [SNYA^{+}21]. containing [CJS^{+}21, RPM^{+}21]. contains [GPEC^{+}23]. content [YSR^{+}21]. context [SH20]. contexts [WDL^{+}20]. contractile [BJAR^{+}21, MBA^{+}22, MHN20, NR22, SCN^{+}23]. contractility [EJBB^{+}20, KSM^{+}21b, TNC^{+}20, WLM^{+}20, ZGR^{+}22]. contraction [KST^{+}23, vLEM^{+}20]. contributes [GNML^{+}20, HKK^{+}20, KTT^{+}22, LKMM^{+}23, MSB^{+}21, MLL^{+}20]. contribution [HHT^{+}20, SNP^{+}22]. control [Alm21, BSC22, BLU21, BHK20, CSG22, CFV^{+}21, CVMB^{+}23, CSS20, DDD^{+}20, DLZ^{+}20, EM22, FHM^{+}22, FBR^{+}21, GGFBR^{+}22, HGK20, HH21, KRH^{+}20, LAH^{+}21, MSC^{+}20, MBA^{+}22, ML22, MSJ20, MMSP20, NvGK20, OMK^{+}22, PK23, PSS^{+}20, SBBJ21, TSL^{+}20, TB20a, ZGR^{+}22]. controlling [APL^{+}21, MP22i]. controls [BHS^{+}21, CDD^{+}22, CFD^{+}20, CHPF^{+}21a, CHPF^{+}21b, CG21, DOA^{+}22, EJBB^{+}20, EM20, FGBD^{+}21, GDB^{+}20, GMIC^{+}20, HSW^{+}22, KCP^{+}21, KAH^{+}21, LWG^{+}22, MBG^{+}23, OZW^{+}21, PRB^{+}20, PAS^{+}22, RZN^{+}22, SFC^{+}23, SCK^{+}19, SCK^{+}23, STY^{+}20, SSF^{+}22, VDC^{+}20, WAK^{+}20, ZJH22, ZTL^{+}23]. converge [AGW^{+}20]. Convergence [ZXY^{+}23]. conversion [BHS^{+}21, VBG^{+}22]. cooperation [CLC^{+}21]. cooperative [SLH^{+}20b]. cooperatively [CYH^{+}21]. coordinate [BB20, LLK^{+}22, MRWK^{+}22, PGW^{+}21]. Coordinated [SRUdC^{+}22, NYN^{+}21]. coordinates [BMS^{+}22, GKM^{+}20, HI21, HDG22, HMSF22, KNA^{+}22, LXJ^{+}23, MDV^{+}21]. Coordinating [AR20]. coordination [LKW^{+}21, SLES20]. COP [XGD^{+}23]. cope [CNL^{+}21]. COPI [WPCB^{+}21]. COPII [GNML^{+}20, JKL^{+}22, SNYA^{+}21, SLH^{+}20b]. copy [AMG^{+}20]. core [CDD^{+}22, TWH^{+}21, ZVC^{+}21]. Coro1B [KBH^{+}22]. Coro1C [KBH^{+}22]. corona [ARCM20]. coronavirus [JLS^{+}22]. Coronin [SV22]. corpse [SLS^{+}23]. Correction [BJSOS^{+}21, CHPF^{+}21a, Col22a, FIK^{+}20, GVD^{+}20a, KPA^{+}20, KSS^{+}20b, MYK^{+}21, MYK^{+}22, SLL^{+}23, SPT^{+}21, SCK^{+}20a, SCK^{+}23, hYKO^{+}20a, hYKO^{+}21, YSC^{+}21, DKCT21, FOR^{+}20]. correlative [vdBdHLK22]. cortex [HW22, LYL^{+}23, MSB^{+}21, OMK^{+}22]. Cortical [vLEM^{+}20, BG22, yLHW^{+}20, DYW^{+}20, GKM^{+}20, IHBP^{+}23, McC21, MLS^{+}22, OCLB21, SvDSW^{+}20]. corticogenesis [HYL^{+}20]. cotransport [BS20a]. countertransport [KSN^{+}22]. couple [SBV^{+}20]. Coupled [dDFGP^{+}21, SKF^{+}23, TJAG^{+}21]. couples [DYW^{+}20, EJBB^{+}20, HSL^{+}20, PGD^{+}20]. Coupling [HGN^{+}21, MMSP20, NGG^{+}20, BG22, FPZ^{+}22, RCA^{+}23]. CoV [SCK^{+}20a, WCC^{+}23, MNvdS^{+}20, SCK^{+}20b]. covalently [UTR^{+}23]. COVID [CS21b, CS21c, CS21d]. COVID-19 [CS21b, CS21c, CS21d]. CP110 [SYQ^{+}22]. CPC [AGH^{+}22, WDJ^{+}21]. CPT1C [CFD^{+}20]. CRAC [ZCD^{+}21]. crawls [BD20]. create [CBC^{+}20]. creates [SSHC21]. Crippling [MNvdS^{+}20]. CRISPR [FHM^{+}20, LHS^{+}22, YSR^{+}21]. CRISPR-Cas12a-assisted [FHM^{+}20]. CRISPRi [KSM^{+}21a]. cristae [BWEHS21]. Critical [TRHS23, YKK^{+}20, LNY^{+}22]. Cross [VOR^{+}21, DdCVT22, EMEZ^{+}20, GLM^{+}22]. cross-linking [DdCVT22, EMEZ^{+}20]. cross-presentation [GLM^{+}22]. Cross-talk [VOR^{+}21]. crossing [DCRDC^{+}22]. crosslinker [SHBF^{+}20]. crosstalk [JMKS^{+}23, PAS^{+}22]. crowding [GNML^{+}20]. crucial [BBP^{+}20]. Crumbs [SLP^{+}22]. Cryo [LLLR20, BMF^{+}23, FSC22, NBI^{+}22, PMB^{+}20, SMM^{+}21, GSP^{+}20]. cryo-electron [BMF^{+}23, NBI^{+}22, PMB^{+}20]. Cryo-EM [LLLR20, SMM^{+}21, GSP^{+}20]. cryo-ET [FSC22]. cryotomography [GVA20]. cryptic [OHY^{+}20]. crystalline [RGK^{+}22, WC22]. CSPP1 [vdBVS^{+}23]. CTLs [FGBD^{+}21]. cue [LDE^{+}22]. Cul5 [DHTP22]. Cullin5 [LDH^{+}21]. curb [GLM^{+}22]. curvature [CSD22, MMKM21]. curved [GOR^{+}20]. cut [HL21]. Cuylen [MP22h]. Cuylen-Haering [MP22h]. Cvm1 [BCM^{+}22]. cyanobacteria [ABB^{+}22]. cycle [AMMK^{+}22, GMC^{+}20, IWI^{+}21, JIBK23, MBG^{+}23, XYG^{+}23]. Cyclin [JMB^{+}20, DOA^{+}22, HLGD20, JMC^{+}20, STS21]. Cyclin-dependent [SWN^{+}22]. cycling [ESB^{+}21]. Cyk4 [SRK22]. CYRI [Kin21, LYP^{+}21]. CYRI-A [LYP^{+}21]. cytokinesis [BJAR^{+}21, Hic22, MSC^{+}20, SCN^{+}23, STY^{+}20]. cytoneme [JRGH21]. Cytonemes [WBH^{+}21, WPM21]. cytoplasm [CAS23b, NR22]. cytoplasmic [FY20, GSL^{+}23, MHN20, SNP^{+}22]. cytoprotective [SSR^{+}22]. Cytoskeletal [Pro20, ALPH20, LZT^{+}23, PLG^{+}23, PVYJ^{+}21, PAS^{+}22, YKSC^{+}22]. cytoskeleton [BCC^{+}21, GM23, LDH^{+}21, MYM^{+}21, POL^{+}20, SJL^{+}22, SCL^{+}21]. cytoskeletons [BG22]. cytosolic [CLL^{+}21a]. cytotoxicity [DSY^{+}22, LSOM23].

D [WCL^{+}23]. D1 [STS21]. D54 [LLBC^{+}20]. DAD [GCL^{+}21]. Dam1 [FPZ^{+}22]. damage [CBJ^{+}21, CW23, DSB22, ITM^{+}21, JWB^{+}22, JFM^{+}22, MRL^{+}21, MFC^{+}20, SGW^{+}20]. damaged [vdBVS^{+}23]. DAPLE [MHGM22]. DarT [DSB22]. DarT-mediated [DSB22]. Dbnl [HMT^{+}21]. death [ABB^{+}22, DRC^{+}20, MRA20, OMI22, Ove21, TWT20, ZWJ22]. Decoding [MP23c]. Deep [DES^{+}23, GMD^{+}23]. DeepContact [LYL^{+}22]. defects [HKK^{+}20, MH22, RLV^{+}20]. deficiency [KNiY^{+}21, RLV^{+}20, VTL^{+}20]. deficient [DMR^{+}20]. defines [GVA20, GM23, LLBC^{+}20, MRG^{+}20, SNYA^{+}21]. Defining [PKH^{+}20]. deforms [MOK^{+}22]. degeneration [Hök22, KMD20, LPMA^{+}22]. degradation [JTM^{+}23, LGB^{+}21, OHHR23, OCLB21, PFPB^{+}20, PE22, SMK20, SSF^{+}22, TSL^{+}20, VGO^{+}23, XZJ^{+}21, ZS21, ZLW23, ZPG^{+}23, ZDGB^{+}22, ZRO^{+}23, ZVL^{+}23]. degradative [LFD^{+}21, VOR^{+}21]. Delaying [HL21]. delivered [PFS^{+}22]. delivers [KB21]. delivery [CMT^{+}21, GLGL^{+}21, MRD21, WPS22, ZBM^{+}22, ZLJ^{+}22]. delta [BJPH^{+}20]. delta-catenin [BJPH^{+}20]. demarcates [GCW^{+}23]. Dendrite [eSG23, BJPH^{+}20, HKK^{+}20, OYJJ23]. dendrites [BS20b, KAH^{+}21]. Dendritic [LAH^{+}21, BS20b, GLM^{+}22, PLG^{+}23, SPKP22, WHE^{+}22, YCC^{+}21]. density [BMS^{+}22, FLJ^{+}22, PCZ^{+}23]. Deorphanizing [KSS^{+}20a]. dependence [MMDK^{+}22]. dependent [ANRS^{+}20, ABB^{+}22, AANLL^{+}20, AII^{+}21, BD20, CCV^{+}21, CLL^{+}21b, CH22, DZA^{+}22, DLZ^{+}20, DCG^{+}23, GKFR20, HCWX^{+}22, HVPM20, HDW^{+}21, HCRMTC23, INM^{+}21, JMKS^{+}23, KST^{+}23, KWdB^{+}20, KBB^{+}23, LMRG20, LLA^{+}21, LLK^{+}21, LYS^{+}20, LLY22, MVM20, MC21, OKH^{+}20, PMB^{+}22, PHMD20, PGW^{+}21, PCGB20, PSP^{+}21, SBEB20, SFWB21, STY^{+}20, SWN^{+}22, YZW^{+}20, ZRO^{+}23, ZSJE20]. depends [CYU^{+}21, PRMF^{+}23, WPS22]. dephosphorylates [FAMQW22, QLC^{+}20]. dephosphorylation [BCdS22]. depletion [LWL^{+}23]. deploys [HYX^{+}20]. Depolarization [IvCD^{+}21]. depolymerization [SHGG21]. deposition [AANLL^{+}20]. derived [BLZ^{+}21, CMM^{+}20, ESX^{+}20, GLM^{+}22, MYK^{+}20, MYK^{+}21, MYK^{+}22, WCC^{+}23]. Design [IHBP^{+}23, dCS^{+}21]. despite [SdRVH^{+}21]. destabilization [ESB^{+}21]. destruction [KSWC22, NKS^{+}21]. detachment [PGH^{+}23]. detail [Bog21]. detect [dCS^{+}21]. Detection [MAW^{+}22, WBR^{+}20]. determines [DTG23, SNP^{+}22]. detrimental [CWZ^{+}20]. detyrosination [FOR^{+}20, LSOM23, RRCS^{+}23]. Deubiquitinases [CM21]. deubiquitylase [CHPF^{+}21a, CHPF^{+}21b]. deubiquitylation [JFM^{+}22]. Developing [ACPR21, LLC^{+}20, LJT^{+}22, LLX^{+}21, WKX^{+}21]. development [BJPH^{+}20, GKM^{+}20, GMIC^{+}20, JBV^{+}20, KKN^{+}21, LW20a, NBI^{+}22, RWSZ^{+}20, SCK^{+}19, SCK^{+}23, SCB^{+}20]. Developmental [KWGR23, YMAS20, Let20]. developmentally [CVMB^{+}23]. Dia1 [HDG22].

Different [TSP21, WDL^{+}20, RRCS^{+}23]. Differential [CML20, LL22, VDC^{+}20, CFV^{+}21]. differentially [XHF^{+}20]. differentiation [HDG22, KKPH^{+}21, MFC^{+}20, PDW^{+}20, SKF^{+}23]. diffuses [PCZ^{+}23]. diffusion [CLR^{+}20, STvT23]. dilute [ITB^{+}23]. dimer [SYW^{+}20]. dimerization [YZY^{+}20]. dimers [WMS^{+}21]. diminish [BJR^{+}21]. dimmer [Sea21]. diphosphatase [BBP^{+}20]. Direct [TTM^{+}21, CYL^{+}20, JGN^{+}20, WDL^{+}20]. directed [CDD^{+}22, RBL22, SLP^{+}22]. directing [TEH^{+}20]. directly [FER^{+}23]. directs [LGL^{+}23, WDJ^{+}21, ZMMM^{+}20]. Disagreement [JJ23]. disassembly [AHQ20, KSWC22, LDE^{+}22, MTR^{+}20, SBV^{+}20, YLH^{+}21, ZBY^{+}21]. disassociation [LWL^{+}23]. DISCO [GGA21]. Discoidin [NR22]. Discrete [MRH^{+}23, BTF^{+}20, BDD20, CEM^{+}20]. discriminate [DCS^{+}20]. disease [CKW^{+}22, DRZ^{+}23, KPG20, MH22, SDD^{+}22, PGDD21, TF20]. diseases [HKK^{+}20]. Dishevelled [BP22, KSWC22]. disinhibition [HKK^{+}20]. disjunction [AHQ20]. displaces [VHPP^{+}20]. display [YPM^{+}21]. disposal [RG23]. disrupting [FAHZ21]. disruption [WJW^{+}22]. disrupts [MPKB^{+}20]. Dissecting [FHM^{+}22]. dissection [ZHHJ22]. dissipate [LSD20b]. dissolution [RSB^{+}23]. distal [KRHP^{+}21, VHPP^{+}20]. distance [DY21]. Distinct [LRM^{+}20, CLH21, FC21, NBC^{+}21, PKC^{+}22, RCF^{+}22, WDL^{+}20, WRG23, WDRRF^{+}23, YMAS20]. distribute [WPM21]. distribution [CYU^{+}21, DdCVT22, LLC^{+}20, LWD^{+}21, PKH^{+}20, ZMS^{+}20, vdBdHLK22]. divergent [HYX^{+}20]. diverse [VTL^{+}20]. division [BWA^{+}23, CCH^{+}21, MDB^{+}20, OMK^{+}22, RSWP20, SPRWB20, Tev20]. divisions [FAMQW22]. DLY [BSC22]. DMV [JLS^{+}22]. DNA [ABM^{+}23, CWZ^{+}20, CBJ^{+}21, CBS^{+}21, ITM^{+}21, JFM^{+}22, KSP^{+}21, LLA^{+}21, MRL^{+}21, MSH^{+}20, MV20, MFC^{+}20, MMC20, PDW^{+}20, SSHC21, SBBJ21, SGW^{+}20]. DNA-PK-AKT [MRL^{+}21]. DNase [PZWW21]. do [Col22a, Col22b, SMD^{+}21]. docking [SJL^{+}22]. Does [BW23, SNYA^{+}21]. domain [BS20a, CMM^{+}20, CPW^{+}23, CJS^{+}21, DLZ^{+}20, FWP^{+}20, HSSK20, SYW^{+}20, YZY^{+}20, ZLS^{+}21, ZVM^{+}20]. domains [SWT^{+}22]. Don [BW20]. dopaminergic [JMKS^{+}23, KJ23]. Dorothy [MP22b]. Double [MS23, KMJ^{+}23, WCC^{+}23]. Double-checking [MS23]. double-membrane [WCC^{+}23]. double-strand [KMJ^{+}23]. downregulating [BZD^{+}21]. downstream [AHvR^{+}20, KMD20, RLK^{+}20]. DPYSL2 [ASK^{+}22]. DRG [LYS^{+}20, GKFR20]. drink [Kin21]. drive [DJI^{+}21, HLB^{+}22, JMY^{+}23, KHV^{+}22, SHGG21, SWT^{+}22, SLH^{+}20b, ZXW^{+}20]. driven [AANLL^{+}20, SLL^{+}21, SLL^{+}23, VGO^{+}23]. drives [AKN^{+}22, CAS23b, GLGL^{+}21, HCB^{+}23, JMC^{+}20, KST^{+}23, KMJ^{+}23, LC20, LDE^{+}22, MS20, NTA^{+}21, OYS^{+}22, OCLB21, PSP^{+}21, RPM^{+}21, RGP^{+}22, VFL20, WLBS20, WKC^{+}22, WZK^{+}23, YCC^{+}21, ZGR^{+}22, vLEM^{+}20]. driving [Kin21, LSG^{+}22]. droplet [Cas21, CYR^{+}21, CEM^{+}20, DZA^{+}20, Goo20, GMCO^{+}22, HAW^{+}22, RE20, ZHW^{+}21, ZDM^{+}22]. droplets [CT20, DZA^{+}20, DY21, ITB^{+}23, MYT^{+}21, RGK^{+}22, SOT^{+}21, WC22]. Drosophila [BCC^{+}21, DdCVT22, FY20, KWGR23, LTL^{+}20, MdCT23, PKD^{+}20, PMSO^{+}23, SLES20]. DRP [CLL^{+}21b]. DRP-1-dependent [CLL^{+}21b]. Drp1 [OCB^{+}21]. Drp1-mediated [OCB^{+}21]. Ds [TB20a]. Dscam2 [OKH^{+}20]. Dual [SdRVH^{+}21, LLK^{+}21]. duct [BED^{+}21]. duplication [CVMB^{+}23, IWI^{+}21, PKD^{+}20, PSC^{+}20, VDC^{+}20]. duration [LAH^{+}21]. during [AMFW^{+}21, BCC^{+}21, BCWM21, BHK20, CS21b, CS21d, CWAT20, CLR^{+}20, DPM^{+}20, DHTP22, EM20, FAMQW22, FGBD^{+}21, GMIC^{+}20, HHT^{+}20, Hic22, HYL^{+}20, JLS^{+}22, JWB^{+}22, LMS^{+}21, LNY^{+}22, LDH^{+}21, MTR^{+}20, MRWK^{+}22, MYM^{+}21, MTW^{+}23, NBC^{+}21, OYJJ23, PVYJ^{+}21, RLS^{+}20, SRUdC^{+}22, SGL^{+}23, SCN^{+}23, SCK^{+}19, SCK^{+}23, STY^{+}20, SMC^{+}20, TP20, VCS^{+}22, VV23, WXM22, WAK^{+}20, hYKO^{+}20a, hYKO^{+}20b, hYKO^{+}21, ZLW23]. dynactin [KRS21, dAC^{+}22]. Dynamic [Kin21, DSB22, Gui21, MSJ20, RGP^{+}22, ZMW^{+}22]. dynamically [MBA^{+}22]. Dynamics [HSU^{+}20, AH20b, ABM^{+}23, BPF^{+}21, BMM^{+}20, CDD^{+}22, DTG23, DES^{+}23, EM22, GMB^{+}20, JIBK23, JCL^{+}23, KBH^{+}22, LAH^{+}21, LDH^{+}21, LMJ^{+}20, LSD^{+}21, MRL^{+}21, MRWK^{+}22, NVPP20, PGH^{+}23, PMSO^{+}23, PPG21, PLL^{+}20, SLP^{+}22, STY^{+}20, SMC^{+}20, WXM22, WH22, WKC^{+}22, YPM^{+}21]. dynamin [LHL^{+}23]. dynamin-2 [LHL^{+}23]. Dynamin2 [LMM^{+}23]. dynein [ARCM20, CCV^{+}21, CGCR^{+}22, yLHW^{+}20, DCRDC^{+}22, GSL^{+}23, KRS21, KKP^{+}21, QZX23, BSC22, dAC^{+}22, SRK22]. dynein-2 [DCRDC^{+}22]. dynein-mediated [yLHW^{+}20]. dyneins [BOW^{+}22]. Dyrk1a [LNY^{+}22]. dysfunction [CFK^{+}22, IMR^{+}23, SLH^{+}20a]. dysfunctional [BC23]. dysplasia [KNiY^{+}21]. dysregulation [VTL^{+}20]. dystrophin [AZR^{+}22].

E-cadherin [HVPM20]. E-catenin [SMS^{+}20]. E-Syt1 [LM23, SvVV^{+}23]. E3 [BMM^{+}20, DMR^{+}20, LSD^{+}21, PE22, SSF^{+}22, TSL^{+}20]. E4orf4 [DRC^{+}20]. early [CCFN^{+}20, MYK^{+}20, MYK^{+}21, MYK^{+}22, O’D22, RWSZ^{+}20, SCK^{+}20a, SCK^{+}20b, ZLJ^{+}22]. earmark [SNN20]. Easy [LM21]. Eating [Yam21, GG20]. EB1 [KMW20]. ebb [ASC20]. ECM [AANLL^{+}20, MMDK^{+}22, PFPB^{+}20]. Ecm29 [LLC^{+}20]. Ecm29-mediated [LLC^{+}20]. Ect2 [MLS^{+}22, SRK22]. Ect2/Cyk4/Mklp1 [SRK22]. ectodomain [GSP^{+}20]. ectopic [MKO^{+}21]. educate [CKR^{+}20]. Effector [ZLS^{+}21, EMEZ^{+}20, MAW^{+}22, PCZ^{+}23, WHE^{+}22, XZJ^{+}21]. Effector-mediated [ZLS^{+}21]. effectors [CCV^{+}21]. effects [KSM^{+}21b]. efferocytosis [RG23]. efficacy [WAOS^{+}21]. efficiency [LAH^{+}21]. Efficient [DF22, KMW20, SBBJ21]. EGF [CHZ^{+}20]. EGFR [LGB^{+}21, NTA^{+}21, SWS21b]. EGFR-mediated [NTA^{+}21]. EGFR-RAS-MAPK [SWS21b]. egress [RCA^{+}21, dCTOG^{+}20]. eIF6 [WI22]. Elda [MP22c]. Electron [GVA20, BMF^{+}23, GMD^{+}23, LYL^{+}22, NBI^{+}22, PMB^{+}20, RMM^{+}21]. electrostatic [GCL^{+}21]. elegans [CSG22, DPM^{+}20, HČK^{+}20, JBV^{+}20, LGL^{+}23, LMJ^{+}20, RCH^{+}20, TP20].

Elimination [AMFW^{+}21]. elongated [KSS^{+}20b, KSS^{+}20c]. elongating [UIS^{+}22]. elongation [RMA21, YMAS20]. Elvan [MP22d]. embryo [JBV^{+}20, MS20]. embryonic [JRGH21]. embryos [CSG22]. Emergence [ALPH20]. enables [FDA21, HRS^{+}20]. enclosing [DG22]. encode [SLM23]. encoded [KLC^{+}20, dCS^{+}21]. encounter [HL21]. end [FAHZ21, RDL^{+}20, SHGG21, WRG23]. endo [GCL^{+}21]. endo-plasma [GCL^{+}21]. endocytic [BSH^{+}22, CSD22, EMY^{+}22, MLQ^{+}21, YLH^{+}21, dDFGP^{+}21]. endocytosis [CMM^{+}20, EM20, GMIC^{+}20, KBB^{+}23, LHL^{+}23, LWG^{+}22, MC21, MTW^{+}23, PHMD20, PGW^{+}21, TOL^{+}20, ZSJE20]. endogenous [BGM^{+}21, vdBdHLK22]. endogenously [WDRRF^{+}23]. endolysosomal [BLU21, RCS22]. Endomembranes [FDSR22, DG22]. Endophilin [YCC^{+}21]. Endoplasmic [CSM^{+}21, AAR^{+}21, BBP^{+}20, GCS^{+}20, GMB^{+}20, SPT^{+}09, SPT^{+}21, SLM23, WMS^{+}21, ZHW^{+}21, ZDM^{+}22]. endorecycling [SWS21b]. Endos [LKW^{+}21]. Endosomal [MH22, PFS^{+}22, GLGL^{+}21, JDKK^{+}22, KKN^{+}21, LLY22, OKH^{+}20, SV22, VBG^{+}22, WME22, ZXY^{+}23, vdBdHLK22]. endosome [BLZ^{+}21, HSW^{+}22, HMSF22, KSN^{+}22, MYK^{+}20, MYK^{+}21, MYK^{+}22, PWW^{+}20, RCM^{+}23b, RBL22, Sea21, SV22, WR22, YLH^{+}21]. endosome-associated [RBL22, YLH^{+}21]. endosome-derived [BLZ^{+}21, MYK^{+}20, MYK^{+}21]. endosome-to-cell [Sea21]. endosome-to-TGN [RCM^{+}23b]. endosomes [LCM22, MVM20, O’D22, WR22, ZLJ^{+}22]. Endothelial [LWL^{+}23, CFV^{+}21, CKM^{+}20, KPM^{+}22]. ends [Gui21, Sir23, TSP21, vdBVS^{+}23]. energy [RZN^{+}22]. enforced [BRD^{+}21]. engage [GLM^{+}22]. engagement [IWI^{+}21, NMO^{+}22]. engages [SKX^{+}23]. Engineered [LRB^{+}22, SHLS22, FHM^{+}22, TB20a]. enhance [JCL^{+}23, WHE^{+}22]. enhanced [MRL^{+}21]. enhances [EZB^{+}20]. Enhancing [WZG22]. enough [ITB^{+}23]. enriched [RSB^{+}23]. enrichment [KKZ^{+}22, MWSX23]. ensheathing [FDSR22]. ensure [IWI^{+}21, JMB^{+}20, MKO^{+}21, YLH^{+}21, Zar20]. ensures [CSOG^{+}20, FCCH21, HGK20, PK23, RSB^{+}23, ZHW^{+}21]. Entosis [AHvR^{+}20, BDS^{+}21]. entrocortin [RFL20]. entry [AMMK^{+}22, DOA^{+}22, LDE^{+}22]. envelope [DNVP23, Köh21, KAS^{+}22, LSD^{+}20a, LW20b, LD20, ML22, PSS^{+}20, PRMF^{+}23, PSP^{+}21, TTM^{+}21, WLBS20]. envelopes [SMD^{+}21]. environment [ZAR^{+}21]. enzymatic [WCL^{+}23]. enzyme [JBV^{+}20]. EpCAM [HSF^{+}23, VRSN23]. EPH [KSM^{+}21b]. EPH/EPHRIN [KSM^{+}21b]. EPHecting [McC21]. EPHRIN [KSM^{+}21b]. epidermal [BHS^{+}21, NTA^{+}21]. epidermis [MBG^{+}23]. epigenetic [BHS^{+}21, CD21]. epigenetically [DCK^{+}20]. epigenomic [BDH^{+}21]. epithelial [AR20, BRB^{+}20, CHS^{+}22, DDD^{+}20, DCS^{+}20, DYW^{+}20, FBR^{+}21, GY20, MDV^{+}21, OHY^{+}20, PAS^{+}22, QLC^{+}20, SLP^{+}22, VRSN23, WB20, vLEM^{+}20, vdGM22]. epithelial-to-neural [AR20]. epithelium [HDG22, SLES20]. epithelium-to-neural [SLES20]. EPLIN [GSC^{+}20, LDH^{+}21]. Eps15 [EMY^{+}22]. Eps15/Pan1p [EMY^{+}22]. ER-bound [LHS^{+}22]. ER-derived [WCC^{+}23]. ER-lipid [DZA^{+}20]. ER-lysosome [HCWX^{+}22]. ER-mitochondria [CCH^{+}21, SvVV^{+}23]. ER-phagy [WJL^{+}23]. ERAD [TSL^{+}20]. ErbB4 [AVC^{+}22]. ERdj8 [hYKO^{+}20a, hYKO^{+}20b, hYKO^{+}21]. Erg1 [FUBS22]. ERK7 [OHHR23]. ERM [RCA^{+}21, ZLS^{+}21]. ERM-guided [RCA^{+}21]. Ernst [TB20b]. erosion [VZQ^{+}21]. error [DKCT21, FOR^{+}20, RFL20]. error-free [RFL20]. escape [CWAT20, MP22a, PFS^{+}22]. ESCRT [LMRG20, LSD^{+}20a, TTM^{+}21, WLBS20, YZW^{+}20]. ESCRT-dependent [YZW^{+}20]. ESCRT-III [WLBS20]. ESCRT-III-dependent [LMRG20]. ESCRTs [LD20]. essential [CSD22, CLZ^{+}20, JTM^{+}23, JLS^{+}22, PSA^{+}23, PE22]. establish [CEM^{+}20]. establishes [PPB^{+}21]. establishment [CSG22]. esters [MYT^{+}21, RCF^{+}22]. estrogen [ANRS^{+}20]. estrogen-dependent [ANRS^{+}20]. Eukaryotic [KPM^{+}22]. evasion [AMG^{+}20]. even [PCZ^{+}23]. eviction [SPRWB20]. Evidence [DPM^{+}20]. EVL [PLG^{+}23]. evoked [BS20b]. evolutionarily [SSR^{+}22]. Evolutionary [WRG23]. Evolving [CS20]. exchange [BRB^{+}20]. excitatory [LLC^{+}20]. exclusion [Tev20]. exit [CBC^{+}20, DF22, GCNL21, MTR^{+}20, SNYA^{+}21, WJW^{+}22, WMS^{+}21]. Exocyst [RLK^{+}20, MRH^{+}23, PSA^{+}23, SKX^{+}23]. exocytosis [PWW^{+}20, RCM^{+}23a]. exon [GPEC^{+}23]. Exosomal [MNC20, AANLL^{+}20, WAK^{+}20]. exosome [VBG^{+}22]. exosomes [LMRG20]. Expanded [FER^{+}23]. expansion [BRD^{+}21, FFZ^{+}22, SHD^{+}21, SPT^{+}09, SPT^{+}21, SLD^{+}21]. Exploring [MRA20]. export [DZA^{+}22, HVPM20, LHS^{+}22]. expression [AZR^{+}22, BCWM21, HCL^{+}21, KVG^{+}20, LLK^{+}22, SRUdC^{+}22, TRJ^{+}20, WHA20]. extend [AH20b, LMJ^{+}20, WB20, WBH^{+}21, XDY^{+}22]. Extracellular [STS21, BSH^{+}22, GKFR20, GKM^{+}20, ICMM20, JKL^{+}22, LYS^{+}20, MBW22, RPM^{+}21, SMK20, WDB^{+}21]. ExTrack [STvT23]. extraordinary [VM21]. extravasation [SMC^{+}20]. extrinsic [KBN^{+}21]. extrusion [AHvR^{+}20, KTT^{+}22].

F [ARCM20, APL^{+}21, MLS^{+}22, YLH^{+}21]. F-actin [MLS^{+}22, APL^{+}21, YLH^{+}21]. F-actin/mitochondria [APL^{+}21]. F508 [HVPM20]. facilitate [OYS^{+}22, PGH^{+}23, WCC^{+}23]. facilitates [AH20a, CLL^{+}21b, CWX^{+}21, JMB^{+}20, PTS^{+}22, QZX23, RFL20, WLM^{+}20]. facilitating [DHB^{+}21]. factor [KPM^{+}22, ZJDR22]. Factoring [WI22]. Factors [LSD^{+}20a, BTF^{+}20, BOW^{+}22, WDL^{+}20, WHA20, YMAS20]. FAM134B [WJL^{+}23]. FAM134B-mediated [WJL^{+}23]. FAM19A [KSS^{+}20a]. Fam20C [HBS^{+}20]. family [MVM20]. Farquhar [SSB20]. fascin [PLL^{+}20]. fashion [HCRMTC23]. fast [MV20]. fast-tracks [MV20]. faster [RMA21]. Fat [FER^{+}23, SHD^{+}21]. fate [BHS^{+}21, DCK^{+}20, MP22i, ZPG^{+}23]. father [WMA^{+}23]. Fbp17/RacC [LYL^{+}23]. Fbxo42 [BZD^{+}21]. features [CLH^{+}20]. feedback [FCHM20]. feedforward [LJJ^{+}21]. feeds [BD20]. feet [GY20]. female [TP20, WLM^{+}21]. FER [LGB^{+}21]. ferritin [OYS^{+}22]. ferritinophagy [WZ22]. ferroptosis [ABB^{+}22, Gan21, RCM^{+}23a]. fertilization [BW23, MYM^{+}21, RCH^{+}20, SSZL21]. FFAT [KHB^{+}22, WME22]. FG [CPC^{+}20, Dor20]. FG-nucleoporins [CPC^{+}20, Dor20]. FGD1 [ZMMM^{+}20]. FGD1/CDC42 [ZMMM^{+}20]. FGF2 [LSG^{+}22, WZG22]. FHL2 [BPF^{+}21]. FIB [LSS^{+}23, MSX^{+}21]. FIB-SEM [LSS^{+}23, MSX^{+}21]. fibers [FSZ^{+}22, KST^{+}23, LSD20b, SvDSW^{+}20]. fibrillar [AKN^{+}22]. fibrinogen [LM21, WAOS^{+}21]. fibroblast [JML^{+}21]. Fibroblasts [HCRMTC23]. fibronectin [AKN^{+}22, BJSOS^{+}20, BJSOS^{+}21, HCRMTC23]. fibronectin-associated [AKN^{+}22]. fibrotic [CHZ^{+}20]. fidelity [CSOG^{+}20, FMY^{+}21, INM^{+}21, KHV^{+}22, LZC^{+}20, MKO^{+}21, PK23, Zar20]. Filament [PMB^{+}22, BG22, CVMB^{+}23, FLJ^{+}22, Gui21, GM23, SHGG21, Sir23]. filamentous [GC22, PMB^{+}20]. filaments [MTCL^{+}23]. Filamin [SJL^{+}22]. Filling [HH22]. filopodia [CJS^{+}21, DJI^{+}21, HRB^{+}21, JGN^{+}20, LC20, PLG^{+}23, PLL^{+}20]. filopodial [BMM^{+}20]. filter [PHT^{+}23]. final [HL21]. finds [CD21]. Fine [McW23, MC21, AFB^{+}20, GL20, LLW^{+}21, ZMW^{+}22]. Fine-tune [McW23]. fine-tuned [ZMW^{+}22]. fine-tunes [AFB^{+}20, GL20, LLW^{+}21]. Fine-tuning [MC21]. FIP200 [SYW^{+}20]. firehose [PH20]. Fis1 [WKC^{+}22]. FISHing [MP22g]. Fission [AGW^{+}20, MSC^{+}20, KSN^{+}22, SV22, WME22, ZJH22]. FIT2 [BBP^{+}20, CYR^{+}21]. flagella [ATS^{+}21]. flagellum [ATS^{+}21]. flashes [VCS^{+}22]. flat [HAL^{+}23]. flies [O’D20b]. FLN [SJL^{+}22]. FLN-2 [SJL^{+}22]. flow [ASC20, MVM20]. flows [IHBP^{+}23]. fluorescence [LQS23, WBR^{+}20]. flux [AAF^{+}20, GOR^{+}20]. fly [AR20, EJBB^{+}20, HKK^{+}20]. FMNL2 [PLL^{+}20]. FMR1 [WAK^{+}20]. FMRP [RFL20]. focal [GMB^{+}20, JKL^{+}22, RRBW^{+}21, Tan23, WZtM^{+}20]. focuses [DLK^{+}21]. foe [SLM20]. folding [SLM23, WB21]. Follicle [MdCT23]. Food [HI21]. Force [KBB^{+}23, ALC^{+}20, BJR^{+}21, LSD20b, SvDSW^{+}20, WZtM^{+}20, WI22]. force-independent [ALC^{+}20]. force-responsive [SvDSW^{+}20]. forces [DPM^{+}20, MBA^{+}22, ME21, MP22i]. fork [DMR^{+}20, RDW^{+}20]. forks [MYC^{+}23]. form [ABB^{+}22, ACPR21, LLK^{+}22, MYT^{+}21, Ped22, RMA21, SMD^{+}21, SOT^{+}21]. formation [AKN^{+}22, AII^{+}21, BBPS23, CJS^{+}21, CG21, DJI^{+}21, DY21, EBZC^{+}21, FSZ^{+}22, FLJ^{+}22, GBBT^{+}22, GSC^{+}20, HAW^{+}22, JLS^{+}22, JMC^{+}20, LYP^{+}21, LLX^{+}21, MKO^{+}21, MPVD^{+}21, OTOF21, OHY^{+}20, SRK22, SMHH^{+}20, SWT^{+}22, TRJ^{+}20, WTS^{+}21, WCC^{+}23, hYKO^{+}20a, hYKO^{+}20b, hYKO^{+}21, ZMMM^{+}20, ZXW^{+}20, ZFZ^{+}23]. formin [Sir23, Sir23]. forming [CYR^{+}21, JKZ^{+}22, MLS^{+}22]. fortifies [KKZ^{+}22]. Fps1 [LL22]. Fps1-mediated [LL22]. fractionation [UZS^{+}23]. fragment [LLLR20]. fragmentation [HRS^{+}20]. fragments [WLBS20]. free [MLvdL^{+}21, RFL20]. freely [PCZ^{+}23]. friend [SLM20, VM21]. fucose [SNP^{+}22]. Function [HLGD20, Tev20, AGH^{+}22, BJPH^{+}20, BDK21, BNV^{+}23, CSD22, CLH21, DSG21, FRO^{+}20, FC21, GNL^{+}20, GVD^{+}20a, GVD^{+}20b, GM23, HHT^{+}20, KAH^{+}21, LLK^{+}22, LGS22, ML22, MRG^{+}20, Ped22, WESR22, WTU^{+}21, ZLJ^{+}22]. Functional [HESH^{+}22]. functionality [RCS22]. functioning [ZDM^{+}22]. functions [DACG^{+}21, DCG^{+}23, KKP^{+}21, KSP^{+}21, LMM^{+}23, NVPP20, PBPBS22, SFWB21, VOR^{+}21, WDB^{+}21, WHE^{+}22, YW21]. fundamentals [GH20]. FUNDC1 [CCH^{+}21]. furrow [SRK22]. FUS [CHZ^{+}20, LLA^{+}21]. FUS-dependent [LLA^{+}21]. fuse [RCH^{+}20]. Fusion [FCT^{+}20, AGW^{+}20, BW23, BNV^{+}23, BSC^{+}23, CMN^{+}22, JMY^{+}23, LML^{+}21, MS20, MWF^{+}23, MMKM21, SSZL21, WLBS20, XZJ^{+}21]. FXR1 [SCB^{+}20]. Fyn [CDLZ^{+}22].

G [MHS^{+}20, TJAG^{+}21, YZY^{+}20]. G0 [AMMK^{+}22]. G1 [BTF^{+}20]. G1/G0 [AMMK^{+}22]. G1/S [BTF^{+}20]. G3BP [KPA^{+}16, KPA^{+}20]. GABA [LLC^{+}20, Let20, SIP^{+}23]. Gaia [MP21a]. GAK [HSU^{+}20]. galectin [ZTL^{+}23]. galectin-3 [ZTL^{+}23]. galectin-3/Lrp1 [ZTL^{+}23]. Gamete [SSZL21]. GAP [WZZ^{+}23, vdGM22, HMSF22]. gaps [HH22]. garbage [Let20]. GARP [OYJJ23, eSG23]. GAS2L1 [AHQ20]. GBF1 [NMO^{+}22]. GCN5 [AZR^{+}22]. GDP [SNP^{+}22]. GDP-fucose [SNP^{+}22]. GDPGP1 [SSO^{+}20, SLM20]. GDPGP1/mcp [SSO^{+}20]. GDPGP1/mcp-1 [SSO^{+}20]. GEF [WLW^{+}22]. Gene [KVG^{+}20, BCWM21, CVMB^{+}23, HCL^{+}21, LLK^{+}22, RDW^{+}20, SRUdC^{+}22, SSO^{+}20, WHA20]. general [UZS^{+}23]. generate [RRCS^{+}23, ZAR^{+}21]. generates [DRW^{+}23]. generation [BJR^{+}21, MWSX23, MP22b, TG21, WZtM^{+}20]. genes [FHM^{+}20, KJ23]. genetic [BGM^{+}21]. Genetically [KLC^{+}20, dCS^{+}21]. genome [BZD20, KSM^{+}21a, Mar21, ME21]. geometries [WBH^{+}21]. germ [HRB^{+}21, MHN20, ME21, SBBJ21]. germline [BCWM21]. GET [FUBS22, MOS^{+}22]. Get1 [CLC^{+}21]. Get1/2 [CLC^{+}21]. gets [RG23]. Giant [CYL^{+}20, GPES21]. Giantin [SBL^{+}21]. Gilgamesh [LTL^{+}20]. Gish [LTL^{+}20]. Gist [SSB20]. gland [MND^{+}20]. GlcNAc [YM21]. Glial [LCB^{+}23, CVT^{+}21, KNiY^{+}21]. glioblastoma [KCP^{+}21]. Glo3 [XGD^{+}23]. global [BCWM21]. globally [Zar20]. glucose [BPF^{+}21]. GLUT4 [CCFN^{+}20, LHL^{+}23]. glycine [RCDMM20]. glycocalyx [BW20]. glycogen [SSO^{+}20, SLM20]. glycolytic [CFK^{+}22]. glycoprotein [LLW^{+}20, TJAG^{+}21]. glycosylation [SNP^{+}22]. Glypican [HRB^{+}21]. Glypicans [WPM21]. Go [Yam21, ASC20, BP22, WC22, Col22a, Col22b]. Godinho [O’D20a]. goes [MP21b]. Golgi [GVD^{+}20a, Bur21, CJK^{+}22, GPL^{+}21, GVD^{+}20b, HSW^{+}22, LKMM^{+}23, Low21, MWSX23, NSB^{+}21, OYJJ23, PFPB^{+}20, PBPBS22, SBV^{+}20, TML22, WHN^{+}21, WPCB^{+}21, XGD^{+}23, Yam21, ZS21, ZXY^{+}23]. Golgi-associated [SBV^{+}20]. GOLPH3 [Low21, WPCB^{+}21]. GOLPH3L [WPCB^{+}21]. good [VRSN23]. GORASPs [GVD^{+}20a, GVD^{+}20b]. governed [YLH^{+}22]. governs [PMSO^{+}23, TJAG^{+}21, hYKO^{+}20a, hYKO^{+}20b, hYKO^{+}21].

GP130 [TJAG^{+}21]. GPCR [CPS^{+}22]. GPCRs [CWKP23, SNN20]. GPI [ARM23, LWZ^{+}23, TWY^{+}22]. GPI-anchored [LWZ^{+}23]. Gq [CPS^{+}22]. Gq-GPCR [CPS^{+}22]. Grabocka [MP22c]. Gradient [WPS22, EMEZ^{+}20]. gradients [GPW^{+}22]. GRAF2 [HVPM20]. Granular [Bog21]. granule [BVYW20, KPA^{+}16, KPA^{+}20, MYK^{+}20, MYK^{+}21, MYK^{+}22, POL^{+}20]. granules [FMY^{+}21, FPMS^{+}21, JWB^{+}22, LFF^{+}22, MMSP20, MP22c, PTS^{+}22, YPM^{+}21]. GRASP55 [ZS21]. GRASP65 [ZS21]. GRASPing [Bur21]. greater [VRSN23]. Greatwall [LKW^{+}21]. groom [Ver21]. groove [LLLR20]. growing [FAHZ21, vdBVS^{+}23]. growth [FER^{+}23, GC22, KCP^{+}21, MLS20, PKC^{+}22, RBL22, SHBF^{+}20, SWN^{+}22, ZHW^{+}21]. GSK3 [LHL^{+}23]. GSK3 [KHB^{+}22]. GTP [AII^{+}21]. GTP-dependent [AII^{+}21]. GTPase [BLU21, KRS21, LD20, SFC^{+}23, VBG^{+}22]. GTPases [HHGR21, RLK^{+}20, WDRRF^{+}23]. guidance [BMM^{+}20]. guide [BKR^{+}22]. guided [RCA^{+}21, WJW^{+}22]. GxcM [LYL^{+}23]. GxcM-Fbp17 [LYL^{+}23]. GxcM-Fbp17/RacC-WASP [LYL^{+}23].

H [MAW^{+}22, SGN^{+}20]. H-zone [SGN^{+}20]. H1 [CBS^{+}21]. Haering [MP22h]. hair [LLW^{+}20]. hair-like [LLW^{+}20]. hairpin [FUBS22]. Hands [GY20]. Haspin [HHT^{+}20, PKY^{+}20]. Hatched [ME21]. HDAC6 [ORCT^{+}20]. head [HGK20]. health [DRZ^{+}23, KPG20, PK23, Pie20, TF20]. heart [BWEHS21, LJT^{+}22]. heat [CLL^{+}21b, FAS^{+}21, SSR^{+}22]. heavy [MLL^{+}20, SCN^{+}23, TSL^{+}20]. Hec1 [INM^{+}21]. Hedgehog [MKD^{+}21, AT21, DSLP20, FDG^{+}21, LLW^{+}21, LSD^{+}21, PRB^{+}20]. height [WB20]. helices [ZY21]. helix [CT20]. help [NR22, YKSC^{+}22]. helps [Kin21]. Hematopoietic [BCS^{+}21, Dus21, HZN^{+}21, LD21]. Hemicentin [GKRL^{+}23]. Hemicentin-mediated [GKRL^{+}23]. hemichannels [KIV^{+}20]. Hemidesmosomes [WZtM^{+}20]. Heparan [ICMM20, SMK20]. Hepatocyte [BBM^{+}23, BRD^{+}21]. herniations [TTM^{+}21]. herpesvirus [CAS23b, ZFZ^{+}23]. heterogeneity [LW20a, NGG^{+}20, SRW^{+}21]. Heteromer [GM23]. heterotrimeric [MHS^{+}20]. hide [MP22f]. hierarchies [VLdRADJ22]. High [BDH^{+}21, FMY^{+}21, KHFK^{+}20, LYL^{+}22, RMM^{+}21, WAOS^{+}21, YSR^{+}21, PCZ^{+}23]. High-content [YSR^{+}21]. high-density [PCZ^{+}23]. High-efficacy [WAOS^{+}21]. High-fidelity [FMY^{+}21]. High-precision [RMM^{+}21]. High-speed [KHFK^{+}20]. High-throughput [BDH^{+}21, LYL^{+}22]. Hippo [FER^{+}23, DYW^{+}20, JMC^{+}20, RSWP20]. HIV [MWF^{+}23]. HIV-1 [MWF^{+}23]. HLH [LMJ^{+}20]. HLH-30 [LMJ^{+}20]. hnRNP [TRJ^{+}20]. home [CD21, Low21]. homeostasis [BBP^{+}20, BCM^{+}22, CSM^{+}21, IMR^{+}23, LTL^{+}20, ZJH22, ZDM^{+}22]. homologous [MSH^{+}20]. Homophilic [LXJ^{+}23]. Hongyuan [Cas21]. Horizontal [DRZ^{+}23]. host [NMO^{+}22]. HSPG [ZVC^{+}21]. Human [BSC^{+}23, JMY^{+}23, MTCL^{+}23, ANRS^{+}20, Bez22, BDD20, CCFN^{+}20, CZTL21, CMN^{+}22, CPS^{+}22, ESH^{+}23, FLW^{+}23, JMKS^{+}23, MFC^{+}20, MPVD^{+}21, SMFC^{+}22, TG21, VZQ^{+}21, WTU^{+}21, ZHHJ22]. Huntington [CKW^{+}22, MH22]. hybrid [CBS^{+}21]. hybrids [SSHC21]. hydrolase [WZK^{+}23]. Hydroxylated [HSSK20]. Hyperstabilization [BEM^{+}23]. hypervariable [KBN^{+}21]. hypo [OMI22]. hypo-osmotic [OMI22]. hypoxia [CCH^{+}21]. hypoxia-induced [CCH^{+}21].

I-band [SGN^{+}20]. I-mediated [KTT^{+}22]. iASPP [MSB^{+}21]. Identification [HČK^{+}20]. identifies [DSMB20]. identity [ML22, TMG^{+}21]. ides [PE22]. IFN [SPKP22]. IFN- [SPKP22]. IFT [DCRDC^{+}22, PL22]. IGF1 [LJJ^{+}21]. Igf2 [KKPH^{+}21]. II [yLHW^{+}20, HCRMTC23, IvCD^{+}21, PKY^{+}20, UIS^{+}22, ZLJ^{+}22]. III [LMRG20, WLBS20]. IL1R [DACG^{+}21]. ILEE [LZT^{+}23]. Image [KSM^{+}21a, SHA20]. Image-based [KSM^{+}21a, SHA20]. images [LZT^{+}23, VLdRADJ22]. imaging [CBS^{+}21, FLW^{+}23, KHFK^{+}20, LYL^{+}22, NGG^{+}20, SSHC21, UIS^{+}22, VVW^{+}23, WDRRF^{+}23, YSR^{+}21]. imaging-based [YSR^{+}21]. immediate [ZS21]. immune [AMG^{+}20, CW23, DSG21, LJJ^{+}21, MP21b, MP22a, MP23b, NS20, RS22, WH22, WHE^{+}22]. immunity [LGS22]. immunological [ACPR21, BB20, LAH^{+}21, WM20]. impact [ZS21]. Impaired [CKW^{+}22]. impairs [FOR^{+}20]. impart [EM22]. impedes [DZA^{+}22]. import [AAR^{+}21, MOS^{+}20, XDY^{+}22, YTH^{+}20, YLH^{+}22]. important [LLA^{+}21, PRB^{+}20, SCB^{+}20]. importin [EMEZ^{+}20]. importing [CKW^{+}22]. inactivates [FFZ^{+}22]. inactivation [TKK^{+}20, WPS22]. inactive [CBS^{+}21, SWT^{+}22]. INAVA [CLL^{+}21a]. INCENP [PZ21]. incompletely [PDW^{+}20]. increase [LPMA^{+}22]. increasing [BTF^{+}20, McC21]. indent [ITB^{+}23]. Independent [BS20a, AMMK^{+}22, ALC^{+}20, BRB^{+}20, BWA^{+}23, DSY^{+}22, KTT^{+}22, LSD^{+}20a, OCB^{+}21, SMHH^{+}20, SFWB21, SOT^{+}21, TNC^{+}20, VFL20, WPS22, WMS^{+}20]. independently [MYT^{+}21, OZW^{+}21, PKC^{+}22, SPT^{+}09, SPT^{+}21, SFO^{+}21, TNLPF20]. Individual [LSD20b]. induce [AHQ20, DRC^{+}20, KMD20]. Induced [CSD22, CCH^{+}21, CLZ^{+}20, CPS^{+}22, FCT^{+}20, FIK^{+}05, FIK^{+}20, HBDC^{+}20, Hök22, Ike20, ITM^{+}21, JTM^{+}23, KIV^{+}20, KHFK^{+}20, OMI22, RZN^{+}22, SHD^{+}21, TRJ^{+}20, YSC^{+}21, ZLW23, YSC^{+}02]. induces [BZD20, DSB22, EMEZ^{+}20, FWP^{+}20, IvCD^{+}21, JKZ^{+}22, NYN^{+}21, NMO^{+}22, SPKP22]. inducible [WHA20]. inducing [RDW^{+}20]. Induction [MPVD^{+}21, HCB^{+}23, STS21]. inductive [RWSZ^{+}20]. infected [MWF^{+}23]. infection [JLS^{+}22, MWF^{+}23, SCK^{+}20b, SCK^{+}20a]. inflammasome [DHB^{+}21, SLH^{+}20a]. Inflammasomes [MNC20]. inflammation [HTL^{+}21, WM20, WAK^{+}20, ZRO^{+}23]. influences [BW20]. influx [Hök22]. information [SLM23]. ingression [SLP^{+}22]. inheritance [MSJ20, OCLB21]. inhibiting [YSC^{+}02, YSC^{+}21, ZLW23]. Inhibition [EZB^{+}20, FIK^{+}05, HGN^{+}21, HTL^{+}21, HCB^{+}23, Kin21, KST^{+}22, MVM20, SNL^{+}22, TOL^{+}20, FIK^{+}20]. inhibitor [CMM^{+}20]. inhibits [CFV^{+}21, FBR^{+}21, RDL^{+}20, ZLS^{+}21]. initial [AH20a, TOL^{+}20]. initiate [CEM^{+}20, KRHP^{+}21, LZZ^{+}21]. initiates [AT21, FCT^{+}20, RWSZ^{+}20]. initiation [BDR20, KPM^{+}22, LLA^{+}21, QZX23]. initiations [CWX^{+}21].

Innate [MP21b]. Inner [MSJ20, CMT^{+}21, GBBT^{+}22, MOK^{+}22, MP23b, OCB^{+}21, SOT^{+}21, TNLPF20]. inositol [DWA^{+}22]. Inp1 [HHD^{+}20, KWdB^{+}20]. Inp1-dependent [KWdB^{+}20]. INPP5B [DWA^{+}22]. ins [WR22]. insertion [BDH^{+}21, CLC^{+}21, KLB^{+}22]. insights [AHLR22, KWV^{+}23, LGS22, YZY^{+}20]. instability [BZD20, Gui21]. insulin [Bog21, GSP^{+}20, LJJ^{+}21, LFF^{+}22, SHD^{+}21]. insulin-induced [SHD^{+}21]. insulin/IGF1 [LJJ^{+}21]. intact [GVA20]. integrated [ZMW^{+}22]. integrates [BKR^{+}22]. integration [AMMK^{+}22]. Integrin [GGFBR^{+}22, BJSOS^{+}20, BJSOS^{+}21, CLH^{+}20, GDB^{+}20, HAL^{+}23, KST^{+}22, KBB^{+}23, LMS^{+}21, LYP^{+}21, MMDK^{+}22, SPKP22, SMC^{+}20, WXM22]. Integrin-based [GGFBR^{+}22]. integrins [LRL^{+}20, ZAK^{+}22]. integrity [AZR^{+}22, BED^{+}21, CWAT20, DDD^{+}20, GMIC^{+}20, MPKB^{+}20, MDV^{+}21, SCK^{+}19, SCK^{+}23, ZHW^{+}21, ZLJ^{+}23, vdGM22]. interacting [CHS^{+}22, FPZ^{+}22, GMB^{+}20]. interaction [AKN^{+}22, APL^{+}21, CYL^{+}20, DF22, FIK^{+}05, FIK^{+}20, LXJ^{+}23, MSX^{+}21, RLS^{+}20, XGD^{+}23, dAC^{+}22, vdGM22]. interactions [GCL^{+}21, HH21, ITM^{+}21, KGVK^{+}23, MDB^{+}20, PVYJ^{+}21, PPG21, PSP^{+}21, RCF^{+}22, SvDSW^{+}20, SLH^{+}20b, TPM^{+}21, ZXW^{+}20]. interactome [CLH21]. interacts [ASK^{+}22, CGCR^{+}22, DRC^{+}20, HLB^{+}22, LLX^{+}21, XYG^{+}23, YLC^{+}21]. interface [KHV^{+}22, LKMM^{+}23]. Interferon [RDW^{+}20]. Interferon-stimulated [RDW^{+}20]. interlocks [VOR^{+}21]. intermediate [BG22]. internalization [LMM^{+}23]. interneurons [GKM^{+}20]. Interphase [LDE^{+}22, IWI^{+}21, MHN20]. interplay [BG21]. interpretation [BGM^{+}21]. Intersection [ZMMM^{+}20]. Interviewing [CS21a]. intra [CJK^{+}22, LPMA^{+}22, PFPB^{+}20, TML22, WPCB^{+}21]. intra-axonal [LPMA^{+}22]. intra-Golgi [CJK^{+}22, PFPB^{+}20, TML22, WPCB^{+}21]. Intracellular [KMK21, LMS^{+}21, FSC22, JMC^{+}20, LLBC^{+}20, MP23c, NvGK20, SBL^{+}21, TF20]. intrachromosomal [WJW^{+}22]. intraflagellar [DSLP20]. intraphagosomal [WZK^{+}23]. intrinsic [GKM^{+}20, KBN^{+}21, NS20]. invadopodia [AO20, KLCM^{+}23, PZWW21, SPS^{+}20, ZMMM^{+}20]. invadosomes [VOR^{+}21]. invaginations [WBH^{+}21]. invasion [AO20, CKM^{+}20, FRO^{+}20, FBR^{+}21, JCL^{+}23, KIV^{+}20, PWW^{+}20, SFC^{+}23]. invasive [LYP^{+}21]. invasiveness [KCP^{+}21, RPM^{+}21]. invertebrates [Pro20]. investigators [CS21c]. involved [OTOF21, RBL22]. involves [KWdB^{+}20]. involving [SGW^{+}20]. ion [CSM^{+}21]. IPO11 [MOS^{+}20]. IQGAP [TRHS23]. IRE1 [GCS^{+}20, GLM^{+}22, LGS22, ZLW23, HYX^{+}20]. IRE1-induced [ZLW23]. IRF8 [DHB^{+}21]. IRF8-mediated [DHB^{+}21]. iron [ABB^{+}22]. iron-dependent [ABB^{+}22]. IRSp53 [FLJ^{+}22]. ISG15 [MV20]. isotype [WM23]. isotypes [NBC^{+}21]. Ist1 [LCM22]. Ist2 [WYL21]. IV [GKRL^{+}23]. IZUMO1 [BW23, BNV^{+}23].

J [FAS^{+}21]. J-protein [FAS^{+}21]. JAK1 [ASK^{+}22]. JAM [KST^{+}22].

JAM-A-tetraspanin- [KST^{+}22]. JIP3 [CGCR^{+}22]. JNK [HBDC^{+}20]. Joachim [TB20b]. join [ME21]. jointly [PSS^{+}20]. Jou [MP23a]. journey [SS22]. Judith [MP23b]. junction [CLL^{+}21a, ESB^{+}21, HSF^{+}23, OHY^{+}20, PVYJ^{+}21, VCS^{+}22]. junctional [MHGM22, MDV^{+}21]. junctions [BRB^{+}20, CHS^{+}22, ORCT^{+}20, SMS^{+}20, YKSC^{+}22]. juxtaposed [GKRL^{+}23].

K63 [JFM^{+}22]. Karyopherin [KKZ^{+}22]. karyotypic [SRW^{+}21]. KASH5 [GSL^{+}23]. Katanin [JBV^{+}20, SCL^{+}21]. KDM5A [KSP^{+}21]. keep [MA20, MP22h, MYM^{+}21]. Keeping [GH20]. keeps [Low21]. key [GGA21, KWV^{+}23]. kidney [BED^{+}21]. KIF13A [GLGL^{+}21]. KIF14 [PRB^{+}20]. Kif18a [SMD^{+}21]. KIF1A [BJR^{+}21, HH21]. KIF4 [WMM^{+}23]. KIF4A [PCGB20]. KIF5A [FPMS^{+}21]. KIF5A/KLC1 [FPMS^{+}21]. killer [CKR^{+}20, POL^{+}20]. Killing [Tai22, FGBD^{+}21]. kinase [BHK20, BDD20, CRZ^{+}21, CSS20, DLZ^{+}20, LZC^{+}20, OZW^{+}21, PGW^{+}21, TNC^{+}20, TSP21, VHPP^{+}20, WB20, WYG^{+}20, ZMW^{+}22, ZBY^{+}21]. kinase-independent [TNC^{+}20]. kinases [HL21, LRB^{+}22, MC21, PKY^{+}20, ZLJ^{+}22]. kinectin [GMB^{+}20]. kinectin-1 [GMB^{+}20]. Kinesin [KMW20, BJR^{+}21, CGCR^{+}22, CPW^{+}23, NVPP20, HLB^{+}22, PPB^{+}21, QZX23]. Kinesin- [KMW20]. kinesin-1 [CGCR^{+}22, HLB^{+}22, QZX23]. Kinesin-13 [PPB^{+}21]. kinesin-3 [BJR^{+}21]. kinesin-like [NVPP20]. kinetics [STvT23]. Kinetochore [CSS20, SKN^{+}21, ARCM20, BDD20, CRZ^{+}21, CWN^{+}23, DKCT21, GOR^{+}20, HLGD20, KMW20, KHV^{+}22, LSD20b, PGH^{+}23, RCA^{+}23, RSB^{+}23, SWS^{+}21a, VVW^{+}23]. Kinetochore-bound [SKN^{+}21]. kinetochore-fibers [LSD20b]. kinetochore-microtubule [ARCM20]. kinetochores [BDT^{+}22]. Kip2 [CPW^{+}23]. KLC1 [FPMS^{+}21]. knockouts [PBPBS22]. Kulathu [MP23c]. Kv1 [KGVK^{+}23].

L [SLS^{+}23]. lab [CS21b, CS21d]. Label [SLD^{+}21]. Label-retention [SLD^{+}21]. labeling [NGG^{+}20]. Lack [XVW^{+}23, ZMS^{+}20]. LAM [MVM20]. LAM-family [MVM20]. lamellipodia [HCRMTC23, KBH^{+}22, OHY^{+}20]. Lamellipodin [MRWK^{+}22]. lamin [KAS^{+}22]. lamina [ITB^{+}23, TPM^{+}21]. lamins [KST^{+}21]. large [CWAT20, CLR^{+}20, MA20]. late [EMY^{+}22]. laterally [FBR^{+}21]. Lattice [Bak23]. lattices [HAL^{+}23, vdBVS^{+}23]. layered [SGN^{+}20]. layers [HRB^{+}21]. LC3 [FCHM20, HJL^{+}22]. LC3-associated [HJL^{+}22]. LC3B [FWP^{+}20, KJ23]. LC3C [BZC^{+}21]. LD [RGK^{+}22, ZDM^{+}22]. lead [RS22]. leads [AAR^{+}21, AII^{+}21, RLV^{+}20, Tev20, VZQ^{+}21]. leakage [KKZ^{+}22]. learning [DES^{+}23, SHA20]. leave [PF21, SS22]. leaves [Ove21]. lectin [NR22]. Leep1 [YLC^{+}21]. Legionella [MAW^{+}22]. LEM2 [PSS^{+}20]. LEM2/CHMP7 [PSS^{+}20]. length [ATS^{+}21, CVMB^{+}23, KNA^{+}22, SCK^{+}19, SCK^{+}23]. lens [LRL^{+}20]. lesions [KSP^{+}21]. Let [BP20]. lethality [MNvdS^{+}20, NMO^{+}22]. Letting [BP22]. leukocyte [CW23, GPW^{+}22]. level [DZA^{+}20]. levels [EBZC^{+}21, LFF^{+}22, SFC^{+}23, VDC^{+}20]. LGI3 [KGVK^{+}23]. LGI3/2 [KGVK^{+}23]. LI [WXM22]. library [SHA20]. LIC1 [KKP^{+}21]. licenses [JFM^{+}22]. life [MNvdS^{+}20, MRA20, SSB20]. lifespan [AH20b, LMJ^{+}20, XDY^{+}22]. ligand [ESH^{+}23, GSP^{+}20]. ligand-saturated [GSP^{+}20]. ligands [KSS^{+}20a, WPM21]. ligase [AHY^{+}21, DMR^{+}20, LSD^{+}21, PE22, SSF^{+}22, ZCL^{+}22, ZSJE20]. ligases [BMM^{+}20, TSL^{+}20]. ligation [TPM^{+}21]. light [BSB^{+}21, Dri20, FBVD^{+}22, LSS^{+}23, Tai22]. light-regulated [BSB^{+}21]. LIKE [GCL^{+}21, BS20b, BHK20, CGK^{+}22, CSS20, FC21, GBBT^{+}22, LLW^{+}20, MSC^{+}20, NVPP20, OZW^{+}21, PMSO^{+}23, SWT^{+}22, SCL^{+}21, TWY^{+}22, WYG^{+}20, WCL^{+}23]. limit [CW23]. limited [MND^{+}20, SWN^{+}22]. limiting [BED^{+}21]. limits [ARCM20, LYP^{+}21, PHT^{+}23, SKS^{+}23]. LINC [GSL^{+}23]. Lineage [LJT^{+}22, WDL^{+}20]. lines [SDD^{+}22]. link [MH22]. linkage [GKRL^{+}23, PVYJ^{+}21]. linked [UTR^{+}23]. linker [AHQ20]. linking [DdCVT22, EMEZ^{+}20]. links [DSLP20, LML^{+}21, LSOM23, MLQ^{+}21, MOK^{+}22, MFC^{+}20, MMDK^{+}22, RSWP20, WB20]. Linton [AO21]. Lipid [GCL^{+}21, WC22, AHLR22, CYR^{+}21, CT20, CEM^{+}20, DZA^{+}20, DY21, Goo20, GMCO^{+}22, HCWX^{+}22, HSW^{+}22, HYX^{+}20, HAW^{+}22, ITB^{+}23, JMY^{+}23, LM23, LHS^{+}22, MYT^{+}21, PKA20, PSS^{+}20, RP21, RE20, RGK^{+}22, SvVV^{+}23, SMM^{+}21, SOT^{+}21, WYG^{+}20, WYL21, ZY21, ZHW^{+}21, ZAR^{+}21, ZDM^{+}22]. lipidation [FWP^{+}20, FCHM20]. lipids [Cas21, CT20, Köh21, LLLR20, PRMF^{+}23, RCM^{+}23a, RCF^{+}22]. lipogenesis [LML^{+}21]. lipolysis [RGK^{+}22]. lipoprotein [BMS^{+}22]. liposomes [JMY^{+}23]. Liquid [CAS23b, NWZ20, PTS^{+}22, ZFZ^{+}23, CPC^{+}20, Dor20, KMK21, LLA^{+}21, RGK^{+}22, WC22, ZVM^{+}20]. LIR [ZRO^{+}23]. LIR-dependent [ZRO^{+}23]. Live [UIS^{+}22, JIBK23, KLC^{+}20, MP21d, MP22d, dCS^{+}21]. live-cell [JIBK23, KLC^{+}20]. lived [BWEHS21]. living [FLW^{+}23]. LMX1B [JMKS^{+}23, KJ23]. LMX1B-autophagy [JMKS^{+}23]. LMX1B-mediated [KJ23]. LNCcation [BDK21]. lncRNA [AMG^{+}20, BDK21]. load [FPZ^{+}22]. loaded [MNC20]. loading [DCRDC^{+}22, WAK^{+}20]. lobe [AR20, SLES20]. Local [BSH^{+}22, NYN^{+}21, NBI^{+}22, UTR^{+}23]. localization [AGH^{+}22, BDK21, CSG22, DDD^{+}20, DHTP22, FAS^{+}21, FY20, MLvdL^{+}21, RFL20, SSF^{+}22, TML22]. localizations [SHA20, WDRRF^{+}23]. localized [HRS^{+}20, LZC^{+}20, SHLS22, Zar20]. localizes [DZA^{+}20, ZCL^{+}22]. locally [LSD20b, ZLS^{+}21, Zar20]. located [Let20]. locomotion [KST^{+}22]. Long [BWEHS21, MTD20, MP22d, GLGL^{+}21, LLLR20, MBG^{+}23, MPVD^{+}21]. Long-lived [BWEHS21]. long-range [MBG^{+}23]. long-term [GLGL^{+}21, MPVD^{+}21]. longer [MP21d]. loop [FCHM20, KIV^{+}20, KTT^{+}22, LJJ^{+}21]. Lose [Set21]. Loss [BZD20, DZA^{+}22, WLM^{+}21, KNiY^{+}21]. lost [DG22]. low [BMS^{+}22, PDW^{+}20, PGH^{+}23]. low-density [BMS^{+}22]. low-tension [PGH^{+}23]. loyal [VM21]. LPHN2 [CFV^{+}21]. Lrp1 [ZTL^{+}23]. LRRK2 [SSF^{+}22]. LTB [SMC^{+}20]. Ltc1 [MVM20]. Ltc1-dependent [MVM20]. LUBAC [SYQ^{+}22]. lumen [PMB^{+}20]. lumenal [CMT^{+}21]. lumina [BRD^{+}21]. luminal [CG21, LRM^{+}20, vdBVS^{+}23]. lung [AMG^{+}20, MTD20, PHAM^{+}20]. LUTI [VGO^{+}23]. LUTI-mediated [VGO^{+}23]. LUZP1 [GSC^{+}20]. lymphatic [DSG^{+}23]. Lymphocyte [RCA^{+}21]. lysate [FMY^{+}21]. Lysine [ALPH20]. lysolipid [PRMF^{+}23]. Lysosomal [HZCX22, ZLJ^{+}23, DCG^{+}23, EEW^{+}22, JWB^{+}22, LHS^{+}22, NGG^{+}20, RCS22, WKC^{+}22, dCTOG^{+}20]. Lysosome [YW21, HCWX^{+}22, KRS21, SNL^{+}22, WCG^{+}22, XZJ^{+}21, YJX^{+}20, ZLJ^{+}23]. Lysosomes [MP22f, AH20b, ATTF20, FCT^{+}20, LFD^{+}21, LFF^{+}22, RCM^{+}23b, ZXY^{+}23]. lytic [POL^{+}20].

M [MP22f, SGL^{+}23]. M2 [Nag23]. M6PR [RCM^{+}23b]. mac [Nag23]. mac-in-touch [Nag23]. machine [SHA20]. machineries [AGW^{+}20, NBI^{+}22, dDFGP^{+}21]. machinery [BSH^{+}22, JKL^{+}22, NR22, VFL20, YTH^{+}20, ZHHJ22]. macroferritinophagy [OYS^{+}22]. macroH2A [KSP^{+}21]. macromolecular [CWAT20]. macrophage [SLS^{+}23]. macrophage-mediated [SLS^{+}23]. macrophages [BG21, EJBB^{+}20, LJT^{+}22, MWF^{+}23, MRWK^{+}22]. macropinocytosis [HCB^{+}23, YLC^{+}21]. macropinosome [LYP^{+}21]. MAD1 [JMB^{+}20, CSOG^{+}20, HLGD20]. main [MdCT23]. maintain [GCNL21, HMT^{+}21, HSF^{+}23, IWI^{+}21, LSD20b, WYL21]. maintaining [AAF^{+}20, RCS22]. maintains [AZR^{+}22, CSM^{+}21, CZTL21, DSG^{+}23, LFD^{+}21, MdCT23, RDL^{+}20, SNL^{+}22, WM23, XGD^{+}23, vdGM22]. maintenance [GMCO^{+}22]. MAIT [LWG^{+}22]. maize [MDB^{+}20]. MakA [JKZ^{+}22]. make [PVYJ^{+}21]. makes [ASC20, Cas22]. Mammalian [BW23, FSC22, FMY^{+}21, FHM^{+}20, KRC^{+}22, LSD20b, OWY^{+}23, PSA^{+}23, UZS^{+}23, WLM^{+}21, YSR^{+}21]. manganese [CKW^{+}22]. manner [SOT^{+}21]. map [EEW^{+}22]. MAP1LC3C [BZC^{+}21]. MAPK [GC22, SWS21b, WTS^{+}21]. MAPK11 [MLQ^{+}21]. MAPK11/14 [MLQ^{+}21]. mapping [TPM^{+}21, Dri20]. MARCH5 [PE22, ZCL^{+}22]. Marilyn [SSB20]. Mark [AO21, Cas22]. marker [JIBK23]. marrow [BCS^{+}21]. mass [ABM^{+}23, DSMB20, NGG^{+}20]. master [EMY^{+}22, WR22, WESR22]. mastigonemes [LLW^{+}20]. MASTL [TNC^{+}20]. Material [WCG^{+}22, MTR^{+}20]. maternal [MYM^{+}21]. mathematical [DES^{+}23]. mating [WPS22]. Matriptase [AHvR^{+}20]. matrix [JKL^{+}22, RS22, WMS^{+}20]. Matsunaga [MP22g]. maturation [CCV^{+}21, CKM^{+}20, FY20, GGA21, HMSF22, KNA^{+}22, MYK^{+}20, MPVD^{+}21, NPdC^{+}21, PHMD20, MYK^{+}21, MYK^{+}22]. maturation-dependent [CCV^{+}21]. mature [WHE^{+}22, YKSC^{+}22, ZBM^{+}22]. may [MDB^{+}20]. MCAK [FOR^{+}20, LSOM23]. McLeod [MH22]. mcp-1 [SSO^{+}20]. me [Bak23]. means [BBM^{+}23, TSP21]. measurement [LQS23]. Mechanical [BGM^{+}21, BBM^{+}23, KPM^{+}22, KWGR23]. mechanics [SMHH^{+}20]. Mechanism [VGK^{+}21, AHY^{+}21, AHvR^{+}20, BD20, CVMB^{+}23, KSS^{+}20a, KTT^{+}22, MSR^{+}20]. Mechanisms [RP21, BJR^{+}21, ML22, ME21, NBC^{+}21]. Mechanistic [AGH^{+}22, ZHHJ22]. mechanoactivation [JML^{+}21]. mechanoresponse [MDV^{+}21]. mechanosensing [MMDK^{+}22, SGL^{+}23]. Mechanosensitive [DSG^{+}23, LRL^{+}20, VCS^{+}22]. mechanosensitivity [GPEC^{+}23]. mechanosensory [SCL^{+}21]. mechanotransducers [DCS^{+}20]. mechanotransduction [BB20]. mediate [ASK^{+}22, BW23, CJK^{+}22, HVPM20, HHGR21, KPA^{+}16, KPA^{+}20, KSP^{+}21, LMS^{+}21, MTCL^{+}23, RWSZ^{+}20, YMH^{+}20, YKSC^{+}22]. mediated [AGH^{+}22, AAR^{+}21, AHQ20, BJSOS^{+}20, BJSOS^{+}21, yLHW^{+}20, CMM^{+}20, CHZ^{+}20, CSOG^{+}20, DCRDC^{+}22, DHB^{+}21, DSB22, EZB^{+}20, GKRL^{+}23, GCW^{+}23, HSU^{+}20, HAW^{+}22, IvCD^{+}21, KKPH^{+}21, KTT^{+}22, KJ23, LLC^{+}20, LJJ^{+}21, LL22, LM23, LGL^{+}23, NTA^{+}21, NSB^{+}21, OTOF21, OCB^{+}21, PWW^{+}20, PHMD20, PAS^{+}22, SKX^{+}23, SLS^{+}23, SNN20, SGW^{+}20, VGK^{+}21, VZQ^{+}21, WJW^{+}22, WJL^{+}23, WLBS20, VGO^{+}23, ZLS^{+}21, ZDGB^{+}22, ZTL^{+}23, ZVL^{+}23]. mediates [CCFN^{+}20, CGK^{+}22, CKM^{+}20, DLZ^{+}20, EBZC^{+}21, FY20, HCL^{+}21, Hök22, HRB^{+}21, ITM^{+}21, KMK21, LSX^{+}22, MDB^{+}20, MOS^{+}22, MOS^{+}20, RCM^{+}23b, ZFZ^{+}23, dCTOG^{+}20]. mediating [SJL^{+}22]. meeting [MS20]. meiosis [CLR^{+}20, DPM^{+}20, OCLB21, STY^{+}20, TP20, VV23, WLM^{+}21, ZJH22]. Meiotic [KWV^{+}23, VGO^{+}23, BHK20, CML20, FAMQW22, GSL^{+}23]. melanocyte [BS20b]. melanogaster [MdCT23]. melanosomes [ZLJ^{+}22]. Membrane [AO20, HZN^{+}21, MS20, SRK22, SPT^{+}09, AHY^{+}21, BLU21, BSC^{+}23, CSD22, CMT^{+}21, CSM^{+}21, CWAT20, CWKP23, CLC^{+}21, CM21, DLZ^{+}20, DCG^{+}23, EYC^{+}20, FUBS22, FCT^{+}20, FC21, FWP^{+}20, GKRL^{+}23, GCL^{+}21, GCW^{+}23, GBBT^{+}22, HSW^{+}22, HHD^{+}20, JWB^{+}22, KSN^{+}22, KWdB^{+}20, KMK21, LML^{+}21, LCM22, LM23, LYL^{+}22, MTCL^{+}23, MND^{+}20, MLS20, MWSX23, MMKM21, OWY^{+}23, OCB^{+}21, PCZ^{+}23, PZWW21, PPG21, PHT^{+}23, RCH^{+}20, RBL22, SLP^{+}22, SOT^{+}21, SBV^{+}20, SWS21b, TNLPF20, UTR^{+}23, VBG^{+}22, WHN^{+}21, WLBS20, WESR22, WCC^{+}23, WCL^{+}23, YZW^{+}20, ZMS^{+}20, ZLJ^{+}23, ZSJE20, SPT^{+}21]. membrane-anchored [AHY^{+}21]. Membrane-bound [HZN^{+}21, KMK21, PZWW21]. membranes [JDKK^{+}22, LLLR20, MOK^{+}22, MP22e, TTM^{+}21]. mentor [VM21]. mesenchymal [DCS^{+}20]. messages [MP23c]. metabolic [HCB^{+}23, PKH^{+}20, SRUdC^{+}22]. metabolism [LL22, LKMM^{+}23, RZN^{+}22, SSO^{+}20, SLM20, TMG^{+}21, WYG^{+}20, WYL21, ZWJ22]. metabolite [BVYW20, EM22]. metaphase [KMW20, PKY^{+}20, PGH^{+}23]. metastasis [Cas23a, CKR^{+}20, SFC^{+}23, SPS^{+}20]. metastasis-promoting [CKR^{+}20]. method [TPM^{+}21, UZS^{+}23]. metrics [FBVD^{+}22]. MICAL1 [HVPM20]. mice [SMD^{+}21]. MICOS [TNLPF20]. microautophagy [YZW^{+}20]. microenvironment [NS20]. microferritinophagy [OYS^{+}22]. microglia [DRW^{+}23, MP22b]. Micron [YKSC^{+}22]. Micron-scale [YKSC^{+}22]. micronuclear [SMD^{+}21]. Micronuclei [SMD^{+}21]. micropatterning [LM21, WAOS^{+}21]. microscope [LQS23]. microscopy [FBVD^{+}22, GMD^{+}23, LYL^{+}22, LSS^{+}23, MLvdL^{+}21, RMM^{+}21, SLD^{+}21, VLdRADJ22, WBR^{+}20, vdBdHLK22]. Microtubule [FDA21, Mer21, TG21, ARCM20, AII^{+}21, BWA^{+}23, CRZ^{+}21, CPW^{+}23, CVT^{+}21, DdCVT22, DKCT21, EMEZ^{+}20, FAHZ21, FPZ^{+}22, GOR^{+}20, HBDC^{+}20, HCRMTC23, JBV^{+}20, KHV^{+}22, LDE^{+}22, MRL^{+}21, MDB^{+}20, MSX^{+}21, NYN^{+}21, OZW^{+}21, ORCT^{+}20, PGH^{+}23, PMB^{+}20, PPB^{+}21, RCA^{+}23, RGP^{+}22, SHBF^{+}20, SKN^{+}21, STY^{+}20, VFL20, ZHHJ22, ZVL^{+}23, vdBVS^{+}23]. microtubule- [HCRMTC23]. Microtubule-associated [TG21]. microtubule-independent [VFL20]. microtubule-nucleating [OZW^{+}21]. microtubule-severing [JBV^{+}20]. Microtubules [KRH^{+}20, MA20, RMA21, TEH^{+}20, CYL^{+}20, EYC^{+}20, FSC22, Gar21, Hic22, JIBK23, KLC^{+}20, MW21, RRCS^{+}23, RLS^{+}20, WDJ^{+}21]. microvascular [LWL^{+}23]. microvilli [BEM^{+}23]. Mid1 [MSC^{+}20]. Mid51 [WKC^{+}22]. Mid51/Fis1 [WKC^{+}22]. midbody [HL21, Hic22, HESH^{+}22]. midbrain [JMKS^{+}23]. midgut [LTL^{+}20]. migrasome [FSZ^{+}22]. migrating [KRH^{+}20, LYL^{+}23]. migration [ASK^{+}22, BCC^{+}21, BKR^{+}22, CDD^{+}22, DHTP22, EJBB^{+}20, FSZ^{+}22, FRO^{+}20, GGFBR^{+}22, GY20, LMS^{+}21, LYP^{+}21, LDH^{+}21, MGM22, OHY^{+}20, RCA^{+}21, RS22, TEH^{+}20, WXM22, WHE^{+}22, XVW^{+}23, YLC^{+}21, ZAK^{+}22]. MIM [PLG^{+}23]. MIM/MTSS1 [PLG^{+}23]. mimics [CPC^{+}20]. Mind [MP22e]. Minibrain [PGW^{+}21]. ministacks [TML22]. miR [TMG^{+}21]. miR-146 [TMG^{+}21]. miRNA [WAK^{+}20]. miRNAs [MNC20]. Miro [GSLH^{+}21]. misaligned [FDSR22]. misinsertion [PHT^{+}23]. mislocalized [MOS^{+}22]. missegregation [DG22, FDSR22]. mitigate [LW20b]. Mitochondria [DRZ^{+}23, TF20, ASC20, APL^{+}21, BPF^{+}21, BC23, CL21, CCH^{+}21, DSB22, EBZC^{+}21, ESX^{+}20, GSLH^{+}21, GMCO^{+}22, Ike20, MOS^{+}22, SvVV^{+}23]. Mitochondrial [AH20b, CFK^{+}22, CWZ^{+}20, Gan21, HRS^{+}20, IMR^{+}23, LM23, LMJ^{+}20, AGW^{+}20, BPF^{+}21, BWEHS21, CGK^{+}22, CCH^{+}21, CLL^{+}21b, DRZ^{+}23, ESX^{+}20, GCW^{+}23, HAW^{+}22, LML^{+}21, LGL^{+}23, MRG^{+}20, SvVV^{+}23, TNLPF20, WKC^{+}22, XDY^{+}22, YKK^{+}20, ZJDR22]. mitochondrial-derived [ESX^{+}20]. Mitoguardin [HAW^{+}22]. Mitoguardin-2 [HAW^{+}22]. Mitophagy [Ike20, KPG20, GCW^{+}23, McW23, OCB^{+}21, SFWB21, YKK^{+}20]. mitosis [HHT^{+}20, KSS^{+}20b, KSS^{+}20c, NBC^{+}21, RLS^{+}20, RFL20, VHPP^{+}20]. Mitotic [YTH^{+}20, CML20, CSS20, DES^{+}23, DOA^{+}22, DDD^{+}20, FOR^{+}20, FDA21, GCNL21, GNL^{+}20, INM^{+}21, KHV^{+}22, KKP^{+}21, LKW^{+}21, LDE^{+}22, MSB^{+}21, MKO^{+}21, MTR^{+}20, MDV^{+}21, PSP^{+}21, SMHH^{+}20, SKS^{+}23, VZQ^{+}21, WMS^{+}20, Zar20]. mix [LC20]. Mklp1 [SRK22]. MKLP2 [SBEB20]. MKLP2-dependent [SBEB20]. MLF2 [RLV^{+}20]. mmBCFA [ZHW^{+}21]. MMP [PWW^{+}20, SPS^{+}20]. MMP14 [HVPM20]. MMPs [PFPB^{+}20]. modality [SFC^{+}23]. model [Bez22, CT20, SDD^{+}22]. modeling [DES^{+}23]. models [CS20, HKK^{+}20]. modes [DKCT21]. modification [YM21]. modified [LML^{+}21, dCTOG^{+}20]. modulate [GSC^{+}20, WZtM^{+}20]. modulated [YPM^{+}21, ZJDR22]. modulates [FRO^{+}20, GC22, LLC^{+}20, PRMF^{+}23, WTS^{+}21]. Modulation [NS20, TWT20]. modulator [DMR^{+}20]. modulators [CLL^{+}21a]. module [PL22]. Molecular [BJAR^{+}21, SBEB20, MLvdL^{+}21, MSR^{+}20]. molecule [CLL^{+}21a, KHFK^{+}20, LXJ^{+}23, MLvdL^{+}21]. Mon1 [HMSF22]. monitor [UTR^{+}23]. monocytes [ESH^{+}23]. monolayer [RE20]. monopolar [WZZ^{+}23]. monotopic [FUBS22]. morphogenesis [CG21, DDD^{+}20, DYW^{+}20, HSSK20, MBA^{+}22, PVYJ^{+}21, QLC^{+}20, SGL^{+}23]. morphology [AGW^{+}20, DSG21, HZN^{+}21, HAW^{+}22, WMM^{+}23]. morphometrics [BMF^{+}23]. morphometry [KRC^{+}22]. morphotypes [SMFC^{+}22]. MOSPD2 [ZDM^{+}22]. mother [SYQ^{+}22, VHPP^{+}20]. motif [HSW^{+}22, KHB^{+}22]. motile [GVA20]. motility [CCV^{+}21, Cas22, HLB^{+}22, KBH^{+}22, KAH^{+}21, NTA^{+}21, SHBF^{+}20, TNC^{+}20, VFL20]. motor [BJR^{+}21, CPW^{+}23, TEH^{+}20, WKX^{+}21]. motors [FPMS^{+}21, KRS21]. mount [SGW^{+}20]. mouse [BWEHS21, MSX^{+}21, SPL^{+}20]. move [DRZ^{+}23]. movement [BJR^{+}21, FIK^{+}05, FIK^{+}20, LSX^{+}22, WDJ^{+}21, WCL^{+}23]. Moving [BD20, CS21b]. Mps1 [CSOG^{+}20, SKN^{+}21]. Mps1-mediated [CSOG^{+}20]. MR1 [LWG^{+}22, PK22]. MRCK [ZGR^{+}22]. MreB [PMB^{+}22]. mRNA [FY20, MMSP20, PFS^{+}22, PSC^{+}20]. mRNAs [Zar20]. mRNP [CLH21]. MS [DSMB20]. Msp300 [TRJ^{+}20]. Msp300/Nesprin [TRJ^{+}20]. Msp300/Nesprin-1 [TRJ^{+}20]. mt [XDY^{+}22]. MT1 [PWW^{+}20, SPS^{+}20]. MT1-MMP [PWW^{+}20, SPS^{+}20]. MTCH2 [LML^{+}21]. mtDNA [DSB22, HCWX^{+}22, KMJ^{+}23, RPM^{+}21]. mtDNA-containing [RPM^{+}21]. mtDNA-dependent [HCWX^{+}22]. mTOR [JWB^{+}22]. mTORC1 [BMS^{+}22, DSG^{+}23, RRBW^{+}21, ZWJ22]. mTORC2 [OMI22]. MTSS1 [PLG^{+}23]. multi [VVW^{+}23, ZJDR22]. multi-color [VVW^{+}23]. multi-factor [ZJDR22]. multi-span [ZJDR22]. multiciliogenesis [LNY^{+}22]. multiorganelle [HH22]. Multiple [BB20, WHE^{+}22, BCM^{+}22, BDD20]. multiplexed [CLH^{+}20]. Multivalent [PVYJ^{+}21, ZXW^{+}20, SLH^{+}20b]. muscle [AZR^{+}22, LHL^{+}23, LLX^{+}21, SCB^{+}20]. muscular [HGG^{+}23, RH23]. musketeers [JJ23]. must [Ver21]. mutant [SMD^{+}21, VGK^{+}21]. mutation [BJSOS^{+}20, CWZ^{+}20, YZY^{+}20, BJSOS^{+}21]. mutations [BJR^{+}21]. Mutual [JMC^{+}20]. MVB [SJL^{+}22]. Myc [YSC^{+}02, YSC^{+}21, SLL^{+}21, SLL^{+}23]. MYC-driven [SLL^{+}21, SLL^{+}23]. MyD88 [DACG^{+}21]. Myddosome [DACG^{+}21]. myelinated [BS20a, KGVK^{+}23, MPKB^{+}20]. Myelination [DRW^{+}23, HCL^{+}21, WKX^{+}21]. myofibers [FCCH21]. myogenic [KKPH^{+}21]. myosin [yLHW^{+}20, DLK^{+}21, HCRMTC23, YKSC^{+}22]. myotubularin [AAF^{+}20]. myotubularin-related [AAF^{+}20].

N [HMT^{+}21, RDL^{+}20, SFC^{+}23, SBL^{+}21]. N-cadherin [SFC^{+}23]. N-end [RDL^{+}20]. N-terminal [SBL^{+}21]. NAD [SRUdC^{+}22]. Naegleria [BD20, VFL20]. Naips [DHB^{+}21]. Nan [MP21b]. Nanoscale [CWN^{+}23, CSD22, LQS23, PFS^{+}22, WDRRF^{+}23]. Nanoscopy [SGN^{+}20]. nanostructure [VVW^{+}23]. nanovesicles [LMS^{+}21]. nascent [ABM^{+}23]. natural [CKR^{+}20, POL^{+}20]. navigation [GPW^{+}22]. Navigator [SHBF^{+}20]. Navigator-1 [SHBF^{+}20]. NBAS [WLW^{+}22]. NBR1 [RKLJ22]. NCAM [HYL^{+}20]. NCOA4 [OYS^{+}22, WZ22]. NDC80 [SKS^{+}23, FPZ^{+}22, RCA^{+}23, SKN^{+}21]. Ndc80-Cdt1-Ska1 [RCA^{+}23]. near [FDG^{+}21]. neat [MP22h]. necessary [CFK^{+}22]. Necl [FIK^{+}20, FIK^{+}05]. Necl-5 [FIK^{+}20, FIK^{+}05]. Necroptosis [Pie20, SdCS^{+}22]. necroptotic [KMD20]. nectin [FIK^{+}05, FIK^{+}20]. nectin-3 [FIK^{+}05, FIK^{+}20]. NEDD1 [CWX^{+}21]. needed [ARM23]. negatively [GCS^{+}20]. Nek2 [AHQ20, VHPP^{+}20]. Nek2-mediated [AHQ20]. Nem1 [CEM^{+}20]. nematode [YMH^{+}20]. neo [CD21]. neocentromere [MPVD^{+}21]. nerves [LCB^{+}23]. Nesprin [GL20]. Nesprin-1 [TRJ^{+}20]. Nesprin-2G [GL20]. Nesprins [DCS^{+}20]. network [CVT^{+}21, GMD^{+}23, GCNL21, OYJJ23, RBL22, WKC^{+}22]. networks [BB20, EYC^{+}20, Sir23, WBR^{+}20]. Neur [AR20]. neural [AR20, Bez22, CPS^{+}22, GMD^{+}23, HYL^{+}20, SLES20, STS21, WDL^{+}20]. Neuralized [CGK^{+}22, SLES20]. Neuralized-like [CGK^{+}22]. neurexin [KSS^{+}20a]. neurites [KWGR23]. NEURL4 [CGK^{+}22]. neurodegenerative [HKK^{+}20]. neuroepithelial [LXJ^{+}23]. neurofascin [AH20a]. neurofascin-186 [AH20a]. neuroinflammatory [KMD20]. neuromuscular [ORCT^{+}20]. neuron [LCB^{+}23, YMAS20]. Neuronal [HDW^{+}21, MPKB^{+}20, RCS22, SLM20, BSB^{+}21, JMKS^{+}23, KKN^{+}21, SSO^{+}20, Tar21, TEH^{+}20, WMM^{+}23, WKX^{+}21]. neurons [GKFR20, KJ23, LYS^{+}20, PLG^{+}23, PPB^{+}21, RCM^{+}23a, STS21, XZJ^{+}21]. neurotransmitter [LLK^{+}21]. Neutral [CT20, RCF^{+}22]. neutrophil [BKR^{+}22, SS22, SMC^{+}20]. next [MP22b]. NF [HKK^{+}20]. NF- [HKK^{+}20]. niche [LD21]. NKCC1b [MPKB^{+}20]. NLRC4 [DHB^{+}21]. NLRP3 [SLH^{+}20a]. NMDA [PPG21]. nociceptor [IvCD^{+}21]. node [BJAR^{+}21]. nodes [OMK^{+}22, TRHS23]. noisy [STvT23]. nomenclature [BOW^{+}22]. Non [AMMK^{+}22, Cas22, RGP^{+}22, AT21, HJL^{+}22, JKZ^{+}22]. Non-canonical [Cas22, AT21, HJL^{+}22, JKZ^{+}22]. Non-catalytic [RGP^{+}22]. Non-G1 [AMMK^{+}22]. Non-G1/G0 [AMMK^{+}22]. Noncanonical [WYG^{+}20, BZC^{+}21, SGW^{+}20]. noncanonically [NVPP20]. noncoding [MTD20, WTS^{+}21]. nonlytic [dCTOG^{+}20]. nonspecifically [CLR^{+}20]. nonstop [BCC^{+}21]. nonvesicular [RP21]. NOT-LIKE-DAD [GCL^{+}21]. Notch1 [RWSZ^{+}20]. Notch2 [RWSZ^{+}20]. Novel [BJPH^{+}20, RE20, BNV^{+}23, CJS^{+}21, DMR^{+}20, ESH^{+}23, HČK^{+}20, LLK^{+}21, RCH^{+}20, WI22]. novo [NPdC^{+}21]. Nox [CW23]. NPC [VV23]. NPHP [PL22]. Nrf2 [CW23]. NRG3 [AVC^{+}22]. NS1 [CLZ^{+}20]. NS1-induced [CLZ^{+}20]. NUCKS1 [MSH^{+}20]. Nuclear [ANRS^{+}20, DOA^{+}22, Köh21, LW20b, RSB^{+}23, SOT^{+}21, CL21, CPC^{+}20, CMT^{+}21, CHZ^{+}20, CSOG^{+}20, DNVP23, DdCVT22, DSMB20, DY21, DRC^{+}20, GBBT^{+}22, GG20, HDW^{+}21, JMB^{+}20, KKZ^{+}22, KVG^{+}20, KWV^{+}23, KST^{+}21, KAS^{+}22, KYR^{+}22, LSD^{+}20a, LD20, ML22, MOS^{+}20, MOK^{+}22, MP22g, PSS^{+}20, PRMF^{+}23, PSP^{+}21, RLV^{+}20, SMHH^{+}20, TTM^{+}21, TKK^{+}20, TPM^{+}21, WLBS20, YLH^{+}22, YM21]. Nuclear-enriched [RSB^{+}23]. nucleate [Gar21]. nucleating [OZW^{+}21]. nucleation [AII^{+}21, Mer21, RMA21]. nuclei [AMFW^{+}21]. Nucleobindin [PFPB^{+}20]. Nucleobindin-1 [PFPB^{+}20]. nucleocytoplasmic [KKZ^{+}22]. nucleoli [FMY^{+}21]. Nucleoplasmic [KAS^{+}22]. nucleoporin [TKK^{+}20]. nucleoporins [CPC^{+}20, Dor20]. nucleosome [dCS^{+}21]. nucleotide [BWA^{+}23, PMB^{+}22]. nucleus [EBZC^{+}21, ITB^{+}23, MOK^{+}22]. NuMA [SRK22, SMHH^{+}20]. NuMA/dynein [SRK22]. NuMA1 [AH20a, TOL^{+}20]. Numb [FFZ^{+}22]. number [AMG^{+}20, ZJH22]. numbers [BTF^{+}20]. Nup188 [VDC^{+}20]. nurse [AMFW^{+}21]. nutrient [HI21, LRL^{+}20]. nutrients [CFD^{+}20].

O [YM21]. O-GlcNAc [YM21]. Object [WBR^{+}20]. objects [CLR^{+}20]. occupancy [KHV^{+}22]. occurs [WMS^{+}21]. octamer [MTCL^{+}23]. octamer-based [MTCL^{+}23]. OFF [BSC22]. offer [PF21]. old [MMC20]. oligodendrocyte [HCL^{+}21]. oligomer [DACG^{+}21]. oligomeric [SMM^{+}21]. oligomerization [WKC^{+}22]. oligomers [AII^{+}21]. Om14 [ZJDR22]. Oncogenic [SKF^{+}23, NS20, YSC^{+}02, YSC^{+}21]. one [Cas21, MS20, RCDMM20]. one-carbon [RCDMM20]. one-cell [MS20]. only [XHF^{+}20]. onset [ZVL^{+}23]. onto [FWP^{+}20]. oocyte [MdCT23, MP22d, WDJ^{+}21]. Oocytes [SGW^{+}20, AMFW^{+}21, CLR^{+}20, RDL^{+}20, SPL^{+}20]. oogenesis [AMFW^{+}21]. OPA1 [YZY^{+}20]. Open [Bak23]. oppose [BBM^{+}23]. opposed [HAL^{+}23]. Opposing [WDB^{+}21]. optic [AR20, SLES20]. optimize [NBC^{+}21]. optimized [NvGK20]. OPTN [YKK^{+}20]. Optogenetic [SdCS^{+}22, NvGK20, TB20a]. orchestrate [GMCO^{+}22]. orchestrates [FSZ^{+}22, MHGM22]. organ [LLK^{+}22]. organellar [BMF^{+}23]. organelle [CGCR^{+}22, MRA20, MSX^{+}21, SLH^{+}20a, WESR22, dCTOG^{+}20]. organelles [KMK21, NGG^{+}20]. organization [BJAR^{+}21, BDT^{+}22, CWN^{+}23, GVD^{+}20a, GVD^{+}20b, JRGH21, KSM^{+}21b, KWV^{+}23, KYR^{+}22, PMB^{+}22, PMSO^{+}23, Pro20, SMS^{+}20, SYW^{+}20, SGN^{+}20]. organize [CVT^{+}21, MTCL^{+}23, SV22]. organized [MKO^{+}21]. organizer [ZVC^{+}21]. organizes [CYR^{+}21, KLB^{+}22, LLW^{+}20, NKS^{+}21]. organoids [Bez22, CPS^{+}22]. Ori [MP22e]. orientation [KMW20]. Origin [SNP^{+}22, LJT^{+}22]. ORP10 [KSN^{+}22, WME22]. ORP5 [DZA^{+}20, GMCO^{+}22, RE20]. ORP8 [GMCO^{+}22]. ORP9 [WME22]. ortholog [PMSO^{+}23]. oscillation [INM^{+}21]. Osh6 [WYL21]. Osh6/7 [WYL21]. osmotic [OMI22]. Osteoclast [Bak23, ZTL^{+}23]. osteoclast-mediated [ZTL^{+}23]. other [MNvdS^{+}20]. our [MP22i]. outer [GCW^{+}23, LML^{+}21, MOK^{+}22, OCB^{+}21, RSB^{+}23]. outgrowth [WZZ^{+}23]. outs [WR22]. outward [LSX^{+}22]. over-elongated [KSS^{+}20b, KSS^{+}20c]. overcome [PL22]. overexpression [SRW^{+}21]. oxygen [CGBMC20, VTL^{+}20, VOR^{+}21]. oxysterol [FDG^{+}21].

P [LSG^{+}22, PCZ^{+}23, DZA^{+}20, MRWK^{+}22, RE20, XYG^{+}23]. P-bodies [XYG^{+}23]. p120 [WMA^{+}23]. P120catenin [EM20]. P2 [MWSX23]. p53 [VTL^{+}20]. p60 [SCL^{+}21]. p60-like [SCL^{+}21]. p97 [JTM^{+}23]. p97/VCP [JTM^{+}23]. pachytene [XYG^{+}23]. paclitaxel [Hök22]. pair [BMM^{+}20]. PAK1 [ESB^{+}21, MSC^{+}20]. pan [KSS^{+}20a]. pan-neurexin [KSS^{+}20a]. Pan1p [EMY^{+}22]. pancreatic [HPO^{+}23]. pandemic [CS21a, CS21b, CS21c, CS21d]. PAR-2 [CSG22]. Par3 [DRC^{+}20]. Par6 [DLZ^{+}20]. Par6-dependent [DLZ^{+}20]. paracrine [SIP^{+}23]. paradox [Tan23]. Paradoxical [EE22]. Parallel [DTG23]. Parameter [MLvdL^{+}21]. Parameter-free [MLvdL^{+}21]. PARK23 [HCWX^{+}22]. Parkin [OCB^{+}21, SFWB21]. Parkin-independent [OCB^{+}21, SFWB21]. Parkinson [PGDD21]. PARP1 [WMM^{+}23]. partially [MTW^{+}23, XHF^{+}20]. particle [SMM^{+}21, STvT23]. PARts [BCdS22]. pass [Col22a, Col22b]. passage [ACPR21]. paternal [MYM^{+}21, SdRVH^{+}21]. Pathogenic [BJR^{+}21, YZY^{+}20]. pathologies [PHAM^{+}20]. pathway [AT21, BCWM21, CCFN^{+}20, CCV^{+}21, EMY^{+}22, FUBS22, FDG^{+}21, FER^{+}23, GC22, LSD^{+}20a, LM23, MYK^{+}20, MYK^{+}21, MYK^{+}22, MOS^{+}22, OKH^{+}20, PSA^{+}23, PZ21, RDL^{+}20, RCDMM20, SFWB21, SCK^{+}20a, SCK^{+}20b, WMM^{+}23, WCG^{+}22]. pathways [BSB^{+}21, CJK^{+}22, EEW^{+}22, KPG20, MLQ^{+}21, MTD20, TWT20]. pattern [JMC^{+}20, MLS^{+}22]. pattern-forming [MLS^{+}22]. patterning [GMIC^{+}20, vLEM^{+}20]. Pavarotti [DNVP23, NVPP20]. Paxillin [XVW^{+}23]. PCM [CYH^{+}21]. PCNT [WMS^{+}20]. Pcp1 [ZJH22]. Pcp1/pericentrin [ZJH22]. PCR [FHM^{+}20]. PD [WXM22, XHF^{+}20]. PD-1 [XHF^{+}20]. PD-LI [WXM22]. PDIA3 [TSL^{+}20]. PEAK1 [ZAK^{+}22]. Peln1 [WJW^{+}22]. Peln1-mediated [WJW^{+}22]. penetration [SBV^{+}20]. Penman [Ped22]. peptide [BZC^{+}21, CMM^{+}20]. peptides [GLM^{+}22]. Per1 [KKPH^{+}21]. Per1/Per2 [KKPH^{+}21]. Per2 [KKPH^{+}21]. Perera [MP22f]. Pericentrin [HLB^{+}22, ZJH22]. pericentriolar [WMS^{+}20]. period [KGVK^{+}23]. peripheral [LCB^{+}23]. PERK [LM23, SvVV^{+}23, ZLW23]. PERK/E [LM23]. PERK/E-Syt1 [LM23]. permeability [CFV^{+}21, CPC^{+}20]. peroxisomal [YTH^{+}20]. Peroxisome [KWdB^{+}20, MRA20, BWK^{+}21, KHB^{+}22, PE22]. peroxisomes [GSLH^{+}21, HHD^{+}20, ZCL^{+}22]. Persistent [CBJ^{+}21, WXM22]. pervasive [SSHC21]. Pex14p [YTH^{+}20]. Pex3 [HHD^{+}20]. Pex30 [FC21]. Pex30-like [FC21]. pexophagy [PE22, ZCL^{+}22]. phagocytes [LFD^{+}21]. phagocytosis [EJBB^{+}20, HJL^{+}22, MRWK^{+}22, VFL20, ZGR^{+}22]. Phagosome [LFD^{+}21, WZK^{+}23]. phagy [WJL^{+}23]. pharmacological [TMG^{+}21]. Phase [CYU^{+}21, NKS^{+}21, BP22, BTF^{+}20, CAS23b, Dor20, GWR^{+}21, KSWC22, LLA^{+}21, NWZ20, OYS^{+}22, PTS^{+}22, WCG^{+}22, WC22, ZPŠS21, ZVM^{+}20, ZFZ^{+}23]. phase-separated [WCG^{+}22]. phases [KHFK^{+}20, RGK^{+}22]. phenotype [ESH^{+}23]. phenotypes [KSM^{+}21a]. phenotypic [LSS^{+}23]. Phollow [WH22]. phosphatase [AAF^{+}20, DWA^{+}22, FHM^{+}22, RSB^{+}23]. phosphatases [CSG22, CSS20, LLW^{+}21, MC21]. phosphate [HHGR21, RCA^{+}21]. phosphatidic [TB20a, TTM^{+}21]. phosphatidylethanolamine [TWY^{+}22]. phosphatidylinositol [HHGR21, PKH^{+}20, ZMS^{+}20, ZLJ^{+}22]. phosphatidylserine [LWD^{+}21, WYG^{+}20]. phospho [BJPH^{+}20, GDB^{+}20, KHV^{+}22, RGP^{+}22]. phospho-occupancy [KHV^{+}22]. phospho-switch [BJPH^{+}20, GDB^{+}20, RGP^{+}22]. phosphoinositol [WH22]. phospholipase [TB20a]. Phospholipid [ASC20, DTG23, EBZC^{+}21, OTOF21, WYL21]. phospholipids [PH20]. Phosphoregulation [GMC^{+}20, Tan23]. phosphorylated [CWX^{+}21]. phosphorylates [JCL^{+}23, LHL^{+}23, MSC^{+}20]. Phosphorylation [JBV^{+}20, KKP^{+}21, LGB^{+}21, MLQ^{+}21, PSP^{+}21, AAR^{+}21, AHQ20, CHZ^{+}20, DSY^{+}22, FFZ^{+}22, HBS^{+}20, HBDC^{+}20, INM^{+}21, KHB^{+}22, LNY^{+}22, LRB^{+}22, SKN^{+}21, SKS^{+}23, SWN^{+}22, XVW^{+}23, YTH^{+}20, ZRO^{+}23, ZAK^{+}22]. Phosphorylation-dependent [PSP^{+}21]. photoreceptor [HSSK20]. physical [DACG^{+}21]. physiological [JMY^{+}23, PHAM^{+}20, VTL^{+}20]. PI [CS21a, Dri20, DZA^{+}20, LSG^{+}22, PCZ^{+}23, RE20]. PI3 [WB20]. PI3K [CW23, EZB^{+}20, FCHM20, MHS^{+}20, OKH^{+}20]. PI3K-calcium-Nox [CW23]. PI3K-dependent [OKH^{+}20]. PI3K-WIPI2 [FCHM20]. PI4P [JDKK^{+}22, KSN^{+}22]. PI4P/PS [KSN^{+}22]. pictures [SSB20]. Pigino [MP21a]. pigmentosa [ZLW23]. PIM1 [JCL^{+}23]. Pin1 [KKP^{+}21]. Ping [XYG^{+}23]. Ping-pong [XYG^{+}23]. PINK1 [RPM^{+}21, SFWB21]. Pink1-dependent [SFWB21]. PIP [YLC^{+}21]. pipeline [BMF^{+}23]. PIPN [LPMA^{+}22]. piRNA [XYG^{+}23]. pits [CDLZ^{+}22, CS20, MLL^{+}20, Smy22]. pituitary [AFB^{+}20]. pivotal [JML^{+}21]. pivoting [FDA21]. PK [MRL^{+}21]. PKA [IvCD^{+}21]. PKA-II [IvCD^{+}21]. PKC [LGB^{+}21]. PKD2 [LLW^{+}20]. PKR [ZMW^{+}22]. Placing [O’D20a]. Plan [SS22]. Planar [NYN^{+}21, HW22, MHS^{+}20]. plane [MDB^{+}20]. Plant [ZBM^{+}22, KB22, MP22g]. plant-specific [KB22]. plaques [MLL^{+}20]. plasma [CSM^{+}21, FCT^{+}20, GCL^{+}21, HHD^{+}20, KWdB^{+}20, MWSX23, MMKM21, PCZ^{+}23, RBL22, SWS21b, UTR^{+}23, WCL^{+}23, ZMS^{+}20, ZSJE20]. plasmacytoid [SPKP22]. plasticity [BDT^{+}22, PKC^{+}22, YCC^{+}21]. Plastin [HGG^{+}23, RH23]. platforms [Smy22, TF20]. play [BP20]. players [RE20]. plays [SMHH^{+}20]. PLC [ZPŠS21]. Plectin [BG22, PAS^{+}22]. Plectin-mediated [PAS^{+}22]. PLK4 [CWX^{+}21, NPdC^{+}21]. PLK4-phosphorylated [CWX^{+}21]. ploidy [ZJH22]. pluralist [MW21]. Pluripotency [JRGH21, WJW^{+}22]. Pluripotent [PDW^{+}20, MMC20, VZQ^{+}21]. PM [WYL21]. PML [JTM^{+}23]. PML-RARA [JTM^{+}23]. podosome [CKM^{+}20]. podosomes [PZWW21]. points [KLB^{+}22]. Polar [yLHW^{+}20]. polarisome [DLK^{+}21]. polarity [BCdS22, BKR^{+}22, CSG22, DRC^{+}20, GMC^{+}20, HMT^{+}21, HW22, KNiY^{+}21, MHS^{+}20, MHGM22, MGM22, MdCT23, NYN^{+}21, PPB^{+}21, Tar21]. polarization [FGBD^{+}21, IHBP^{+}23]. polarize [BCS^{+}21]. polarized [LD21, LYL^{+}23]. polarizes [BCC^{+}21]. pole [RVNS21]. poles [CYH^{+}21, MKO^{+}21]. policy [GPES21]. pollen [GCL^{+}21]. Polo [OZW^{+}21, BP20, BHK20, CSS20]. Polo-like [OZW^{+}21, BHK20, CSS20]. Poly [KSP^{+}21, CYU^{+}21]. polybasic [DLZ^{+}20]. polymerase [CBJ^{+}21, FLW^{+}23, UIS^{+}22]. polymerization [CFK^{+}22, CPW^{+}23]. polymers [LLW^{+}20]. polyphosphoinositide [Dri20]. polyploid [GNL^{+}20]. polyposis [EYC^{+}20]. pombe [VVW^{+}23]. pong [XYG^{+}23]. pool [FFZ^{+}22, PCGB20, SHD^{+}21, VGO^{+}23]. pooled [KSM^{+}21a, YSR^{+}21]. pore [CPC^{+}20, GBBT^{+}22, GG20, JMB^{+}20, JKZ^{+}22, KKZ^{+}22, KWV^{+}23, KST^{+}21, LW20b, RLV^{+}20, SMM^{+}21, TKK^{+}20, YM21]. pore-forming [JKZ^{+}22]. pores [CSOG^{+}20]. portals [CBC^{+}20]. position [CS21a, PHMD20]. position-dependent [PHMD20]. positioning [MSB^{+}21, MDB^{+}20, NBC^{+}21]. positive [FCHM20]. positives [MP23a]. post [XGD^{+}23]. post-Golgi [XGD^{+}23]. postsynaptic [AVC^{+}22, FLJ^{+}22, KKN^{+}21]. potential [PDW^{+}20]. potentiation [GLGL^{+}21]. power [Dus21]. powers [DCRDC^{+}22]. PP1 [CSG22]. PP2A [BZD^{+}21, LKW^{+}21]. PP2A-B56 [BZD^{+}21]. PP6 [SKS^{+}23]. PPM1F [GDB^{+}20]. PQLC2 [ATTF20]. PR [CYU^{+}21]. Prdm16 [HPO^{+}23]. pre [PSC^{+}20]. pre-mRNA [PSC^{+}20]. preassemble [MTW^{+}23]. precision [RMM^{+}21]. Precursor [ESH^{+}23, GPL^{+}21, SHD^{+}21]. precursors [LWZ^{+}23]. predictor [KRC^{+}22]. predicts [BJAR^{+}21]. premature [OHHR23]. presence [SdRVH^{+}21]. presentation [GLM^{+}22, LWG^{+}22, TJAG^{+}21]. preserves [HAW^{+}22, XDY^{+}22]. pressure [BBM^{+}23, MP23a]. presynaptic [GPL^{+}21, PMSO^{+}23, WDB^{+}21]. prevent [FBR^{+}21]. Preventing [Ver21]. prevents [HCWX^{+}22, LJJ^{+}21, MKO^{+}21, OYJJ23, PHAM^{+}20, ZLW23]. Prickle1 [HW22]. primary [DSG21, GSC^{+}20, MND^{+}20, MSX^{+}21, SIP^{+}23, YMAS20]. ′ − COP [XGD^{+}23].primes[CW23].priming[LAH^{+}21].primordial[MHN20].principles[IHBP^{+}23, WTU^{+}21, BLU21].Prion[SWT^{+}22].Prion − like[SWT^{+}22].prior[RSB^{+}23, VHPP^{+}20, ZBM^{+}22].pro[HMT^{+}21, WDL^{+}20].pro − N − cadherin[HMT^{+}21].pro − neural[WDL^{+}20].probe[CBS^{+}21, UIS^{+}22, UTR^{+}23].Probing[ZMS^{+}20].procentriole[MSR^{+}20].process[CLL^{+}21b, hYKO^{+}20a, hYKO^{+}20b, hYKO^{+}21].processes[PKC^{+}22, VTL^{+}20].processing[HMT^{+}21, SBL^{+}21].processive[RCA^{+}23].procollagen[SBL^{+}21].produces[LPMA^{+}22].production[RPM^{+}21, SPKP22].Productive[MWF^{+}23].profibrotic[PHAM^{+}20].profiles[RRCS^{+}23].profilin2[HYL^{+}20].profiling[ABM^{+}23, BDH^{+}21, MKD^{+}21, SMFC^{+}22].progenitor[BHS^{+}21, HYL^{+}20, Tar21].progenitors[BCS^{+}21].prognostications[MRA20].program[BGM^{+}21, HYX^{+}20].Programmed[OCLB21].Programming[MMC20].programs[DCS^{+}20].progranulin[DF22].progression[BDR20, BW20, HPO^{+}23, LKW^{+}21, SDD^{+}22, RDW^{+}20].progressively[RMA21].proinflammatory[ESH^{+}23].projections[CG21].proliferation[EE22, FIK^{+}05, FIK^{+}20, HSL^{+}20, LXJ^{+}23, MFC^{+}20].proline[ZVM^{+}20].proline − rich[ZVM^{+}20].Prolonged[KSS^{+}20c, KSS^{+}20b].prolongs[HTL^{+}21].promote[AZR^{+}22, CYH^{+}21, CBJ^{+}21, CM21, DG22, ESX^{+}20, FRO^{+}20, FDSR22, HDW^{+}21, HMT^{+}21, JCL^{+}23, LHL^{+}23, LPMA^{+}22, LLX^{+}21, MDB^{+}20, MMKM21, PSC^{+}20, PLG^{+}23, PWW^{+}20, QLC^{+}20, RPM^{+}21, SMD^{+}21, SvDSW^{+}20, STS21, TG21, VCS^{+}22, WCC^{+}23].promoters[BDH^{+}21].promotes[AMMK^{+}22, BWK^{+}21, BZD^{+}21, BED^{+}21, BDS^{+}21, CCH^{+}21, CRZ^{+}21, CPW^{+}23, CLC^{+}21, CHZ^{+}20, CJS^{+}21, ESB^{+}21, FAHZ21, FLJ^{+}22, FPMS^{+}21, GOR^{+}20, GBBT^{+}22, HBDC^{+}20, INM^{+}21, KSWC22, LKW^{+}21, LCB^{+}23, LSG^{+}22, MYK^{+}20, MYK^{+}21, MRL^{+}21, MYK^{+}22, MSH^{+}20, MWSX23, MYC^{+}23, PCGB20, RCA^{+}21, SPS^{+}20, SJL^{+}22, SRW^{+}21, TNC^{+}20, TOL^{+}20, WHN^{+}21, WXM22, WZZ^{+}23, XZJ^{+}21, XVW^{+}23, YSC^{+}02, YSC^{+}21, ZPŠS21, ZVM^{+}20, ZBY^{+}21].promoting[AH20a, CKR^{+}20, HTL^{+}21, LSD^{+}20a, LSX^{+}22, OMI22, PFPB^{+}20, SYQ^{+}22].prompts[WKX^{+}21, ZVL^{+}23].pronuclear[MS20].pronuclei[RCH^{+}20].proper[Tev20].properties[CPC^{+}20, MTR^{+}20, UTR^{+}23, WCG^{+}22].prophase[BHK20, CLR^{+}20].prosaposin[DF22].prostate[Cas23a, SDD^{+}22].protease[TSL^{+}20].proteasomal[LLC^{+}20].proteasome[WB21].proteasomes[Let20].protect[ML22, RCM^{+}23a].protects[CWAT20, OHHR23, ZLJ^{+}23].Protein[BRB^{+}20, BNV^{+}23, CDD^{+}22, CGK^{+}22, CWAT20, CWKP23, CLC^{+}21, CJS^{+}21, DRC^{+}20, EEW^{+}22, FUBS22, FHM^{+}22, FAS^{+}21, GCW^{+}23, GC22, HCWX^{+}22, HSL^{+}20, HSW^{+}22, HBS^{+}20, HW22, ITM^{+}21, JJ23, KNA^{+}22, LLBC^{+}20, LMM^{+}23, LRM^{+}20, LLK^{+}22, MHS^{+}20, MAW^{+}22, MRD21, NVPP20, NGG^{+}20, NMO^{+}22, PCZ^{+}23, PPG21, PGD^{+}20, PHT^{+}23, RSB^{+}23, SMK20, SHA20, SPT^{+}09, SPT^{+}21, SMHH^{+}20, Sir23, SLM23, TTM^{+}21, TJAG^{+}21, WYG^{+}20, WLW^{+}22, VGO^{+}23, WAK^{+}20, ZPG^{+}23, ZJDR22, ESH^{+}23].proteins[ALPH20, AHLR22, BWEHS21, CGK^{+}22, CNL^{+}21, CYR^{+}21, CM21, DNVP23, DJI^{+}21, DSMB20, FC21, HH22, IHBP^{+}23, ICMM20, KSS^{+}20a, LCM22, LWZ^{+}23, MOS^{+}22, SSR^{+}22, TEH^{+}20, TWY^{+}22, TG21, UTR^{+}23, WAOS^{+}21, YMH^{+}20, vdBdHLK22].proteolysis[HSF^{+}23, NTA^{+}21, SGW^{+}20].Proteolytic[ZTL^{+}23, WB21].proteome[HDW^{+}21, RGK^{+}22].proteomics[MKD^{+}21].proteostasis[MBW22, SNL^{+}22].Protrudin[AO20, PWW^{+}20].Protrudin − mediated[PWW^{+}20].protrudinG[AO20].protrusion[EYC^{+}20].protrusions[CDD^{+}22, SWS21b].protrusive[MBA^{+}22].provide[BBM^{+}23, RVNS21, TWY^{+}22].provides[GMC^{+}20, KWV^{+}23].proximity[CLH^{+}20, TNLPF20, TPM^{+}21].pruning[KWGR23].PS[KSN^{+}22].PtdIns[MWSX23, MRWK^{+}22].PTEN[QLC^{+}20].public[GPES21].pulling[DPM^{+}20, TP20].pulls[BG22].Pumping[Gal23].purified[CBS^{+}21].purinergic[KIV^{+}20].pushing[TP20].pVHL[ZDGB^{+}22].pVHL − mediated[ZDGB^{+}22].pyroptosis[SdCS^{+}22].

Q [TRJ^{+}20]. QPCT/L [SLS^{+}23]. Quality [FBVD^{+}22, HGK20, ML22, PK23, MMSP20, SBBJ21]. quantification [LYL^{+}22, MLvdL^{+}21, MAW^{+}22]. Quantifying [BMF^{+}23]. Quantitative [BBPS23, vdBdHLK22, LZT^{+}23, UZS^{+}23]. quiescence [AMMK^{+}22].

R [GPES21]. RAB [WLW^{+}22, BLU21, KCP^{+}21]. RAB-8 [WLW^{+}22]. Rab11 [CH22, ESB^{+}21, WDB^{+}21]. Rab18 [GMB^{+}20]. Rab2 [GPL^{+}21]. Rab35 [CG21]. Rab40 [LDH^{+}21]. Rab40b [DHTP22]. Rab40b/Cul5 [DHTP22]. Rab5 [HMSF22]. Rab5-GAP [HMSF22]. Rab7 [XZJ^{+}21]. Rab8 [HVPM20]. Rab8/10/11 [HVPM20]. Rac1 [BED^{+}21, ESB^{+}21, Kin21]. Rac1-PAK1 [ESB^{+}21]. RacC-WASP [LYL^{+}23]. RAD54 [MSH^{+}20]. Radial [WLM^{+}20, CVT^{+}21, KNiY^{+}21]. radiation [FCT^{+}20]. radiation-induced [FCT^{+}20]. raft [KHFK^{+}20]. Rag [HHGR21]. Range [MND^{+}20, MBG^{+}23]. RanGTP [EMEZ^{+}20, MYM^{+}21]. Rap2 [Cas22, DHTP22]. Rapid [ABM^{+}23, ZS21, CFK^{+}22]. rapidly [KAS^{+}22]. RARA [JTM^{+}23]. RAS [SWS21b]. Rashomon [HLGD20]. rate [BJAR^{+}21, FGBD^{+}21, LMM^{+}23]. ratio [DSMB20]. RB [KYR^{+}22]. Rcd4 [PKD^{+}20]. Rcr1 [ZSJE20]. reactions [SNP^{+}22]. Reactive [CGBMC20, Bez22, VOR^{+}21]. reactivity [CPS^{+}22]. ready [Ver21]. Real [FLW^{+}23]. Real-time [FLW^{+}23]. reality [WBR^{+}20]. rear [WXM22]. rearrangements [BZD20]. rebalances [MBA^{+}22]. rebuilding [CLL^{+}21b]. REC [KMJ^{+}23]. receptor [AFB^{+}20, BMS^{+}22, CFD^{+}20, CLC^{+}21, CHZ^{+}20, DWA^{+}22, GLGL^{+}21, GSP^{+}20, ICMM20, LSD^{+}21, RKLJ22, TRHS23, TJAG^{+}21, VGK^{+}21, WZ22]. receptors [BEM^{+}23, PPG21]. Recognition [SSHC21, AHY^{+}21, HH21, JFM^{+}22]. recombination [KMJ^{+}23, MSH^{+}20]. Reconstitution [CMN^{+}22, ZHHJ22, FMY^{+}21]. reconstitutions [WTU^{+}21]. reconstruction [LLLR20, MSX^{+}21]. recruit [CYH^{+}21, SKX^{+}23]. recruited [BDD20, CDLZ^{+}22]. recruitment [HHGR21, KST^{+}23, KSP^{+}21, MSR^{+}20, MYC^{+}23, PGD^{+}20, WDJ^{+}21, ZPG^{+}23]. recruits [ATTF20, BP22, SvVV^{+}23, WYL21]. recycles [SPS^{+}20]. Recycling [LCM22, CJK^{+}22, CKW^{+}22, LLY22, LGB^{+}21, MLQ^{+}21, MH22, PFS^{+}22, Sea21, XGD^{+}23, YLH^{+}21]. reduce [LMM^{+}23]. Reduced [LJJ^{+}21, BRB^{+}20]. REEP4 [GBBT^{+}22]. reference [SHA20]. refractory [KGVK^{+}23]. regenerates [LFD^{+}21]. regenerating [MBG^{+}23]. regeneration [GSB^{+}20, NTA^{+}21]. region [KBN^{+}21, SGL^{+}23]. regions [BDD20, FPZ^{+}22, FER^{+}23]. regrowth [YMAS20]. regulate [AGW^{+}20, BMM^{+}20, CDLZ^{+}22, CFD^{+}20, CCV^{+}21, CGCR^{+}22, CT20, DNVP23, DSLP20, GKM^{+}20, HZN^{+}21, KBH^{+}22, KGVK^{+}23, KKP^{+}21, Let20, LWD^{+}21, LZZ^{+}21, MHS^{+}20, MYC^{+}23, MMDK^{+}22, NVPP20, PLG^{+}23, WCG^{+}22, WLW^{+}22, XHF^{+}20, YLC^{+}21, YJX^{+}20, ZCL^{+}22]. Regulated [ARM23, AFB^{+}20, FUBS22, MTR^{+}20, PSS^{+}20, ABB^{+}22, BSB^{+}21, CVMB^{+}23, HCL^{+}21, JWB^{+}22, LCM22, RFL20]. regulates [AAF^{+}20, ATS^{+}21, APL^{+}21, DCK^{+}20, DDD^{+}20, DSY^{+}22, DHB^{+}21, DWA^{+}22, DZA^{+}20, DY21, DHTP22, FAS^{+}21, GCNL21, GSB^{+}20, GPL^{+}21, GCS^{+}20, GMB^{+}20, HSSK20, HBS^{+}20, HDW^{+}21, HYL^{+}20, JBV^{+}20, JRGH21, KSN^{+}22, KPM^{+}22, KKN^{+}21, KSM^{+}21b, KLCM^{+}23, KST^{+}22, LL22, LTL^{+}20, LLK^{+}21, LFF^{+}22, LYL^{+}23, LDH^{+}21, LD20, LSD^{+}21, ORCT^{+}20, OHY^{+}20, PFPB^{+}20, PKC^{+}22, PLL^{+}20, POL^{+}20, RE20, SMS^{+}20, SRK22, SHBF^{+}20, SKN^{+}21, SSO^{+}20, SYQ^{+}22, SPL^{+}20, SMC^{+}20, Tar21, WMM^{+}23, WJL^{+}23, YTH^{+}20, YZW^{+}20, ZAK^{+}22]. Regulating [KHB^{+}22, DOA^{+}22, EE22, SFC^{+}23, XZJ^{+}21]. Regulation [GC22, JKL^{+}22, NSB^{+}21, PPG21, YW21, ZWJ22, AANLL^{+}20, BVYW20, BSH^{+}22, CML20, ESB^{+}21, Gan21, GWR^{+}21, HW22, KKPH^{+}21, KBN^{+}21, KBB^{+}23, LYP^{+}21, MBW22, MRH^{+}23, PHMD20, PRB^{+}20, PPB^{+}21, SKS^{+}23, WYG^{+}20, WM23, WHA20, ZSJE20, ZTL^{+}23, ZCD^{+}21]. regulator [AMG^{+}20, EMY^{+}22, HJL^{+}22, JML^{+}21, WESR22]. regulators [BZC^{+}21, LHS^{+}22]. regulatory [DJI^{+}21, WJL^{+}23]. reinforce [KLB^{+}22]. reinforces [CH22]. related [AAF^{+}20]. relationships [KST^{+}21, VLdRADJ22]. relative [DKCT21]. relaxation [yLHW^{+}20]. release [BSB^{+}21, CRZ^{+}21, CSOG^{+}20, GKFR20, HSF^{+}23, JMB^{+}20, LLK^{+}21, LYS^{+}20, PGD^{+}20, POL^{+}20, VRSN23, ZCD^{+}21]. releases [SPL^{+}20]. Relief [ALC^{+}20]. relies [WMS^{+}20]. remarkable [WM20]. remodel [JDKK^{+}22]. remodeling [CLZ^{+}20, GLGL^{+}21, KWV^{+}23, LD20, MHN20, MYC^{+}23, OYJJ23, PLG^{+}23, eSG23, VCS^{+}22, WKX^{+}21]. removal [yLHW^{+}20, SLS^{+}23, SYQ^{+}22, SNN20]. reorganization [DSB22, VV23]. reovirus [dCTOG^{+}20]. repair [CSM^{+}21, HSF^{+}23, HRS^{+}20, KMJ^{+}23, LLA^{+}21, MRL^{+}21, MSH^{+}20, MWSX23]. replication [ABM^{+}23, CWZ^{+}20, CLZ^{+}20, ITM^{+}21, MV20, MYC^{+}23, MMC20, PDW^{+}20, RDW^{+}20, WCC^{+}23]. Reply [LYS^{+}20]. representation [MP21c]. repress [XYG^{+}23]. represses [BCWM21]. repression [VGO^{+}23]. repressor [ZRO^{+}23]. reproducibility [LVMFL20]. reprogrammed [PDW^{+}20]. repurposed [RLS^{+}20]. request [Köh21]. require [GKFR20, LYS^{+}20, MGM22]. required [BCM^{+}22, CSM^{+}21, CG21, DCG^{+}23, Gar21, HCL^{+}21, HRS^{+}20, KMW20, LZC^{+}20, PSC^{+}20, SHD^{+}21, SGL^{+}23, SBL^{+}21, SBBJ21, TRJ^{+}20]. requirement [CML20, PKD^{+}20, WM23]. requires [BZC^{+}21, DF22, FUBS22, SLES20, SPRWB20, WCL^{+}23]. rescue [HBDC^{+}20]. rescues [HGG^{+}23, RH23]. research [CS21b, GPES21, O’D20a]. researchers [MP22b]. resetting [VGO^{+}23]. reshapes [MHN20]. resident [LJT^{+}22]. residents [Low21]. resilience [CW23, JMKS^{+}23]. resistance [TMG^{+}21]. resolution [LFD^{+}21, WZK^{+}23]. resolved [DRW^{+}23, MKD^{+}21]. Resorb [Bak23]. resorption [ZTL^{+}23]. respiration [SvVV^{+}23]. response [ATTF20, BCS^{+}21, EBZC^{+}21, FAS^{+}21, GPW^{+}22, GCS^{+}20, HYX^{+}20, JJ23, MKD^{+}21, MFC^{+}20, PKA20, SPT^{+}09, SGW^{+}20, YCC^{+}21, ZMW^{+}22, SPT^{+}21]. responses [GSB^{+}20, ITM^{+}21, KPM^{+}22, LLC^{+}20]. responsive [MBW22, SSO^{+}20, SvDSW^{+}20]. restrain [FER^{+}23]. restrict [CWZ^{+}20, GSC^{+}20]. restricted [GCW^{+}23]. restricts [SV22]. results [KSS^{+}20b, KSS^{+}20c]. resurfacing [AFB^{+}20]. retention [AH20a, AVC^{+}22, KWdB^{+}20, SLD^{+}21]. Reticular [HAL^{+}23]. Reticulon [CWAT20, GBBT^{+}22, PMSO^{+}23]. Reticulon-like [GBBT^{+}22, PMSO^{+}23]. Reticulons [WCC^{+}23]. reticulum [AAR^{+}21, BBP^{+}20, CSM^{+}21, GCS^{+}20, GMB^{+}20, SPT^{+}09, SPT^{+}21, SLM23, WMS^{+}21, ZHW^{+}21, ZDM^{+}22]. retina [LXJ^{+}23]. retinal [DYW^{+}20, WDL^{+}20]. retinitis [ZLW23]. Retinyl [MYT^{+}21]. retraction [FSZ^{+}22, WXM22]. retrieval [LLY22, RCM^{+}23b]. retrograde [BS20a, DCRDC^{+}22, KRS21, MYK^{+}20, MYK^{+}21, MYK^{+}22]. retromer [LLY22, SPS^{+}20, WDB^{+}21]. retromer-dependent [LLY22]. reveal [KST^{+}21, LHS^{+}22, SHLS22, WTU^{+}21]. revealed [CLH^{+}20, PBPBS22]. reveals [AMG^{+}20, BSB^{+}21, BBPS23, CVMB^{+}23, CLH21, CMN^{+}22, EEW^{+}22, KRC^{+}22, KHFK^{+}20, LLLR20, MVM20, NGG^{+}20, NBI^{+}22, PMB^{+}20, SMM^{+}21, SGN^{+}20, WDRRF^{+}23, ZMS^{+}20, ZS21, vdBdHLK22]. rewires [MTD20]. Reynolds [WMA^{+}23]. RFWD3 [DMR^{+}20, MYC^{+}23]. RGA [MLS^{+}22]. RGA-3 [MLS^{+}22]. RGA-3/4 [MLS^{+}22]. Rga6 [WZZ^{+}23]. RGD [BJSOS^{+}21, BJSOS^{+}20]. RGE [BJSOS^{+}21, BJSOS^{+}20]. RGS [HH22]. RHGF [KST^{+}23]. RHGF-1 [KST^{+}23]. Rho [MLS^{+}22, RLK^{+}20]. Rho/Cdc42 [RLK^{+}20]. RhoA [SKX^{+}23, VCS^{+}22, ZLS^{+}21]. rhodopsin [ZLW23]. RhoGAP19D [FBR^{+}21]. Rhotekin [YLH^{+}21]. Ribbon [LLK^{+}22]. ribose [KSP^{+}21]. ribosomal [CHPF^{+}21a, CHPF^{+}21b, LLK^{+}22]. Ribosome [LW20a, HGK20, PK23]. ribosylation [CGK^{+}22]. rich [TTM^{+}21, ZVM^{+}20]. rigid [ITB^{+}23]. rigidity [MSB^{+}21, MMDK^{+}22]. RIM [PGD^{+}20]. RIM-binding [PGD^{+}20]. ring [BJAR^{+}21, MHN20, Mer21, SCN^{+}23]. rings [WLM^{+}20]. RIPK1 [HTL^{+}21]. RNA [CBS^{+}21, FLW^{+}23, FPMS^{+}21, MTD20, RFL20, SPRWB20, SSHC21, Tev20, UIS^{+}22, WTS^{+}21, WAK^{+}20, WLM^{+}21, ZPG^{+}23]. RNase [CBS^{+}21]. RNF17 [XYG^{+}23]. Robust [PGH^{+}23, FBVD^{+}22, JMB^{+}20, KST^{+}23, LM21, LSD20b, PVYJ^{+}21]. rod [CJS^{+}21]. Role [SCK^{+}20b, ANRS^{+}20, BWA^{+}23, BBPS23, JGN^{+}20, LHS^{+}22, SMHH^{+}20, VDC^{+}20, WI22, YKK^{+}20, SCK^{+}20a]. roles [EE22, LRM^{+}20]. rounding [LDE^{+}22, MSB^{+}21, MDV^{+}21]. routes [ZXY^{+}23]. routine [FBVD^{+}22]. RTKN [YLH^{+}21]. RTKN-1 [YLH^{+}21]. RTKN-1/Rhotekin [YLH^{+}21]. RUFY1 [RCM^{+}23b]. rule [RDL^{+}20]. run [GPW^{+}22, SS22]. rupture [DRC^{+}20, ITB^{+}23, KAS^{+}22]. Rushika [MP22f].

s [HLGD20, BTF^{+}20]. S. [FDA21]. S1PR1 [AAR^{+}21]. S9.6 [SSHC21]. sabers [Tai22]. SAC1 [CFD^{+}20]. Sachihiro [MP22g]. safeguards [DdCVT22]. Sara [MP22h, MP22i]. SARAF [ZCD^{+}21]. sarcomeric [SGN^{+}20]. Sarm1 [LPMA^{+}22, KMD20]. SARMful [Pie20]. SARS [SCK^{+}20a, MNvdS^{+}20, SCK^{+}20b, WCC^{+}23].

SARS-CoV- [SCK^{+}20a, MNvdS^{+}20, SCK^{+}20b]. SARS-CoV-2 [WCC^{+}23]. Sas4 [RSWP20]. saturated [GSP^{+}20]. scaffold [KNA^{+}22, KRS21]. scaffolds [TRHS23]. scale [YKSC^{+}22]. scaling [KRC^{+}22]. scanning [HGK20]. SCAP [WHN^{+}21]. Scar [YLC^{+}21]. Scar/WAVE [YLC^{+}21]. Scc1 [SPL^{+}20]. Scc1-cohesin [SPL^{+}20]. SCF [BZD^{+}21, HZN^{+}21]. SCF-Fbxo42 [BZD^{+}21]. Schafer [MP22b]. Schizosaccharomyces [VVW^{+}23]. science [MP21c]. scientist [VM21]. scramblases [LWD^{+}21]. screening [KSM^{+}21a]. screens [LSS^{+}23, LHS^{+}22, YSR^{+}21]. Sculpting [MP22b]. sculpts [SCL^{+}21]. SDF1 [BCS^{+}21]. sealing [LD20]. seals [SFO^{+}21]. search [GPW^{+}22, SS22]. search-and-run [SS22]. Sec14 [WYG^{+}20]. Sec14-like [WYG^{+}20]. second [O’D22]. secretase [WMS^{+}21]. secrete [HCRMTC23]. secretion [AFB^{+}20, Bog21, CGBMC20, HBS^{+}20, LSG^{+}22, VBG^{+}22, WZG22, WLW^{+}22]. Secretory [SNL^{+}22, CCFN^{+}20, JKL^{+}22, LFF^{+}22, MYK^{+}20, MYK^{+}21, MYK^{+}22, PTS^{+}22, PSA^{+}23, SCK^{+}20a, SCK^{+}20b, WR22, ZXY^{+}23]. secures [CHS^{+}22]. seed [OWY^{+}23]. segment [AH20a, TOL^{+}20]. segmentation [GMD^{+}23]. segregase [JTM^{+}23]. segregation [BDT^{+}22, CML20, CWZ^{+}20, CBJ^{+}21, CSOG^{+}20, DSB22, LZC^{+}20, MS23, SPL^{+}20, TP20]. Seipin [RCF^{+}22, MYT^{+}21, SOT^{+}21, CEM^{+}20, DY21]. seipin-independent [SOT^{+}21]. seizure [WMM^{+}23]. selected [LRB^{+}22]. selection [ACPR21]. Selective [BZC^{+}21, AHY^{+}21, CWKP23, DSB22, HH21, IHBP^{+}23, MOK^{+}22, NSB^{+}21, OCLB21, RKLJ22, WAK^{+}20]. selectively [CMT^{+}21, KKP^{+}21, Yam21]. selectivity [JRGH21, PHT^{+}23]. self [JRGH21, MP23b]. self-immune [MP23b]. self-organization [JRGH21]. SEM [LSS^{+}23, MSX^{+}21]. Semi [LQS23]. Semi-automated [LQS23]. senescence [BZD20]. senescence-associated [BZD20]. Senescent [RG23, SLS^{+}23, BG21]. Sensing [CFD^{+}20, HI21]. sensitive [CBS^{+}21]. sensitivity [PRMF^{+}23]. sensitization [IvCD^{+}21]. sensor [GLM^{+}22, HYX^{+}20, KLC^{+}20, LLK^{+}21, SLH^{+}20a, WHN^{+}21]. sensors [dCS^{+}21]. sensory [KWGR23]. separated [WCG^{+}22]. separation [BP22, CYU^{+}21, CAS23b, Dor20, GWR^{+}21, KSWC22, LNY^{+}22, LLA^{+}21, NWZ20, NKS^{+}21, OYS^{+}22, PTS^{+}22, ZPŠS21, ZVM^{+}20, ZFZ^{+}23]. SEPT9 [FRO^{+}20]. Septin [SKX^{+}23, CVMB^{+}23, GM23, KRS21, POL^{+}20]. Septin-mediated [SKX^{+}23]. Septin2 [CKM^{+}20]. septins [CYR^{+}21, MTCL^{+}23]. seq [BDH^{+}21]. sequences [SLM23]. Sequential [CCV^{+}21, CJK^{+}22]. sequentially [ZLJ^{+}22]. sequestration [RLV^{+}20, SSR^{+}22]. serine [RCDMM20]. service [MRD21]. severe [KNiY^{+}21]. severing [JBV^{+}20, OCB^{+}21]. Sfi1 [RVNS21]. SFPQ [FPMS^{+}21]. SFPQ-RNA [FPMS^{+}21]. Sgo1 [AGH^{+}22]. Sgo1-mediated [AGH^{+}22]. shape [BC23, DYW^{+}20, FBR^{+}21, KRH^{+}20, MA20, MP22e, ZLS^{+}21]. shaped [SYW^{+}20]. shapes [DLK^{+}21, GPW^{+}22, JMKS^{+}23, Sir23]. Shaping [PKA20, CS20, FSZ^{+}22, KTT^{+}22]. Shared [PBPBS22]. sheets [PAS^{+}22]. Sheldon [Ped22]. SHH [MND^{+}20]. shields [YLH^{+}21]. SHIP164 [HSW^{+}22]. shock [FAS^{+}21, SSR^{+}22]. short [Ike20]. show [BDT^{+}22]. SHP1 [XHF^{+}20]. SHP2 [XHF^{+}20]. shuttle [AMFW^{+}21]. side [TML22, vdBVS^{+}23]. side-averaging [TML22]. Sidekick [MBA^{+}22]. SidK [MAW^{+}22]. Signal [FGBD^{+}21, SLM23, DSMB20, KHFK^{+}20, LWG^{+}22, RCA^{+}21]. Signaling [ZMW^{+}22, BS20b, BDS^{+}21, CDLZ^{+}22, CHZ^{+}20, CBJ^{+}21, DACG^{+}21, DSG^{+}23, DYW^{+}20, DSLP20, DWA^{+}22, FHM^{+}22, GL20, GCS^{+}20, HCWX^{+}22, HRS^{+}20, HRB^{+}21, HTL^{+}21, JMC^{+}20, JKL^{+}22, KMD20, LJJ^{+}21, LDE^{+}22, LYL^{+}23, LLW^{+}21, MHS^{+}20, MLQ^{+}21, MND^{+}20, MBG^{+}23, MP21b, NS20, PRB^{+}20, PPB^{+}21, RH23, RWSZ^{+}20, RSWP20, SIP^{+}23, SHLS22, SKF^{+}23, Smy22, SWS21b, TF20, TNC^{+}20, TB20a, TRHS23, WTS^{+}21, WMM^{+}23, XHF^{+}20, YSC^{+}02, YSC^{+}21, ZMMM^{+}20, ZPŠS21, ZDGB^{+}22, ZGR^{+}22]. signalosome [KSWC22]. signals [AMMK^{+}22, GKM^{+}20, MRH^{+}23]. silence [ME21]. Silencing [BP20, BHS^{+}21, PRMF^{+}23]. Similarities [BG21, DSG21]. simple [FBVD^{+}22]. simply [BD20]. simultaneous [WRG23]. Single [SMM^{+}21, BDH^{+}21, FWP^{+}20, KHFK^{+}20, MLvdL^{+}21, NGG^{+}20, STvT23, XVW^{+}23, YMH^{+}20]. single-cell [BDH^{+}21, XVW^{+}23, YMH^{+}20]. single-membrane [FWP^{+}20]. single-molecule [KHFK^{+}20, MLvdL^{+}21]. Single-particle [SMM^{+}21, STvT23]. SIR [PRMF^{+}23]. SIRT7 [WJL^{+}23]. Sis1 [FAS^{+}21, KB21]. sister [RDL^{+}20, SWS^{+}21a]. Site [LSS^{+}23, EBZC^{+}21, KBN^{+}21, PHMD20, RVNS21, SMFC^{+}22, SNYA^{+}21]. site-specific [SMFC^{+}22]. sites [AGW^{+}20, AO20, BCM^{+}22, CCH^{+}21, CBC^{+}20, DCG^{+}23, FC21, GMCO^{+}22, KSN^{+}22, KAS^{+}22, KWdB^{+}20, LYL^{+}22, PWW^{+}20, PGD^{+}20, RLS^{+}20, TNLPF20, UIS^{+}22, VBG^{+}22, WHN^{+}21, dDFGP^{+}21]. situ [NBI^{+}22, PMB^{+}20]. size [DACG^{+}21, GCNL21, OZW^{+}21, OMK^{+}22, SKS^{+}23, hYKO^{+}20a, hYKO^{+}20b, hYKO^{+}21]. Ska1 [RCA^{+}23]. skeletal [LLX^{+}21]. skin [NTA^{+}21]. Slik [DDD^{+}20]. SLX4 [ITM^{+}21]. SMAD3 [ZDGB^{+}22]. Smad4 [HPO^{+}23]. Small [CLL^{+}21a, ITB^{+}23, KHV^{+}22, WTS^{+}21, SFC^{+}23, SSR^{+}22, VBG^{+}22]. Small-molecule [CLL^{+}21a]. SMC3 [RDL^{+}20]. SMGL [WLW^{+}22]. SMGL-1 [WLW^{+}22]. SMGL-1/NBAS [WLW^{+}22]. SMLM [VVW^{+}23]. smoothened [DSLP20, LSD^{+}21]. SNAP [Tar21]. SNAP23 [KNiY^{+}21]. SNARE [BLZ^{+}21, CWKP23, Tar21]. snubs [MRD21]. SNX [HH22, LC20]. SNX-RGS [HH22]. SNX13 [LHS^{+}22]. SNX27 [MLQ^{+}21, SPS^{+}20]. SNX9 [JGN^{+}20, LC20]. soaps [MP22h]. Sod1 [VGO^{+}23]. software [LSS^{+}23]. solely [BJSOS^{+}20, BJSOS^{+}21]. solute [HZCX22]. somatostatin [AFB^{+}20, GKM^{+}20]. Song [O’D20b]. SORLA [SHD^{+}21]. sorting [AANLL^{+}20, GNML^{+}20, HDG22, LMRG20, LRM^{+}20, PKA20, RCM^{+}23b, WPCB^{+}21]. sorts [BLZ^{+}21]. source [Hic22]. span [ZJDR22]. Spatial [BHK20, PHMD20, EM22, KST^{+}21, MSC^{+}20, MRH^{+}23, RRBW^{+}21]. spatially [CLH21, GCW^{+}23]. Spatiotemporal [LKW^{+}21, TB20a, ZGR^{+}22, GCNL21, KLCM^{+}23, WHA20]. spatiotemporally [FAMQW22]. SPB [ZJH22]. Specialist [MW21]. specialized [LWG^{+}22]. species [CGBMC20, VOR^{+}21]. Specific [HH21, CBS^{+}21, KST^{+}21, KB22, LWZ^{+}23, PKD^{+}20, PBPBS22, SMFC^{+}22, UIS^{+}22]. specification [HYL^{+}20]. specificity [GMC^{+}20]. specifies [HPO^{+}23, LLY22]. specify [WDL^{+}20]. speckle [DSMB20, KVG^{+}20, LQS23]. Speckler [LQS23]. Spectrin [DYW^{+}20, SCN^{+}23]. spectrometry [ABM^{+}23, DSMB20, NGG^{+}20]. spectrometry-based [ABM^{+}23]. spectrum [EEW^{+}22, WPCB^{+}21]. speed [KHFK^{+}20]. sperm [BW23, BNV^{+}23]. spermathecal [KST^{+}23]. spermatogenesis [FY20]. SPG11 [HHGR21]. SPG12 [PMSO^{+}23]. SPG15 [HHGR21]. sphingolipid [BCM^{+}22, HSSK20, LKMM^{+}23]. sphingomyelin [OMI22]. sphingosine [RCA^{+}21]. sphingosine- [RCA^{+}21]. SPIN [BSC22]. spinal [HGG^{+}23, RH23]. spindle [BP20, CYH^{+}21, DES^{+}23, EMEZ^{+}20, FDA21, GNL^{+}20, HESH^{+}22, JMB^{+}20, KRC^{+}22, KMW20, LSD20b, MSB^{+}21, MKO^{+}21, NBC^{+}21, RVNS21, SBEB20, SMHH^{+}20, SKS^{+}23, WMS^{+}20, WLM^{+}21]. spindle-independent [SMHH^{+}20]. spindles [SdRVH^{+}21]. Spindly [dAC^{+}22]. spine [BS20b]. spine-like [BS20b]. spines [YCC^{+}21]. spliceosome [MGM22]. splicing [MLL^{+}20, PSC^{+}20, SCB^{+}20]. spontaneous [MPVD^{+}21]. spores [WZZ^{+}23]. sprouting [YMAS20]. squad [SMK20]. squamous [MFC^{+}20]. SREBF2 [HCL^{+}21]. SREBF2-regulated [HCL^{+}21]. stability [DMR^{+}20, GSC^{+}20, KNA^{+}22, LLX^{+}21, ORCT^{+}20, WLM^{+}20]. stabilization [ZVL^{+}23]. stabilize [BP22, RLS^{+}20]. stabilizes [ARCM20, GOR^{+}20, SCN^{+}23, SWS^{+}21a, vdBVS^{+}23]. stabilizing [ZBY^{+}21]. Stable [MSJ20, ATS^{+}21, JIBK23, SMD^{+}21, WMS^{+}21]. stages [EMY^{+}22]. stalled [DMR^{+}20, MYC^{+}23]. stalling [CHPF^{+}21a, CHPF^{+}21b]. Starting [CS21d]. starvation [ATTF20, BCWM21]. starved [ME21]. state [CPC^{+}20, CKR^{+}20, JRGH21, PMB^{+}22, SKF^{+}23, KB21]. staying [Dus21]. STED [WDRRF^{+}23]. Stem [BDR20, AR20, BHS^{+}21, DCK^{+}20, Dus21, FFZ^{+}22, HZN^{+}21, JRGH21, LD21, LW20a, MMC20, NTA^{+}21, PDW^{+}20, SLES20, STS21, TMG^{+}21, VZQ^{+}21]. step [GPW^{+}22]. stereocilia [KLB^{+}22]. Sterol [FDG^{+}21, MVM20, OYJJ23, dDFGP^{+}21]. STIL [SWN^{+}22]. stimulated [RDW^{+}20]. stimulates [TKK^{+}20]. STING [FWP^{+}20, HCWX^{+}22, RZN^{+}22]. stings [RS22]. Stochastic [DJI^{+}21, CYL^{+}20]. stoichiometry [CWN^{+}23]. stops [Kin21]. storage [AFB^{+}20, PTS^{+}22]. straight [AII^{+}21]. Straightening [Gar21]. strand [KMJ^{+}23]. stratified [HDG22]. strength [DKCT21, FGBD^{+}21, MSJ20, OKH^{+}20]. strengthens [GKRL^{+}23]. Stress [GG20, HBDC^{+}20, HYX^{+}20, JWB^{+}22, MBW22, MP22c, YPM^{+}21, AMMK^{+}22, BVYW20, CNL^{+}21, CLL^{+}21b, FMY^{+}21, GCS^{+}20, GLM^{+}22, ITM^{+}21, KPA^{+}16, KPA^{+}20, KST^{+}23, LW20b, MLQ^{+}21, MMSP20, MP21d, OMI22, PKA20, RZN^{+}22, SLL^{+}21, SLL^{+}23, SPT^{+}09, SPT^{+}21, SSO^{+}20, ZMW^{+}22]. Stress-buffering [MP22c]. Stress-induced [HBDC^{+}20, RZN^{+}22]. Stress-responsive [MBW22, SSO^{+}20]. Stressed [Col22b, Col22a]. Stressed-out [Col22b, Col22a]. stressors [WB21]. striatal [CKW^{+}22]. striated [SvDSW^{+}20]. stringency [GNML^{+}20]. STRIPAK [DDD^{+}20]. stripping [ARCM20]. stromal [BCS^{+}21]. Structural [AHLR22, YZY^{+}20, CWN^{+}23, RLK^{+}20, WLM^{+}20, YCC^{+}21]. structurally [KSS^{+}20b, KSS^{+}20c]. structure [GSP^{+}20, LSD20b, RCH^{+}20, SPL^{+}20, ZS21]. structures [BS20b, Bog21, GMD^{+}23, MLvdL^{+}21, WRG23]. STX17 [RZN^{+}22]. Subcellular [FAS^{+}21, KSM^{+}21a, PKH^{+}20, SHA20, WAOS^{+}21, ZMS^{+}20]. subcomplexes [TNLPF20]. subdomain [LLY22]. subdomains [CEM^{+}20]. subsequently [MTW^{+}23]. subset [CWKP23, MOS^{+}20]. substrate [BKR^{+}22]. subtypes [WDL^{+}20]. Subunit [CLC^{+}21, KKP^{+}21]. subunits [KPA^{+}16, KPA^{+}20]. successful [GGA21]. sufficient [DJI^{+}21, JMY^{+}23]. sugarcoat [BW20]. sulfate [ICMM20, SMK20]. SUMOylation [Mar21, PKY^{+}20, PSP^{+}21, SWS^{+}21a]. super [BLZ^{+}21, MLvdL^{+}21]. super-complex [BLZ^{+}21]. super-structures [MLvdL^{+}21]. SuperPlots [LVMFL20]. Superresolution [TWH^{+}21]. supply [BPF^{+}21]. support [FPZ^{+}22, MNvdS^{+}20, TEH^{+}20, VBG^{+}22, WZK^{+}23]. supported [RRBW^{+}21]. suppress [CSS20, SLS^{+}23]. suppresses [CYL^{+}20, OKH^{+}20, OMI22, ZDGB^{+}22]. suppressing [FOR^{+}20]. suppression [LSOM23, RG23]. suppressive [AHvR^{+}20]. suppressor [GSC^{+}20, YLH^{+}22]. supramolecular [YKSC^{+}22]. Surf4 [DF22]. surface [BMF^{+}23, HGG^{+}23, LCM22, Sea21, TJAG^{+}21]. surfaces [vLEM^{+}20]. surprising [ZMS^{+}20]. surrounding [UTR^{+}23]. surveillance [Köh21]. Surveilling [MP21d]. survey [FSC22]. survival [EE22, FPMS^{+}21, KKN^{+}21, LCB^{+}23, SSO^{+}20]. Susana [O’D20a]. susceptibility [WMM^{+}23]. sustained [VCS^{+}22]. SUV39H2 [BHS^{+}21]. SVBP [RRCS^{+}23]. Svf1 [LKMM^{+}23]. switch [BJPH^{+}20, GDB^{+}20, Let20, MLL^{+}20, RGP^{+}22, Sea21]. switches [DKCT21, ZPG^{+}23]. Synapse [LD21, ACPR21, BB20, LAH^{+}21, SHLS22, TRJ^{+}20, WTS^{+}21, WH22]. Synapses [Alm21, DSG21, ZVC^{+}21]. Synaptic [KAH^{+}21, AVC^{+}22, BSH^{+}22, GLGL^{+}21, OKH^{+}20, PKC^{+}22, PGD^{+}20, SHLS22]. synaptojanin [PGW^{+}21]. synaptonemal [BZD^{+}21, HČK^{+}20, ZXW^{+}20]. Syncrip [TRJ^{+}20]. Syncrip/hnRNP [TRJ^{+}20]. syncytium [DdCVT22]. syndecan [ZVC^{+}21]. syndrome [MH22]. synergistically [HZN^{+}21, LMJ^{+}20, ZAR^{+}21]. synthase [WCL^{+}23]. synthase-like [WCL^{+}23]. synthases [FDG^{+}21]. synthesis [HSL^{+}20, PSS^{+}20, TWY^{+}22, dDFGP^{+}21]. synthetic [MNvdS^{+}20, NMO^{+}22]. system [BSB^{+}21, LLK^{+}21, WB21]. systematic [PBPBS22]. systems [WHA20]. Syt1 [LM23, SvVV^{+}23].

t [BW20, ACPR21, BEM^{+}23, BMS^{+}22, LLX^{+}21, MWF^{+}23, RWSZ^{+}20, XHF^{+}20, ZPŠS21]. T-tubule [LLX^{+}21]. tagged [WDRRF^{+}23]. tagging [FHM^{+}20]. tail [CLC^{+}21, CM21, MOS^{+}22, PK22]. tail-anchored [CLC^{+}21, CM21, MOS^{+}22]. tale [BSC22, TP20]. Talin [CJS^{+}21, ALC^{+}20, AKN^{+}22, GPEC^{+}23]. talin-1 [GPEC^{+}23]. talk [VOR^{+}21]. TANGLED1 [MDB^{+}20]. target [GCL^{+}21, WBH^{+}21]. targeted [BDH^{+}21, EYC^{+}20]. targeting [BHK20, CH22, DLZ^{+}20, EEW^{+}22, FUBS22, LFF^{+}22, LZZ^{+}21, MMSP20, RMM^{+}21]. targets [GG20, LRB^{+}22]. TAT1 [RGP^{+}22]. Tau [ZVM^{+}20]. taxol [LSOM23]. TBC1D18 [HMSF22]. TBK1 [ZRO^{+}23]. TDP [DSY^{+}22, GWR^{+}21, HCL^{+}21]. TDP-43 [DSY^{+}22, GWR^{+}21, HCL^{+}21]. teach [MMC20]. tearing [KWGR23]. Teasing [DSG21]. techniques [DES^{+}23]. Telomerase [PHAM^{+}20]. Telomere [VZQ^{+}21, PRMF^{+}23]. Temporal [EM22, BHK20, HYL^{+}20]. Tensin3 [AKN^{+}22, ZAK^{+}22]. Tension [CRZ^{+}21, KST^{+}23, DOA^{+}22, DYW^{+}20, GL20, McC21, MMKM21, PGH^{+}23, PAS^{+}22]. Tension-dependent [KST^{+}23]. term [GLGL^{+}21, MPVD^{+}21]. Terminal [YMH^{+}20, BZC^{+}21, CMM^{+}20, SYW^{+}20, SBL^{+}21]. terminals [WDB^{+}21]. terminus [CMN^{+}22, RVNS21]. tether [KHB^{+}22, ZDM^{+}22]. tethering [MRH^{+}23, RLK^{+}20]. tethers [HH22, HHD^{+}20]. tetraspanin [KST^{+}22]. tetraspanins [LMRG20]. Tex19.1 [RDL^{+}20]. Tex2 [DCG^{+}23]. -catenin [vdGM22]. -tubulins [MW21]. 51 [LMS^{+}21]. -secretase [WMS^{+}21]. -tubulin [BWA^{+}23, TG21]. TFE3 [YJX^{+}20]. TFEB [WCG^{+}22, YJX^{+}20]. TFEB/TFE3 [YJX^{+}20]. TGF [LCB^{+}23, ZDGB^{+}22]. TGF- [ZDGB^{+}22]. TGN [RCM^{+}23b]. Thank [WME22]. their [LD21, RCF^{+}22, WDRRF^{+}23, ZS21, ZBM^{+}22]. them [Kin21]. theta [CBJ^{+}21]. Think [Zar20, GY20]. thought [HI21]. Three [FPZ^{+}22, JJ23, VLdRADJ22]. three-color [VLdRADJ22]. threshold [DACG^{+}21]. thrombopoietin [VGK^{+}21]. throughout [JIBK23]. throughput [BDH^{+}21, LYL^{+}22]. Thy [Bak23]. Tian [MP21d]. tight [HSF^{+}23, VCS^{+}22]. Time [MKD^{+}21, Cas21, CS21c, FLW^{+}23, GH20, O’D20b]. Time-resolved [MKD^{+}21]. timely [SWS^{+}21a]. TIMP [ESH^{+}23]. TIMP-1 [ESH^{+}23]. TIP [SHBF^{+}20, DLK^{+}21, FPZ^{+}22, BDH^{+}21]. tip-coupling [FPZ^{+}22]. TIP-seq [BDH^{+}21]. tips [CPW^{+}23, GOR^{+}20, LGB^{+}21]. Tissue [PKD^{+}20, SLES20, GSB^{+}20, KST^{+}23, LTL^{+}20, LJT^{+}22, MBA^{+}22, MWF^{+}23, SGL^{+}23]. tissue-resident [LJT^{+}22]. Tissue-wide [SLES20]. TKS5 [ZMMM^{+}20]. TLN1 [GPEC^{+}23]. TLNRD1 [CJS^{+}21]. TMEM11 [GCW^{+}23, McW23]. TMEM41B [JLS^{+}22, LWD^{+}21]. TMEM55 [DCG^{+}23]. TMEM55-dependent [DCG^{+}23]. TNIP1 [ZRO^{+}23]. together [BG22, HMSF22]. TOM1L1 [CDLZ^{+}22]. tomography [BMF^{+}23, NBI^{+}22, PMB^{+}20]. Too [DG22]. tool [LRB^{+}22]. toolbox [LZT^{+}23, MRG^{+}20, NvGK20]. tools [FBVD^{+}22, SHLS22]. Topoisomerase [PKY^{+}20, SBBJ21]. Topological [CLH^{+}20]. topology [BKR^{+}22]. TORC1 [CLH21, TKK^{+}20, YZW^{+}20]. TORC1-mediated [LGL^{+}23]. Torsin [RLV^{+}20]. touch [Nag23]. toxicity [CYU^{+}21, SSF^{+}22]. toxin [JKZ^{+}22]. Toxoplasma [OHHR23]. TPX2 [SKS^{+}23]. tracing [LJT^{+}22]. track [GH20, MRG^{+}20]. Tracking [Cas21, WPS22]. tracks [MV20, STvT23]. traffic [BSH^{+}22, CCFN^{+}20, HSW^{+}22, HSU^{+}20, WDB^{+}21]. trafficking [BLU21, BSB^{+}21, CFD^{+}20, FCCH21, KKN^{+}21, LMS^{+}21, LRM^{+}20, MYK^{+}20, MYK^{+}21, MYK^{+}22, PFPB^{+}20, WLM^{+}20, WESR22, YMH^{+}20]. TRAIL [BDS^{+}21, Ove21]. trains [PL22]. traits [WM20]. trajectory [HPO^{+}23]. trans [AVC^{+}22, GPL^{+}21, OYJJ23, ZXY^{+}23]. trans-Golgi [GPL^{+}21, OYJJ23, ZXY^{+}23]. trans-synaptic [AVC^{+}22]. transbilayer [KHFK^{+}20]. transcription [ANRS^{+}20, BTF^{+}20, CZTL21, DHB^{+}21, HDW^{+}21, KJ23, SLL^{+}21, SLL^{+}23, UIS^{+}22]. transcriptional [HYX^{+}20, JML^{+}21]. Transcytosis [AVC^{+}22]. transduction [KHFK^{+}20]. transfer [DTG23, DRZ^{+}23, HCWX^{+}22, HSW^{+}22, HAW^{+}22, MOS^{+}22]. transformation [YSC^{+}02, YSC^{+}21]. transient [VGO^{+}23]. transients [BS20b, GKFR20, LYS^{+}20]. transition [AR20, DCRDC^{+}22, DCS^{+}20, GVA20, LSX^{+}22, PL22, SKX^{+}23, SLES20, STvT23]. transitions [SRUdC^{+}22, dAC^{+}22]. translation [AH20b, APL^{+}21, LGL^{+}23, LMJ^{+}20, MMSP20]. Translational [GWR^{+}21]. translocation [CHZ^{+}20, DOA^{+}22, HGG^{+}23, OTOF21, WZG22]. Transmembrane [OTOF21, AHY^{+}21, ZY21]. transmission [CWZ^{+}20]. transport [AHLR22, BS20a, BLZ^{+}21, CGCR^{+}22, CBC^{+}20, DSLP20, FPMS^{+}21, HRB^{+}21, HZCX22, KRS21, LLBC^{+}20, LL22, LM23, LRL^{+}20, NvGK20, OMI22, QZX23, RP21, RCS22, SvVV^{+}23, SBEB20, TML22, WPCB^{+}21, YM21, dDFGP^{+}21]. transporter [ZLJ^{+}23]. transporters [WYL21]. transportin [YLH^{+}22]. transportin-1 [YLH^{+}22]. Traub [AO21]. TRCky [MRD21]. treatment [PHAM^{+}20]. triacylglycerol [MYT^{+}21]. tricellular [CHS^{+}22, SFO^{+}21]. Tricellulin [CHS^{+}22, SFO^{+}21, vdGM22]. Tricellulin/ [vdGM22]. Tricellulin/-catenin [vdGM22]. tricks [MMC20]. trigger [DACG^{+}21]. Triggered [MWSX23]. triggers [ALC^{+}20, CFK^{+}22, CGBMC20, ESH^{+}23, KRS21, LWL^{+}23, NPdC^{+}21, RGK^{+}22, VTL^{+}20]. Triglyceride [RGK^{+}22]. TRIM1 [SSF^{+}22]. TRIM37 [MKO^{+}21]. TRIOBP [KLB^{+}22]. trip [SS22]. TrkB [HGG^{+}23, RH23]. TRPA1 [LYS^{+}20, GKFR20]. TRPA1-dependent [LYS^{+}20, GKFR20]. Trpml [EJBB^{+}20]. TRPV4 [VOR^{+}21]. Truly [CD21]. TSA [DSMB20]. TSA-MS [DSMB20]. tube [SCK^{+}19, SCK^{+}23]. TUBGCP6 [PSC^{+}20]. tubular [BLZ^{+}21]. tubule [CYR^{+}21, LLX^{+}21]. tubule-forming [CYR^{+}21]. tubules [JDKK^{+}22, PFS^{+}22, PF21]. Tubulin [FOR^{+}20, LSOM23, NBC^{+}21, WM23, AII^{+}21, BWA^{+}23, RMA21, TG21]. tubulins [MW21]. tubulogenesis [YMH^{+}20]. Tumbleweed [DNVP23]. Tumor [LLBC^{+}20, YLH^{+}22, AHvR^{+}20, DRC^{+}20, GSC^{+}20, JCL^{+}23, LGS22, NS20]. tumor-intrinsic [NS20]. tumor-suppressive [AHvR^{+}20]. tumorigenesis [EE22, SMD^{+}21]. tune [McW23]. tuned [ZMW^{+}22]. tunes [AFB^{+}20, GL20, LLW^{+}21]. tuning [EM20, KBH^{+}22, KBB^{+}23, MC21, WRG23]. tunnels [PF21]. TuRC [WTU^{+}21]. turnover [GC22, KBH^{+}22, LSX^{+}22, NGG^{+}20, NSB^{+}21, VDC^{+}20]. Tweaking [Bez22]. Twinfilin [SHGG21]. twist [PK22]. Two [ME21, CJK^{+}22, GPW^{+}22, LLW^{+}21, MRH^{+}23]. two-step [GPW^{+}22]. type [GKRL^{+}23, SBL^{+}21, DLK^{+}21, ZLJ^{+}22]. Tyramide [DSMB20]. tyrosinated [KLC^{+}20]. Tyrosine [AAR^{+}21, FHM^{+}22, LWG^{+}22]. tyrosine-based [LWG^{+}22].

ubiquitin [AHY^{+}21, BMM^{+}20, Ike20, LSD^{+}21, WB21, YZW^{+}20, ZSJE20, DSLP20, SNN20, ZCL^{+}22]. ubiquitin- [YZW^{+}20]. ubiquitin-proteasome [WB21]. ubiquitinated [LCM22]. ubiquitinates [SSF^{+}22]. ubiquitination [OCB^{+}21, YKK^{+}20, dCS^{+}21]. Ubiquitous [PZWW21]. Ubiquitylation [DHTP22, Cas22]. UBR [TSL^{+}20]. ULK [SYW^{+}20]. ULK1 [LZZ^{+}21]. Ultrastructural [MMKM21, vdBdHLK22]. ultrastructure [BMF^{+}23, DRW^{+}23]. unaffected [BJSOS^{+}20, BJSOS^{+}21]. unconventional [CGBMC20, LSG^{+}22, PK22, WZG22, WLW^{+}22]. underlie [AVC^{+}22, MTR^{+}20, dAC^{+}22]. underlies [CKW^{+}22]. unexpected [BSB^{+}21]. unfolded [JJ23, SPT^{+}09, SPT^{+}21, TSL^{+}20]. Ungewickell [TB20b]. unguided [LZT^{+}23]. uniform [DdCVT22]. Union [KB21]. unique [ZAR^{+}21]. unit [RCA^{+}23]. unite [WPM21]. units [ZXW^{+}20]. universal [HJL^{+}22]. unleash [McW23]. unravel [VV23]. Unraveling [VVW^{+}23]. Untangling [HHT^{+}20]. untethering [WKC^{+}22]. until [Ver21]. unusual [KSS^{+}20a]. unveil [MP22f]. update [Nag23]. upon [MVM20, SNL^{+}22]. UPR [LGL^{+}23, XDY^{+}22]. upregulate [LWZ^{+}23]. upregulation [SLS^{+}23]. UPS [MRD21]. uptake [LYP^{+}21]. use [FBVD^{+}22, Set21]. using [ABM^{+}23, BMF^{+}23, BD20, CPS^{+}22, LM21, LSS^{+}23, MAW^{+}22, Tai22, UIS^{+}22, VVW^{+}23, WAOS^{+}21]. USP10 [KPA^{+}20, KPA^{+}16]. USP19 [CCH^{+}21]. USP20 [CM21]. USP20/33 [CM21]. USP22 [BCC^{+}21]. USP22/nonstop [BCC^{+}21]. USP9X [CHPF^{+}21a, CHPF^{+}21b]. utilize [YMAS20]. UVSSA [SLL^{+}21, SLL^{+}23].

v [BJSOS^{+}20, BJSOS^{+}21, KST^{+}22, DLK^{+}21, FWP^{+}20, HJL^{+}22, IvCD^{+}21, LGL^{+}23, CWKP23]. V-ATPase [FWP^{+}20, HJL^{+}22, LGL^{+}23]. V-ATPase/TORC1-mediated [LGL^{+}23]. v-Class [BJSOS^{+}20, BJSOS^{+}21]. v-SNARE [CWKP23]. V. [JKZ^{+}22]. vacuolar [BCM^{+}22, EEW^{+}22, MAW^{+}22]. vacuole [KAH^{+}21, LW20b, YZW^{+}20, ZBM^{+}22]. Vaishnavi [MP21c]. valves [DSG^{+}23]. VAMP4 [LFF^{+}22]. VAPB [KHB^{+}22]. variability [LVMFL20, SMM^{+}21]. varicosities [CVT^{+}21]. vascular [CFV^{+}21, GMIC^{+}20]. VASH1 [RRCS^{+}23]. VASH2 [RRCS^{+}23]. VASP [MRWK^{+}22]. vault [WTS^{+}21]. Vav [PKC^{+}22]. VCAM [HZN^{+}21]. VCAM-1 [HZN^{+}21]. VCC [SMM^{+}21]. VCP [JTM^{+}23]. VE [EM20, GMIC^{+}20]. VE-cadherin [EM20, GMIC^{+}20]. versatile [MLS^{+}22]. versus [CML20, MSJ20]. vertebrates [Pro20]. vertices [vdGM22]. Vesicle [CWKP23, BSH^{+}22, GPL^{+}21, PGD^{+}20, SJL^{+}22, WDB^{+}21, WPS22, YMH^{+}20]. Vesicle-associated [CWKP23]. vesicles [CMT^{+}21, CH22, FWP^{+}20, GNML^{+}20, LLBC^{+}20, OWY^{+}23, PF21, RPM^{+}21, STS21, WPCB^{+}21, WCC^{+}23]. via [BVYW20, CCH^{+}21, CLL^{+}21b, DSY^{+}22, FWP^{+}20, FER^{+}23, GDB^{+}20, GSLH^{+}21, GLGL^{+}21, HRB^{+}21, HYL^{+}20, HCB^{+}23, JKL^{+}22, KHB^{+}22, LJJ^{+}21, LYL^{+}23, MHS^{+}20, MC21, OCB^{+}21, PKY^{+}20, PRB^{+}20, RCF^{+}22, RZN^{+}22, RSWP20, SKN^{+}21, SLS^{+}23, WMM^{+}23, WZtM^{+}20, ZLJ^{+}23]. vinculin [ALC^{+}20]. Viral [NMO^{+}22, CLZ^{+}20, MP21b, MP22a, SBV^{+}20]. Virus [Nag23]. viruses [MNvdS^{+}20]. visceral [SHD^{+}21]. visible [KYR^{+}22]. Visionary [Ped22]. visualize [JIBK23]. visualized [PFS^{+}22]. Visualizing [TML22]. vivo [GPW^{+}22, GVD^{+}20b, LRB^{+}22, SMFC^{+}22, Tan23, XVW^{+}23, GVD^{+}20a]. VMP1 [JLS^{+}22, LWD^{+}21]. volume [Gal23, GMD^{+}23, RMM^{+}21]. Volumetric [KRC^{+}22]. Vps13 [TWY^{+}22, AHLR22, LLLR20, DTG23]. Vps13-like [TWY^{+}22]. VPS13C [HCWX^{+}22]. VPS13C/PARK23 [HCWX^{+}22]. VPS13D [BWK^{+}21, GSLH^{+}21, SFWB21]. vulnerability [CKW^{+}22, NMO^{+}22]. wait [Ver21]. Want [PF21]. Wapl [SPL^{+}20]. WASH [DNVP23]. WASP [BKR^{+}22, LYL^{+}23, RS22]. water [HZCX22]. wave [AR20, Mar21, YLC^{+}21]. way [CNL^{+}21, CS21a, RG23]. Wbox2 [CMM^{+}20]. WBP11 [PSC^{+}20]. WD40 [FWP^{+}20]. WDR44 [HVPM20]. WDR60 [DCRDC^{+}22]. WDR60-mediated [DCRDC^{+}22]. WDR91 [LLY22, XZJ^{+}21]. weak [ZXW^{+}20]. web [YMH^{+}20]. weights [Cas23a]. well [BW23]. Werb [SH20]. wetting [KMK21]. which [CJS^{+}21]. Who [BLU21]. whole [KSM^{+}21a, MSX^{+}21]. whole-genome [KSM^{+}21a]. Wickström [MP22i]. wide [SLES20]. width [KRC^{+}22]. William [GPES21]. WIPI2 [FCHM20]. within [FGBD^{+}21, JFM^{+}22, PCZ^{+}23, PMB^{+}20]. without [Mer21]. Wnt [GL20, HRB^{+}21, KSWC22, PPB^{+}21, WPM21, YSC^{+}02, YSC^{+}21]. Wnt/ [GL20]. Wnts [MHS^{+}20]. women [MP21c]. workflow [RMM^{+}21]. wound [GSB^{+}20]. wounded [FCCH21]. Wwp1 [LSD^{+}21].

X [SWT^{+}22]. Xbp1 [AMMK^{+}22]. XCTK2 [EMEZ^{+}20]. XK [CKW^{+}22]. XMAP215 [FAHZ21]. XPF [ITM^{+}21].

Y635 [ZAK^{+}22]. Yan [MP21b, O’D20b]. Yang [Cas21]. Ye [MP21d]. Yeast [LKMM^{+}23, CEM^{+}20, CWN^{+}23, Col22a, Col22b, DLK^{+}21, EEW^{+}22, HHD^{+}20, KWdB^{+}20, LCM22, MSC^{+}20, MMKM21, NBC^{+}21, O’D22, OCLB21, RVNS21, WPS22, ZJH22]. Yogesh [MP23c]. Yorkie [SCK^{+}19, SCK^{+}23]. YY1 [AZR^{+}22].

zapERtrap [BSB^{+}21]. Zc3h10 [APL^{+}21]. zebrafish [GSB^{+}20, LRM^{+}20, SMFC^{+}22, TEH^{+}20]. Zena [SH20]. Zhong [MP22a]. Zika [CLZ^{+}20]. ZIP [LJJ^{+}21]. ZIP-10 [LJJ^{+}21]. ZIP-10/bZIP-mediated [LJJ^{+}21]. ZNF416 [JML^{+}21]. zone [DCRDC^{+}22, LSX^{+}22, PL22, SKX^{+}23, SGN^{+}20]. ZRANB3 [MYC^{+}23]. zygotes [SdRVH^{+}21].

References

Allen:2020:CMR

[AAF^{+}20] Elizabeth A. Allen, Clelia Amato, Tina M. Fortier, Panagiotis Velentzas, Will Wood, and Eric H. Baehrecke. A conserved myotubularin-related phosphatase regulates autophagy by maintaining autophagic flux. Journal of Cell Biology, 219(11): e201909073, November 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

             org/jcb/article/219/11/e201909073/152081/A-conserved-
             myotubularin-related-phosphatase.

Albacete-Albacete:2020:EDD

[AANLL^{+}20] Lucas Albacete-Albacete, Inmaculada Navarro-Lérida, Juan Antonio López, Inés Martín-Padura, Alma M. Astudillo, Alessia Ferrarini, Michael Van-Der-Heyden, Jesús Balsinde, Gertraud Orend, Jesús Vázquez, and Miguel Ángel del Pozo. ECM deposition is driven by caveolin-1–dependent regulation of exosomal biogenesis and cargo sorting. Journal of Cell Biology, 219(11):e202006178, November 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/219/11/e202006178/211453/ECM-
deposition-is-driven-by-caveolin-1-dependent.

Anwar:2021:TPS

[AAR^{+}21] Mumtaz Anwar, Md Ruhul Amin, Vijay Avin Balaji Ragunathrao, Jacob Matsche, Andrei Karginov, Richard D. Minshall, Gary C. H. Mo, Yulia Komarova, and Dolly Mehta. Tyrosine phosphorylation of S1PR1 leads to chaperone BiPmediated import to the endoplasmic reticulum. Journal of Cell Biology, 220(12):e202006021, December 6, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/12/
e202006021/212707/Tyrosine-phosphorylation-of-S1PR1-
leads-to.

Aguilera:2022:CFI

[ABB^{+}22] Anabella Aguilera, Federico Berdun, Carlos Bartoli, Charlotte Steelheart, Matías Alegre, Hülya Bayir, Yulia Y. Tyurina, Valerian E. Kagan, Graciela Salerno, Gabriela Pagnussat, and María Victoria Martin. C-ferroptosis is an iron-dependent form of regulated cell death in cyanobacteria. Journal of Cell Biology, 221(2):e201911005, February 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/221/2/e201911005/212878/C-
ferroptosis-is-an-iron-dependent-form-of.

Ashour:2023:RPD

[ABM^{+}23] Mohamed E. Ashour, Andrea K. Byrum, Alice Meroni, Jun Xia, Saurabh Singh, Roberto Galletto, Susan M. Rosenberg, Alessandro Vindigni, and Nima Mosammaparast. Rapid profiling of DNA replication dynamics using mass spectrometrybased analysis of nascent DNA. Journal of Cell Biology, 222(4): e202207121, April 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/222/4/e202207121/213875/Rapid-profiling-
of-DNA-replication-dynamics-using.

Allam:2021:DCF

[ACPR21] Amr H. Allam, Mirren Charnley, Kim Pham, and Sarah M. Russell. Developing T cells form an immunological synapse for passage through the β-selection checkpoint. Journal of Cell Biology, 220(3):e201908108, March 1, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/3/
e201908108/211687/Developing-T-cells-form-an-immunological-
synapse.

Alshafie:2020:RRS

[AFB^{+}20] Walaa Alshafie, Vincent Francis, Klaudia Bednarz, Yingzhou Edward Pan, Thomas Stroh, and Peter S. McPherson. Regulated resurfacing of a somatostatin receptor storage compartment fine-tunes pituitary secretion. Journal of Cell Biology, 219(1):e201904054, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Abad:2022:MBS

[AGH^{+}22] Maria Alba Abad, Tanmay Gupta, Michael A. Hadders, Amanda Meppelink, J. Pepijn Wopken, Elizabeth Blackburn, Juan Zou, Anjitha Gireesh, Lana Buzuk, David A. Kelly, Toni McHugh, Juri Rappsilber, Susanne M. A. Lens, and A. Arockia Jeyaprakash. Mechanistic basis for Sgo1-mediated centromere localization and function of the CPC. Journal of Cell Biology, 221 (8):e202108156, August 1, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/221/8/e202108156/213318/Mechanistic-
basis-for-Sgo1-mediated-centromere.

Abrisch:2020:FFM

[AGW^{+}20] Robert G. Abrisch, Samantha C. Gumbin, Brett Taylor Wisniewski, Laura L. Lackner, and Gia K. Voeltz. Fission and fusion machineries converge at ER contact sites to regulate mitochondrial morphology. Journal of Cell Biology, 219(4): e201911122, April 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/4/e201911122/133843/Fission-and-
fusion-machineries-converge-at-ER.

Abouelezz:2020:NFA

[AH20a] Amr Abouelezz and Pirta Hotulainen. NuMA1 facilitates the assembly of the axon initial segment by promoting the retention of neurofascin-186. Journal of Cell Biology, 219(2):e201911139, February 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/219/2/e201911139/133623/NuMA1-facilitates-the-
assembly-of-the-axon-initial.

Ali:2020:MTD

[AH20b] Levi Ali and Cole M. Haynes. Mitochondrial translation, dynamics, and lysosomes combine to extend lifespan. Journal of Cell Biology, 219(6):e202005084, June 1, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/6/
e202005084/151796/Mitochondrial-translation-dynamics-
and-lysosomes.

Adlakha:2022:SBI

[AHLR22] Jyoti Adlakha, Zhouping Hong, PeiQi Li, and Karin M. Reinisch. Structural and biochemical insights into lipid transport by VPS13 proteins. Journal of Cell Biology, 221(5):e202202030, May 2, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/5/e202202030/213104/Structural-and-biochemical-insights-
into-lipid.

Au:2020:NMG

[AHQ20] Franco K. C. Au, Bill K. T. Hau, and Robert Z. Qi. Nek2mediated GAS2L1 phosphorylation and centrosome-linker disassembly induce centrosome disjunction. Journal of Cell Biology, 219(5):e201909094, May 4, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/219/5/e201909094/151657/Nek2-
mediated-GAS2L1-phosphorylation-and.

Armistead:2020:EAC

[AHvR^{+}20] Joy Armistead, Julia Hatzold, Anna van Roye, Evelin Fahle, and Matthias Hammerschmidt. Entosis and apical cell extrusion constitute a tumor-suppressive mechanism downstream of Matriptase. Journal of Cell Biology, 219(2):e201905190, February 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/2/
e201905190/132730/Entosis-and-apical-cell-extrusion-
constitute-a.

Arines:2021:STR

[AHY^{+}21] Felichi Mae Arines, Aaron Jeremy Hamlin, Xi Yang, Yun-Yu Jennifer Liu, and Ming Li. A selective transmembrane recognition mechanism by a membrane-anchored ubiquitin ligase adaptor. Journal of Cell Biology, 220(1):e202001116, January 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/1/e202001116/211632/A-selective-
transmembrane-recognition-mechanism-by.

Ayukawa:2021:GDF

[AII^{+}21] Rie Ayukawa, Seigo Iwata, Hiroshi Imai, Shinji Kamimura, Masahito Hayashi, Kien Xuan Ngo, Itsushi Minoura, Seiichi Uchimura, Tsukasa Makino, Mikako Shirouzu, Hideki Shigematsu, Ken Sekimoto, Benoît Gigant, and Etsuko Muto. GTP-dependent formation of straight tubulin oligomers leads to microtubule nucleation. Journal of Cell Biology, 220(4):??, April 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/220/4/e202007033/211760/GTP-
dependent-formation-of-straight-tubulin.

Atherton:2022:TIT

[AKN^{+}22] Paul Atherton, Rafaella Konstantinou, Suat Peng Neo, Emily Wang, Eleonora Balloi, Marina Ptushkina, Hayley Bennett, Kath Clark, Jayantha Gunaratne, David Critchley, Igor Barsukov, Edward Manser, and Christoph Ballestrem. Tensin3 interaction with talin drives the formation of fibronectin-associated fibrillar adhesions. Journal of Cell Biology, 221(10):e202107022, October 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/221/10/e202107022/213452/Tensin3-interaction-
with-talin-drives-the.

Atherton:2020:RTA

[ALC^{+}20] Paul Atherton, Franziska Lausecker, Alexandre Carisey, Andrew Gilmore, David Critchley, Igor Barsukov, and Christoph Ballestrem. Relief of talin autoinhibition triggers a forceindependent association with vinculin. Journal of Cell Biology, 219(1):e201903134, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Almeida:2021:SAU

[Alm21] Claudia Guimas Almeida. Synapses have autophagy under control. Journal of Cell Biology, 220(6):e202105008, June 7, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/6/
e202105008/212106/Synapses-have-autophagy-under-controlSynapses-
have.

A:2020:LAC

[ALPH20] Mu A., Casey J. Latario, Laura E. Pickrell, and Henry N. Higgs. Lysine acetylation of cytoskeletal proteins: Emergence of an actin code. Journal of Cell Biology, 219(12):e202006151, December 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/219/12/e202006151/211455/Lysine-acetylation-
of-cytoskeletal-proteins.

Ali-Murthy:2021:ENC

[AMFW^{+}21] Zehra Ali-Murthy, Richard D. Fetter, Wanpeng Wang, Bin Yang, Loic A. Royer, and Thomas B. Kornberg. Elimination of nurse cell nuclei that shuttle into oocytes during oogenesis. Journal of Cell Biology, 220(7):e202012101, July 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
7/e202012101/212051/Elimination-of-nurse-cell-nuclei-
that-shuttle-into.

Athie:2020:ACN

[AMG^{+}20] Alejandro Athie, Francesco P. Marchese, Jovanna González, Teresa Lozano, Ivan Raimondi, Prasanna Kumar Juvvuna, Amaya Abad, Oskar Marin-Bejar, Jacques Serizay, Dannys Martínez, Daniel Ajona, Maria Jose Pajares, Juan Sandoval, Luis M. Montuenga, Chandrasekhar Kanduri, Juan J. Lasarte, and Maite Huarte. Analysis of copy number alterations reveals the lncRNA ALAL-1 as a regulator of lung cancer immune evasion. Journal of Cell Biology, 219(9):e201908078, September 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/219/9/e201908078/152052/Analysis-of-copy-
number-alterations-reveals-the.

Arguello-Miranda:2022:CCI

[AMMK^{+}22] Orlando Argüello-Miranda, Ashley J. Marchand, Taylor Kennedy, Marielle A. X. Russo, and Jungsik Noh. Cell cycle–independent integration of stress signals by Xbp1 promotes non-G1/G0 quiescence entry. Journal of Cell Biology, 221(1):e202103171, January 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/1/
e202103171/212720/Cell-cycle-independent-integration-
of-stress.

Acuna:2020:NRH

[ANRS^{+}20] Luciana I. Gómez Acuña, Ezequiel Nazer, Santiago A. Rodríguez-Seguí, Berta Pozzi, Valeria Buggiano, Luciano E. Marasco, Eneritz Agirre, Cody He, Mariano Alló, and Alberto R. Kornblihtt. Nuclear role for human Argonaute-1 as an estrogen-dependent transcription coactivator. Journal of Cell Biology, 219(9):e201908097, September 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/9/
e201908097/151953/Nuclear-role-for-human-Argonaute-1-
as-an-estrogen.

Arora:2020:PPI

[AO20] Amita Arora and Vesa M. Olkkonen. Protrudin in protrudinG invadopodia: Membrane contact sites and cell invasion. Journal of Cell Biology, 219(8):e202006146, August 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/8/
e202006146/151960/Protrudin-in-protrudinG-invadopodia-
Membrane.

Aridor:2021:LMT

[AO21] Meir Aridor and David J. Owen. Linton Mark Traub (1962– 2020). Journal of Cell Biology, 220(2):e202011169, February 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/2/
e202011169/211653/Linton-Mark-Traub-1962-2020-Linton-
Mark-Traub-1962.

Audano:2021:ZRA

[APL^{+}21] Matteo Audano, Silvia Pedretti, Simona Ligorio, Francesco Gualdrini, Sara Polletti, Marta Russo, Serena Ghisletti, Camilla Bean, Maurizio Crestani, Donatella Caruso, Emma De Fabiani, and Nico Mitro. Zc3h10 regulates adipogenesis by controlling translation and F-actin/mitochondria interaction. Journal of Cell Biology, 220(3):e202003173, March 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
3/e202003173/211779/Zc3h10-regulates-adipogenesis-by-
controlling.

Ambrosini:2020:NAB

[AR20] Arnaud Ambrosini and Katja Röper. “Neur”al brain wave: Coordinating epithelial-to-neural stem cell transition in the fly optic lobe. Journal of Cell Biology, 219(11):e202009040, November 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/
11/e202009040/211466/Neur-al-brain-wave-Coordinating-
epithelial-to.

Auckland:2020:CFS

[ARCM20] Philip Auckland, Emanuele Roscioli, Helena Louise Elvidge Coker, and Andrew D. McAinsh. CENP-F stabilizes kinetochoremicrotubule attachments and limits dynein stripping of corona cargoes. Journal of Cell Biology, 219(5):e201905018, May 4, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/5/
e201905018/151576/CENP-F-stabilizes-kinetochore-microtubule.

Aguilera-Romero:2023:GAR

[ARM23] Auxiliadora Aguilera-Romero and Manuel Muñiz. GPI anchors: Regulated as needed. Journal of Cell Biology, 222(5): e202303097, May 1, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/222/5/e202303097/214027/GPI-anchors-
Regulated-as-neededGPI-anchors.

Acoba:2020:PEF

[ASC20] Michelle Grace Acoba, Nanami Senoo, and Steven M. Claypool. Phospholipid ebb and flow makes mitochondria go. Journal of Cell Biology, 219(8):e202003131, August 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/8/
e202003131/151918/Phospholipid-ebb-and-flow-makes-mitochondria.

AbuRmaileh:2022:DIJ

[ASK^{+}22] Areej Abu Rmaileh, Balakrishnan Solaimuthu, Anees Khatib, Shirel Lavi, Mayur Tanna, Arata Hayashi, Michal Ben Yosef, Michal Lichtenstein, Nir Pillar, and Yoav D. Shaul. DPYSL2 interacts with JAK1 to mediate breast cancer cell migration. Journal of Cell Biology, 221(7):e202106078, July 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/7/
e202106078/213220/DPYSL2-interacts-with-JAK1-to-mediate-
breast.

Akhshi:2021:NCH

[AT21] Tara Akhshi and William S. Trimble. A non-canonical Hedgehog pathway initiates ciliogenesis and autophagy. Journal of Cell Biology, 220(1):e202004179, January 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/220/1/e202004179/211568/A-
non-canonical-Hedgehog-pathway-initiates.

Atkins:2021:CRF

[ATS^{+}21] Madison Atkins, Jiří Týč, Shahaan Shafiq, Manu Ahmed, Eloïse Bertiaux, Artur Leonel De Castro Neto, Jack Sunter, Philippe Bastin, Samuel Dale Dean, and Sue Vaughan. CEP164C regulates flagellum length in stable flagella. Journal of Cell Biology, 220(1):e202001160, January 4, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/1/
e202001160/211523/CEP164C-regulates-flagellum-length-
in-stable.

Amick:2020:PRC

[ATTF20] Joseph Amick, Arun Kumar Tharkeshwar, Gabriel Talaia, and Shawn M. Ferguson. PQLC2 recruits the C9orf72 complex to lysosomes in response to cationic amino acid starvation. Journal of Cell Biology, 219(1):e201906076, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Ahmad:2022:TTS

[AVC^{+}22] Tanveer Ahmad, Detlef Vullhorst, Rituparna Chaudhuri, Carlos M. Guardia, Nisha Chaudhary, Irina Karavanova, Juan S. Bonifacino, and Andres Buonanno. Transcytosis and transsynaptic retention by postsynaptic ErbB4 underlie axonal accumulation of NRG3. Journal of Cell Biology, 221(7): e202110167, July 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/221/7/e202110167/213222/Transcytosis-and-
trans-synaptic-retention-by.

Addicks:2022:GMM

[AZR^{+}22] Gregory C. Addicks, Hongbo Zhang, Dongryeol Ryu, Goutham Vasam, Alexander E. Green, Philip L. Marshall, Sonia Patel, Baeki E. Kang, Doyoun Kim, Elena Katsyuba, Evan G. Williams, Jean-Marc Renaud, Johan Auwerx, and Keir J. Menzies. GCN5 maintains muscle integrity by acetylating YY1 to promote dystrophin expression. Journal of Cell Biology, 221 (2):e202104022, February 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/221/2/e202104022/212962/GCN5-
maintains-muscle-integrity-by-acetylating-YY1.

Bakiri:2023:OTL

[Bak23] Latifa Bakiri. Open thy lattice osteoclast, resorb me! Journal of Cell Biology, 222(4):e202302033, April 3, 2023. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/222/4/
e202302033/213945/Open-Thy-Lattice-Osteoclast-Resorb-
me-Open-Thy.

Blumenthal:2020:MAN

[BB20] Daniel Blumenthal and Janis K. Burkhardt. Multiple actin networks coordinate mechanotransduction at the immunological synapse. Journal of Cell Biology, 219(2):e201911058, February 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/2/
e201911058/133650/Multiple-actin-networks-coordinate.

Bebelman:2023:HAB

[BBM^{+}23] Maarten P. Bebelman, Matthew J. Bovyn, Carlotta M. Mayer, Julien Delpierre, Ronald Naumann, Nuno P. Martins, Alf Honigmann, Yannis Kalaidzidis, Pierre A. Haas, and Marino Zerial. Hepatocyte apical bulkheads provide a mechanical means to oppose bile pressure. Journal of Cell Biology, 222(4):e202208002, April 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
222/4/e202208002/213840/Hepatocyte-apical-bulkheads-
provide-a-mechanical.

Becuwe:2020:FAC

[BBP^{+}20] Michel Becuwe, Laura M. Bond, Antonio F. M. Pinto, Sebastian Boland, Niklas Mejhert, Shane D. Elliott, Marcelo Cicconet, Morven M. Graham, Xinran N. Liu, Olga Ilkayeva, Alan Saghatelian, Tobias C. Walther, and Robert V. Farese, Jr. FIT2 is an acyl–coenzyme A diphosphatase crucial for endoplasmic reticulum homeostasis. Journal of Cell Biology, 219(10): e202006111, October 5, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/219/10/e202006111/152082/FIT2-is-an-acyl-
coenzyme-A-diphosphatase-crucial.

Broadbent:2023:QAA

[BBPS23] David G. Broadbent, Carlo Barnaba, Gloria I. Perez, and Jens C. Schmidt. Quantitative analysis of autophagy reveals the role of ATG9 and ATG2 in autophagosome formation. Journal of Cell Biology, 222(7):e202210078, July 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/7/
e202210078/214070/Quantitative-analysis-of-autophagy-
reveals-the.

Bear:2023:CBS

[BC23] Rachel M. Bear and Tamara Caspary. Cilia bent out of shape over dysfunctional astrocyte mitochondria. Journal of Cell Biology, 222(1):e202211123, January 2, 2023. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/222/1/
e202211123/213762/Cilia-bent-out-of-shape-over-dysfunctional.

Badmos:2021:DUN

[BCC^{+}21] Hammed Badmos, Neville Cobbe, Amy Campbell, Richard Jackson, and Daimark Bennett. Drosophila USP22/nonstop polarizes the actin cytoskeleton during collective border cell migration. Journal of Cell Biology, 220(7):e202007005, July 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/7/
e202007005/212101/Drosophila-USP22-nonstop-polarizes-
the-actin.

Barros-Carvalho:2022:BCP

[BCdS22] André Barros-Carvalho and Eurico Morais de Sá. Balancing cell polarity PARts through dephosphorylation. Journal of Cell Biology, 221(10):e202208008, October 3, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/10/
e202208008/213477/Balancing-cell-polarity-PARts-through.

Bisinski:2022:CCM

[BCM^{+}22] Daniel D. Bisinski, Inês Gomes Castro, Muriel Mari, Stefan Walter, Florian Fröhlich, Maya Schuldiner, and Ayelén González Montoro. Cvm1 is a component of multiple vacuolar contact sites required for sphingolipid homeostasis. Journal of Cell Biology, 221(8):e202103048, August 1, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/221/8/e202103048/213309/Cvm1-
is-a-component-of-multiple-vacuolar-contact.

Bessy:2021:HPP

[BCS^{+}21] Thomas Bessy, Adrian Candelas, Benoit Souquet, Khansa Saadallah, Alexandre Schaeffer, Benoit Vianay, Damien Cuvelier, Samy Gobaa, Cecilia Nakid-Cordero, Julien Lion, Jean-Christophe Bories, Nuala Mooney, Thierry Jaffredo, Jerome Larghero, Laurent Blanchoin, Lionel Faivre, Stephane Brunet, and Manuel Théry. Hematopoietic progenitors polarize in contact with bone marrow stromal cells in response to SDF1. Journal of Cell Biology, 220(11):e202005085, November 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/11/
e202005085/212662/Hematopoietic-progenitors-polarize-
in-contact-with.

Belew:2021:GCC

[BCWM21] Mezmur D. Belew, Emilie Chien, Matthew Wong, and W. Matthew Michael. A global chromatin compaction pathway that represses germline gene expression during starvation. Journal of Cell Biology, 220(9):e202009197, September 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/9/
e202009197/212349/A-global-chromatin-compaction-pathway-
that.

Baum:2020:MSN

[BD20] Buzz Baum and Gautam Dey. Moving simply: Naegleria crawls and feeds using an ancient Arp2/3-dependent mechanism. Journal of Cell Biology, 219(11):e202009031, November 2, 2020.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/11/
e202009031/211468/Moving-simply-Naegleria-crawls-and-
feeds-using-an.

Broad:2020:ABK

[BDD20] Amanda J. Broad, Keith F. DeLuca, and Jennifer G. DeLuca. Aurora B kinase is recruited to multiple discrete kinetochore and centromere regions in human cells. Journal of Cell Biology, 219(3):e201905144, March 2, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/3/
e201905144/133701/Aurora-B-kinase-is-recruited-to-multiple-
discrete.

Bartlett:2021:HTS

[BDH^{+}21] Daniel A. Bartlett, Vishnu Dileep, Tetsuya Handa, Yasuyuki Ohkawa, Hiroshi Kimura, Steven Henikoff, and David M. Gilbert. High-throughput single-cell epigenomic profiling by targeted insertion of promoters (TIP-seq). Journal of Cell Biology, 220(12):e202103078, December 6, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/12/
e202103078/212821/High-throughput-single-cell-epigenomic-
profiling.

Bridges:2021:LLL

[BDK21] Mary Catherine Bridges, Amanda C. Daulagala, and Antonis Kourtidis. LNCcation: lncRNA localization and function. Journal of Cell Biology, 220(2):e202009045, February 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/2/
e202009045/211695/LNCcation-lncRNA-localization-and-
functionlncRNA.

Bajaj:2020:SCC

[BDR20] Jeevisha Bajaj, Emily Diaz, and Tannishtha Reya. Stem cells in cancer initiation and progression. Journal of Cell Biology, 219(1):e201911053, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Bozkurt:2021:TSP

[BDS^{+}21] Emir Bozkurt, Heiko Düssmann, Manuela Salvucci, Brenton L. Cavanagh, Sandra Van Schaeybroeck, Daniel B. Longley, Seamus J. Martin, and Jochen H. M. Prehn. TRAIL signaling promotes entosis in colorectal cancer. Journal of Cell Biology, 220(11):e202010030, November 1, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/11/
e202010030/212649/TRAIL-signaling-promotes-entosis-in-
colorectal.

Brusini:2022:COK

[BDT^{+}22] Lorenzo Brusini, Nicolas Dos Santos Pacheco, Eelco C. Tromer, Dominique Soldati-Favre, and Mathieu Brochet. Composition and organization of kinetochores show plasticity in apicomplexan chromosome segregation. Journal of Cell Biology, 221(9): e202111084, September 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/221/9/e202111084/213421/Composition-
and-organization-of-kinetochores-show.

Bock:2021:RPK

[BED^{+}21] Fabian Bock, Bertha C. Elias, Xinyu Dong, Diptiben V. Parekh, Glenda Mernaugh, Olga M. Viquez, Anjana Hassan, Venkateswara Rao Amara, Jiageng Liu, Kyle L. Brown, Andrew S. Terker, Manuel Chiusa, Leslie S. Gewin, Agnes B. Fogo, Cord H. Brakebusch, Ambra Pozzi, and Roy Zent. Rac1 promotes kidney collecting duct integrity by limiting actomyosin activity. Journal of Cell Biology, 220(11):e202103080, November 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
11/e202103080/212704/Rac1-promotes-kidney-collecting-
duct-integrity-by.

Beppler:2023:HCM

[BEM^{+}23] Casey Beppler, John Eichorst, Kyle Marchuk, En Cai, Carlos A. Castellanos, Venkataraman Sriram, Kole T. Roybal, and Matthew F. Krummel. Hyperstabilization of T cell microvilli contacts by chimeric antigen receptors. Journal of Cell Biology, 222(3):e202205118, March 6, 2023. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/222/3/
e202205118/213760/Hyperstabilization-of-T-cell-microvilli-
contacts.

Bezzi:2022:TNO

[Bez22] Paola Bezzi. Tweaking neural organoids to model human reactive astrocytes. Journal of Cell Biology, 221(4):e202202026, April 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/4/e202202026/213083/Tweaking-neural-organoids-to-
model-human-reactive.

Behmoaras:2021:SIB

[BG21] Jacques Behmoaras and Jesús Gil. Similarities and interplay between senescent cells and macrophages. Journal of Cell Biology, 220(2):e202010162, February 1, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/2/
e202010162/211634/Similarities-and-interplay-between-
senescent-cells.

Broussard:2022:PPI

[BG22] Joshua A. Broussard and Kathleen J. Green. Plectin pulls it together, coupling the cortical actin and intermediate filament cytoskeletons. Journal of Cell Biology, 221(3):e202201054, March 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/3/
e202201054/213012/Plectin-pulls-it-together-coupling-
the-cortical.

Bhide:2021:MCA

[BGM^{+}21] Sourabh Bhide, Denisa Gombalova, Gregor Mönke, Johannes Stegmaier, Valentyna Zinchenko, Anna Kreshuk, Julio M. Belmonte, and Maria Leptin. Mechanical competition alters the cellular interpretation of an endogenous genetic program. Journal of Cell Biology, 220(11):e202104107, November 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/11/
e202104107/212605/Mechanical-competition-alters-the-
cellular.

Brandt:2020:STC

[BHK20] James N. Brandt, Katarzyna A. Hussey, and Yumi Kim. Spatial and temporal control of targeting Polo-like kinase during meiotic prophase. Journal of Cell Biology, 219(11): e202006094, November 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/11/e202006094/152136/Spatial-and-
temporal-control-of-targeting-Polo.

Balmer:2021:SES

[BHS^{+}21] Pierre Balmer, William V. J. Hariton, Beyza S. Sayar, Vidhya Jagannathan, Arnaud Galichet, Tosso Leeb, Petra Roosje, and Eliane J. Müller. SUV39H2 epigenetic silencing controls fate conversion of epidermal stem and progenitor cells. Journal of Cell Biology, 220(4):??, April 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/4/e201908178/211810/SUV39H2-epigenetic-silencing-
controls-fate.

Bellingham-Johnstun:2021:MOC

[BJAR^{+}21] Kimberly Bellingham-Johnstun, Erica Casey Anders, John Ravi, Christina Bruinsma, and Caroline Laplante. Molecular organization of cytokinesis node predicts the constriction rate of the contractile ring. Journal of Cell Biology, 220(3):e202008032, March 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/3/
e202008032/211718/Molecular-organization-of-cytokinesis-
node.

Baumert:2020:NPS

[BJPH^{+}20] Ryan Baumert, Hong Ji, Adriana Paulucci-Holthauzen, Aaron Wolfe, Cari Sagum, Louis Hodgson, Jyothi Arikkath, Xiaojiang Chen, Mark T. Bedford, M. Neal Waxham, and Pierre D. Mc-Crea. Novel phospho-switch function of delta-catenin in dendrite development. Journal of Cell Biology, 219(11):e201909166, November 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/219/11/e201909166/152151/Novel-phospho-switch-
function-of-delta-catenin-in.

Budaitis:2021:PMK

[BJR^{+}21] Breane G. Budaitis, Shashank Jariwala, Lu Rao, Yang Yue, David Sept, Kristen J. Verhey, and Arne Gennerich. Pathogenic mutations in the kinesin-3 motor KIF1A diminish force generation and movement through allosteric mechanisms. Journal of Cell Biology, 220(4):??, April 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/

220/4/e202004227/211720/Pathogenic-mutations-in-the-
kinesin-3-motor-KIF1A.

Benito-Jardon:2020:TVC

[BJSOS^{+}20] María Benito-Jardón, Nico Strohmeyer, Sheila Ortega-Sanchís, Mitasha Bharadwaj, Markus Moser, Daniel J. Müller, Reinhard Fässler, and Mercedes Costell. α v-class integrin binding to fibronectin is solely mediated by RGD and unaffected by an RGE mutation. Journal of Cell Biology, 219(12):e202004198, December 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/12/e202004198/211518/v-Class-integrin-binding-to-

fibronectin-is-solely. See correction [BJSOS^{+}21].

Benito-Jardon:2021:CTV

[BJSOS^{+}21] María Benito-Jardón, Nico Strohmeyer, Sheila Otega-Sanchís, Mitasha Bharadwaj, Markus Moser, Daniel J. Müller, Reinhard Fässler, and Mercedes Costell. Correction: α v-Class integrin binding to fibronectin is solely mediated by RGD and unaffected by an RGE mutation. Journal of Cell Biology, 220(1):??, January 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/1/jcb.20200419812072020c/211584/
Correction-v-Class-integrin-binding-to-fibronectin. See

[BJSOS^{+}20].

Brunetti:2022:WIS

[BKR^{+}22] Rachel M. Brunetti, Gabriele Kockelkoren, Preethi Raghavan, George R. R. Bell, Derek Britain, Natasha Puri, Sean R. Collins, Manuel D. Leonetti, Dimitrios Stamou, and Orion D. Weiner. WASP integrates substrate topology and cell polarity to guide neutrophil migration. Journal of Cell Biology, 221(2):e202104046, February 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/221/2/e202104046/212945/WASP-
integrates-substrate-topology-and-cell.

Borchers:2021:WCP

[BLU21] Ann-Christin Borchers, Lars Langemeyer, and Christian Ungermann. Who’s in control? Principles of Rab GTPase activation in endolysosomal membrane trafficking and beyond. Journal of Cell Biology, 220(9):e202105120, September 6, 2021.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/9/
e202105120/212549/Who-s-in-control-Principles-of-Rab-
GTPase.

Bowman:2021:BAS

[BLZ^{+}21] Shanna L. Bowman, Linh Le, Yueyao Zhu, Dawn C. Harper, Anand Sitaram, Alexander C. Theos, Elena V. Sviderskaya, Dorothy C. Bennett, Graça Raposo-Benedetti, David J. Owen, Megan K. Dennis, and Michael S. Marks. A BLOC-1-AP-3 super-complex sorts a cis-SNARE complex into endosomederived tubular transport carriers. Journal of Cell Biology, 220 (7):e202005173, July 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/7/e202005173/212016/A-BLOC-1-AP-
3-super-complex-sorts-a-cis-SNARE.

Barad:2023:QOU

[BMF^{+}23] Benjamin A. Barad, Michaela Medina, Daniel Fuentes, R. Luke Wiseman, and Danielle A. Grotjahn. Quantifying organellar ultrastructure in cryo-electron tomography using a surface morphometrics pipeline. Journal of Cell Biology, 222(4): e202204093, April 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/222/4/e202204093/213869/Quantifying-
organellar-ultrastructure-in-cryo.

Boyer:2020:PEU

[BMM^{+}20] Nicholas P. Boyer, Laura E. McCormick, Shalini Menon, Fabio L. Urbina, and Stephanie L. Gupton. A pair of E3 ubiquitin ligases compete to regulate filopodial dynamics and axon guidance. Journal of Cell Biology, 219(1):e201902088, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Bonacina:2022:LDL

[BMS^{+}22] Fabrizia Bonacina, Annalisa Moregola, Monika Svecla, David Coe, Patrizia Uboldi, Sara Fraire, Simona Beretta, Giangiacomo Beretta, Fabio Pellegatta, Alberico Luigi Catapano, Federica M. Marelli-Berg, and Giuseppe Danilo Norata. The low-density lipoprotein receptor–mTORC1 axis coordinates CD8 + T cell activation. Journal of Cell Biology, 221(11):e202202011, November 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/11/e202202011/213488/The-low-density-lipoprotein-
receptor-mTORC1-axis.

Brukman:2023:NFS

[BNV^{+}23] Nicolas G. Brukman, Kohdai P. Nakajima, Clari Valansi, Kateryna Flyak, Xiaohui Li, Tetsuya Higashiyama, and Benjamin Podbilewicz. A novel function for the sperm adhesion protein IZUMO1 in cell–cell fusion. Journal of Cell Biology, 222(2):e202207147, February 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/222/2/e202207147/213693/A-
novel-function-for-the-sperm-adhesion-protein.

Bogan:2021:GDT

[Bog21] Jonathan S. Bogan. Granular detail of β cell structures for insulin secretion. Journal of Cell Biology, 220(2):e202012082, February 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/220/2/e202012082/211670/Granular-detail-
of-cell-structures-for-insulin.

Braschi:2022:CND

[BOW^{+}22] Bryony Braschi, Heymut Omran, George B. Witman, Gregory J. Pazour, K. Kevin Pfister, Elspeth A. Bruford, and Stephen M. King. Consensus nomenclature for dyneins and associated assembly factors. Journal of Cell Biology, 221(2):e202109014, February 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/2/
e202109014/212950/Consensus-nomenclature-for-dyneins-
and-associated.

Benzi:2020:SSA

[BP20] Giorgia Benzi and Simonetta Piatti. Silencing the spindle assembly checkpoint: Let’s play Polo! Journal of Cell Biology, 219(12):e202010053, December 7, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/12/
e202010053/211545/Silencing-the-spindle-assembly-checkpoint-
Let-s.

Babcock:2022:LGD

[BP22] Rachel L. Babcock and Kevin Pruitt. Letting go: Dishevelled phase separation recruits Axin to stabilize βcatenin. Journal of Cell Biology, 221(12):e202211001, December 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/221/12/e202211001/213691/Letting-go-
Dishevelled-phase-separation-recruits.

Basu:2021:FAM

[BPF^{+}21] Himanish Basu, Gulcin Pekkurnaz, Jill Falk, Wei Wei, Morven Chin, Judith Steen, and Thomas L. Schwarz. FHL2 anchors mitochondria to actin and adapts mitochondrial dynamics to glucose supply. Journal of Cell Biology, 220(10):e201912077, October 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/10/e201912077/212527/FHL2-anchors-mitochondria-to-
actin-and-adapts.

Bartle:2020:PER

[BRB^{+}20] Emily I. Bartle, Tejeshwar C. Rao, Reena R. Beggs, William F. Dean, Tara M. Urner, Andrew P. Kowalczyk, and Alexa L. Mattheyses. Protein exchange is reduced in calcium-independent epithelial junctions. Journal of Cell Biology, 219(6):e201906153, June 1, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/6/e201906153/151763/Protein-exchange-is-reduced-
in-calcium-independent.

Belicova:2021:AEH

[BRD^{+}21] Lenka Belicova, Urska Repnik, Julien Delpierre, Elzbieta Gralinska, Sarah Seifert, José Ignacio Valenzuela, Hernán Andrés Morales-Navarrete, Christian Franke, Helin Räägel, Evgeniya Shcherbinina, Tatiana Prikazchikova, Victor Koteliansky, Martin Vingron, Yannis L. Kalaidzidis, Timofei Zatsepin, and Marino Zerial. Anisotropic expansion of hepatocyte lumina enforced by apical bulkheads. Journal of Cell Biology, 220(10): e202103003, October 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/10/e202103003/212522/Anisotropic-
expansion-of-hepatocyte-lumina.

Bekku:2020:IAT

[BS20a] Yoko Bekku and James L. Salzer. Independent anterograde transport and retrograde cotransport of domain components of myelinated axons. Journal of Cell Biology, 219(6): e201906071, June 1, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/6/e201906071/151658/Independent-
anterograde-transport-and-retrograde.

Belote:2020:CTM

[BS20b] Rachel L. Belote and Sanford M. Simon. Ca^{2+} transients in melanocyte dendrites and dendritic spine-like structures evoked by cell-to-cell signaling. Journal of Cell Biology, 219(1): e201902014, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Bourke:2021:ZLR

[BSB^{+}21] Ashley M. Bourke, Samantha L. Schwartz, Aaron B. Bowen, Mason S. Kleinjan, Christina S. Winborn, Dean J. Kareemo, Amos Gutnick, Thomas L. Schwarz, and Matthew J. Kennedy. zapERtrap: a light-regulated ER release system reveals unexpected neuronal trafficking pathways. Journal of Cell Biology, 220(9): e202103186, September 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/9/e202103186/212461/zapERtrap-A-
light-regulated-ER-release-system.

Barbosa:2022:SDT

[BSC22] João Barbosa, Claudio E. Sunkel, and Carlos Conde. SPIN(DLY)- OFF: a tale of conformational change to control DYNEIN. Journal of Cell Biology, 221(11):e202209063, November 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/11/
e202209063/213524/SPIN-DLY-OFF-A-tale-of-conformational-
change-to.

Bryce:2023:HAC

[BSC^{+}23] Samantha Bryce, Maureen Stolzer, Daniel Crosby, Ruijin Yang, Dannie Durand, and Tina H. Lee. Human atlastin-3 is a constitutive ER membrane fusion catalyst. Journal of Cell Biology, 222 (7):e202211021, July 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/222/7/e202211021/214058/Human-atlastin-3-
is-a-constitutive-ER-membrane.

Blanchette:2022:LRE

[BSH^{+}22] Cassandra R. Blanchette, Amy L. Scalera, Kathryn P. Harris, Zechuan Zhao, Erica C. Dresselhaus, Kate Koles, Anna Yeh, Julia K. Apiki, Bryan A. Stewart, and Avital A. Rodal. Local regulation of extracellular vesicle traffic by the synaptic endocytic machinery. Journal of Cell Biology, 221(5): e202112094, May 2, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/221/5/e202112094/213086/Local-regulation-
of-extracellular-vesicle-traffic.

Black:2020:GTF

[BTF^{+}20] Labe Black, Sylvain Tollis, Guo Fu, Jean-Bernard Fiche, Savanna Dorsey, Jing Cheng, Ghada Ghazal, Stephen Notley, Benjamin Crevier, Jeremy Bigness, Marcelo Nollmann, Mike Tyers, and Catherine Ann Royer. G1/S transcription factors assemble in increasing numbers of discrete clusters through G1 phase. Journal of Cell Biology, 219(9): e202003041, September 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/219/9/e202003041/151997/G1-S-
transcription-factors-assemble-in-increasing.

Burd:2021:GCA

[Bur21] Christopher G. Burd. GRASPing for consensus about the Golgi apparatus. Journal of Cell Biology, 220(5):e202103117, May 3, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
5/e202103117/211980/GRASPing-for-consensus-about-the-
Golgi.

Begovich:2020:CMR

[BVYW20] Kyle Begovich, Anthony Q. Vu, Gene Yeo, and James E. Wilhelm. Conserved metabolite regulation of stress granule assembly via AdoMet. Journal of Cell Biology, 219(8):e201904141, August 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/8/
e201904141/151916/Conserved-metabolite-regulation-of-
stress-granule.

Buffone:2020:DSI

[BW20] Alexander Buffone, Jr. and Valerie M. Weaver. Don t sugarcoat it: How glycocalyx composition influences cancer progression. Journal of Cell Biology, 219(1):e201910070, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Bianchi:2023:MFD

[BW23] Enrica Bianchi and Gavin J. Wright. Mammalian fertilization: Does sperm IZUMO1 mediate fusion as well as adhesion? Journal of Cell Biology, 222(2):e202301035, February 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/2/
e202301035/213812/Mammalian-fertilization-Does-sperm-
IZUMO1-mediate.

Berman:2023:NBI

[BWA^{+}23] Adi Y. Berman, Michal Wieczorek, Amol Aher, Paul Dominic B. Olinares, Brian T. Chait, and Tarun M. Kapoor. A nucleotide binding–independent role for -tubulin in microtubule capping and cell division. Journal of Cell Biology, 222 (3):e202204102, March 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/222/3/e202204102/213828/A-nucleotide-
binding-independent-role-for-tubulin.

Bomba-Warczak:2021:LLM

[BWEHS21] Ewa Bomba-Warczak, Seby L. Edassery, Timothy J. Hark, and Jeffrey N. Savas. Long-lived mitochondrial cristae proteins in mouse heart and brain. Journal of Cell Biology, 220(9):e202005193, September 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/220/9/e202005193/212469/Long-
lived-mitochondrial-cristae-proteins-in-mouse.

Baldwin:2021:VPP

[BWK^{+}21] Heather A. Baldwin, Chunxin Wang, Gil Kanfer, Hetal V. Shah, Antonio Velayos-Baeza, Marija Dulovic-Mahlow, Norbert Brüggemann, Allyson Anding, Eric H. Baehrecke, Dragan Maric, William A. Prinz, and Richard J. Youle. VPS13D promotes peroxisome biogenesis. Journal of Cell Biology, 220(5): e202001188, May 3, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/220/5/e202001188/212018/VPS13D-promotes-
peroxisome-biogenesisVPS13D-and.

Bischoff:2021:SML

[BZC^{+}21] Megan E. Bischoff, Yuanwei Zang, Johnson Chu, Adam D. Price, Birgit Ehmer, Nicholas J. Talbot, Michael J. Newbold, Anurag Paul, Jun-Lin Guan, David R. Plas, Jarek Meller, and Maria F. Czyzyk-Krzeska. Selective MAP1LC3C (LC3C) autophagy requires noncanonical regulators and the C-terminal peptide. Journal of Cell Biology, 220(7):e202004182, July 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/7/
e202004182/212107/Selective-MAP1LC3C-LC3C-autophagy-
requires.

Baumann:2020:LCI

[BZD20] Claudia Baumann, Xiangyu Zhang, and Rabindranath De La Fuente. Loss of CBX2 induces genome instability and senescence-associated chromosomal rearrangements. Journal of Cell Biology, 219(11):e201910149, November 2, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/11/
e201910149/152063/Loss-of-CBX2-induces-genome-instability-
and.

Barbosa:2021:SFP

[BZD^{+}21] Pedro Barbosa, Liudmila Zhaunova, Simona Debilio, Verdiana Steccanella, Van Kelly, Tony Ly, and Hiroyuki Ohkura. SCFfbxo42 promotes synaptonemal complex assembly by downregulating PP2A-b56. Journal of Cell Biology, 220(2):e202009167, February 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/220/2/e202009167/211645/SCF-Fbxo42-promotes-
synaptonemal-complex-assembly.

Casadio:2021:HYT

[Cas21] Melina Casadio. Hongyuan Yang: Tracking lipids, one droplet at a time. Journal of Cell Biology, 220(5):e202103075, May 3, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/5/
e202103075/212008/Hongyuan-Yang-Tracking-lipids-one-
droplet-at-a.

Caswell:2022:NCU

[Cas22] Patrick T. Caswell. Non-canonical ubiquitylation makes its mark on Rap2 and cell motility. Journal of Cell Biology, 221(4):e202203029, April 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/221/4/e202203029/213094/Non-
canonical-ubiquitylation-makes-its-mark-on.

Caswell:2023:AWB

[Cas23a] Patrick T. Caswell. ARF3 weights the balance for prostate cancer metastasis. Journal of Cell Biology, 222(4):e202303037, April 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
222/4/e202303037/213942/ARF3-weights-the-balance-for-
prostate-cancer.

Chung:2023:LLP

[CAS23b] Woo-Chang Chung, Jin-Hyun Ahn, and Moon Jung Song. Liquid–liquid phase separation drives herpesvirus assembly in the cytoplasm. Journal of Cell Biology, 222(1):e202211015, January 2, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/222/1/e202211015/213772/Liquid-liquid-phase-
separation-drives-herpesvirus.

Chowdhury:2020:ASA

[CBC^{+}20] Sudeshna Roy Chowdhury, Chumki Bhattacharjee, Jason C. Casler, Bhawik Kumar Jain, Benjamin S. Glick, and Dibyendu Bhattacharyya. ER arrival sites associate with ER exit sites to create bidirectional transport portals. Journal of Cell Biology, 219(4):e201902114, April 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/219/4/e201902114/133830/ER-
arrival-sites-associate-with-ER-exit-sites-to.

Clay:2021:PDD

[CBJ^{+}21] Delisa E. Clay, Heidi S. Bretscher, Erin A. Jezuit, Korie B. Bush, and Donald T. Fox. Persistent DNA damage signaling and DNA polymerase theta promote broken chromosome segregation. Journal of Cell Biology, 220(12):e202106116, December 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
12/e202106116/212679/Persistent-DNA-damage-signaling-
and-DNA-polymerase.

Crossley:2021:CIP

[CBS^{+}21] Magdalena P. Crossley, Joshua R. Brickner, Chenlin Song, Su Mon Thin Zar, Su S. Maw, Frédéric Chédin, Miaw-Sheue Tsai, and Karlene A. Cimprich. Catalytically inactive, purified RNase H1: a specific and sensitive probe for RNA–DNA hybrid imaging. Journal of Cell Biology, 220(9):e202101092, September 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
9/e202101092/212458/Catalytically-inactive-purified-
RNase-H1-A.

Camus:2020:CCM

[CCFN^{+}20] Stéphane M. Camus, Marine D. Camus, Carmen Figueras-Novoa, Gaelle Boncompain, L. Amanda Sadacca, Christopher Esk, Anne Bigot, Gwyn W. Gould, Dimitrios Kioumourtzoglou, Franck Perez, Nia J. Bryant, Shaeri Mukherjee, and Frances M. Brodsky. CHC22 clathrin mediates traffic from early secretory compartments for human GLUT4 pathway biogenesis. Journal of Cell Biology, 219(1):e201812135, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Chai:2021:UPH

[CCH^{+}21] Peiyuan Chai, Yiru Cheng, Chuyi Hou, Lei Yin, Donghui Zhang, Yingchun Hu, Qingzhou Chen, Pengli Zheng, Junlin Teng, and Jianguo Chen. USP19 promotes hypoxia-induced mitochondrial division via FUNDC1 at ER-mitochondria contact sites. Journal of Cell Biology, 220(7):e202010006, July 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/7/
e202010006/212091/USP19-promotes-hypoxia-induced-mitochondrial.

Cason:2021:SDE

[CCV^{+}21] Sydney E. Cason, Peter J. Carman, Claire Van Duyne, Juliet Goldsmith, Roberto Dominguez, and Erika L. F. Holzbaur. Sequential dynein effectors regulate axonal autophagosome motility in a maturation-dependent pathway. Journal of Cell Biology, 220(7):e202010179, July 5, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/7/
e202010179/212171/Sequential-dynein-effectors-regulate-
axonal.

Carty:2021:TEC

[CD21] Ben L. Carty and Elaine M. Dunleavy. Truly epigenetic: a centromere finds a “neo” home. Journal of Cell Biology, 220(3): e202101027, March 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/220/3/e202101027/211780/Truly-epigenetic-
A-centromere-finds-a-neo.

Campisi:2022:CAP

[CDD^{+}22] Daniele Campisi, Laurence Desrues, Kléouforo-Paul Dembélé, Alexandre Mutel, Renaud Parment, Pierrick Gandolfo, Hélène Castel, and Fabrice Morin. The core autophagy protein ATG9A controls dynamics of cell protrusions and directed migration. Journal of Cell Biology, 221(3):e202106014, March 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/
3/e202106014/213022/The-core-autophagy-protein-ATG9A-
controls-dynamics.

Cabral-Dias:2022:FTR

[CDLZ^{+}22] Rebecca Cabral-Dias, Stefanie Lucarelli, Karolina Zak, Sadia Rahmani, Gurjeet Judge, John Abousawan, Laura F. DiGiovanni, Dafne Vural, Karen E. Anderson, Michael G. Sugiyama, Gizem Genc, Wanjin Hong, Roberto J. Botelho, Gregory D. Fairn, Peter K. Kim, and Costin N. Antonescu. Fyn and TOM1L1 are recruited to clathrin-coated pits and regulate Akt signaling. Journal of Cell Biology, 221(4):e201808181, April 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/4/e201808181/213045/Fyn-and-TOM1L1-are-recruited-
to-clathrin-coated.

Choudhary:2020:SNE

[CEM^{+}20] Vineet Choudhary, Ola El Atab, Giulia Mizzon, William A. Prinz, and Roger Schneiter. Seipin and Nem1 establish discrete ER subdomains to initiate yeast lipid droplet biogenesis. Journal of Cell Biology, 219(7):e201910177, July 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/7/
e201910177/151718/Seipin-and-Nem1-establish-discrete-
ER-subdomains.

Casas:2020:SNC

[CFD^{+}20] Maria Casas, Rut Fadó, José Luis Domínguez, Aina Roig, Moena Kaku, Shigeru Chohnan, Montse Solé, Mercedes Unzeta, Alfredo Jesús Miñano-Molina, José Rodríguez-Álvarez, Eamonn James Dickson, and Núria Casals. Sensing of nutrients by CPT1C controls SAC1 activity to regulate AMPA receptor trafficking. Journal of Cell Biology, 219(10):e201912045, October 5, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/10/e201912045/152088/Sensing-of-
nutrients-by-CPT1C-controls-SAC1.

Chakrabarti:2022:MDT

[CFK^{+}22] Rajarshi Chakrabarti, Tak Shun Fung, Taewook Kang, Pieti W. Elonkirjo, Anu Suomalainen, Edward J. Usherwood, and Henry N. Higgs. Mitochondrial dysfunction triggers actin polymerization necessary for rapid glycolytic activation. Journal of Cell Biology, 221(11):e202201160, November 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/11/
e202201160/213462/Mitochondrial-dysfunction-triggers-
actin.

Camillo:2021:LIV

[CFV^{+}21] Chiara Camillo, Nicola Facchinello, Giulia Villari, Giulia Mana, Noemi Gioelli, Chiara Sandri, Matteo Astone, Dora Tortarolo, Fabiana Clapero, Dafne Gays, Roxana E. Oberkersch, Marco Arese, Luca Tamagnone, Donatella Valdembri, Massimo M. Santoro, and Guido Serini. LPHN2 inhibits vascular permeability by differential control of endothelial cell adhesion. Journal of Cell Biology, 220(11):e202006033, November 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/220/11/e202006033/212665/LPHN2-inhibits-
vascular-permeability-by.

Cozmescu:2021:RCF

[CG21] Claudiu Andrei Cozmescu and Paul Gissen. Rab35 controls formation of luminal projections required for bile canalicular morphogenesis. Journal of Cell Biology, 220(10):e202108047, October 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/

220/10/e202108047/212638/Rab35-controls-formation-of-
luminal-projections.

Cruz-Garcia:2020:ROS

[CGBMC20] David Cruz-Garcia, Nathalie Brouwers, Vivek Malhotra, and Amy J. Curwin. Reactive oxygen species triggers unconventional secretion of antioxidants and Acb1. Journal of Cell Biology, 219(4):e201905028, April 6, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/4/
e201905028/151570/Reactive-oxygen-species-triggers-unconventional.

Celestino:2022:JID

[CGCR^{+}22] Ricardo Celestino, José B. Gama, Artur F. Castro-Rodrigues, Daniel J. Barbosa, Helder Rocha, Ennio A. d’Amico, Andrea Musacchio, Ana Xavier Carvalho, João H. Morais-Cabral, and Reto Gassmann. JIP3 interacts with dynein and kinesin-1 to regulate bidirectional organelle transport. Journal of Cell Biology, 221(8):e202110057, August 1, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/221/8/e202110057/213353/JIP3-
interacts-with-dynein-and-kinesin-1-to.

Cardamone:2022:NLP

[CGK^{+}22] Maria Dafne Cardamone, Yuan Gao, Julian Kwan, Vanessa Hayashi, Megan Sheeran, Junxiang Xu, Justin English, Joseph Orofino, Andrew Emili, and Valentina Perissi. Neuralizedlike protein 4 (NEURL4) mediates ADP-ribosylation of mitochondrial proteins. Journal of Cell Biology, 221(3):e202101021, March 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/3/e202101021/213006/Neuralized-like-protein-4-NEURL4-
mediates-ADP.

Chen:2022:AAD

[CH22] Wei Chen and Bing He. Actomyosin activity-dependent apical targeting of Rab11 vesicles reinforces apical constriction. Journal of Cell Biology, 221(6):e202103069, June 6, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/221/6/e202103069/213118/Actomyosin-activity-
dependent-apical-targeting-of.

Clancy:2021:CDU

[CHPF^{+}21a] Anne Clancy, Claire Heride, Adán Pinto-Fernández, Hannah Elcocks, Andreas Kallinos, Katherine J. Kayser-Bricker, Weiping Wang, Victoria Smith, Simon Davis, Shawn Fessler, Crystal McKinnon, Marie Katz, Tim Hammonds, Neil P. Jones, Jonathan O’Connell, Bruce Follows, Steven Mischke, Justin A. Caravella, Stephanos Ioannidis, Christopher Dinsmore, Sunkyu Kim, Axel Behrens, David Komander, Benedikt M. Kessler, Sylvie Urbé, and Michael J. Clague. Correction: The deubiquitylase USP9X controls ribosomal stalling. Journal of Cell Biology, 220(3):??, March 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/220/3/jcb.20200421102102021c/
211799/Correction-The-deubiquitylase-USP9X-controls.

See [CHPF^{+}21b].

Clancy:2021:DUC

[CHPF^{+}21b] Anne Clancy, Claire Heride, Adán Pinto-Fernández, Hannah Elcocks, Andreas Kallinos, Katherine J. Kayser-Bricker, Weiping Wang, Victoria Smith, Simon Davis, Shawn Fessler, Crystal McKinnon, Marie Katz, Tim Hammonds, Neil P. Jones, Jonathan O’Connell, Bruce Follows, Steven Mischke, Justin A. Caravella, Stephanos Ioannidis, Christopher Dinsmore, Sunkyu Kim, Axel Behrens, David Komander, Benedikt M. Kessler, Sylvie Urbé, and Michael J. Clague. The deubiquitylase USP9X controls ribosomal stalling. Journal of Cell Biology, 220(3):e202004211, March 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:/

/rupress.org/jcb/article/220/3/e202004211/211735/The-
deubiquitylase-USP9X-controls-ribosomal. See correction

[CHPF^{+}21a].

Cho:2022:TSE

[CHS^{+}22] Yuma Cho, Daichi Haraguchi, Kenta Shigetomi, Kenji Matsuzawa, Seiichi Uchida, and Junichi Ikenouchi. Tricellulin secures the epithelial barrier at tricellular junctions by interacting with actomyosin. Journal of Cell Biology, 221(4):e202009037, April 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/4/
e202009037/213005/Tricellulin-secures-the-epithelial-
barrier-at.

Chiusa:2020:ERM

[CHZ^{+}20] Manuel Chiusa, Wen Hu, Jozef Zienkiewicz, Xiwu Chen, Ming-Zhi Zhang, Raymond C. Harris, Roberto M. Vanacore, Jennifer A. Bentz, Giuseppe Remuzzi, Ariela Benigni, Agnes B. Fogo, Wentian Luo, Stavroula Mili, Matthew H. Wilson, Roy Zent, Jacek Hawiger, and Ambra Pozzi. EGF receptor– mediated FUS phosphorylation promotes its nuclear translocation and fibrotic signaling. Journal of Cell Biology, 219(9): e202001120, September 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/9/e202001120/151955/EGF-receptor-
mediated-FUS-phosphorylation-promotes.

Casler:2022:CAM

[CJK^{+}22] Jason C. Casler, Natalie Johnson, Adam H. Krahn, Areti Pantazopoulou, Kasey J. Day, and Benjamin S. Glick. Clathrin adaptors mediate two sequential pathways of intra-Golgi recycling. Journal of Cell Biology, 221(1):e202103199, January 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/1/
e202103199/212747/Clathrin-adaptors-mediate-two-sequential-
pathways.

Cowell:2021:TRD

[CJS^{+}21] Alana R. Cowell, Guillaume Jacquemet, Abhimanyu K. Singh, Lorena Varela, Anna S. Nylund, York-Christoph Ammon, David G. Brown, Anna Akhmanova, Johanna Ivaska, and Benjamin T. Goult. Talin rod domain–containing protein 1 (TL-NRD1) is a novel actin-bundling protein which promotes filopodia formation. Journal of Cell Biology, 220(9):e202005214, September 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/220/9/e202005214/212472/Talin-rod-domain-
containing-protein-1-TLNRD1-is-a.

Collins:2020:SMP

[CKM^{+}20] Kerrie B. Collins, Hojin Kang, Jacob Matsche, Jennifer E. Klomp, Jalees Rehman, Asrar B. Malik, and Andrei V. Karginov. Septin2 mediates podosome maturation and endothelial cell invasion associated with angiogenesis. Journal of Cell Biology, 219(2):e201903023, February 3, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/2/
e201903023/133523/Septin2-mediates-podosome-maturation-
and.

Chan:2020:CCE

[CKR^{+}20] Isaac S. Chan, Hildur Knútsdóttir, Gayathri Ramakrishnan, Veena Padmanaban, Manisha Warrier, Juan Carlos Ramirez, Matthew Dunworth, Hao Zhang, Elizabeth M. Jaffee, Joel S. Bader, and Andrew Josef Ewald. Cancer cells educate natural killer cells to a metastasis-promoting cell state. Journal of Cell Biology, 219(9):e202001134, September 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/9/e202001134/151934/Cancer-cells-educate-natural-
killer-cells-to-a.

Chhetri:2022:IXR

[CKW^{+}22] Gaurav Chhetri, Yuting Ke, Ping Wang, Muhammad Usman, Yan Li, Ellen Sapp, Jing Wang, Arabinda Ghosh, Md Ariful Islam, Xiaolong Wang, Adel Boudi, Marian Di-Figlia, and Xueyi Li. Impaired XK recycling for importing manganese underlies striatal vulnerability in Huntington’s disease. Journal of Cell Biology, 221(10):e202112073, October 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/221/10/e202112073/213461/Impaired-XK-
recycling-for-importing-manganese.

Casler:2021:CBB

[CL21] Jason C. Casler and Laura L. Lackner. Cnm1: a bridge between mitochondria and nuclear ER. Journal of Cell Biology, 220(11): e202109021, November 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/11/e202109021/212718/Cnm1-A-bridge-
between-mitochondria-and-nuclear.

Chio:2021:SCG

[CLC^{+}21] Un Seng Chio, Yumeng Liu, SangYoon Chung, Woo Jun Shim, Sowmya Chandrasekar, Shimon Weiss, and Shu ou Shan. Subunit cooperation in the Get1/2 receptor promotes tail-anchored membrane protein insertion. Journal of Cell Biology, 220(11): e202103079, November 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/11/e202103079/212681/Subunit-cooperation-
in-the-Get1-2-receptor.

Chastney:2020:TFI

[CLH^{+}20] Megan R. Chastney, Craig Lawless, Jonathan D. Humphries, Stacey Warwood, Matthew C. Jones, David Knight, Claus Jorgensen, and Martin J. Humphries. Topological features of integrin adhesion complexes revealed by multiplexed proximity biotinylation. Journal of Cell Biology, 219(8):e202003038, August 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/
8/e202003038/151896/Topological-features-of-integrin-
adhesion.

Chang:2021:ATI

[CLH21] Yeonji Chang, Gyubum Lim, and Won-Ki Huh. Analysis of the TORC1 interactome reveals a spatially distinct function of TORC1 in mRNP complexes. Journal of Cell Biology, 220 (4):??, April 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/220/4/e201912060/211781/Analysis-of-the-
TORC1-interactome-reveals-a.

Chang:2021:SMM

[CLL^{+}21a] Denis Chang, Phi Luong, Qian Li, Jamie LeBarron, Michael Anderson, Lee Barrett, and Wayne I. Lencer. Small-molecule modulators of INAVA cytosolic condensate and cell–cell junction assemblies. Journal of Cell Biology, 220(9):e202007177, September 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/9/
e202007177/212462/Small-molecule-modulators-of-INAVA-
cytosolic.

Chen:2021:AFM

[CLL^{+}21b] Yanfang Chen, Romane Leboutet, Céline Largeau, Siham Zentout, Christophe Lefebvre, Agnès Delahodde, Emmanuel Culetto, and Renaud Legouis. Autophagy facilitates mitochondrial rebuilding after acute heat stress via a DRP-1-dependent process. Journal of Cell Biology, 220(4):??, April 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/4/
e201909139/211895/Autophagy-facilitates-mitochondrial-
rebuilding.

Colin:2020:ADO

[CLR^{+}20] Alexandra Colin, Gaëlle Letort, Nitzan Razin, Maria Almonacid, Wylie Ahmed, Timo Betz, Marie-Emilie Terret, Nir S. Gov, Raphaël Voituriez, Zoher Gueroui, and Marie-Hélène Verlhac. Active diffusion in oocytes nonspecifically centers large objects during prophase I and meiosis I. Journal of Cell Biology, 219(3): e201908195, March 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/219/3/e201908195/133616/Active-diffusion-
in-oocytes-nonspecifically.

Ci:2020:ZNI

[CLZ^{+}20] Yali Ci, Zhong-Yu Liu, Na-Na Zhang, Yuqiang Niu, Yang Yang, Caimin Xu, Wei Yang, Cheng-Feng Qin, and Lei Shi. Zika NS1-induced ER remodeling is essential for viral replication. Journal of Cell Biology, 219(2):e201903062, February 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/2/
e201903062/133534/Zika-NS1-induced-ER-remodeling-is-
essential-for.

Culver:2021:DUP

[CM21] Jacob A. Culver and Malaiyalam Mariappan. Deubiquitinases USP20/33 promote the biogenesis of tail-anchored membrane proteins. Journal of Cell Biology, 220(5):e202004086, May 3, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
5/e202004086/211933/Deubiquitinases-USP20-33-promote-
the-biogenesis-of.

Cairo:2020:DRB

[CML20] Gisela Cairo, Anne M. MacKenzie, and Soni Lacefield. Differential requirement for Bub1 and Bub3 in regulation of meiotic versus mitotic chromosome segregation. Journal of Cell Biology, 219(4):e201909136, April 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/4/e201909136/133770/Differential-
requirement-for-Bub1-and-Bub3-in.

Chen:2020:WCT

[CMM^{+}20] Zhiming Chen, Rosa E. Mino, Marcel Mettlen, Peter Michaely, Madhura Bhave, Dana Kim Reed, and Sandra L. Schmid.

Wbox2: a clathrin terminal domain–derived peptide inhibitor of clathrin-mediated endocytosis. Journal of Cell Biology, 219(9): e201908189, September 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/9/e201908189/151850/Wbox2-A-clathrin-
terminal-domain-derived-peptide.

Crosby:2022:RHA

[CMN^{+}22] Daniel Crosby, Melissa R. Mikolaj, Sarah B. Nyenhuis, Samantha Bryce, Jenny E. Hinshaw, and Tina H. Lee. Reconstitution of human atlastin fusion activity reveals autoinhibition by the C terminus. Journal of Cell Biology, 221(2):e202107070, February 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/
2/e202107070/212879/Reconstitution-of-human-atlastin-
fusion-activity.

Chandra:2021:ASC

[CMT^{+}21] Sunandini Chandra, Philip J. Mannino, David J. Thaller, Nicholas R. Ader, Megan C. King, Thomas J. Melia, and C. Patrick Lusk. Atg39 selectively captures inner nuclear membrane into lumenal vesicles for delivery to the autophagosome. Journal of Cell Biology, 220(12):e202103030, December 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/12/e202103030/212738/Atg39-selectively-captures-
inner-nuclear-membrane.

Carosi:2021:APW

[CNL^{+}21] Julian M. Carosi, Thanh N. Nguyen, Michael Lazarou, Sharad Kumar, and Timothy J. Sargeant. ATG8ylation of proteins: a way to cope with cell stress? Journal of Cell Biology, 220(11): e202108120, November 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/11/e202108120/212716/ATG8ylation-
of-proteins-A-way-to-cope-with-cell.

Coller:2022:CSY

[Col22a] Hilary A. Coller. Correction: Stressed-out yeast do not pass GO. Journal of Cell Biology, 221(2):e202109111, February 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/

2/jcb.20211103201052022c/212971/Correction-Stressed-
out-yeast-do-not-pass-GO.

Coller:2022:SYD

[Col22b] Hilary A. Coller. Stressed-out yeast do not pass GO. Journal of Cell Biology, 221(1):e202111032, January 3, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/1/
e202111032/212929/Stressed-out-yeast-do-not-pass-GOStressed-
out.

Celetti:2020:LSF

[CPC^{+}20] Giorgia Celetti, Giulia Paci, Joana Caria, Virginia VanDelinder, George Bachand, and Edward A. Lemke. The liquid state of FGnucleoporins mimics permeability barrier properties of nuclear pore complexes. Journal of Cell Biology, 219(1):e201907157, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Cvetkovic:2022:AGG

[CPS^{+}22] Caroline Cvetkovic, Rajan Patel, Arya Shetty, Matthew K. Hogan, Morgan Anderson, Nupur Basu, Samira Aghlara-Fotovat, Srivathsan Ramesh, Debosmita Sardar, Omid Veiseh, Michael E. Ward, Benjamin Deneen, Philip J. Horner, and Robert Krencik. Assessing Gq-GPCR–induced human astrocyte reactivity using bioengineered neural organoids. Journal of Cell Biology, 221(4):e202107135, April 4, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/4/
e202107135/212997/Assessing-Gq-GPCR-induced-human-astrocyte.

Chen:2023:MDK

[CPW^{+}23] Xiuzhen Chen, Didier Portran, Lukas A. Widmer, Marcel M. Stangier, Mateusz P. Czub, Dimitris Liakopoulos, Jörg Stelling, Michel O. Steinmetz, and Yves Barral. The motor domain of the kinesin Kip2 promotes microtubule polymerization at microtubule tips. Journal of Cell Biology, 222(7):e202110126, July 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/
7/e202110126/214052/The-motor-domain-of-the-kinesin-
Kip2-promotes.

Chen:2021:TPK

[CRZ^{+}21] Geng-Yuan Chen, Fioranna Renda, Huaiying Zhang, Alper Gokden, Daniel Z. Wu, David M. Chenoweth, Alexey Khodjakov, and Michael A. Lampson. Tension promotes kinetochore– microtubule release by Aurora B kinase. Journal of Cell Biology, 220(6):e202007030, June 7, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/6/e202007030/212027/Tension-promotes-
kinetochore-microtubule-release.

Chen:2020:EMA

[CS20] Zhiming Chen and Sandra L. Schmid. Evolving models for assembling and shaping clathrin-coated pits. Journal of Cell Biology, 219(9):e202005126, September 7, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/9/
e202005126/152007/Evolving-models-for-assembling-and-
shaping.

Casadio:2021:IPP

[CS21a] Melina Casadio and Dan Simon. Interviewing for a PI position — the pandemic way. Journal of Cell Biology, 220(4):??, April 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/4/
e202102149/211902/Interviewing-for-a-PI-position-the-
pandemic.

Casadio:2021:MRL

[CS21b] Melina Casadio and Dan Simon. Moving a research lab during the COVID-19 pandemic. Journal of Cell Biology, 220(1): e202012016, January 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/220/1/e202012016/211608/Moving-a-research-
lab-during-the-COVID-19.

Casadio:2021:NIT

[CS21c] Melina Casadio and Dan Simon. New investigators in the time of the COVID-19 pandemic. Journal of Cell Biology, 220(2):e202012131, February 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:/

/rupress.org/jcb/article/220/2/e202012131/211703/New-
investigators-in-the-time-of-the-COVID-19.

Casadio:2021:SLD

[CS21d] Melina Casadio and Dan Simon. Starting a lab during the COVID-19 pandemic. Journal of Cell Biology, 220(3): e202101156, March 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/220/3/e202101156/211819/Starting-a-lab-during-
the-COVID-19.

Cail:2022:INM

[CSD22] Robert C. Cail, Cyna R. Shirazinejad, and David G. Drubin. Induced nanoscale membrane curvature bypasses the essential endocytic function of clathrin. Journal of Cell Biology, 221(7): e202109013, July 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/221/7/e202109013/213203/Induced-nanoscale-
membrane-curvature-bypasses-the.

Calvi:2022:PPC

[CSG22] Ida Calvi, Françoise Schwager, and Monica Gotta. PP1 phosphatases control PAR-2 localization and polarity establishment in C. elegans embryos. Journal of Cell Biology, 221 (10):e202201048, October 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/221/10/e202201048/213453/PP1-
phosphatases-control-PAR-2-localization-and.

Chandra:2021:ERM

[CSM^{+}21] Goutam Chandra, Sen Chandra Sreetama, Davi A. G. Mázala, Karine Charton, Jack H. VanderMeulen, Isabelle Richard, and Jyoti K. Jaiswal. Endoplasmic reticulum maintains ion homeostasis required for plasma membrane repair. Journal of Cell Biology, 220(5):e202006035, May 3, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
5/e202006035/211873/Endoplasmic-reticulum-maintains-
ion-homeostasis.

Cunha-Silva:2020:MMR

[CSOG^{+}20] Sofia Cunha-Silva, Mariana Osswald, Jana Goemann, João Barbosa, Luis M. Santos, Pedro Resende, Tanja Bange, Cristina Ferrás, Claudio E. Sunkel, and Carlos Conde. Mps1mediated release of Mad1 from nuclear pores ensures the fidelity of chromosome segregation. Journal of Cell Biology, 219(3):e201906039, March 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/219/3/e201906039/133569/Mps1-
mediated-release-of-Mad1-from-nuclear-pores.

Cordeiro:2020:KPS

[CSS20] Marilia H. Cordeiro, Richard J. Smith, and Adrian T. Saurin. Kinetochore phosphatases suppress autonomous Polo-like kinase 1 activity to control the mitotic checkpoint. Journal of Cell Biology, 219(12):e202002020, December 7, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/12/
e202002020/211513/Kinetochore-phosphatases-suppress-
autonomous-Polo.

Chorlay:2020:NLR

[CT20] Aymeric Chorlay and Abdou Rachid Thiam. Neutral lipids regulate amphipathic helix affinity for model lipid droplets. Journal of Cell Biology, 219(4):e201907099, April 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/4/
e201907099/133864/Neutral-lipids-regulate-amphipathic-
helix-affinity.

Cannon:2023:GDS

[CVMB^{+}23] Kevin S. Cannon, Jose M. Vargas-Muniz, Neil Billington, Ian Seim, Joanne Ekena, James R. Sellers, and Amy. S. Gladfelter. A gene duplication of a septin reveals a developmentally regulated filament length control mechanism. Journal of Cell Biology, 222(3):e202204063, March 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/222/3/e202204063/213866/A-
gene-duplication-of-a-septin-reveals-a.

Coquand:2021:COA

[CVT^{+}21] Laure Coquand, Guiliana Soraya Victoria, Alice Tata, Jacopo Amerigo Carpentieri, Jean-Baptiste Brault, Fabien Guimiot, Vincent Fraisier, and Alexandre D. Baffet. CAM-SAPs organize an acentrosomal microtubule network from basal varicosities in radial glial cells. Journal of Cell Biology, 220(8): e202003151, August 2, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/220/8/e202003151/212175/CAMSAPs-organize-
an-acentrosomal-microtubule.

Clemente:2023:PCN

[CW23] Giuliana D. Clemente and Helen Weavers. A PI3K-calcium-Nox axis primes leukocyte Nrf2 to boost immune resilience and limit collateral damage. Journal of Cell Biology, 222 (6):e202203062, June 5, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/222/6/e202203062/213986/A-
PI3K-calcium-Nox-axis-primes-leukocyte-Nrf2-to.

Chen:2020:RPI

[CWAT20] Yu-Jie Chen, Jeffrey M. Williams, Peter Arvan, and Billy Tsai. Reticulon protects the integrity of the ER membrane during ER escape of large macromolecular protein complexes. Journal of Cell Biology, 219(2):e201908182, February 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/
2/e201908182/133556/Reticulon-protects-the-integrity-
of-the-ER.

Chen:2023:VAM

[CWKP23] Hao Chen, Zara Y. Weinberg, G. Aditya Kumar, and Manojkumar A. Puthenveedu. Vesicle-associated membrane protein 2 is a cargo-selective v-SNARE for a subset of GPCRs. Journal of Cell Biology, 222(7):e202207070, July 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
222/7/e202207070/214003/Vesicle-associated-membrane-
protein-2-is-a-cargo.

Cieslinski:2023:NSO

[CWN^{+}23] Konstanty Cieslinski, Yu-Le Wu, Lisa Nechyporenko, Sarah Janice Hörner, Duccio Conti, Michal Skruzny, and Jonas Ries. Nanoscale structural organization and stoichiometry of the budding yeast kinetochore. Journal of Cell Biology, 222(4): e202209094, April 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/222/4/e202209094/213833/Nanoscale-structural-
organization-and.

Chi:2021:PPN

[CWX^{+}21] Wangfei Chi, Gang Wang, Guangwei Xin, Qing Jiang, and Chuanmao Zhang. PLK4-phosphorylated NEDD1 facilitates cartwheel assembly and centriole biogenesis initiations. Journal of Cell Biology, 220(1):e202002151, January 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/220/1/e202002151/211633/PLK4-phosphorylated-
NEDD1-facilitates-cartwheel.

Chen:2020:MDS

[CWZ^{+}20] Zhe Chen, Zong-Heng Wang, Guofeng Zhang, Christopher K. E. Bleck, Dillon J. Chung, Grey P. Madison, Eric Lindberg, Christian Combs, Robert S. Balaban, and Hong Xu. Mitochondrial DNA segregation and replication restrict the transmission of detrimental mutation. Journal of Cell Biology, 219(7):e201905160, July 6, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/7/
e201905160/151740/Mitochondrial-DNA-segregation-and-
replication.

Chinen:2021:CPC

[CYH^{+}21] Takumi Chinen, Kaho Yamazaki, Kaho Hashimoto, Ken Fujii, Koki Watanabe, Yutaka Takeda, Shohei Yamamoto, Yuka Nozaki, Yuki Tsuchiya, Daisuke Takao, and Daiju Kitagawa. Centriole and PCM cooperatively recruit CEP192 to spindle poles to promote bipolar spindle assembly. Journal of Cell Biology, 220(2):e202006085, February 1, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/2/
e202006085/211683/Centriole-and-PCM-cooperatively-recruit-
CEP192-to.

Chen:2020:GAB

[CYL^{+}20] Keyu Chen, Rui Yang, Yubing Li, Jin Chuan Zhou, and Mingjie Zhang. Giant ankyrin-B suppresses stochastic collateral axon branching through direct interaction with microtubules. Journal of Cell Biology, 219(8):e201910053, August 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/219/8/e201910053/151929/Giant-ankyrin-B-
suppresses-stochastic-collateral.

Chen:2021:FOL

[CYR^{+}21] Fang Chen, Bing Yan, Jie Ren, Rui Lyu, Yanfang Wu, Yuting Guo, Dong Li, Hong Zhang, and Junjie Hu. FIT2 organizes lipid droplet biogenesis with ER tubule-forming proteins and septins. Journal of Cell Biology, 220(5):e201907183, May 3, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/5/e201907183/211999/FIT2-organizes-lipid-droplet-
biogenesis-with-ER.

Chen:2021:PST

[CYU^{+}21] Chen Chen, Yoshiaki Yamanaka, Koji Ueda, Peiying Li, Tamami Miyagi, Yuichiro Harada, Sayaka Tezuka, Satoshi Narumi, Masahiro Sugimoto, Masahiko Kuroda, Yuhei Hayamizu, and Kohsuke Kanekura. Phase separation and toxicity of C9orf72 poly(PR) depends on alternate distribution of arginine. Journal of Cell Biology, 220(11):e202103160, November 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/11/
e202103160/212626/Phase-separation-and-toxicity-of-C9orf72-
poly-PR.

Chen:2021:CTM

[CZTL21] Yujue Chen, Qian Zhang, Zhen Teng, and Hong Liu. Centromeric transcription maintains centromeric cohesion in human cells. Journal of Cell Biology, 220(7):e202008146, July 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/7/
e202008146/212009/Centromeric-transcription-maintains-
centromeric.

dAmico:2022:CTS

[dAC^{+}22] Ennio A. d’Amico, Misbha Ud Din Ahmad, Verena Cmentowski, Mathias Girbig, Franziska Müller, Sabine Wohlgemuth, Andreas Brockmeyer, Stefano Maffini, Petra Janning, Ingrid R. Vetter, Andrew P. Carter, Anastassis Perrakis, and Andrea Musacchio. Conformational transitions of the spindly adaptor underlie its interaction with Dynein and Dynactin. Journal of Cell Biology, 221(11):e202206131, November 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/11/
e202206131/213466/Conformational-transitions-of-the-
Spindly-adaptor.

Deliz-Aguirre:2021:MOS

[DACG^{+}21] Rafael Deliz-Aguirre, Fakun Cao, Fenja H. U. Gerpott, Nichanok Auevechanichkul, Mariam Chupanova, YeVin Mun, Elke Ziska, and Marcus J. Taylor. MyD88 oligomer size functions as a physical threshold to trigger IL1R myddosome signaling. Journal of Cell Biology, 220(7):e202012071, July 5, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/7/
e202012071/212080/MyD88-oligomer-size-functions-as-a-
physical.

Du:2023:TRL

[DCG^{+}23] Yuanjiao Du, Weiping Chang, Lei Gao, Lin Deng, and Wei-Ke Ji. Tex2 is required for lysosomal functions at TMEM55-dependent ER membrane contact sites. Journal of Cell Biology, 222(4): e202205133, April 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/222/4/e202205133/213838/Tex2-is-required-
for-lysosomal-functions-at-TMEM55.

Dattoli:2020:AAC

[DCK^{+}20] Anna Ada Dattoli, Ben L. Carty, Antje M. Kochendoerfer, Conall Morgan, Annie E. Walshe, and Elaine M. Dunleavy. Asymmetric assembly of centromeres epigenetically regulates stem cell fate. Journal of Cell Biology, 219(4):e201910084, April 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/4/
e201910084/133868/Asymmetric-assembly-of-centromeres-
epigenetically.

De-Castro:2022:WMD

[DCRDC^{+}22] Ana R. G. De-Castro, Diogo R. M. Rodrigues, Maria J. G. De-Castro, Neide Vieira, Cármen Vieira, Ana X. Carvalho, Reto Gassmann, Carla M. C. Abreu, and Tiago J. Dantas. WDR60-mediated dynein-2 loading into cilia powers retrograde IFT and transition zone crossing. Journal of Cell Biology, 221(1):e202010178, January 3, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/1/
e202010178/212746/WDR60-mediated-dynein-2-loading-into-
cilia-powers.

Dejardin:2020:NMD

[DCS^{+}20] Théophile Déjardin, Pietro Salvatore Carollo, François Sipieter, Patricia M. Davidson, Cynthia Seiler, Damien Cuvelier, Bruno Cadot, Cecile Sykes, Edgar R. Gomes, and Nicolas Borghi. Nesprins are mechanotransducers that discriminate epithelial–mesenchymal transition programs. Journal of Cell Biology, 219(10):e201908036, October 5, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/10/
e201908036/152020/Nesprins-are-mechanotransducers-that-
discriminate.

dosSantosPassos:2021:DGE

[dCS^{+}21] Carolina dos Santos Passos, Yun-Seok Choi, Christopher D. Snow, Tingting Yao, and Robert E. Cohen. Design of genetically encoded sensors to detect nucleosome ubiquitination in live cells. Journal of Cell Biology, 220(4):??, April 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/4/
e201911130/211785/Design-of-genetically-encoded-sensors-
to-detect.

deCastro:2020:MLO

[dCTOG^{+}20] Isabel Fernández de Castro, Raquel Tenorio, Paula Ortega-González, Jonathan J. Knowlton, Paula F. Zamora, Christopher H. Lee, José J. Fernández, Terence S. Dermody, and Cristina Risco. A modified lysosomal organelle mediates nonlytic egress of reovirus. Journal of Cell Biology, 219(7):e201910131, July 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/219/7/e201910131/151728/A-modified-lysosomal-
organelle-mediates-nonlytic.

Deshpande:2022:AMC

[DdCVT22] Ojas Deshpande, Jorge de Carvalho, Diana V. Vieira, and Ivo A. Telley. Astral microtubule cross-linking safeguards uniform nuclear distribution in the Drosophila syncytium. Journal of Cell Biology, 221(1):e202007209, January 3, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/1/
e202007209/212810/Astral-microtubule-cross-linking-safeguards.

DeJamblinne:2020:SRS

[DDD^{+}20] Camille Valérie De Jamblinne, Barbara Decelle, Mehrnoush Dehghani, Mathieu Joseph, Neera Sriskandarajah, Kévin Leguay, Basile Rambaud, Sébastien Lemieux, Philippe P. Roux, David R.

Hipfner, and Sébastien Carréno. STRIPAK regulates Slik localization to control mitotic morphogenesis and epithelial integrity. Journal of Cell Biology, 219(11):e201911035, November 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/11/
e201911035/152107/STRIPAK-regulates-Slik-localization-
to-control.

delDedo:2021:CSS

[dDFGP^{+}21] Javier Encinar del Dedo, Isabel María Fernández-Golbano, Laura Pastor, Paula Meler, Cristina Ferrer-Orta, Elena Rebollo, and Maria Isabel Geli. Coupled sterol synthesis and transport machineries at ER–endocytic contact sites. Journal of Cell Biology, 220(10):e202010016, October 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/10/
e202010016/212484/Coupled-sterol-synthesis-and-transport-
machineries.

Dang:2023:DLT

[DES^{+}23] David Dang, Christoforos Efstathiou, Dijue Sun, Haoran Yue, Nishanth R. Sastry, and Viji M. Draviam. Deep learning techniques and mathematical modeling allow 3D analysis of mitotic spindle dynamics. Journal of Cell Biology, 222(5):e202111094, May 1, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
222/5/e202111094/213913/Deep-learning-techniques-and-
mathematical-modeling.

Devireddy:2022:EPE

[DF22] Swathi Devireddy and Shawn M. Ferguson. Efficient progranulin exit from the ER requires its interaction with prosaposin, a Surf4 cargo. Journal of Cell Biology, 221(2):e202104044, February 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/
2/e202104044/212919/Efficient-progranulin-exit-from-
the-ER-requires.

Donker:2022:TCC

[DG22] Lisa Donker and Susana A. Godinho. Too close for comfort? Endomembranes promote missegregation by enclosing lost chromosomes. Journal of Cell Biology, 221(6): e202204114, June 6, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/221/6/e202204114/213206/Too-
close-for-comfort-Endomembranes-promote.

Dong:2021:BRN

[DHB^{+}21] Xingchen Dong, Xiangming Hu, Yan Bao, Guo Li, Xiao dong Yang, James M. Slauch, and Lin-Feng Chen. Brd4 regulates NLRC4 inflammasome activation by facilitating IRF8mediated transcription of Naips. Journal of Cell Biology, 220 (3):e202005148, March 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/220/3/e202005148/211747/Brd4-
regulates-NLRC4-inflammasome-activation-by.

Duncan:2022:URC

[DHTP22] Emily D. Duncan, Ke-Jun Han, Margaret A. Trout, and Rytis Prekeris. Ubiquitylation by Rab40b/Cul5 regulates Rap2 localization and activity during cell migration. Journal of Cell Biology, 221(4):e202107114, April 4, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/4/
e202107114/213068/Ubiquitylation-by-Rab40b-Cul5-regulates-
Rap2.

Dobramysl:2021:SCA

[DJI^{+}21] Ulrich Dobramysl, Iris Katharina Jarsch, Yoshiko Inoue, Hanae Shimo, Benjamin Richier, Jonathan R. Gadsby, Julia Mason, Alicja Szałapak, Pantelis Savvas Ioannou, Guilherme Pereira Correia, Astrid Walrant, Richard Butler, Edouard Hannezo, Benjamin D. Simons, and Jennifer L. Gallop. Stochastic combinations of actin regulatory proteins are sufficient to drive filopodia formation. Journal of Cell Biology, 220(4):??, April 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
4/e202003052/211897/Stochastic-combinations-of-actin-
regulatory.

Doodhi:2021:ABS

[DKCT21] Harinath Doodhi, Taciana Kasciukovic, Lesley Clayton, and Tomoyuki U. Tanaka. Aurora B switches relative strength of kinetochore–microtubule attachment modes for error correction. Journal of Cell Biology, 220(6):e202011117, June 7, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/6/
e202011117/211981/Aurora-B-switches-relative-strength-
of-kinetochore.

Dunkler:2021:TVM

[DLK^{+}21] Alexander Dünkler, Marcin Leda, Jan-Michael Kromer, Joachim Neller, Thomas Gronemeyer, Andrew B. Goryachev, and Nils Johnsson. Type V myosin focuses the polarisome and shapes the tip of yeast cells. Journal of Cell Biology, 220 (5):e202006193, May 3, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/220/5/e202006193/211845/Type-
V-myosin-focuses-the-polarisome-and-shapes.

Dong:2020:PDA

[DLZ^{+}20] Wei Dong, Juan Lu, Xuejing Zhang, Yan Wu, Kaela Lettieri, Gerald R. Hammond, and Yang Hong. A polybasic domain in aPKC mediates Par6-dependent control of membrane targeting and kinase activity. Journal of Cell Biology, 219(7): e201903031, July 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/7/e201903031/151883/A-polybasic-
domain-in-aPKC-mediates-Par6-dependent.

Duan:2020:ELR

[DMR^{+}20] Haohui Duan, Sarah Mansour, Rachel Reed, Margaret K. Gillis, Benjamin Parent, Ben Liu, Zsofia Sztupinszki, Nicolai Birkbak, Zoltan Szallasi, Andrew E. H. Elia, Judy E. Garber, and Shailja Pathania. E3 ligase RFWD3 is a novel modulator of stalled fork stability in BRCA2-deficient cells. Journal of Cell Biology, 219(6):e201908192, June 1, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/219/6/e201908192/151752/E3-
ligase-RFWD3-is-a-novel-modulator-of-stalled.

Davidson:2023:CPP

[DNVP23] Kerri A. Davidson, Mitsutoshi Nakamura, Jeffrey M. Verboon, and Susan M. Parkhurst. Centralspindlin proteins Pavarotti and Tumbleweed along with WASH regulate nuclear envelope budding. Journal of Cell Biology, 222(8):e202211074, August 7, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/8/

e202211074/214102/Centralspindlin-proteins-Pavarotti-
and-Tumbleweed.

Dantas:2022:NTC

[DOA^{+}22] Margarida Dantas, Andreia Oliveira, Paulo Aguiar, Helder Maiato, and Jorge G. Ferreira. Nuclear tension controls mitotic entry by regulating cyclin B1 nuclear translocation. Journal of Cell Biology, 221(12):e202205051, December 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/221/12/e202205051/213539/Nuclear-tension-
controls-mitotic-entry-by.

Dormann:2020:FNC

[Dor20] Dorothee Dormann. FG-nucleoporins caught in the act of liquid liquid phase separation. Journal of Cell Biology, 219(1): e201910211, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Danlasky:2020:EAP

[DPM^{+}20] Brennan M. Danlasky, Michelle T. Panzica, Karen P. McNally, Elizabeth Vargas, Cynthia Bailey, Wenzhe Li, Ting Gong, Elizabeth S. Fishman, Xueer Jiang, and Francis J. McNally. Evidence for anaphase pulling forces during C. elegans meiosis. Journal of Cell Biology, 219(12):e202005179, December 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/12/
e202005179/211469/Evidence-for-anaphase-pulling-forces-
during-C.

Dziengelewski:2020:APE

[DRC^{+}20] Claire Dziengelewski, Marc-Antoine Rodrigue, Alexia Caillier, Kévin Jacquet, Marie-Chloé Boulanger, Jonathan Bergeman, Margit Fuchs, Herman Lambert, Patrick Laprise, Darren E. Richard, François Bordeleau, Marc-Étienne Huot, and Josée N. Lavoie. Adenoviral protein E4orf4 interacts with the polarity protein Par3 to induce nuclear rupture and tumor cell death. Journal of Cell Biology, 219(4):e201805122, April 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/4/
e201805122/151580/Adenoviral-protein-E4orf4-interacts-
with-the.

Drin:2020:MPI

[Dri20] Guillaume Drin. MapPIng PI inside cells brings new light to polyphosphoinositide biology. Journal of Cell Biology, 219(3): e202001185, March 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/219/3/e202001185/133823/MapPIng-PI-inside-
cells-brings-new-light-to.

Djannatian:2023:MGA

[DRW^{+}23] Minou Djannatian, Swathi Radha, Ulrich Weikert, Shima Safaiyan, Christoph Wrede, Cassandra Deichsel, Georg Kislinger, Agata Rhomberg, Torben Ruhwedel, Douglas S. Campbell, Tjakko van Ham, Bettina Schmid, Jan Hegermann, Wiebke Möbius, Martina Schifferer, and Mikael Simons. Myelination generates aberrant ultrastructure that is resolved by microglia. Journal of Cell Biology, 222(3):e202204010, March 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/3/
e202204010/213804/Myelination-generates-aberrant-ultrastructure-
that.

Dong:2023:MMH

[DRZ^{+}23] Lan-Feng Dong, Jakub Rohlena, Renata Zobalova, Zuzana Nahacka, Anne-Marie Rodriguez, Michael V. Berridge, and Jiri Neuzil. Mitochondria on the move: Horizontal mitochondrial transfer in disease and health. Journal of Cell Biology, 222(3): e202211044, March 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/222/3/e202211044/213873/Mitochondria-
on-the-move-Horizontal-mitochondrial.

Dua:2022:DMM

[DSB22] Nitish Dua, Akshaya Seshadri, and Anjana Badrinarayanan. DarT-mediated mtDNA damage induces dynamic reorganization and selective segregation of mitochondria. Journal of Cell Biology, 221(10):e202205104, October 3, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/10/
e202205104/213451/DarT-mediated-mtDNA-damage-induces-
dynamic.

Douanne:2021:TFM

[DSG21] Tiphaine Douanne, Jane C. Stinchcombe, and Gillian M. Griffiths. Teasing out function from morphology: Similarities between primary cilia and immune synapses. Journal of Cell Biology, 220(6):e202102089, June 7, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/6/
e202102089/212075/Teasing-out-function-from-morphology-
Similarities.

Demir:2023:MMS

[DSG^{+}23] Cansaran Saygili Demir, Amélie Sabine, Muyun Gong, Olivier Dormond, and Tatiana V. Petrova. Mechanosensitive mTORC1 signaling maintains lymphatic valves. Journal of Cell Biology, 222(6):e202207049, June 5, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/222/6/e202207049/214013/Mechanosensitive-
mTORC1-signaling-maintains.

Desai:2020:ULS

[DSLP20] Paurav B. Desai, Michael W. Stuck, Bo Lv, and Gregory J. Pazour. Ubiquitin links smoothened to intraflagellar transport to regulate Hedgehog signaling. Journal of Cell Biology, 219(7): e201912104, July 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/219/7/e201912104/151798/Ubiquitin-links-
smoothened-to-intraflagellar.

Dopie:2020:TSA

[DSMB20] Joseph Dopie, Michael J. Sweredoski, Annie Moradian, and Andrew S. Belmont. Tyramide signal amplification mass spectrometry (TSA-MS) ratio identifies nuclear speckle proteins. Journal of Cell Biology, 219(9):e201910207, September 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/9/
e201910207/151914/Tyramide-signal-amplification-mass-
spectrometry.

Deng:2022:CRT

[DSY^{+}22] Xue Deng, Xing Sun, Wenkai Yue, Yongjia Duan, Rirong Hu, Kai Zhang, Jiangxia Ni, Jihong Cui, Qiangqiang Wang, Yelin Chen, Ang Li, and Yanshan Fang. CHMP2B regulates TDP-43 phosphorylation and cytotoxicity independent of autophagy via CK1. Journal of Cell Biology, 221(1): e202103033, January 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/221/1/e202103033/212740/CHMP2B-regulates-
TDP-43-phosphorylation-and.

Dabrowski:2023:PPT

[DTG23] Rahel Dabrowski, Susanna Tulli, and Martin Graef. Parallel phospholipid transfer by Vps13 and Atg2 determines autophagosome biogenesis dynamics. Journal of Cell Biology, 222(7): e202211039, July 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/222/7/e202211039/214069/Parallel-phospholipid-
transfer-by-Vps13-and-Atg2.

Dustin:2021:SPH

[Dus21] Michael L. Dustin. The staying power of hematopoietic stem cells. Journal of Cell Biology, 220(11):e202109019, November 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/11/e202109019/212709/The-staying-
power-of-hematopoietic-stem-cellsThe.

Droubi:2022:IPI

[DWA^{+}22] Alaa Droubi, Connor Wallis, Karen E. Anderson, Saifur Rahman, Aloka de Sa, Taufiq Rahman, Len R. Stephens, Philip T. Hawkins, and Martin Lowe. The inositol 5-phosphatase INPP5B regulates B cell receptor clustering and signaling. Journal of Cell Biology, 221(9):e202112018, September 5, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/9/
e202112018/213365/The-inositol-5-phosphatase-INPP5B-
regulates-B-cell.

Du:2021:SRF

[DY21] Ximing Du and Hongyuan Yang. Seipin regulates the formation of nuclear lipid droplets from a distance. Journal of Cell Biology, 220(1):e202011166, January 4, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/1/
e202011166/211595/Seipin-regulates-the-formation-of-
nuclear-lipid.

Deng:2020:SCC

[DYW^{+}20] Hua Deng, Limin Yang, Pei Wen, Huiyan Lei, Paul Blount, and Duojia Pan. Spectrin couples cell shape, cortical tension, and Hippo signaling in retinal epithelial morphogenesis. Journal of Cell Biology, 219(4):e201907018, April 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/4/e201907018/133846/Spectrin-couples-cell-shape-
cortical-tension-and.

Du:2020:OLE

[DZA^{+}20] Ximing Du, Linkang Zhou, Yvette Celine Aw, Hoi Yin Mak, Yanqing Xu, James Rae, Wenmin Wang, Armella Zadoorian, Sarah E. Hancock, Brenna Osborne, Xiang Chen, Jia-Wei Wu, Nigel Turner, Robert G. Parton, Peng Li, and Hongyuan Yang. ORP5 localizes to ER-lipid droplet contacts and regulates the level of PI(4)P on lipid droplets. Journal of Cell Biology, 219 (1):e201905162, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Dai:2022:LAI

[DZA^{+}22] Jin Dai, Gui Zhang, Rama A. Alkhofash, Betlehem Mekonnen, Sahana Saravanan, Bin Xue, Zhen-Chuan Fan, Ewelina Betleja, Douglas G. Cole, Peiwei Liu, and Karl Lechtreck. Loss of ARL13 impedes BBSome-dependent cargo export from Chlamydomonas cilia. Journal of Cell Biology, 221(10):e202201050, October 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/
10/e202201050/213429/Loss-of-ARL13-impedes-BBSome-dependent-
cargo.

Eisenberg-Bord:2021:CMN

[EBZC^{+}21] Michal Eisenberg-Bord, Naama Zung, Javier Collado, Layla Drwesh, Emma J. Fenech, Amir Fadel, Nili Dezorella, Yury S. Bykov, Doron Rapaport, Ruben Fernandez-Busnadiego, and Maya Schuldiner. Cnm1 mediates nucleus–mitochondria contact site formation in response to phospholipid levels. Journal of Cell Biology, 220(11):e202104100, November 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/11/
e202104100/212719/Cnm1-mediates-nucleus-mitochondria-
contact-site.

Eskandari:2022:PRC

[EE22] Ebrahim Eskandari and Connie J. Eaves. Paradoxical roles of caspase-3 in regulating cell survival, proliferation, and tumorigenesis. Journal of Cell Biology, 221(6):e202201159, June 6, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/6/e202201159/213213/Paradoxical-roles-of-caspase-
3-in-regulating-cell.

Eising:2022:LBM

[EEW^{+}22] Sebastian Eising, Bianca Esch, Mike Wälte, Prado Vargas Duarte, Stefan Walter, Christian Ungermann, Maria Bohnert, and Florian Fröhlich. A lysosomal biogenesis map reveals the cargo spectrum of yeast vacuolar protein targeting pathways. Journal of Cell Biology, 221(4):e202107148, April 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/4/
e202107148/213011/A-lysosomal-biogenesis-map-reveals-
the-cargo.

Edwards-Jorquera:2020:TCA

[EJBB^{+}20] Sandra Sofía Edwards-Jorquera, Floris Bosveld, Yohanns A. Bellaïche, Ana-María Lennon-Duménil, and Álvaro Glavic. Trpml controls actomyosin contractility and couples migration to phagocytosis in fly macrophages. Journal of Cell Biology, 219(3):e201905228, March 2, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/3/
e201905228/133603/Trpml-controls-actomyosin-contractility-
and.

Etienne-Manneville:2020:PTV

[EM20] Sandrine Etienne-Manneville. P120catenin tuning of VEcadherin endocytosis controls collective cell behavior during angiogenesis. Journal of Cell Biology, 219(5):e202003005, May 4, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/5/
e202003005/151707/P120catenin-tuning-of-VE-cadherin-
endocytosis.

Eckel-Mahan:2022:TSM

[EM22] Kristin Eckel-Mahan. Temporal and spatial metabolite dynamics impart control in adipogenesis. Journal of Cell Biology, 221(12): e202210021, December 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/221/12/e202210021/213702/Temporal-and-
spatial-metabolite-dynamics-impart.

Ems-McClung:2020:RIE

[EMEZ^{+}20] Stephanie C. Ems-McClung, Mackenzie Emch, Stephanie Zhang, Serena Mahnoor, Lesley N. Weaver, and Claire E. Walczak. RanGTP induces an effector gradient of XCTK2 and importin α /β for spindle microtubule cross-linking. Journal of Cell Biology, 219(2):e201906045, February 3, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/2/
e201906045/133528/RanGTP-induces-an-effector-gradient-
of-XCTK2-and.

Enshoji:2022:EPM

[EMY^{+}22] Mariko Enshoji, Yoshiko Miyano, Nao Yoshida, Makoto Nagano, Minami Watanabe, Mayumi Kunihiro, Daria E. Siekhaus, Junko Y. Toshima, and Jiro Toshima. Eps15/Pan1p is a master regulator of the late stages of the endocytic pathway. Journal of Cell Biology, 221(10):e202112138, October 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/10/
e202112138/213415/Eps15-Pan1p-is-a-master-regulator-
of-the-late.

Erasmus:2021:RPR

[ESB^{+}21] Jennifer C. Erasmus, Kasia Smolarczyk, Helena Brezovjakova, Noor F. Mohd-Naim, Encarnación Lozano, Karl Matter, and Vania M. M. Braga. Rac1-PAK1 regulation of Rab11 cycling promotes junction destabilization. Journal of Cell Biology, 220(6):e202002114, June 7, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/220/6/e202002114/212034/Rac1-
PAK1-regulation-of-Rab11-cycling-promotes.

Shin:2023:DRA

[eSG23] Grace Ji eun Shin and Wesley B. Grueber. Dendrite remodeling according to GARP. Journal of Cell Biology, 222(1):e202211072, January 2, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/222/1/e202211072/213775/Dendrite-remodeling-
according-to-GARPDendrite.

Eckfeld:2023:TNL

[ESH^{+}23] Celina Eckfeld, Benjamin Schoeps, Daniel Häußler, Julian Frädrich, Felix Bayerl, Jan Philipp Böttcher, Percy Knolle, Simone Heisz, Olga Prokopchuk, Hans Hauner, Enkhtsetseg Munkhbaatar, Ihsan Ekin Demir, Chris D. Hermann, and Achim Krüger. TIMP-1 is a novel ligand of Amyloid Precursor Protein and triggers a proinflammatory phenotype in human monocytes. Journal of Cell Biology, 222(2):e202206095, February 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
222/2/e202206095/213799/TIMP-1-is-a-novel-ligand-of-
Amyloid-Precursor.

English:2020:MCP

[ESX^{+}20] Alyssa M. English, Max-Hinderk Schuler, Tianyao Xiao, Benoît Kornmann, Janet M. Shaw, and Adam L. Hughes. ER– mitochondria contacts promote mitochondrial-derived compartment biogenesis. Journal of Cell Biology, 219(12):e202002144, December 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/219/12/e202002144/211482/ER-mitochondria-
contacts-promote-mitochondrial.

Efimova:2020:BAN

[EYC^{+}20] Nadia Efimova, Changsong Yang, Jonathan X. Chia, Ning Li, Christopher J. Lengner, Kristi L. Neufeld, and Tatyana M. Svitkina. Branched actin networks are assembled on microtubules by adenomatous polyposis coli for targeted membrane protrusion. Journal of Cell Biology, 219(9):e202003091, September 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/9/e202003091/151902/Branched-actin-networks-are-
assembled-on.

Endicott:2020:ICP

[EZB^{+}20] S. Joseph Endicott, Zachary J. Ziemba, Logan J. Beckmann, Dennis N. Boynton, Jr., and Richard A. Miller. Inhibition of class I PI3K enhances chaperone-mediated autophagy. Journal of Cell Biology, 219(12):e202001031, December 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/12/e202001031/211459/Inhibition-of-class-I-PI3K-
enhances-chaperone.

Farmer:2021:XPM

[FAHZ21] Veronica Farmer, Göker Arpağ, Sarah L. Hall, and Marija Zanic. XMAP215 promotes microtubule catastrophe by disrupting the growing microtubule end. Journal of Cell Biology, 220(10):e202012144, October 4, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/10/
e202012144/212518/XMAP215-promotes-microtubule-catastrophe-
by.

Feng:2022:CSD

[FAMQW22] Wenzhi Feng, Orlando Argüello-Miranda, Suhong Qian, and Fei Wang. Cdc14 spatiotemporally dephosphorylates Atg13 to activate autophagy during meiotic divisions. Journal of Cell Biology, 221(5):e202107151, May 2, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/5/
e202107151/213046/Cdc14-spatiotemporally-dephosphorylates-
Atg13-to.

Feder:2021:SLJ

[FAS^{+}21] Zoë A. Feder, Asif Ali, Abhyudai Singh, Joanna Krakowiak, Xu Zheng, Vytas P. Bindokas, Donald Wolfgeher, Stephen J. Kron, and David Pincus. Subcellular localization of the J-protein Sis1 regulates the heat shock response. Journal of Cell Biology, 220(1):e202005165, January 4, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/1/
e202005165/211600/Subcellular-localization-of-the-J-
protein-Sis1.

Fic:2021:RIC

[FBR^{+}21] Weronika Fic, Rebecca Bastock, Francesco Raimondi, Erinn Los, Yoshiko Inoue, Jennifer L. Gallop, Robert B. Russell, and Daniel St Johnston. RhoGAP19D inhibits Cdc42 laterally to control epithelial cell shape and prevent invasion. Journal of Cell Biology, 220(4):??, April 5, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/4/
e202009116/211832/RhoGAP19D-inhibits-Cdc42-laterally-
to-control.

Faklaris:2022:QAL

[FBVD^{+}22] Orestis Faklaris, Leslie Bancel-Vallée, Aurélien Dauphin, Baptiste Monterroso, Perrine Frère, David Geny, Tudor Manoliu, Sylvain de Rossi, Fabrice P. Cordelières, Damien Schapman, Roland Nitschke, Julien Cau, and Thomas Guilbert. Quality assessment in light microscopy for routine use through simple tools and robust metrics. Journal of Cell Biology, 221(11):e202107093, November 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/221/11/e202107093/213512/Quality-assessment-
in-light-microscopy-for-routine.

Ferreira:2021:PLP

[FC21] Joana Veríssimo Ferreira and Pedro Carvalho. Pex30-like proteins function as adaptors at distinct ER membrane contact sites. Journal of Cell Biology, 220(10):e202103176, October 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
10/e202103176/212563/Pex30-like-proteins-function-as-
adaptors-at.

Foltz:2021:AET

[FCCH21] Steven J. Foltz, Yuan Yuan Cui, Hyojung J. Choo, and H. Criss Hartzell. ANO5 ensures trafficking of annexins in wounded myofibers. Journal of Cell Biology, 220(3):e202007059, March 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/3/e202007059/211717/ANO5-ensures-trafficking-of-
annexins-in-wounded.

Fracchiolla:2020:PWP

[FCHM20] Dorotea Fracchiolla, Chunmei Chang, James H. Hurley, and Sascha Martens. A PI3K-WIPI2 positive feedback loop allosterically activates LC3 lipidation in autophagy. Journal of Cell Biology, 219(7):e201912098, July 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/219/7/e201912098/151802/A-
PI3K-WIPI2-positive-feedback-loop-allosterically.

Ferranti:2020:FLP

[FCT^{+}20] Charles S. Ferranti, Jin Cheng, Chris Thompson, Jianjun Zhang, Jimmy A. Rotolo, Salma Buddaseth, Zvi Fuks, and Richard N. Kolesnick. Fusion of lysosomes to plasma membrane initiates radiation-induced apoptosis. Journal of Cell Biology, 219(4):e201903176, April 6, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/4/
e201903176/133857/Fusion-of-lysosomes-to-plasma-membrane-
initiates.

Fong:2021:MPE

[FDA21] Kimberly K. Fong, Trisha N. Davis, and Charles L. Asbury. Microtubule pivoting enables mitotic spindle assembly in S. cerevisiae. Journal of Cell Biology, 220(3):e202007193, March 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/3/e202007193/211686/Microtubule-pivoting-enables-
mitotic-spindle.

Findakly:2021:SOS

[FDG^{+}21] Sarah Findakly, Vikas Daggubati, Galo Garcia III, Sydney A. LaStella, Abrar Choudhury, Cecilia Tran, Amy Li, Pakteema Tong, Jason Q. Garcia, Natasha Puri, Jeremy F. Reiter, Libin Xu, and David R. Raleigh. Sterol and oxysterol synthases near the ciliary base activate the Hedgehog pathway. Journal of Cell Biology, 220(1):e202002026, January 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/220/1/e202002026/211576/Sterol-and-oxysterol-
synthases-near-the-ciliary.

Ferrandiz:2022:EPC

[FDSR22] Nuria Ferrandiz, Laura Downie, Georgina P. Starling, and Stephen J. Royle. Endomembranes promote chromosome missegregation by ensheathing misaligned chromosomes. Journal of Cell Biology, 221(6):e202203021, June 6, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/6/
e202203021/213189/Endomembranes-promote-chromosome-missegregation-
by.

Fulford:2023:EDB

[FER^{+}23] Alexander D. Fulford, Leonie Enderle, Jannette Rusch, Didier Hodzic, Maxine V. Holder, Alex Earl, Robin Hyunseo Oh, Nicolas Tapon, and Helen McNeill. Expanded directly binds conserved regions of fat to restrain growth via the hippo pathway. Journal of Cell Biology, 222(5):e202204059, May 1, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/5/
e202204059/214041/Expanded-directly-binds-conserved-
regions-of-Fat.

Filippone:2022:API

[FFZ^{+}22] Maria Grazia Filippone, Stefano Freddi, Silvia Zecchini, Silvia Restelli, Ivan Nicola Colaluca, Giovanni Bertalot, Salvatore Pece, Daniela Tosoni, and Pier Paolo Di Fiore. Aberrant phosphorylation inactivates numb in breast cancer causing expansion of the stem cell pool. Journal of Cell Biology, 221(12):e202112001, December 5, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/12/
e202112001/213525/Aberrant-phosphorylation-inactivates-
Numb-in.

Frazer:2021:SSC

[FGBD^{+}21] Gordon L. Frazer, Christian M. Gawden-Bone, Nele M. G. Dieckmann, Yukako Asano, and Gillian M. Griffiths. Signal strength controls the rate of polarization within CTLs during killing. Journal of Cell Biology, 220(10):e202104093, October 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/10/
e202104093/212498/Signal-strength-controls-the-rate-
of-polarization.

Fueller:2020:CCA

[FHM^{+}20] Julia Fueller, Konrad Herbst, Matthias Meurer, Krisztina Gubicza, Bahtiyar Kurtulmus, Julia D. Knopf, Daniel Kirrmaier, Benjamin C. Buchmuller, Gislene Pereira, Marius K. Lemberg, and Michael Knop. CRISPR-Cas12a-assisted PCR tagging of mammalian genes. Journal of Cell Biology, 219(6):e201910210, June 1, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/6/e201910210/151766/CRISPR-Cas12a-assisted-PCR-tagging-
of-mammalian.

Fauser:2022:DPT

[FHM^{+}22] Jordan Fauser, Vincent Huyot, Jacob Matsche, Barbara N. Szynal, Yuri Alexeev, Pradeep Kota, and Andrei V. Karginov.

Dissecting protein tyrosine phosphatase signaling by engineered chemogenetic control of its activity. Journal of Cell Biology, 221(8):e202111066, August 1, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/8/
e202111066/213352/Dissecting-protein-tyrosine-phosphatase-
signaling.

Fujito:2005:ICM

[FIK^{+}05] Tsutomu Fujito, Wataru Ikeda, Shigeki Kakunaga, Yukiko Minami, Mihoko Kajita, Yasuhisa Sakamoto, Morito Monden, and Yoshimi Takai. Inhibition of cell movement and proliferation by cell–cell contact-induced interaction of Necl-5 with nectin-3. Journal of Cell Biology, 171(1):165–??, October 2005. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/171/1/165. See correction

[FIK^{+}20].

Fujito:2020:CIC

[FIK^{+}20] Tsutomu Fujito, Wataru Ikeda, Shigeki Kakunaga, Yukiko Minami, Mihoko Kajita, Yasuhisa Sakamoto, Morito Monden, and Yoshimi Takai. Correction: Inhibition of cell movement and proliferation by cell–cell contact-induced interaction of Necl-5 with nectin-3. Journal of Cell Biology, 219(6):??, June 1, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/6/
jcb.20050109005122020c/151782/Correction-Inhibition-
of-cell-movement-and. See [FIK^{+}05].

Feng:2022:IPP

[FLJ^{+}22] Zhe Feng, Suho Lee, Bowen Jia, Tao Jian, Eunjoon Kim, and Mingjie Zhang. IRSp53 promotes postsynaptic density formation and actin filament bundling. Journal of Cell Biology, 221(8): e202105035, August 1, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/221/8/e202105035/213346/IRSp53-promotes-
postsynaptic-density-formation-and.

Fu:2023:RTI

[FLW^{+}23] Yujuan Fu, Yaxin Liu, Tanye Wen, Jie Fang, Yalong Chen, Ziying Zhou, Xinyi Gu, Hao Wu, Jinghao Sheng, Zhengping Xu, Wei Zou, and Baohui Chen. Real-time imaging of RNA polymerase I activity in living human cells. Journal of Cell Biology, 222(1):e202202110, January 2, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/222/1/e202202110/213608/Real-
time-imaging-of-RNA-polymerase-I-activity-in.

Freibaum:2021:HFR

[FMY^{+}21] Brian D. Freibaum, James Messing, Peiguo Yang, Hong Joo Kim, and J. Paul Taylor. High-fidelity reconstitution of stress granules and nucleoli in mammalian cellular lysate. Journal of Cell Biology, 220(3):e202009079, March 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/220/3/e202009079/211726/High-
fidelity-reconstitution-of-stress-granules.

Ferreira:2020:TTD

[FOR^{+}20] Luísa T. Ferreira, Bernardo Orr, Girish Rajendraprasad, António J. Pereira, Carolina Lemos, Joana T. Lima, Clàudia Guasch Boldú, Jorge G. Ferreira, Marin Barisic, and Helder Maiato. α-tubulin detyrosination impairs mitotic error correction by suppressing MCAK centromeric activity. Journal of Cell Biology, 219(4):e201910064, April 6, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/4/
e201910064/133849/Tubulin-detyrosination-impairs-mitotic-
error.

Fukuda:2021:BTS

[FPMS^{+}21] Yusuke Fukuda, Maria F. Pazyra-Murphy, Elizabeth S. Silagi, Ozge E. Tasdemir-Yilmaz, Yihang Li, Lillian Rose, Zoe C. Yeoh, Nicholas E. Vangos, Ezekiel A. Geffken, Hyuk-Soo Seo, Guillaume Adelmant, Gregory H. Bird, Loren D. Walensky, Jarrod A. Marto, Sirano Dhe-Paganon, and Rosalind A. Segal. Binding and transport of SFPQ-RNA granules by KIF5A/KLC1 motors promotes axon survival. Journal of Cell Biology, 220(1): e202005051, January 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/1/e202005051/211577/Binding-and-
transport-of-SFPQ-RNA-granules-by.

Flores:2022:TIR

[FPZ^{+}22] Rachel L. Flores, Zachary E. Peterson, Alex Zelter, Michael Riffle, Charles L. Asbury, and Trisha N. Davis. Three interacting regions of the Ndc80 and Dam1 complexes support microtubule tip-coupling under load. Journal of Cell Biology, 221(5): e202107016, May 2, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/221/5/e202107016/213102/Three-interacting-
regions-of-the-Ndc80-and-Dam1.

Farrugia:2020:CBM

[FRO^{+}20] Aaron J. Farrugia, Javier Rodríguez, Jose L. Orgaz, María Lucas, Victoria Sanz-Moreno, and Fernando Calvo. CDC42EP5/BORG3 modulates SEPT9 to promote actomyosin function, migration, and invasion. Journal of Cell Biology, 219(9):e201912159, September 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/219/9/e201912159/152021/CDC42EP5-BORG3-modulates-
SEPT9-to-promote.

Foster:2022:CSM

[FSC22] Helen E. Foster, Camilla Ventura Santos, and Andrew P. Carter. A cryo-ET survey of microtubules and intracellular compartments in mammalian axons. Journal of Cell Biology, 221(2):e202103154, February 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/221/2/e202103154/212894/A-
cryo-ET-survey-of-microtubules-and-intracellular.

Fan:2022:CMO

[FSZ^{+}22] Changyuan Fan, Xuemeng Shi, Kaikai Zhao, Linbo Wang, Kun Shi, Yan-Jun Liu, Hui Li, Baohua Ji, and Yaming Jiu. Cell migration orchestrates migrasome formation by shaping retraction fibers. Journal of Cell Biology, 221(4):e202109168, April 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/4/e202109168/213015/Cell-migration-orchestrates-
migrasome-formation-by.

Farkas:2022:RTM

[FUBS22] Ákos Farkas, Henning Urlaub, Katherine E. Bohnsack, and Blanche Schwappach. Regulated targeting of the monotopic hairpin membrane protein Erg1 requires the GET pathway. Journal of Cell Biology, 221(6):e202201036, June 6, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/221/6/e202201036/213228/Regulated-targeting-
of-the-monotopic-hairpin.

Fischer:2020:SIL

[FWP^{+}20] Tara D. Fischer, Chunxin Wang, Benjamin S. Padman, Michael Lazarou, and Richard J. Youle. STING induces LC3B lipidation onto single-membrane vesicles via the V-ATPase and ATG16L1-WD40 domain. Journal of Cell Biology, 219(12):e202009128, December 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/219/12/e202009128/211541/STING-induces-LC3B-
lipidation-onto-single-membrane.

Fingerhut:2020:MLM

[FY20] Jaclyn M. Fingerhut and Yukiko M. Yamashita. mRNA localization mediates maturation of cytoplasmic cilia in Drosophila spermatogenesis. Journal of Cell Biology, 219(9):e202003084, September 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/219/9/e202003084/151973/mRNA-localization-mediates-
maturation-of.

Galbraith:2023:PV

[Gal23] Catherine G. Galbraith. Pumping up the volume. Journal of Cell Biology, 222(2):e202212042, February 6, 2023. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/222/2/
e202212042/213825/Pumping-up-the-volumeASEM-An-automated.

Gan:2021:MRF

[Gan21] Boyi Gan. Mitochondrial regulation of ferroptosis. Journal of Cell Biology, 220(9):e202105043, September 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/9/
e202105043/212523/Mitochondrial-regulation-of-ferroptosisFerroptosis.

Gardner:2021:SRN

[Gar21] Melissa K. Gardner. Straightening up is required to nucleate new microtubules. Journal of Cell Biology, 220(4):??, April 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/4/e202102123/211886/Straightening-up-is-required-
to-nucleate-new.

Golchoubian:2022:RLR

[GBBT^{+}22] Banafsheh Golchoubian, Andreas Brunner, Helena Bragulat-Teixidor, Annett Neuner, Busra A. Akarlar, Nurhan Ozlu, and Anne-Lore Schlaitz. Reticulon-like REEP4 at the inner nuclear membrane promotes nuclear pore complex formation. Journal of Cell Biology, 221(2):e202101049, February 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/2/
e202101049/212893/Reticulon-like-REEP4-at-the-inner-
nuclear-membrane.

Gonzalez:2022:RCP

[GC22] Beatriz González and Paul J. Cullen. Regulation of Cdc42 protein turnover modulates the filamentous growth MAPK pathway. Journal of Cell Biology, 221(12):e202112100, December 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/12/e202112100/213675/Regulation-of-Cdc42-protein-
turnover-modulates-the.

Gilles:2021:LAE

[GCL^{+}21] Laurine M. Gilles, Andrea R. M. Calhau, Veronica La Padula, Nathanaël M. A. Jacquier, Claire Lionnet, Jean-Pierre Martinant, Peter M. Rogowsky, and Thomas Widiez. Lipid anchoring and electrostatic interactions target NOT-LIKE-DAD to pollen endo-plasma membrane. Journal of Cell Biology, 220(10): e202010077, October 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/220/10/e202010077/212519/Lipid-anchoring-
and-electrostatic-interactions.

Gihana:2021:MEN

[GCNL21] Gabriel M. Gihana, Arthur A. Cross-Najafi, and Soni Lacefield. The mitotic exit network regulates the spatiotemporal activity of Cdc42 to maintain cell size. Journal of Cell Biology, 220(1): e202001016, January 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/220/1/e202001016/211575/The-mitotic-exit-
network-regulates-the.

Grey:2020:INR

[GCS^{+}20] Michael J. Grey, Eva Cloots, Mariska S. Simpson, Nicole LeDuc, Yevgeniy V. Serebrenik, Heidi De Luca, Delphine De Sutter, Phi Luong, Jay R. Thiagarajah, Adrienne W. Paton, James C. Paton, Markus A. Seeliger, Sven Eyckerman, Sophie Janssens, and Wayne I. Lencer. IRE1β negatively regulates IRE1α signaling in response to endoplasmic reticulum stress. Journal of Cell Biology, 219(2):e201904048, February 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/219/2/e201904048/133656/IRE1-
negatively-regulates-IRE1-signaling-in.

Gok:2023:OMM

[GCW^{+}23] Mehmet Oguz Gok, Olivia M. Connor, Xun Wang, Cameron J. Menezes, Claire B. Llamas, Prashant Mishra, and Jonathan R. Friedman. The outer mitochondrial membrane protein TMEM11 demarcates spatially restricted BNIP3/BNIP3L-mediated mitophagy. Journal of Cell Biology, 222(4):e202204021, April 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/
4/e202204021/213874/The-outer-mitochondrial-membrane-
protein-TMEM11.

Grimm:2020:PCI

[GDB^{+}20] Tanja M. Grimm, Nina I. Dierdorf, Karin Betz, Christoph Paone, and Christof R. Hauck. PPM1F controls integrin activity via a conserved phospho-switch. Journal of Cell Biology, 219(12):e202001057, December 7, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/12/
e202001057/211512/PPM1F-controls-integrin-activity-via-
a-conserved.

Gross:2020:SEA

[GG20] Angelina Sarah Gross and Martin Graef. Stress eating: Autophagy targets nuclear pore complexes. Journal of Cell Biology, 219(7):e202006007, July 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/7/e202006007/151894/Stress-eating-
Autophagy-targets-nuclear-pore.

Gaudin:2021:DKS

[GGA21] Noémie Gaudin, Paula Martin Gil, and Juliette Azimzadeh. DISCO is key to successful centriole maturation. Journal of Cell Biology, 220(9):e202107033, September 6, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/9/
e202107033/212560/DISCO-is-key-to-successful-centriole.

Guevara-Garcia:2022:IBA

[GGFBR^{+}22] Amaris Guevara-Garcia, Laure Fourel, Ingrid Bourrin-Reynard, Adria Sales, Christiane Oddou, Mylène Pezet, Olivier Rossier, Paul Machillot, Line Chaar, Anne-Pascale Bouin, Gregory Giannone, Olivier Destaing, Catherine Picart, and Corinne Albiges-Rizo. Integrin-based adhesion compartmentalizes ALK3 of the BMPRII to control cell adhesion and migration. Journal of Cell Biology, 221(12):e202107110, December 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/221/12/e202107110/213529/Integrin-based-
adhesion-compartmentalizes-ALK3-of.

Gliech:2020:KTT

[GH20] Colin R. Gliech and Andrew J. Holland. Keeping track of time: The fundamentals of cellular clocks. Journal of Cell Biology, 219(11):e202005136, November 2, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/11/
e202005136/152075/Keeping-track-of-time-The-fundamentals-
of-cellular.

Gebhardt:2020:TDC

[GKFR20] Lisa A. Gebhardt, Tetyana I. Kichko, Michael J. M. Fischer, and Peter W. Reeh. TRPA1-dependent calcium transients and CGRP release in DRG neurons require extracellular calcium. Journal of Cell Biology, 219(6):e201702151, June 1, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/6/
e201702151/151799/TRPA1-dependent-calcium-transients-

and-CGRP. See reply [LYS^{+}20].

Gngrich:2020:ACE

[GKM^{+}20] Christina Göngrich, Favio A. Krapacher, Hermany Munguba, Diana Fernández-Suárez, Annika Andersson, Jens Hjerling-Leffler, and Carlos F. Ibáñez. ALK4 coordinates extracellular and intrinsic signals to regulate development of cortical somatostatin interneurons. Journal of Cell Biology, 219(1):e201905002, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Gianakas:2023:HMT

[GKRL^{+}23] Claire A. Gianakas, Daniel P. Keeley, William Ramos-Lewis, Kieop Park, Ranjay Jayadev, Isabel W. Kenny, Qiuyi Chi, and David R. Sherwood. Hemicentin-mediated type IV collagen assembly strengthens juxtaposed basement membrane linkage. Journal of Cell Biology, 222(1):e202112096, January 2, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
222/1/e202112096/213571/Hemicentin-mediated-type-IV-
collagen-assembly.

Gottardi:2020:NTF

[GL20] Cara J. Gottardi and G. W. Gant Luxton. Nesprin-2G tension fine-tunes Wnt/β-catenin signaling. Journal of Cell Biology, 219(10):e202009042, October 5, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/10/
e202009042/152112/Nesprin-2G-tension-fine-tunes-Wnt-
catenin.

Gutierrez:2021:KDA

[GLGL^{+}21] Yolanda Gutiérrez, Sergio López-García, Argentina Lario, Silvia Gutiérrez-Eisman, Cédric Delevoye, and José A. Esteban. KIF13A drives AMPA receptor synaptic delivery for long-term potentiation via endosomal remodeling. Journal of Cell Biology, 220(6):e202003183, June 7, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/6/
e202003183/212112/KIF13A-drives-AMPA-receptor-synaptic-
delivery-for.

Guttman:2022:ADP

[GLM^{+}22] Ofer Guttman, Adrien Le Thomas, Scot Marsters, David A. Lawrence, Lauren Gutgesell, Iratxe Zuazo-Gaztelu, Jonathan M. Harnoss, Simone M. Haag, Aditya Murthy, Geraldine Strasser, Zora Modrusan, Thomas Wu, Ira Mellman, and Avi Ashkenazi. Antigen-derived peptides engage the ER stress sensor IRE1α to curb dendritic cell cross-presentation. Journal of Cell Biology, 221(6):e202111068, June 6, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/6/
e202111068/213173/Antigen-derived-peptides-engage-the-
ER-stress.

Guler:2023:SCH

[GM23] Gizem Özbaykal Güler and Serge Mostowy. The septin cytoskeleton: Heteromer composition defines filament function. Journal of Cell Biology, 222(3):e202302010, March 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/222/3/
e202302010/213890/The-septin-cytoskeleton-Heteromer-
composition.

Guadagno:2020:RRF

[GMB^{+}20] Noemi Antonella Guadagno, Azzurra Margiotta, Synne Arstad Bjørnestad, Linda Hofstad Haugen, Ingrid Kjos, Xiaochun Xu, Xian Hu, Oddmund Bakke, Felix Margadant, and Cinzia Progida. Rab18 regulates focal adhesion dynamics by interacting with kinectin-1 at the endoplasmic reticulum. Journal of Cell Biology, 219(7):e201809020, July 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/7/
e201809020/151855/Rab18-regulates-focal-adhesion-dynamics-
by.

Gerbich:2020:PPS

[GMC^{+}20] Therese M. Gerbich, Grace A. McLaughlin, Katelyn Cassidy, Scott Gerber, David Adalsteinsson, and Amy S. Gladfelter. Phosphoregulation provides specificity to biomolecular condensates in the cell cycle and cell polarity. Journal of Cell Biology, 219(7):e201910021, July 6, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/7/
e201910021/151764/Phosphoregulation-provides-specificity-
to.

Guyard:2022:OOO

[GMCO^{+}22] Valentin Guyard, Vera Filipa Monteiro-Cardoso, Mohyeddine Omrane, Cécile Sauvanet, Audrey Houcine, Claire Boulogne, Kalthoum Ben Mbarek, Nicolas Vitale, Orestis Faklaris, Naima El Khallouki, Abdou Rachid Thiam, and Francesca Giordano. ORP5 and ORP8 orchestrate lipid droplet biogenesis and maintenance at ER–mitochondria contact sites. Journal of Cell Biology, 221(9):e202112107, September 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/221/9/e202112107/213393/ORP5-
and-ORP8-orchestrate-lipid-droplet-biogenesis.

Gallusser:2023:DNN

[GMD^{+}23] Benjamin Gallusser, Giorgio Maltese, Giuseppe Di Caprio, Tegy John Vadakkan, Anwesha Sanyal, Elliott Somerville, Mihir Sahasrabudhe, Justin O’Connor, Martin Weigert, and Tom Kirchhausen. Deep neural network automated segmentation of cellular structures in volume electron microscopy. Journal of Cell Biology, 222(2):e202208005, February 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/222/2/e202208005/213736/Deep-neural-network-
automated-segmentation-of.

Grimsley-Myers:2020:VCE

[GMIC^{+}20] Cynthia M. Grimsley-Myers, Robin H. Isaacson, Chantel M. Cadwell, Jazmin Campos, Marina S. Hernandes, Kenneth R. Myers, Tadahiko Seo, William Giang, Kathy K. Griendling, and Andrew P. Kowalczyk. VE-cadherin endocytosis controls vascular integrity and patterning during development. Journal of Cell Biology, 219(5):e201909081, May 4, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/219/5/e201909081/151601/VE-
cadherin-endocytosis-controls-vascular.

Goupil:2020:CFB

[GNL^{+}20] Alix Goupil, Maddalena Nano, Gaëlle Letort, Simon Gemble, Frances Edwards, Oumou Goundiam, Delphine Gogendeau, Carole Pennetier, and Renata Basto. Chromosomes function as a barrier to mitotic spindle bipolarity in polyploid cells. Journal of Cell Biology, 219(4):e201908006, April 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/4/
e201908006/133854/Chromosomes-function-as-a-barrier-
to-mitotic.

Gomez-Navarro:2020:CCC

[GNML^{+}20] Natalia Gomez-Navarro, Alejandro Melero, Xiao-Han Li, Jérôme Boulanger, Wanda Kukulski, and Elizabeth A. Miller. Cargo crowding contributes to sorting stringency in COPII vesicles. Journal of Cell Biology, 219(7):e201806038, July 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/7/
e201806038/151777/Cargo-crowding-contributes-to-sorting-
stringency.

Goodman:2020:BLD

[Goo20] Joel M. Goodman. Building the lipid droplet assembly complex. Journal of Cell Biology, 219(7):e202006025, July 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/7/
e202006025/151882/Building-the-lipid-droplet-assembly.

Girao:2020:CBC

[GOR^{+}20] Hugo Girão, Naoyuki Okada, Tony A. Rodrigues, Alexandra O. Silva, Ana C. Figueiredo, Zaira Garcia, Tatiana Moutinho-Santos, Ikuko Hayashi, Jorge E. Azevedo, Sandra Macedo-Ribeiro, and Helder Maiato. CLASP2 binding to curved microtubule tips promotes flux and stabilizes kinetochore attachments. Journal of Cell Biology, 219(2):e201905080, February 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/2/
e201905080/132496/CLASP2-binding-to-curved-microtubule-
tips-promotes.

Gallego-Paez:2023:TCC

[GPEC^{+}23] Lina M. Gallego-Paez, William J. S. Edwards, Manasa Chanduri, Yanyu Guo, Thijs Koorman, Chieh-Yu Lee, Nina Grexa, Patrick Derksen, Jie Yan, Martin A. Schwartz, Jan Mauer, and Benjamin Thomas Goult. TLN1 contains a cancer-associated cassette exon that alters talin-1 mechanosensitivity. Journal of Cell Biology, 222(5):e202209010, May 1, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/222/5/e202209010/213923/TLN1-
contains-a-cancer-associated-cassette-exon.

Gerbi:2021:WRB

[GPES21] Susan A. Gerbi, Robert E. Palazzo, William C. Earnshaw, and William T. Schrader. William R. Brinkley: a giant in biomedical research and public policy. Journal of Cell Biology, 220(10):e202106102, October 4, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/10/
e202106102/212526/William-R-Brinkley-A-giant-in-biomedical-
research.

Gotz:2021:RRP

[GPL^{+}21] Torsten W. B. Götz, Dmytro Puchkov, Veronika Lysiuk, Janine Lützkendorf, Alexander G. Nikonenko, Christine Quentin, Martin Lehmann, Stephan J. Sigrist, and Astrid G. Petzoldt. Rab2 regulates presynaptic precursor vesicle biogenesis at the trans-Golgi. Journal of Cell Biology, 220 (5):e202006040, May 3, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/220/5/e202006040/211946/Rab2-
regulates-presynaptic-precursor-vesicle.

Georgantzoglou:2022:TSS

[GPW^{+}22] Antonios Georgantzoglou, Hugo Poplimont, Hazel A. Walker, Tim Lämmermann, and Milka Sarris. A two-step search and run response to gradients shapes leukocyte navigation in vivo. Journal of Cell Biology, 221(8):e202103207, August 1, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/8/
e202103207/213303/A-two-step-search-and-run-response-
to-gradients.

Golenberg:2020:CRW

[GSB^{+}20] Netta Golenberg, Jayne M. Squirrell, David A. Bennin, Julie Rindy, Paige E. Pistono, Kevin W. Eliceiri, Miriam A. Shelef, Junsu Kang, and Anna Huttenlocher. Citrullination regulates wound responses and tissue regeneration in zebrafish. Journal of Cell Biology, 219(4):e201908164, April 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/4/
e201908164/133858/Citrullination-regulates-wound-responses-
and.

Goncalves:2020:LTS

[GSC^{+}20] João Gonçalves, Amit Sharma, Étienne Coyaud, Estelle M. N. Laurent, Brian Raught, and Laurence Pelletier. LUZP1 and the tumor suppressor EPLIN modulate actin stability to restrict primary cilia formation. Journal of Cell Biology, 219(7): e201908132, July 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/219/7/e201908132/151837/LUZP1-and-the-tumor-
suppressor-EPLIN-modulate.

Garner:2023:MLC

[GSL^{+}23] Kirsten E. L. Garner, Anna Salter, Clinton K. Lau, Manickam Gurusaran, Cécile M. Villemant, Elizabeth P. Granger, Gavin McNee, Philip G. Woodman, Owen R. Davies, Brian E.

Burke, and Victoria J. Allan. The meiotic LINC complex component KASH5 is an activating adaptor for cytoplasmic dynein. Journal of Cell Biology, 222(5):e202204042, May 1, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/5/
e202204042/213963/The-meiotic-LINC-complex-component-
KASH5-is-an.

Guillen-Samander:2021:VBM

[GSLH^{+}21] Andrés Guillén-Samander, Marianna Leonzino, IV Michael G. Hanna, Ni Tang, Hongying Shen, and Pietro De Camilli. VPS13D bridges the ER to mitochondria and peroxisomes via Miro. Journal of Cell Biology, 220(5):e202010004, May 3, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/5/
e202010004/212021/VPS13D-bridges-the-ER-to-mitochondria-
and.

Gutmann:2020:CES

[GSP^{+}20] Theresia Gutmann, Ingmar B. Schäfer, Chetan Poojari, Beate Brankatschk, Ilpo Vattulainen, Mike Strauss, and Ünal Coskun. Cryo-EM structure of the complete and ligand-saturated insulin receptor ectodomain. Journal of Cell Biology, 219(1): e201907210, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

GuillaumeRomet-Lemonne:2021:DIA

[Gui21] Antoine Jégou Guillaume Romet-Lemonne . The dynamic instability of actin filament barbed ends. Journal of Cell Biology, 220(4):??, April 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/4/e202102020/211908/The-dynamic-
instability-of-actin-filament-barbed.

Greenan:2020:ECI

[GVA20] Garrett A. Greenan, Ronald D. Vale, and David A. Agard. Electron cryotomography of intact motile cilia defines the basal body to axoneme transition. Journal of Cell Biology, 219(1): e201907060, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Grond:2020:CFG

[GVD^{+}20a] Rianne Grond, Tineke Veenendaal, Juan M. Duran, Ishier Raote, Johan H. van Es, Sebastiaan Corstjens, Laura Delfgou, Benaissa El Haddouti, Vivek Malhotra, and Catherine Rabouille. Correction: The function of GORASPs in Golgi apparatus organization in vivo. Journal of Cell Biology, 219 (9):??, September 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/9/jcb.20200419106242020c/151901/
Correction-The-function-of-GORASPs-in-Golgi. See [GVD^{+}20b].

Grond:2020:FGG

[GVD^{+}20b] Rianne Grond, Tineke Veenendaal, Juan M. Duran, Ishier Raote, Johan H. van Es, Sebastiaan Corstjens, Laura Delfgou, Benaissa El Haddouti, Vivek Malhotra, and Catherine Rabouille. The function of GORASPs in Golgi apparatus organization in vivo. Journal of Cell Biology, 219(9):e202004191, September 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/
9/e202004191/151880/The-function-of-GORASPs-in-Golgi-

apparatus. See correction [GVD^{+}20a].

Gao:2021:TRB

[GWR^{+}21] Ju Gao, Luwen Wang, Xiaojia Ren, Justin R. Dunn, Ariele Peters, Masaru Miyagi, Hisashi Fujioka, Fangli Zhao, Candice Askwith, Jingjing Liang, and Xinglong Wang. Translational regulation in the brain by TDP-43 phase separation. Journal of Cell Biology, 220(10):e202101019, October 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
10/e202101019/212594/Translational-regulation-in-the-
brain-by-TDP-43.

Gupta:2020:HFC

[GY20] Shafali Gupta and Alpha S. Yap. Hands and feet: Closer than you think in epithelial migration. Journal of Cell Biology, 219(10):e202008069, October 5, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/219/10/e202008069/152110/Hands-
and-feet-Closer-than-you-think-in-epithelial.

Hakanpaa:2023:RAA

[HAL^{+}23] Laura Hakanpää, Amr Abouelezz, An-Sofie Lenaerts, Seyda Culfa, Michael Algie, Jenny Bärlund, Pekka Katajisto, Harvey McMahon, and Leonardo Almeida-Souza. Reticular adhesions are assembled at flat clathrin lattices and opposed by active integrin α5β1. Journal of Cell Biology, 222(8):e202303107, August 7, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/8/
e202303107/214133/Reticular-adhesions-are-assembled-
at-flat-clathrin.

Hong:2022:MML

[HAW^{+}22] Zhouping Hong, Jyoti Adlakha, Neng Wan, Emily Guinn, Fabian Giska, Kallol Gupta, Thomas J. Melia, and Karin M. Reinisch. Mitoguardin-2–mediated lipid transfer preserves mitochondrial morphology and lipid droplet formation. Journal of Cell Biology, 221(12):e202207022, December 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/12/
e202207022/213573/Mitoguardin-2-mediated-lipid-transfer-
preserves.

Henrie:2020:SIP

[HBDC^{+}20] Hélène Henrie, Dalal Bakhos-Douaihy, Isabelle Cantaloube, Antoine Pilon, Maya Talantikite, Virginie Stoppin-Mellet, Anita Baillet, Christian Poüs, and Béatrice Benoit. Stress-induced phosphorylation of CLIP-170 by JNK promotes microtubule rescue. Journal of Cell Biology, 219(7):e201909093, July 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/7/
e201909093/151834/Stress-induced-phosphorylation-of-
CLIP-170-by-JNK.

Hecht:2020:FRP

[HBS^{+}20] Tobias Karl-Heinz Hecht, Birgit Blank, Martin Steger, Victor Lopez, Gisela Beck, Bulat Ramazanov, Matthias Mann, Vincent Tagliabracci, and Julia von Blume. Fam20C regulates protein secretion by Cab45 phosphorylation. Journal of Cell Biology, 219(6):e201910089, June 1, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/6/
e201910089/151783/Fam20C-regulates-protein-secretion-
by-Cab45.

Huang:2023:AID

[HCB^{+}23] Zhentai Huang, Chi-Wei Chen, Raquel Buj, Naveen Kumar Tangudu, Richard S. Fang, Kelly E. Leon, Erika S. Dahl, Erika L.

Varner, Eliana von Krusenstiern, Aidan R. Cole, Nathaniel W. Snyder, and Katherine M. Aird. ATM inhibition drives metabolic adaptation via induction of macropinocytosis. Journal of Cell Biology, 222(1):e202007026, January 2, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:/

/rupress.org/jcb/article/222/1/e202007026/213697/ATM-
inhibition-drives-metabolic-adaptation-via.

Hurlock:2020:INS

[HČK^{+}20] Matthew E. Hurlock, Ivana Čavka, Lisa E. Kursel, Jocelyn Haversat, Matthew Wooten, Zehra Nizami, Rashi Turniansky, Philipp Hoess, Jonas Ries, Joseph G. Gall, Ofer Rog, Simone Köhler, and Yumi Kim. Identification of novel synaptonemal complex components in C. elegans. Journal of Cell Biology, 219(5):e201910043, May 4, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/5/
e201910043/151584/Identification-of-novel-synaptonemal-
complex.

Ho:2021:TMS

[HCL^{+}21] Wan Yun Ho, Jer-Cherng Chang, Kenneth Lim, Amaury Cazenave-Gassiot, Aivi T. Nguyen, Juat Chin Foo, Sneha Muralidharan, Ashley Viera-Ortiz, Sarah J. M. Ong, Jin Hui Hor, Ira Agrawal, Shawn Hoon, Olubankole Aladesuyi Arogundade, Maria J. Rodriguez, Su Min Lim, Seung Hyun Kim, John Ravits, Shi-Yan Ng, Markus R. Wenk, Edward B. Lee, Greg Tucker-Kellogg, and Shuo-Chien Ling. TDP-43 mediates SREBF2regulated gene expression required for oligodendrocyte myelination. Journal of Cell Biology, 220(9):e201910213, September 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
9/e201910213/212536/TDP-43-mediates-SREBF2-regulated-
gene-expression.

Huet-Calderwood:2023:FSF

[HCRMTC23] Clotilde Huet-Calderwood, Felix E. Rivera-Molina, Derek K. Toomre, and David A. Calderwood. Fibroblasts secrete fibronectin under lamellipodia in a microtubule- and myosin II– dependent fashion. Journal of Cell Biology, 222(2):e202204100, February 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/222/2/e202204100/213712/Fibroblasts-secrete-
fibronectin-under-lamellipodia.

Hancock-Cerutti:2022:LLT

[HCWX^{+}22] William Hancock-Cerutti, Zheng Wu, Peng Xu, Narayana Yadavalli, Marianna Leonzino, Arun Kumar Tharkeshwar, Shawn M. Ferguson, Gerald S. Shadel, and Pietro De Camilli. ER-lysosome lipid transfer protein VPS13C/PARK23 prevents aberrant mtDNA-dependent STING signaling. Journal of Cell Biology, 221(7):e202106046, July 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/221/7/e202106046/213259/ER-
lysosome-lipid-transfer-protein-VPS13C-PARK23.

Harmon:2022:DCD

[HDG22] Robert M. Harmon, John Devany, and Margaret L. Gardel. Dia1 coordinates differentiation and cell sorting in a stratified epithelium. Journal of Cell Biology, 221(5):e202101008, May 2, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/
5/e202101008/213092/Dia1-coordinates-differentiation-
and-cell-sorting.

Herbst:2021:NAR

[HDW^{+}21] Wendy A. Herbst, Weixian Deng, James A. Wohlschlegel, Jennifer M. Achiro, and Kelsey C. Martin. Neuronal activity regulates the nuclear proteome to promote activity-dependent transcription. Journal of Cell Biology, 220(12):e202103087, December 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/12/e202103087/212684/Neuronal-activity-regulates-
the-nuclear-proteome.

Hirsch:2022:FMA

[HESH^{+}22] Sophia M. Hirsch, Frances Edwards, Mimi Shirasu-Hiza, Julien Dumont, and Julie C. Canman. Functional midbody assembly in the absence of a central spindle. Journal of Cell Biology, 221(3):e202011085, March 7, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/3/
e202011085/212948/Functional-midbody-assembly-in-the-
absence-of-a.

Hennlein:2023:PRC

[HGG^{+}23] Luisa Hennlein, Hanaa Ghanawi, Florian Gerstner, Eduardo Palominos García, Ezgi Yildirim, Lena Saal-Bauernschubert, Mehri Moradi, Chunchu Deng, Teresa Klein, Silke Appenzeller, Markus Sauer, Michael Briese, Christian Simon, Michael Sendtner, and Sibylle Jablonka. Plastin 3 rescues cell surface translocation and activation of TrkB in spinal muscular atrophy. Journal of Cell Biology, 222(3):e202204113, March 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/3/
e202204113/213794/Plastin-3-rescues-cell-surface-translocation-
and.

Huang:2020:QCR

[HGK20] Haina Huang, Homa Ghalei, and Katrin Karbstein. Quality control of 40S ribosome head assembly ensures scanning competence. Journal of Cell Biology, 219(11):e202004161, November 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/
11/e202004161/152152/Quality-control-of-40S-ribosome-
head-assembly.

Hein:2021:CCI

[HGN^{+}21] Jamin B. Hein, Dimitriya H. Garvanska, Isha Nasa, Arminja N. Kettenbach, and Jakob Nilsson. Coupling of Cdc20 inhibition and activation by BubR1. Journal of Cell Biology, 220 (5):e202012081, May 3, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/5/e202012081/211939/Coupling-of-
Cdc20-inhibition-and-activation-by.

Hummel:2021:SKA

[HH21] Jessica J. A. Hummel and Casper C. Hoogenraad. Specific KIF1A–adaptor interactions control selective cargo recognition. Journal of Cell Biology, 220(10):e202105011, October 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/10/
e202105011/212488/Specific-KIF1A-adaptor-interactions-
control.

Hariri:2022:FGS

[HH22] Hanaa Hariri and W. Mike Henne. Filling in the gaps: SNX-RGS proteins as multiorganelle tethers. Journal of Cell Biology, 221 (5):e202203061, May 2, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/221/5/e202203061/213114/Filling-in-the-gaps-
SNX-RGS-proteins-as.

Hulmes:2020:PIC

[HHD^{+}20] Georgia E. Hulmes, John D. Hutchinson, Noa Dahan, James M. Nuttall, Ellen G. Allwood, Kathryn R. Ayscough, and Ewald H. Hettema. The Pex3–Inp1 complex tethers yeast peroxisomes to the plasma membrane. Journal of Cell Biology, 219(10): e201906021, October 5, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/219/10/e201906021/152119/The-
Pex3-Inp1-complex-tethers-yeast-peroxisomes-to.

Hirst:2021:RGP

[HHGR21] Jennifer Hirst, Geoffrey G. Hesketh, Anne-Claude Gingras, and Margaret S. Robinson. Rag GTPases and phosphatidylinositol 3-phosphate mediate recruitment of the AP-5/SPG11/ SPG15 complex. Journal of Cell Biology, 220(2):e202002075, February 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/220/2/e202002075/211690/Rag-GTPases-and-
phosphatidylinositol-3-phosphate.

Hadders:2020:UCH

[HHT^{+}20] Michael A. Hadders, Sanne Hindriksen, My Anh Truong, Aditya N. Mhaskar, J. Pepijn Wopken, Martijn J. M. Vromans, and Susanne M. A. Lens. Untangling the contribution of Haspin and Bub1 to Aurora B function during mitosis. Journal of Cell Biology, 219(3):e201907087, March 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/3/e201907087/133700/Untangling-the-contribution-
of-Haspin-and-Bub1-to.

Hamidi:2021:FTH

[HI21] Hellyeh Hamidi and Johanna Ivaska. Food for thought: How cell adhesion coordinates nutrient sensing. Journal of Cell Biology, 220(5):e202103128, May 3, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/220/5/e202103128/211983/Food-
for-thought-How-cell-adhesion-coordinates.

Hickson:2022:BSM

[Hic22] Gilles R. X. Hickson. A back-up source of microtubules for the midbody during cytokinesis. Journal of Cell Biology, 221(3): e202201028, March 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/221/3/e202201028/213023/A-back-up-source-
of-microtubules-for-the-midbody.

Hooper:2022:VAU

[HJL^{+}22] Kirsty M. Hooper, Elise Jacquin, Taoyingnan Li, Jonathan M. Goodwin, John H. Brumell, Joanne Durgan, and Oliver Florey. V-ATPase is a universal regulator of LC3-associated phagocytosis and non-canonical autophagy. Journal of Cell Biology, 221 (6):e202105112, June 6, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/221/6/e202105112/213194/V-ATPase-is-
a-universal-regulator-of-LC3.

Han:2020:NTB

[HKK^{+}20] Myeong Hoon Han, Min Jee Kwon, Byung Su Ko, Do Young Hyeon, Davin Lee, Hyung-Jun Kim, Daehee Hwang, and Sung Bae Lee. NF-κB disinhibition contributes to dendrite defects in fly models of neurodegenerative diseases. Journal of Cell Biology, 219(12):e202004107, December 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/12/
e202004107/211484/NF-B-disinhibition-contributes-to-
dendrite-defects.

Hadders:2021:DFC

[HL21] Michael A. Hadders and Susanne M. A. Lens. Delaying the final cut: a close encounter of checkpoint kinases at the midbody. Journal of Cell Biology, 220(2):e202012130, February 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/2/e202012130/211655/Delaying-the-
final-cut-A-close-encounter-of.

Hannaford:2022:PIK

[HLB^{+}22] Matthew R. Hannaford, Rong Liu, Neil Billington, Zachary T. Swider, Brian J. Galletta, Carey J. Fagerstrom, Christian Combs, James R. Sellers, and Nasser M. Rusan. Pericentrin interacts with Kinesin-1 to drive centriole motility. Journal of Cell Biology, 221(9):e202112097, September 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/9/
e202112097/213380/Pericentrin-interacts-with-Kinesin-
1-to-drive.

Houston:2020:RKF

[HLGD20] Jack Houston, Pablo Lara-Gonzalez, and Arshad Desai. Rashomon at the kinetochore: Function(s) of the Mad1–cyclin B1 complex. Journal of Cell Biology, 219(8):e202006006, August 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/8/e202006006/151919/Rashomon-at-the-kinetochore-
Function-s-of-the-Mad1.

Hiragi:2022:TRG

[HMSF22] Shu Hiragi, Takahide Matsui, Yuriko Sakamaki, and Mitsunori Fukuda. TBC1D18 is a Rab5-GAP that coordinates endosome maturation together with Mon1. Journal of Cell Biology, 221(12):e202201114, December 5, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/12/
e202201114/213520/TBC1D18-is-a-Rab5-GAP-that-coordinates-
endosome.

Herrera:2021:DTC

[HMT^{+}21] Antonio Herrera, Anghara Menendez, Blanca Torroba, Andrea Ochoa, and Sebastián Pons. Dbnl and β-catenin promote pro-n-cadherin processing to maintain apico-basal polarity. Journal of Cell Biology, 220(6):e202007055, June 7, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/6/e202007055/212044/Dbnl-and-catenin-promote-pro-
N-cadherin-processing.

Hoke:2022:CIC

[Hök22] Ahmet Höke. cADPR induced calcium influx mediates axonal degeneration caused by paclitaxel. Journal of Cell Biology, 221 (2):e202112021, February 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/221/2/e202112021/212955/cADPR-induced-
calcium-influx-mediates-axonal.

Hurwitz:2023:ABP

[HPO^{+}23] Eric Hurwitz, Parash Parajuli, Seval Ozkan, Celine Prunier, Thien Ly Nguyen, Deanna Campbell, Creighton Friend, Allyn Austin Bryan, Ting-Xuan Lu, Steven Christopher Smith, Mohammed Shawkat Razzaque, Keli Xu, and Azeddine Atfi. Antagonism between Prdm16 and Smad4 specifies the trajectory and progression of pancreatic cancer. Journal of Cell Biology, 222(4):e202203036, April 3, 2023. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/222/4/
e202203036/213901/Antagonism-between-Prdm16-and-Smad4-
specifies-the.

Hu:2021:GMW

[HRB^{+}21] Bo Hu, Juan J. Rodriguez, Anurag Kakkerla Balaraju, Yuanyuan Gao, Nhan T. Nguyen, Heston Steen, Saeb Suhaib, Songhai Chen, and Fang Lin. Glypican 4 mediates Wnt transport between germ layers via signaling filopodia. Journal of Cell Biology, 220(12):e202009082, December 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/12/
e202009082/212673/Glypican-4-mediates-Wnt-transport-
between-germ.

Horn:2020:MFE

[HRS^{+}20] Adam Horn, Shreya Raavicharla, Sonna Shah, Dan Cox, and Jyoti K. Jaiswal. Mitochondrial fragmentation enables localized signaling required for cell repair. Journal of Cell Biology, 219(5):e201909154, May 4, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/5/
e201909154/151605/Mitochondrial-fragmentation-enables-
localized.

Higashi:2023:EPR

[HSF^{+}23] Tomohito Higashi, Akira C. Saito, Yugo Fukazawa, Mikio Furuse, Atsuko Y. Higashi, Masahiro Ono, and Hideki Chiba. EpCAM proteolysis and release of complexed claudin-7 repair and maintain the tight junction barrier. Journal of Cell Biology, 222(1):e202204079, January 2, 2023. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/222/1/

e202204079/213688/EpCAM-proteolysis-and-release-of-complexed-
claudin.

Haneke:2020:CCP

[HSL^{+}20] Katharina Haneke, Johanna Schott, Doris Lindner, Anne Kruse Hollensen, Christian Kroun Damgaard, Cyril Mongis, Michael Knop, Wilhelm Palm, Alessia Ruggieri, and Georg Stoecklin. CDK1 couples proliferation with protein synthesis. Journal of Cell Biology, 219(3):e201906147, March 2, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/3/
e201906147/133706/CDK1-couples-proliferation-with-protein.

Hebbar:2020:HSB

[HSSK20] Sarita Hebbar, Kai Schuhmann, Andrej Shevchenko, and Elisabeth Knust. Hydroxylated sphingolipid biosynthesis regulates photoreceptor apical domain morphogenesis. Journal of Cell Biology, 219(12):e201911100, December 7, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/12/
e201911100/211460/Hydroxylated-sphingolipid-biosynthesis-
regulates.

He:2020:DAG

[HSU^{+}20] Kangmin He, Eli Song, Srigokul Upadhyayula, Song Dang, Raphael Gaudin, Wesley Skillern, Kevin Bu, Benjamin R. Capraro, Iris Rapoport, Ilja Kusters, Minghe Ma, and Tom Kirchhausen. Dynamics of Auxilin 1 and GAK in clathrinmediated traffic. Journal of Cell Biology, 219(3):e201908142, March 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/219/3/e201908142/133624/Dynamics-of-Auxilin-
1-and-GAK-in-clathrin-mediated.

Hanna:2022:SCM

[HSW^{+}22] Michael G. Hanna, Patreece H. Suen, Yumei Wu, Karin M. Reinisch, and Pietro De Camilli. SHIP164 is a chorein motif lipid transfer protein that controls endosome–Golgi membrane traffic. Journal of Cell Biology, 221(6):e202111018, June 6, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/
6/e202111018/213192/SHIP164-is-a-chorein-motif-lipid-
transfer-protein.

Huang:2021:CIP

[HTL^{+}21] Xinyue Huang, Shuixia Tan, Yanxia Li, Shuangyi Cao, Xingyan Li, Heling Pan, Bing Shan, Lihui Qian, and Junying Yuan. Caspase inhibition prolongs inflammation by promoting a signaling complex with activated RIPK1. Journal of Cell Biology, 220(6):e202007127, June 7, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/6/
e202007127/212035/Caspase-inhibition-prolongs-inflammation-
by.

Hasler:2020:GWM

[HVPM20] Safa Lucken-Ardjomande Häsler, Yvonne Vallis, Mathias Pasche, and Harvey T. McMahon. GRAF2, WDR44, and MI-CAL1 mediate Rab8/10/11–dependent export of E-cadherin, MMP14, and CFTR ΔF508. Journal of Cell Biology, 219(5): e201811014, May 4, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/219/5/e201811014/151714/GRAF2-WDR44-and-
MICAL1-mediate-Rab8-10-11.

Huang:2022:CCR

[HW22] Yunyun Huang and Rudolf Winklbauer. Cell cortex regulation by the planar cell polarity protein prickle1. Journal of Cell Biology, 221(7):e202008116, July 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/221/7/e202008116/213195/Cell-
cortex-regulation-by-the-planar-cell-polarity.

Yamamoto:2020:CEG

[hYKO^{+}20a] Yo hei Yamamoto, Ayano Kasai, Hiroko Omori, Tomoe Takino, Munechika Sugihara, Tetsuo Umemoto, Maho Hamasaki, Tomohisa Hatta, Tohru Natsume, Richard I. Morimoto, Ritsuko Arai, Satoshi Waguri, Miyuki Sato, Ken Sato, Shoshana Bar-Nun, Tamotsu Yoshimori, Takeshi Noda, and Kazuhiro Nagata. Correction: ERdj8 governs the size of autophagosomes during the formation process. Journal of Cell Biology, 219(8):e20190312709142020c, August 3, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/8/
e20190312709142020c/152105/Correction-ERdj8-governs-

the-size-of. See [hYKO^{+}20b].

Yamamoto:2020:EGS

[hYKO^{+}20b] Yo hei Yamamoto, Ayano Kasai, Hiroko Omori, Tomoe Takino, Munechika Sugihara, Tetsuo Umemoto, Maho Hamasaki, Tomohisa Hatta, Tohru Natsume, Richard I. Morimoto, Ritsuko Arai, Satoshi Waguri, Miyuki Sato, Ken Sato, Shoshana Bar-Nun, Tamotsu Yoshimori, Takeshi Noda, and Kazuhiro Nagata. ERdj8 governs the size of autophagosomes during the formation process. Journal of Cell Biology, 219(8):e201903127, August 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/8/
e201903127/151832/ERdj8-governs-the-size-of-autophagosomes-

during. See correction [hYKO^{+}20a].

Yamamoto:2021:CEG

[hYKO^{+}21] Yo hei Yamamoto, Ayano Kasai, Hiroko Omori, Tomoe Takino, Munechika Sugihara, Tetsuo Umemoto, Maho Hamasaki, Tomohisa Hatta, Tohru Natsume, Richard I. Morimoto, Ritsuko Arai, Satoshi Waguri, Miyuki Sato, Ken Sato, Shoshana Bar-Nun, Tamotsu Yoshimori, Takeshi Noda, and Kazuhiro Nagata. Correction: ERdj8 governs the size of autophagosomes during the formation process. Journal of Cell Biology, 220(9): e20190312707282021c, September 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/220/9/e20190312707282021c/
212542/Correction-ERdj8-governs-the-size-of.

Huang:2020:NRT

[HYL^{+}20] Rui Huang, De-Juan Yuan, Shao Li, Xue-Song Liang, Yue Gao, Xiao-Yan Lan, Hua-Min Qin, Yu-Fang Ma, Guang-Yin Xu, Melitta Schachner, Vladimir Sytnyk, Johannes Boltze, Quan-Hong Ma, and Shen Li. NCAM regulates temporal specification of neural progenitor cells via profilin2 during corticogenesis. Journal of Cell Biology, 219(1):e201902164, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Ho:2020:SSI

[HYX^{+}20] Nurulain Ho, Wei Sheng Yap, Jiaming Xu, Haoxi Wu, Jhee Hong Koh, Wilson Wen Bin Goh, Bhawana George, Shu Chen Chong, Stefan Taubert, and Guillaume Thibault. Stress sensor Ire1 deploys a divergent transcriptional program in response to lipid bilayer stress. Journal of Cell Biology, 219(7):e201909165, July 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/7/e201909165/151719/Stress-sensor-Ire1-deploys-a-
divergent.

Hu:2022:LSW

[HZCX22] Meiqin Hu, Nan Zhou, Weijie Cai, and Haoxing Xu. Lysosomal solute and water transport. Journal of Cell Biology, 221(11): e202109133, November 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/221/11/e202109133/213536/Lysosomal-solute-
and-water-transportIon-and-water.

Hao:2021:MBS

[HZN^{+}21] Jia Hao, Hao Zhou, Kristen Nemes, Daniel Yen, Winfield Zhao, Charles Bramlett, Bowen Wang, Rong Lu, and Keyue Shen. Membrane-bound SCF and VCAM-1 synergistically regulate the morphology of hematopoietic stem cells. Journal of Cell Biology, 220(10):e202010118, October 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/10/e202010118/212562/Membrane-bound-SCF-and-VCAM-
1-synergistically.

Itakura:2020:HSC

[ICMM20] Eisuke Itakura, Momoka Chiba, Takeshi Murata, and Akira Matsuura. Heparan sulfate is a clearance receptor for aberrant extracellular proteins. Journal of Cell Biology, 219(3): e201911126, March 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/219/3/e201911126/133807/Heparan-sulfate-
is-a-clearance-receptor-for.

Illukkumbura:2023:DPS

[IHBP^{+}23] Rukshala Illukkumbura, Nisha Hirani, Joana Borrego-Pinto, Tom Bland, KangBo Ng, Lars Hubatsch, Jessica McQuade, Robert G. Endres, and Nathan W. Goehring. Design principles for selective polarization of PAR proteins by cortical flows. Journal of Cell Biology, 222(8):e202209111, August 7, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/
8/e202209111/214138/Design-principles-for-selective-
polarization-of.

Ikeda:2020:MIS

[Ike20] Fumiyo Ikeda. Mitophagy is induced by short ubiquitin chains on mitochondria. Journal of Cell Biology, 219(9): e202008031, September 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/9/e202008031/152051/Mitophagy-is-
induced-by-short-ubiquitin-chains-on.

Ignatenko:2023:MDC

[IMR^{+}23] Olesia Ignatenko, Satu Malinen, Sofiia Rybas, Helena Vihinen, Joni Nikkanen, Aleksander Kononov, Eija S. Jokitalo, Gulayse Ince-Dunn, and Anu Suomalainen. Mitochondrial dysfunction compromises ciliary homeostasis in astrocytes. Journal of Cell Biology, 222(1):e202203019, January 2, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/1/
e202203019/213692/Mitochondrial-dysfunction-compromises-
ciliary.

Iemura:2021:COP

[INM^{+}21] Kenji Iemura, Toyoaki Natsume, Kayoko Maehara, Masato T. Kanemaki, and Kozo Tanaka. Chromosome oscillation promotes Aurora A–dependent Hec1 phosphorylation and mitotic fidelity. Journal of Cell Biology, 220(7):e202006116, July 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
7/e202006116/212099/Chromosome-oscillation-promotes-
Aurora-A-dependent.

Ivanovska:2023:SLD

[ITB^{+}23] Irena L. Ivanovska, Michael P. Tobin, Tianyi Bai, Lawrence J. Dooling, and Dennis E. Discher. Small lipid droplets are rigid enough to indent a nucleus, dilute the lamina, and cause rupture. Journal of Cell Biology, 222(8):e202208123, August 7, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/222/8/e202208123/214122/Small-lipid-droplets-
are-rigid-enough-to-indent-a.

Ishimoto:2021:SXM

[ITM^{+}21] Riko Ishimoto, Yota Tsuzuki, Tomoki Matsumura, Seiichiro Kurashige, Kouki Enokitani, Koki Narimatsu, Mitsunori Higa, Nozomi Sugimoto, Kazumasa Yoshida, and Masatoshi Fujita. SLX4–XPF mediates DNA damage responses to replication stress induced by DNA–protein interactions. Journal of Cell Biology, 220(1):e202003148, January 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/220/1/e202003148/211628/SLX4-
XPF-mediates-DNA-damage-responses-to.

Isensee:2021:DIN

[IvCD^{+}21] Jörg Isensee, Marianne van Cann, Patrick Despang, Dioneia Araldi, Katharina Moeller, Jonas Petersen, Achim Schmidtko, Jan Matthes, Jon D. Levine, and Tim Hucho. Depolarization induces nociceptor sensitization by Ca V 1.2-mediated PKA-II activation. Journal of Cell Biology, 220(10):e202002083, October 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/220/10/e202002083/212600/Depolarization-
induces-nociceptor-sensitization-by.

Ito:2021:CCM

[IWI^{+}21] Kei K. Ito, Koki Watanabe, Haruki Ishida, Kyohei Matsuhashi, Takumi Chinen, Shoji Hata, and Daiju Kitagawa. Cep57 and Cep57L1 maintain centriole engagement in interphase to ensure centriole duplication cycle. Journal of Cell Biology, 220(3):e202005153, March 1, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/3/
e202005153/211711/Cep57-and-Cep57L1-maintain-centriole-
engagement-in.

Joly:2020:PMS

[JBV^{+}20] Nicolas Joly, Eva Beaumale, Lucie Van Hove, Lisa Martino, and Lionel Pintard. Phosphorylation of the microtubule-severing AAA+ enzyme Katanin regulates C. elegans embryo development. Journal of Cell Biology, 219(6):e201912037, June 1, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/6/
e201912037/151781/Phosphorylation-of-the-microtubule-
severing-AAA.

Jensen:2023:PPA

[JCL^{+}23] Corbin C. Jensen, Amber N. Clements, Hope Liou, Lauren E. Ball, Jennifer R. Bethard, Paul R. Langlais, Rachel K. Toth, Shailender S. Chauhan, Andrea L. Casillas, Sohail R. Daulat, Andrew S. Kraft, Anne E. Cress, Cindy K. Miranti, Ghassan Mouneimne, Greg C. Rogers, and Noel A. Warfel. PIM1 phosphorylates ABI2 to enhance actin dynamics and promote tumor invasion. Journal of Cell Biology, 222(6):e202208136, June 5, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
222/6/e202208136/214022/PIM1-phosphorylates-ABI2-to-
enhance-actin-dynamics.

Jani:2022:PBR

[JDKK^{+}22] Riddhi Atul Jani, Aurélie Di Cicco, Tal Keren-Kaplan, Silvia Vale-Costa, Daniel Hamaoui, Ilse Hurbain, Feng-Ching Tsai, Mathilde Di Marco, Anne-Sophie Macé, Yueyao Zhu, Maria João Amorim, Patricia Bassereau, Juan S. Bonifacino, Agathe Subtil, Michael S. Marks, Daniel Lévy, Graça Raposo, and Cédric Delevoye. PI4P and BLOC-1 remodel endosomal membranes into tubules. Journal of Cell Biology, 221(11):e202110132, November 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/221/11/e202110132/213508/PI4P-and-BLOC-1-
remodel-endosomal-membranes-into.

Jiang:2022:AKD

[JFM^{+}22] Qinqin Jiang, Martina Foglizzo, Yaroslav I. Morozov, Xuejiao Yang, Arindam Datta, Lei Tian, Vaughn Thada, Weihua Li, Elton Zeqiraj, and Roger A. Greenberg. Autologous K63 deubiquitylation within the BRCA1-A complex licenses DNA damage recognition. Journal of Cell Biology, 221(9):e202111050, September 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/9/e202111050/213388/Autologous-K63-deubiquitylation-
within-the-BRCA1-A.

Jarsch:2020:DRS

[JGN^{+}20] Iris K. Jarsch, Jonathan R. Gadsby, Annalisa Nuccitelli, Julia Mason, Hanae Shimo, Ludovic Pilloux, Bishara Marzook, Claire M. Mulvey, Ulrich Dobramysl, Charles R. Bradshaw, Kathryn S. Lilley, Richard D. Hayward, Tristan J. Vaughan, Claire L. Dobson, and Jennifer L. Gallop. A direct role for SNX9 in the biogenesis of filopodia. Journal of Cell Biology, 219(4):e201909178, April 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/219/4/e201909178/151579/A-
direct-role-for-SNX9-in-the-biogenesis-of.

Jansen:2023:LCM

[JIBK23] Klara I. Jansen, Malina K. Iwanski, Mithila Burute, and Lukas C. Kapitein. A live-cell marker to visualize the dynamics of stable microtubules throughout the cell cycle. Journal of Cell Biology, 222(5):e202106105, May 1, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/222/5/e202106105/213914/A-
live-cell-marker-to-visualize-the-dynamics-of.

Jipa:2023:DAT

[JJ23] András Jipa and Gábor Juhász. Disagreement among the three musketeers of the unfolded protein response. Journal of Cell Biology, 222(5):e202304013, May 1, 2023. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/222/5/
e202304013/214028/Disagreement-among-the-three-musketeers-
of-the.

Jung:2022:RCS

[JKL^{+}22] Juan Jung, Muzamil Majid Khan, Jonathan Landry, Aliaksandr Halavatyi, Pedro Machado, Miriam Reiss, and Rainer Pepperkok. Regulation of the COPII secretory machinery via focal adhesions and extracellular matrix signaling. Journal of Cell Biology, 221(8):e202110081, August 1, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/8/
e202110081/213351/Regulation-of-the-COPII-secretory-
machinery-via.

Jia:2022:VCM

[JKZ^{+}22] Xiaotong Jia, Anastasia Knyazeva, Yu Zhang, Sergio Castro-Gonzalez, Shuhei Nakamura, Lars-Anders Carlson, Tamotsu Yoshimori, Dale P. Corkery, and Yao-Wen Wu. V. cholerae MakA is a cholesterol-binding pore-forming toxin that induces non-canonical autophagy. Journal of Cell Biology, 221(12): e202206040, December 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/221/12/e202206040/213518/V-cholerae-
MakA-is-a-cholesterol-binding-pore.

Ji:2022:VTE

[JLS^{+}22] Mingming Ji, Meng Li, Long Sun, Hongyu Zhao, Ying Li, Lulu Zhou, Zhenni Yang, Xin Zhao, Wenyan Qu, Hanbing Xue, Ze Zheng, Yiming Li, Hongyu Deng, and Yan G. Zhao. VMP1 and TMEM41B are essential for DMV formation during β-coronavirus infection. Journal of Cell Biology, 221(6): e202112081, June 6, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/221/6/e202112081/213207/VMP1-and-TMEM41B-
are-essential-for-DMV-formation.

Jackman:2020:CBC

[JMB^{+}20] Mark Jackman, Chiara Marcozzi, Martina Barbiero, Mercedes Pardo, Lu Yu, Adam L. Tyson, Jyoti S. Choudhary, and Jonathon Pines. Cyclin B1-Cdk1 facilitates MAD1 release from the nuclear pore to ensure a robust spindle checkpoint. Journal of Cell Biology, 219(6):e201907082, June 1, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/
6/e201907082/151606/Cyclin-B1-Cdk1-facilitates-MAD1-
release-from-the.

Jiang:2020:MAB

[JMC^{+}20] Yu-Yang Jiang, Wolfgang Maier, Uzoamaka N. Chukka, Michael Choromanski, Chinkyu Lee, Ewa Joachimiak, Dorota Wloga, Wayland Yeung, Natarajan Kannan, Joseph Frankel, and Jacek Gaertig. Mutual antagonism between Hippo signaling and cyclin E drives intracellular pattern formation. Journal of Cell Biology, 219(9):e202002077, September 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/
9/e202002077/151920/Mutual-antagonism-between-Hippo-
signaling-and.

Jimenez-Moreno:2023:ADL

[JMKS^{+}23] Natalia Jiménez-Moreno, Madhu Kollareddy, Petros Stathakos, Joanna J. Moss, Zuriñe Antón, Deborah K. Shoemark, Richard B. Sessions, Ralph Witzgall, Maeve Caldwell, and Jon D. Lane. ATG8-dependent LMX1B-autophagy crosstalk shapes human midbrain dopaminergic neuronal resilience. Journal of Cell Biology, 222(5):e201910133, May 1, 2023. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/222/5/

e201910133/213994/ATG8-dependent-LMX1B-autophagy-crosstalk-
shapes.

Jones:2021:ZPT

[JML^{+}21] Dakota L. Jones, Jeffrey A. Meridew, Patrick A. Link, Merrick T. Ducharme, Katherine L. Lydon, Kyoung M. Choi, Nunzia Caporarello, Qi Tan, Ana Maria Diaz Espinosa, Yuning Xiong, Jeong-Heon Lee, Zhenqing Ye, Huihuang Yan, Tamas Ordog, Giovanni Ligresti, Xaralabos Varelas, and Daniel J. Tschumperlin. ZNF416 is a pivotal transcriptional regulator of fibroblast mechanoactivation. Journal of Cell Biology, 220 (5):e202007152, May 3, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/5/e202007152/211825/ZNF416-is-a-
pivotal-transcriptional-regulator-of.

Jang:2023:HAS

[JMY^{+}23] Eunhong Jang, Yeojin Moon, So Young Yoon, Joyce Anne R. Diaz, Miriam Lee, Naho Ko, Jongseo Park, Soo Hyun Eom, Changwook Lee, and Youngsoo Jun. Human atlastins are sufficient to drive the fusion of liposomes with a physiological lipid composition. Journal of Cell Biology, 222(4):e202109090, April 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/4/
e202109090/213863/Human-atlastins-are-sufficient-to-
drive-the-fusion.

Junyent:2021:PSR

[JRGH21] Sergi Junyent, Joshua Reeves, Eileen Gentleman, and Shukry J. Habib. Pluripotency state regulates cytoneme selectivity and self-organization of embryonic stem cells. Journal of Cell Biology, 220(4):??, April 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/4/e202005095/211816/Pluripotency-
state-regulates-cytoneme-selectivity.

Jaffray:2023:PVS

[JTM^{+}23] Ellis G. Jaffray, Michael H. Tatham, Barbara Mojsa, Magda Liczmanska, Alejandro Rojas-Fernandez, Yili Yin, Graeme Ball, and Ronald T. Hay. The p97/VCP segregase is essential for arsenic-induced degradation of PML and PML-RARA. Journal of Cell Biology, 222(4):e202201027, April 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/4/
e202201027/213910/The-p97-VCP-segregase-is-essential-
for-arsenic.

Jia:2022:SGM

[JWB^{+}22] Jingyue Jia, Fulong Wang, Zambarlal Bhujabal, Ryan Peters, Michal Mudd, Thabata Duque, Lee Allers, Ruheena Javed, Michelle Salemi, Christian Behrends, Brett Phinney, Terje Johansen, and Vojo Deretic. Stress granules and mTOR are regulated by membrane atg8ylation during lysosomal damage. Journal of Cell Biology, 221(11):e202207091, November 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/221/11/e202207091/213514/Stress-granules-and-
mTOR-are-regulated-by-membrane.

Kulkarni:2021:SAC

[KAH^{+}21] Vineet Vinay Kulkarni, Anip Anand, Jessica Brandt Herr, Christina Miranda, Maria Chalokh Vogel, and Sandra Maday. Synaptic activity controls autophagic vacuole motility and function in dendrites. Journal of Cell Biology, 220(6):e202002084, June 7, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/6/e202002084/211926/Synaptic-activity-controls-autophagic-
vacuole.

Kono:2022:NLC

[KAS^{+}22] Yohei Kono, Stephen A. Adam, Yuko Sato, Karen L. Reddy, Yixian Zheng, Ohad Medalia, Robert D. Goldman, Hiroshi Kimura, and Takeshi Shimi. Nucleoplasmic lamin C rapidly accumulates at sites of nuclear envelope rupture with BAF and cGAS. Journal of Cell Biology, 221(12):e202201024, December 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/221/12/e202201024/213616/Nucleoplasmic-lamin-
C-rapidly-accumulates-at-sites.

Khan:2021:SDS

[KB21] Danish Khan and Onn Brandman. Sis1 delivers the State of the Union. Journal of Cell Biology, 220(1):e202011093, January 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/1/
e202011093/211606/Sis1-delivers-the-State-of-the-UnionSis1-
delivers.

Koestel:2022:PSB

[KB22] Jérôme Koestel and Henri Batoko. A plant-specific bridging adaptor for amphisome biogenesis. Journal of Cell Biology, 221(12):e202210011, December 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/221/12/e202210011/213679/A-
plant-specific-bridging-adaptor-for-amphisome.

Kyumurkov:2023:FTT

[KBB^{+}23] Alexander Kyumurkov, Anne-Pascale Bouin, Mathieu Boissan, Sandra Manet, Francesco Baschieri, Mathilde Proponnet-Guerault, Martial Balland, Olivier Destaing, Myriam Régent-Kloeckner, Claire Calmel, Alice Nicolas, François Waharte, Philippe Chavrier, Guillaume Montagnac, Emmanuelle Planus, and Corinne Albiges-Rizo. Force tuning through regulation of clathrin-dependent integrin endocytosis. Journal of Cell Biology, 222(1):e202004025, January 2, 2023. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/222/1/
e202004025/213549/Force-tuning-through-regulation-of-
clathrin.

King:2022:CCR

[KBH^{+}22] Zayna T. King, Mitchell T. Butler, Max A. Hockenberry, Bhagawat C. Subramanian, Priscila F. Siesser, David M. Graham, Wesley R. Legant, and James E. Bear. Coro1B and Coro1C regulate lamellipodia dynamics and cell motility by tuning branched actin turnover. Journal of Cell Biology, 221(8):e202111126, August 1, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/8/e202111126/213258/Coro1B-and-Coro1C-regulate-lamellipodia-
dynamics.

Kelly:2021:HRA

[KBN^{+}21] Carolyn M. Kelly, Laura J. Byrnes, Niharika Neela, Holger Sondermann, and John P. O’Donnell. The hypervariable region of atlastin-1 is a site for intrinsic and extrinsic regulation. Journal of Cell Biology, 220(11):e202104128, November 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/11/e202104128/212648/The-hypervariable-region-of-
atlastin-1-is-a-site.

Kulasekaran:2021:ARC

[KCP^{+}21] Gopinath Kulasekaran, Mathilde Chaineau, Valerio Emilio Crescenzo Piscopo, Federica Verginelli, Maryam Fotouhi, Martine Girard, Yeman Tang, Rola Dali, Rita Lo, Stefano Stifani, and Peter S. McPherson. An Arf/Rab cascade controls the growth and invasiveness of glioblastoma. Journal of Cell Biology, 220(2):e202004229, February 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/220/2/e202004229/211682/An-
Arf-Rab-cascade-controls-the-growth-and.

Kozar-Gillan:2023:LAI

[KGVK^{+}23] Nina Kozar-Gillan, Atanaska Velichkova, George Kanatouris, Yael Eshed-Eisenbach, Gavin Steel, Martine Jaegle, Eerik Aunin, Elior Peles, Carole Torsney, and Dies N. Meijer. LGI3/2– ADAM23 interactions cluster Kv1 channels in myelinated axons to regulate refractory period. Journal of Cell Biology, 222(4):e202211031, April 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/222/4/e202211031/213903/LGI3-
2-ADAM23-interactions-cluster-Kv1-channels-in.

Kors:2022:RPC

[KHB^{+}22] Suzan Kors, Christian Hacker, Chloe Bolton, Renate Maier, Lena Reimann, Emily J. A. Kitchener, Bettina Warscheid, Joseph L. Costello, and Michael Schrader. Regulating peroxisome–ER contacts via the ACBD5-VAPB tether by FFAT motif phosphorylation and GSK3. Journal of Cell Biology, 221 (3):e202003143, March 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/221/3/e202003143/212956/Regulating-peroxisome-
ER-contacts-via-the-ACBD5.

Koyama-Honda:2020:HSS

[KHFK^{+}20] Ikuko Koyama-Honda, Takahiro K. Fujiwara, Rinshi S. Kasai, Kenichi G. N. Suzuki, Eriko Kajikawa, Hisae Tsuboi, Taka A. Tsunoyama, and Akihiro Kusumi. High-speed single-molecule imaging reveals signal transduction by induced transbilayer raft phases. Journal of Cell Biology, 219(12):e202006125, December 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/12/e202006125/211461/High-speed-single-molecule-
imaging-reveals-signal.

Kucharski:2022:SCP

[KHV^{+}22] Thomas J. Kucharski, Rufus Hards, Sarah E. Vandal, Maria Alba Abad, A. Arockia Jeyaprakash, Edward Kaye, Aymen al Rawi, Tony Ly, Kristina M. Godek, Scott A. Gerber, and Duane A. Compton. Small changes in phospho-occupancy at the kinetochore–microtubule interface drive mitotic fidelity. Journal of Cell Biology, 221(9):e202107107, September 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/9/
e202107107/213364/Small-changes-in-phospho-occupancy-
at-the.

King:2021:DRI

[Kin21] Jason S. King. Dynamic Rac1 inhibition by CYRI helps cells drink, but stops them from driving. Journal of Cell Biology, 220(9):e202108041, September 6, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/9/
e202108041/212561/Dynamic-Rac1-inhibition-by-CYRI-helps-
cells-drink.

Khalil:2020:CII

[KIV^{+}20] Antoine A. Khalil, Olga Ilina, Angela Vasaturo, Jan-Hendrik Venhuizen, Manon Vullings, Victor Venhuizen, Ab Bilos, Carl G. Figdor, Paul N. Span, and Peter Friedl. Collective invasion induced by an autocrine purinergic loop through connexin-43 hemichannels. Journal of Cell Biology, 219(10):e201911120, October 5, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/10/e201911120/152009/Collective-invasion-induced-
by-an-autocrine.

Kournoutis:2023:LCL

[KJ23] Athanasios Kournoutis and Terje Johansen. LC3B is a cofactor for LMX1B-mediated transcription of autophagy genes in dopaminergic neurons. Journal of Cell Biology, 222 (5):e202303008, May 1, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/222/5/e202303008/214011/LC3B-
is-a-cofactor-for-LMX1B-mediated.

Kim:2021:ARE

[KKN^{+}21] Joohyung Kim, Sungdae Kim, Minyeop Nahm, Tsai-Ning Li, Hsin-Chieh Lin, Yeongjin David Kim, Jihye Lee, Chi-Kuang Yao, and Seungbok Lee. ALS2 regulates endosomal trafficking, postsynaptic development, and neuronal survival. Journal of Cell Biology, 220(5):e202007112, May 3, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/220/5/e202007112/211859/ALS2-
regulates-endosomal-trafficking-postsynaptic.

Kumari:2021:PPB

[KKP^{+}21] Amrita Kumari, Chandan Kumar, Rajaiah Pergu, Megha Kumar, Sagar P. Mahale, Neeraj Wasnik, and Sivaram V. S. Mylavarapu. Phosphorylation and Pin1 binding to the LIC1 subunit selectively regulate mitotic dynein functions. Journal of Cell Biology, 220(12):e202005184, December 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/220/12/e202005184/212736/Phosphorylation-and-
Pin1-binding-to-the-LIC1.

Katoku-Kikyo:2021:PPI

[KKPH^{+}21] Nobuko Katoku-Kikyo, Ellen Paatela, Daniel L. Houtz, Britney Lee, Dane Munson, Xuerui Wang, Mohammed Hussein, Jasmeet Bhatia, Seunghyun Lim, Ce Yuan, Yoko Asakura, Atsushi Asakura, and Nobuaki Kikyo. Per1/Per2–Igf2 axis–mediated circadian regulation of myogenic differentiation. Journal of Cell Biology, 220(7):e202101057, July 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/220/7/e202101057/212164/Per1-
Per2-Igf2-axis-mediated-circadian-regulation.

Kalita:2022:KEC

[KKZ^{+}22] Joanna Kalita, Larisa E. Kapinos, Tiantian Zheng, Chantal Rencurel, Anton Zilman, and Roderick Y. H. Lim. Karyopherin enrichment and compensation fortifies the nuclear pore complex against nucleocytoplasmic leakage. Journal of Cell Biology, 221 (3):e202108107, March 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/221/3/e202108107/212986/Karyopherin-
enrichment-and-compensation-fortifies.

Krey:2022:AOT

[KLB^{+}22] Jocelyn F. Krey, Chang Liu, Inna A. Belyantseva, Michael Bateschell, Rachel A. Dumont, Jennifer Goldsmith, Paroma Chatterjee, Rachel S. Morrill, Lev M. Fedorov, Sarah Foster, Jinkyung Kim, Alfred L. Nuttall, Sherri M. Jones, Dongseok Choi, Thomas B. Friedman, Anthony J. Ricci, Bo Zhao, and Peter G. Barr-Gillespie. ANKRD24 organizes TRIOBP to reinforce stereocilia insertion points. Journal of Cell Biology, 221(4):e202109134, April 4, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/4/
e202109134/213014/ANKRD24-organizes-TRIOBP-to-reinforce-
stereocilia.

Kesarwani:2020:GEL

[KLC^{+}20] Shubham Kesarwani, Prakash Lama, Anchal Chandra, P. Purushotam Reddy, A. S. Jijumon, Satish Bodakuntla, Balaji M. Rao, Carsten Janke, Ranabir Das, and Minhajuddin Sirajuddin. Genetically encoded live-cell sensor for tyrosinated microtubules. Journal of Cell Biology, 219(10): e201912107, October 5, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/10/e201912107/152071/Genetically-
encoded-live-cell-sensor-for.

Kreider-Letterman:2023:ARS

[KLCM^{+}23] Gabriel Kreider-Letterman, Abel Castillo, Eike K. Mahlandt, Joachim Goedhart, Agustin Rabino, Silvia Goicoechea, and Rafael Garcia-Mata. ARHGAP17 regulates the spatiotemporal activity of cdc42 at invadopodia. Journal of Cell Biology, 222(2): e202207020, February 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/222/2/e202207020/213782/ARHGAP17-regulates-
the-spatiotemporal-activity-of.

Ko:2020:SAD

[KMD20] Kwang Woo Ko, Jeffrey Milbrandt, and Aaron DiAntonio. SARM1 acts downstream of neuroinflammatory and necroptotic signaling to induce axon degeneration. Journal of Cell Biology, 219(8):e201912047, August 3, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/8/

e201912047/151915/SARM1-acts-downstream-of-neuroinflammatory-
and.

Klucnika:2023:RDR

[KMJ^{+}23] Anna Klucnika, Peiqiang Mu, Jan Jezek, Matthew McCormack, Ying Di, Charles R. Bradshaw, and Hansong Ma. REC drives recombination to repair double-strand breaks in animal mtDNA. Journal of Cell Biology, 222(1):e202201137, January 2, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/1/
e202201137/213677/REC-drives-recombination-to-repair-
double-strand.

Kusumaatmaja:2021:IWM

[KMK21] Halim Kusumaatmaja, Alexander I. May, and Roland L. Knorr. Intracellular wetting mediates contacts between liquid compartments and membrane-bound organelles. Journal of Cell Biology, 220(10):e202103175, October 4, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/10/
e202103175/212595/Intracellular-wetting-mediates-contacts-
between.

Kornakov:2020:EKC

[KMW20] Nikolay Kornakov, Bastian Möllers, and Stefan Westermann. The EB1–kinesin-14 complex is required for efficient metaphase spindle assembly and kinetochore bi-orientation. Journal of Cell Biology, 219(12):e202003072, December 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/12/e202003072/211447/The-EB1-Kinesin-14-complex-
is-required-for.

Karasu:2022:CSP

[KNA^{+}22] Onur Rojhat Karasu, Annett Neuner, Enrico Salvatore Atorino, Gislene Pereira, and Elmar Schiebel. The central scaffold protein CEP350 coordinates centriole length, stability, and maturation. Journal of Cell Biology, 221(12):e202203081, December 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/221/12/e202203081/213625/The-central-
scaffold-protein-CEP350-coordinates.

Kunii:2021:SDC

[KNiY^{+}21] Masataka Kunii, Yuria Noguchi, Shin ichiro Yoshimura, Satoshi Kanda, Tomohiko Iwano, Erda Avriyanti, Nur Atik, Takashi Sato, Ken Sato, Masaharu Ogawa, and Akihiro Harada. SNAP23 deficiency causes severe brain dysplasia through the loss of radial glial cell polarity. Journal of Cell Biology, 220(1):e201910080, January 4, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/1/
e201910080/211605/SNAP23-deficiency-causes-severe-brain-
dysplasia.

Kohler:2021:NEL

[Köh21] Alwin Köhler. Nuclear envelope lipids request border surveillance. Journal of Cell Biology, 220(3):e202101164, March 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
3/e202101164/211788/Nuclear-envelope-lipids-request-
border.

Kedersha:2016:GCU

[KPA^{+}16] Nancy Kedersha, Marc D. Panas, Christopher A. Achorn, Shawn Lyons, Sarah Tisdale, Tyler Hickman, Marshall Thomas, Judy Lieberman, Gerald M. McInerney, Pavel Ivanov, and Paul Anderson. G3BP–Caprin1–USP10 complexes mediate stress granule condensation and associate with 40S subunits. Journal of Cell Biology, 212(7):845–??, March 2016. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL

http://jcb.rupress.org/content/212/7/845. See correction

[KPA^{+}20].

Kedersha:2020:CGC

[KPA^{+}20] Nancy Kedersha, Marc D. Panas, Christopher A. Achorn, Shawn Lyons, Sarah Tisdale, Tyler Hickman, Marshall Thomas, Judy Lieberman, Gerald M. McInerney, Pavel Ivanov, and Paul Anderson. Correction: G3BP Caprin1 USP10 complexes mediate stress granule condensation and associate with 40S subunits. Journal of Cell Biology, 219(1):e20150802809202019c, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). See [KPA^{+}16].

Killackey:2020:MPH

[KPG20] Samuel A. Killackey, Dana J. Philpott, and Stephen E. Girardin. Mitophagy pathways in health and disease. Journal of Cell Biology, 219(11):e202004029, November 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/11/
e202004029/152083/Mitophagy-pathways-in-health-and-diseaseMitophagy.

Keen:2022:EIF

[KPM^{+}22] Adam N. Keen, Luke A. Payne, Vedanta Mehta, Alistair Rice, Lisa J. Simpson, Kar Lai Pang, Armando del Rio Hernandez, John S. Reader, and Ellie Tzima. Eukaryotic initiation factor 6 regulates mechanical responses in endothelial cells. Journal of Cell Biology, 221(2):e202005213, February 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/2/e202005213/212961/Eukaryotic-initiation-factor-
6-regulates.

Kletter:2022:VMR

[KRC^{+}22] Tobias Kletter, Sebastian Reusch, Tommaso Cavazza, Nils Dempewolf, Christian Tischer, and Simone Reber. Volumetric morphometry reveals spindle width as the best predictor of mammalian spindle scaling. Journal of Cell Biology, 221(1): e202106170, January 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/221/1/e202106170/212842/Volumetric-morphometry-
reveals-spindle-width-as.

Kopf:2020:MCC

[KRH^{+}20] Aglaja Kopf, Jörg Renkawitz, Robert Hauschild, Irute Girkontaite, Kerry Tedford, Jack Merrin, Oliver Thorn-Seshold, Dirk Trauner, Hans Häcker, Klaus-Dieter Fischer, Eva Kiermaier, and Michael Sixt. Microtubules control cellular shape and coherence in amoeboid migrating cells. Journal of Cell Biology, 219 (6):e201907154, June 1, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/6/e201907154/151745/Microtubules-
control-cellular-shape-and-coherence.

Kumar:2021:CCB

[KRHP^{+}21] Dhivya Kumar, Addison Rains, Vicente Herranz-Pérez, Quanlong Lu, Xiaoyu Shi, Danielle L. Swaney, Erica Stevenson, Nevan J. Krogan, Bo Huang, Christopher Westlake, Jose Manuel Garcia-Verdugo, Bradley K. Yoder, and Jeremy F. Reiter. A ciliopathy complex builds distal appendages to initiate ciliogenesis. Journal of Cell Biology, 220(9):e202011133, September 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/9/
e202011133/212460/A-ciliopathy-complex-builds-distal-
appendages-to.

Kesisova:2021:SGS

[KRS21] Ilona A. Kesisova, Benjamin P. Robinson, and Elias T. Spiliotis. A septin GTPase scaffold of dynein–dynactin motors triggers retrograde lysosome transport. Journal of Cell Biology, 220(2):e202005219, February 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/220/2/e202005219/211663/A-
septin-GTPase-scaffold-of-dynein-dynactin-motors.

Kanfer:2021:IBP

[KSM^{+}21a] Gil Kanfer, Shireen A. Sarraf, Yaakov Maman, Heather Baldwin, Eunice Dominguez-Martin, Kory R. Johnson, Michael E. Ward, Martin Kampmann, Jennifer Lippincott-Schwartz, and Richard J. Youle. Image-based pooled whole-genome CRISPRi screening for subcellular phenotypes. Journal of Cell Biology, 220(2):e202006180, February 1, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/2/
e202006180/211694/Image-based-pooled-whole-genome-CRISPRi-
screening.

Kindberg:2021:EER

[KSM^{+}21b] Abigail A. Kindberg, Vasudha Srivastava, Jonathon M. Muncie, Valerie M. Weaver, Zev J. Gartner, and Jeffrey O. Bush. EPH/EPHRIN regulates cellular organization by actomyosin contractility effects on cell contacts. Journal of Cell Biology, 220(6):e202005216, June 7, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:/

/rupress.org/jcb/article/220/6/e202005216/211938/EPH-
EPHRIN-regulates-cellular-organization-by.

Kawasaki:2022:PPC

[KSN^{+}22] Asami Kawasaki, Akiko Sakai, Hiroki Nakanishi, Junya Hasegawa, Tomohiko Taguchi, Junko Sasaki, Hiroyuki Arai, Takehiko Sasaki, Michihiro Igarashi, and Fubito Nakatsu. PI4P/PS countertransport by ORP10 at ER–endosome membrane contact sites regulates endosome fission. Journal of Cell Biology, 221(1):e202103141, January 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/221/1/e202103141/212876/PI4P-
PS-countertransport-by-ORP10-at-ER-endosome.

Kumbhar:2021:PAR

[KSP^{+}21] Ramhari Kumbhar, Anthony Sanchez, Jullian Perren, Fade Gong, David Corujo, Frank Medina, Sravan K. Devanathan, Blerta Xhemalce, Andreas Matouschek, Marcus Buschbeck, Bethany A. Buck-Koehntop, and Kyle M. Miller. Poly(ADPribose) binding and macroH2A mediate recruitment and functions of KDM5A at DNA lesions. Journal of Cell Biology, 220(7):e202006149, July 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/220/7/e202006149/212163/Poly-
ADP-ribose-binding-and-macroH2A-mediate.

Khalaj:2020:DFP

[KSS^{+}20a] Anna J. Khalaj, Fredrik H. Sterky, Alessandra Sclip, Jochen Schwenk, Axel T. Brunger, Bernd Fakler, and Thomas C. Südhof. Deorphanizing FAM19A proteins as pan-neurexin ligands with an unusual biosynthetic binding mechanism. Journal of Cell Biology, 219(9):e202004164, September 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/
9/e202004164/151974/Deorphanizing-FAM19A-proteins-as-
pan-neurexin.

Kong:2020:CPM

[KSS^{+}20b] Dong Kong, Natalie Sahabandu, Catherine Sullenberger, Alejandra Vásquez-Limeta, Delgermaa Luvsanjav, Kimberly Lukasik, and Jadranka Loncarek. Correction: Prolonged mitosis results in structurally aberrant and over-elongated centrioles. Journal of Cell Biology, 219(7):??, July 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/219/7/jcb.20191001906192020c/

151881/Correction-Prolonged-mitosis-results-in. See

[KSS^{+}20c].

Kong:2020:PMR

[KSS^{+}20c] Dong Kong, Natalie Sahabandu, Catherine Sullenberger, Alejandra Vásquez-Limeta, Delgermaa Luvsanjav, Kimberly Lukasik, and Jadranka Loncarek. Prolonged mitosis results in structurally aberrant and over-elongated centrioles. Journal of Cell Biology, 219(6):e201910019, June 1, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/6/e201910019/151638/Prolonged-mitosis-results-in-

structurally-aberrant. See correction [KSS^{+}20b].

Kittisopikul:2021:CAR

[KST^{+}21] Mark Kittisopikul, Takeshi Shimi, Meltem Tatli, Joseph Riley Tran, Yixian Zheng, Ohad Medalia, Khuloud Jaqaman, Stephen A. Adam, and Robert D. Goldman. Computational analyses reveal spatial relationships between nuclear pore complexes and specific lamins. Journal of Cell Biology, 220(4):??, April 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/4/e202007082/211786/Computational-analyses-reveal-
spatial.

Kummer:2022:JTT

[KST^{+}22] Daniel Kummer, Tim Steinbacher, Sonja Thölmann, Mariel Flavia Schwietzer, Christian Hartmann, Simone Horenkamp, Sabrina Demuth, Swetha S. D. Peddibhotla, Frauke Brinkmann, Björn Kemper, Jürgen Schnekenburger, Matthias Brandt, Timo Betz, Ivan Liashkovich, Ivan U. Kouzel, Victor Shahin, Nathalie Corvaia, Klemens Rottner, Katsiaryna Tarbashevich, Erez Raz, Lilo Greune, M. Alexander Schmidt, Volker Gerke, and Klaus Ebnet. A JAM-A-tetraspanin-α vβ5 integrin complex regulates contact inhibition of locomotion. Journal of Cell Biology, 221(4):e202105147, April 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/221/4/e202105147/213070/A-
JAM-A-tetraspanin-v-5-integrin-complex-regulates.

Kela:2023:TDR

[KST^{+}23] Shiri Avivi Kela, Kriti Sethi, Pei Yi Tan, Danesha Suresh, Hui Ting Ong, Perla G. Castaneda, Mustafi R. Amin, Tal Laviv, Erin J. Cram, Jan Faix, and Ronen Zaidel-Bar. Tension-dependent RHGF-1 recruitment to stress fibers drives robust spermathecal tissue contraction. Journal of Cell Biology, 222(2):e202203105, February 6, 2023. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/222/2/
e202203105/213784/Tension-dependent-RHGF-1-recruitment-
to-stress.

Kang:2022:DPS

[KSWC22] Kexin Kang, Qiaoni Shi, Xu Wang, and Ye-Guang Chen. Dishevelled phase separation promotes Wnt signalosome assembly and destruction complex disassembly. Journal of Cell Biology, 221(12):e202205069, December 5, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/12/
e202205069/213667/Dishevelled-phase-separation-promotes-
Wnt.

Kinoshita:2022:LEI

[KTT^{+}22] Kazuhisa Kinoshita, Yuko Tsubota, Shoji Tane, Yuuki Aizawa, Ryota Sakata, Kozo Takeuchi, Keishi Shintomi, Tomoko Nishiyama, and Tatsuya Hirano. A loop extrusion–independent mechanism contributes to condensin I-mediated chromosome shaping. Journal of Cell Biology, 221(3):e202109016, March 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/3/e202109016/212966/A-loop-extrusion-independent-
mechanism-contributes.

Kim:2020:GEA

[KVG^{+}20] Jiah Kim, Neha Chivukula Venkata, Gabriela Andrea Hernandez Gonzalez, Nimish Khanna, and Andrew S. Belmont. Gene expression amplification by nuclear speckle association. Journal of Cell Biology, 219(1):e201904046, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Krikken:2020:PRI

[KWdB^{+}20] Arjen M. Krikken, Huala Wu, Rinse de Boer, Damien P. Devos, Tim P. Levine, and Ida J. van der Klei. Peroxisome retention involves Inp1-dependent peroxisome–plasma membrane contact sites in yeast. Journal of Cell Biology, 219(10): e201906023, October 5, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/10/e201906023/152028/Peroxisome-
retention-involves-Inp1-dependent.

Kramer:2023:DPS

[KWGR23] Rafael Krämer, Neele Wolterhoff, Milos Galic, and Sebastian Rumpf. Developmental pruning of sensory neurites by mechanical tearing in Drosophila. Journal of Cell Biology, 222(3):e202205004, March 6, 2023. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/222/3/
e202205004/213805/Developmental-pruning-of-sensory-neurites-
by.

King:2023:MNP

[KWV^{+}23] Grant A. King, Rahel Wettstein, Joseph M. Varberg, Keerthana Chetlapalli, Madison E. Walsh, Ludovic C. J. Gillet, Claudia Hernández-Armenta, Pedro Beltrao, Ruedi Aebersold, Sue L. Jaspersen, Joao Matos, and Elçin Ünal. Meiotic nuclear pore complex remodeling provides key insights into nuclear basket organization. Journal of Cell Biology, 222(2):e202204039, February 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
222/2/e202204039/213751/Meiotic-nuclear-pore-complex-
remodeling-provides.

Krishnan:2022:ARC

[KYR^{+}22] Badri Krishnan, Takaaki Yasuhara, Purva Rumde, Marcello Stanzione, Chenyue Lu, Hanjun Lee, Michael S. Lawrence, Lee Zou, Linda T. Nieman, Ioannis Sanidas, and Nicholas J. Dyson. Active RB causes visible changes in nuclear organization. Journal of Cell Biology, 221(3):e202102144, March 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/
3/e202102144/212957/Active-RB-causes-visible-changes-
in-nuclear.

Leithner:2021:DCA

[LAH^{+}21] Alexander Leithner, Lukas M. Altenburger, Robert Hauschild, Frank P. Assen, Klemens Rottner, Theresia E. B. Stradal, Alba Diz-Muñoz, Jens V. Stein, and Michael Sixt. Dendritic cell actin dynamics control contact duration and priming efficiency at the immunological synapse. Journal of Cell Biology, 220(4): ??, April 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/220/4/e202006081/211749/Dendritic-cell-actin-
dynamics-control-contact.

Lebek:2020:ASM

[LC20] Nadine M. Lebek and Kenneth G. Campellone. Adding SNX to the mix: SNX9 drives filopodia biogenesis. Journal of Cell Biology, 219(4):e202002086, April 6, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/4/
e202002086/151588/Adding-SNX-to-the-mix-SNX9-drives-
filopodia.

Lassetter:2023:GTA

[LCB^{+}23] Alexandria P. Lassetter, Megan M. Corty, Romina Barria, Amy E. Sheehan, Jo Q. Hill, Sue A. Aicher, A. Nicole Fox, and Marc R. Freeman. Glial TGFβ activity promotes neuron survival in peripheral nerves. Journal of Cell Biology, 222(1):e202111053, January 2, 2023. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/222/1/
e202111053/213698/Glial-TGF-activity-promotes-neuron-
survival-in.

Laidlaw:2022:RCS

[LCM22] Kamilla M. E. Laidlaw, Grant Calder, and Chris MacDonald. Recycling of cell surface membrane proteins from yeast endosomes is regulated by ubiquitinated ist1. Journal of Cell Biology, 221(11):e202109137, November 7, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/11/
e202109137/213481/Recycling-of-cell-surface-membrane-
proteins-from.

Lu:2020:CGR

[LD20] Michelle Seiko Lu and David G. Drubin. Cdc42 GTPase regulates ESCRTs in nuclear envelope sealing and ER remodeling. Journal of Cell Biology, 219(8):e201910119, August 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/
8/e201910119/151867/Cdc42-GTPase-regulates-ESCRTs-in-
nuclear-envelope.

Lee:2021:PAH

[LD21] Yeojin Lee and Lei Ding. A polarized anchor for hematopoietic stem cells: Synapse between stem cells and their niche? Journal of Cell Biology, 220(10):e202108031, October 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/220/10/e202108031/212641/A-polarized-anchor-
for-hematopoietic-stem-cells.

Leguay:2022:IMD

[LDE^{+}22] Kévin Leguay, Barbara Decelle, Islam E. Elkholi, Michel Bouvier, Jean-François Côté, and Sébastien Carréno. Interphase microtubule disassembly is a signaling cue that drives cell rounding at mitotic entry. Journal of Cell Biology, 221(6):e202109065, June 6, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/6/e202109065/213183/Interphase-microtubule-disassembly-
is-a-signaling.

Linklater:2021:RCC

[LDH^{+}21] Erik S. Linklater, Emily D. Duncan, Ke-Jun Han, Algirdas Kaupinis, Mindaugas Valius, Traci R. Lyons, and Rytis Prekeris. Rab40–Cullin5 complex regulates EPLIN and actin cytoskeleton dynamics during cell migration. Journal of Cell Biology, 220(7):e202008060, July 5, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/7/
e202008060/212111/Rab40-Cullin5-complex-regulates-EPLIN-
and-actin.

Leterrier:2020:GGA

[Let20] Christophe Leterrier. GABA in, garbage out: AIS-located proteasomes regulate the developmental GABA switch. Journal of Cell Biology, 219(2):e201912006, February 3, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/2/
e201912006/133590/GABA-in-garbage-out-AIS-located-proteasomes.

Lancaster:2021:PRR

[LFD^{+}21] Charlene E. Lancaster, Aaron Fountain, Roaya M. Dayam, Elliott Somerville, Javal Sheth, Vanessa Jacobelli, Alex Somerville, Mauricio R. Terebiznik, and Roberto J. Botelho. Phagosome resolution regenerates lysosomes and maintains the degradative capacity in phagocytes. Journal of Cell Biology, 220(9): e202005072, September 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/9/e202005072/212440/Phagosome-resolution-
regenerates-lysosomes-and.

Li:2022:VRI

[LFF^{+}22] Min Li, Fengping Feng, Han Feng, Pengkai Hu, Yanhong Xue, Tao Xu, and Eli Song. VAMP4 regulates insulin levels by targeting secretory granules to lysosomes. Journal of Cell Biology, 221(10):e202110164, October 3, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/10/
e202110164/213439/VAMP4-regulates-insulin-levels-by-
targeting.

Lonic:2021:PPF

[LGB^{+}21] Ana Lonic, Freya Gehling, Leila Belle, Xiaochun Li, Nicole L. Schieber, Elizabeth V. Nguyen, Gregory J. Goodall, Robert G. Parton, Roger J. Daly, and Yeesim Khew-Goodall. Phosphorylation of PKCδ by FER tips the balance from EGFR degradation to recycling. Journal of Cell Biology, 220(2):e201902073, February 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/220/2/e201902073/211661/Phosphorylation-
of-PKC-by-FER-tips-the-balance.

Li:2023:VAT

[LGL^{+}23] Terytty Yang Li, Arwen W. Gao, Xiaoxu Li, Hao Li, Yasmine J. Liu, Amelia Lalou, Nagammal Neelagandan, Felix Naef, Kristina Schoonjans, and Johan Auwerx. V-ATPase/TORC1-mediated ATFS-1 translation directs mitochondrial UPR activation in C. elegans. Journal of Cell Biology, 222(1):e202205045, January 2, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
222/1/e202205045/213623/V-ATPase-TORC1-mediated-ATFS-
1-translation-directs.

Logue:2022:NII

[LGS22] Susan E. Logue, Adrienne M. Gorman, and Afshin Samali. New insights into IRE1α activation and function in anti-tumor immunity. Journal of Cell Biology, 221(6):e202205019, June 6, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/6/
e202205019/213209/New-insights-into-IRE1-activation-
and-function-in.

Laiman:2023:GPD

[LHL^{+}23] Jessica Laiman, Yen-Jung Hsu, Julie Loh, Wei-Chun Tang, Mei-Chun Chuang, Hui-Kang Liu, Wei-Shun Yang, Bi-Chang Chen, Lee-Ming Chuang, Yi-Cheng Chang, and Ya-Wen Liu. GSK3α phosphorylates dynamin-2 to promote GLUT4 endocytosis in muscle cells. Journal of Cell Biology, 222(2):e202102119, February 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
222/2/e202102119/213725/GSK3-phosphorylates-dynamin-
2-to-promote-GLUT4.

Lu:2022:CSL

[LHS^{+}22] Albert Lu, Frank Hsieh, Bikal R. Sharma, Sydney R. Vaughn, Carlos Enrich, and Suzanne R. Pfeffer. CRISPR screens for lipid regulators reveal a role for ER-bound SNX13 in lysosomal cholesterol export. Journal of Cell Biology, 221(2):e202105060, February 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/2/e202105060/212937/CRISPR-screens-for-lipid-regulators-
reveal-a-role.

Lee:2021:RII

[LJJ^{+}21] Yujin Lee, Yoonji Jung, Dae-Eun Jeong, Wooseon Hwang, Seokjin Ham, Hae-Eun H. Park, Sujeong Kwon, Jasmine M. Ashraf, Coleen T. Murphy, and Seung-Jae V. Lee. Reduced insulin/IGF1 signaling prevents immune aging via ZIP-10/bZIP-mediated feedforward loop. Journal of Cell Biology, 220(5): e202006174, May 3, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/220/5/e202006174/211856/Reduced-insulin-
IGF1-signaling-prevents-immune.

Liu:2022:LTC

[LJT^{+}22] Kuo Liu, Hengwei Jin, Muxue Tang, Shaohua Zhang, Xueying Tian, Mingjun Zhang, Ximeng Han, Xiuxiu Liu, Juan Tang, Wenjuan Pu, Yan Li, Lingjuan He, Zhongzhou Yang, Kathy O. Lui, and Bin Zhou. Lineage tracing clarifies the cellular origin of tissue-resident macrophages in the developing heart. Journal of Cell Biology, 221(6): e202108093, June 6, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/221/6/e202108093/213182/Lineage-tracing-
clarifies-the-cellular-origin-of.

Limar:2023:YSB

[LKMM^{+}23] Sergej Limar, Carolin Körner, Fernando Martínez-Montañés, Viktoriya G. Stancheva, Verena N. Wolf, Stefan Walter, Elizabeth A. Miller, Christer S. Ejsing, Vanesa Viviana Galassi, and Florian Fröhlich. Yeast Svf1 binds ceramides and contributes to sphingolipid metabolism at the ER cis-Golgi interface. Journal of Cell Biology, 222(5):e202109162, May 1, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/5/
e202109162/213932/Yeast-Svf1-binds-ceramides-and-contributes-
to.

Larouche:2021:SCG

[LKW^{+}21] Myreille Larouche, David Kachaner, Peng Wang, Karine Normandin, Damien Garrido, Changfu Yao, Maxime Cormier, Kristen M. Johansen, Jørgen Johansen, and Vincent Archambault. Spatiotemporal coordination of Greatwall–Endos–PP2A promotes mitotic progression. Journal of Cell Biology, 220(6): e202008145, June 7, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/220/6/e202008145/211965/Spatiotemporal-coordination-
of-Greatwall-Endos.

Lee:2022:DMA

[LL22] Jongmin Lee and David E. Levin. Differential metabolism of arsenicals regulates Fps1-mediated arsenite transport. Journal of Cell Biology, 221(3):e202109034, March 7, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/3/
e202109034/212996/Differential-metabolism-of-arsenicals-
regulates.

Levone:2021:FDL

[LLA^{+}21] Brunno R. Levone, Silvia C. Lenzken, Marco Antonaci, Andreas Maiser, Alexander Rapp, Francesca Conte, Stefan Reber, Jonas Mechtersheimer, Antonella E. Ronchi, Oliver Mühlemann, Heinrich Leonhardt, M. Cristina Cardoso, Marc-David Ruepp, and Silvia M. L. Barabino. FUS-dependent liquid–liquid phase separation is important for DNA repair initiation. Journal of Cell Biology, 220(5):e202008030, May 3, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:/

/rupress.org/jcb/article/220/5/e202008030/211877/FUS-
dependent-liquid-liquid-phase-separation-is.

Larocque:2020:TPD

[LLBC^{+}20] Gabrielle Larocque, Penelope J. La-Borde, Nicholas I. Clarke, Nicholas J. Carter, and Stephen J. Royle. Tumor protein D54 defines a new class of intracellular transport vesicles. Journal of Cell Biology, 219(1):e201812044, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Lee:2020:EMP

[LLC^{+}20] Min Lee, Yen-Chen Liu, Chen Chen, Chi-Huan Lu, Shao-Tzu Lu, Tzyy-Nan Huang, Meng-Tsung Hsu, Yi-Ping Hsueh, and Pei-Lin Cheng. Ecm29-mediated proteasomal distribution modulates excitatory GABA responses in the developing brain. Journal of Cell Biology, 219(2):e201903033, February 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/2/
e201903033/133566/Ecm29-mediated-proteasomal-distribution-
modulates.

Li:2021:NDC

[LLK^{+}21] Lei Li, Haowen Liu, Mia Krout, Janet E. Richmond, Yu Wang, Jihong Bai, Saroja Weeratunga, Brett M. Collins, Donovan Ventimiglia, Yi Yu, Jingyao Xia, Jing Tang, Jie Liu, and Zhitao Hu. A novel dual Ca^{2+} sensor system regulates Ca^{2+}- dependent neurotransmitter release. Journal of Cell Biology, 220(4):??, April 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/220/4/e202008121/211787/A-novel-dual-Ca2-
sensor-system-regulates-Ca2.

Loganathan:2022:RBR

[LLK^{+}22] Rajprasad Loganathan, Daniel C. Levings, Ji Hoon Kim, Michael B. Wells, Hannah Chiu, Yifan Wu, Matthew Slattery, and Deborah J. Andrew. Ribbon boosts ribosomal protein gene expression to coordinate organ form and function. Journal of Cell Biology, 221(4):e202110073, April 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/
4/e202110073/213030/Ribbon-boosts-ribosomal-protein-
gene-expression-to.

Li:2020:CER

[LLLR20] PeiQi Li, Joshua Aaron Lees, C. Patrick Lusk, and Karin M. Reinisch. Cryo-EM reconstruction of a VPS13 fragment reveals a long groove to channel lipids between membranes. Journal of Cell Biology, 219(5):e202001161, May 4, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/5/
e202001161/151565/Cryo-EM-reconstruction-of-a-VPS13-
fragment-reveals.

Liu:2020:CPO

[LLW^{+}20] Peiwei Liu, Xiaochu Lou, Jenna L. Wingfield, Jianfeng Lin, Daniela Nicastro, and Karl Lechtreck. Chlamydomonas PKD2 organizes mastigonemes, hair-like glycoprotein polymers on cilia. Journal of Cell Biology, 219(6):e202001122, June 1, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/6/e202001122/151720/Chlamydomonas-PKD2-organizes-
mastigonemes-hair.

Liu:2021:CBT

[LLW^{+}21] Min Liu, Aiguo Liu, Jie Wang, Yansong Zhang, Yajuan Li, Ying Su, and Alan Jian Zhu. Competition between two phosphatases fine-tunes Hedgehog signaling. Journal of Cell Biology, 220(2): e202010078, February 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/2/e202010078/211641/Competition-
between-two-phosphatases-fine-tunes.

Lo:2021:CIB

[LLX^{+}21] Harriet P. Lo, Ye-Wheen Lim, Zherui Xiong, Nick Martel, Charles Ferguson, Nicholas Ariotti, Jean Giacomotto, James Rae, Matthias Floetenmeyer, Shayli Varasteh Moradi, Ya Gao, Vikas A. Tillu, Di Xia, Huang Wang, Samira Rahnama, Susan J. Nixon, Michele Bastiani, Ryan D. Day, Kelly A. Smith, Nathan J. Palpant, Wayne A. Johnston, Kirill Alexandrov, Brett M. Collins, Thomas E. Hall, and Robert G. Parton. Cavin4 interacts with Bin1 to promote T-tubule formation and stability in developing skeletal muscle. Journal of Cell Biology, 220(12):e201905065, December 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/12/e201905065/212693/Cavin4-interacts-with-Bin1-
to-promote-T-tubule.

Liu:2022:WSE

[LLY22] Nan Liu, Kai Liu, and Chonglin Yang. WDR91 specifies the endosomal retrieval subdomain for retromer-dependent recycling. Journal of Cell Biology, 221(12):e202203013, December 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/221/12/e202203013/213515/WDR91-specifies-the-
endosomal-retrieval-subdomain.

Latour:2021:ERM

[LM21] Simon Latour and Alison P. McGuigan. Easy and robust micropatterning using fibrinogen anchors. Journal of Cell Biology, 220(2):e202012105, February 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/220/2/e202012105/211669/Easy-
and-robust-micropatterning-using-fibrinogen.

Leterme:2023:MMB

[LM23] Sébastien Leterme and Morgane Michaud. Mitochondrial membrane biogenesis: a new pathway for lipid transport mediated by PERK/E-Syt1 complex. Journal of Cell Biology, 222(3): e202301132, March 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/222/3/e202301132/213895/Mitochondrial-membrane-
biogenesis-A-new-pathway.

Liu:2020:MTD

[LMJ^{+}20] Yasmine J. Liu, Rebecca L. McIntyre, Georges E. Janssens, Evan G. Williams, Jiayi Lan, Michel van Weeghel, Bauke Schomakers, Henk van der Veen, Nicole N. van der Wel, Pallas Yao, William B. Mair, Ruedi Aebersold, Alyson W. MacInnes, and Riekelt H. Houtkooper. Mitochondrial translation and dynamics synergistically extend lifespan in C. elegans through HLH-30. Journal of Cell Biology, 219(6):e201907067, June 1, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/6/
e201907067/151623/Mitochondrial-translation-and-dynamics.

Labbe:2021:MMO

[LML^{+}21] Katherine Labbé, Shona Mookerjee, Maxence Le Vasseur, Eddy Gibbs, Chad Lerner, and Jodi Nunnari. The modified mitochondrial outer membrane carrier MTCH2 links mitochondrial fusion to lipogenesis. Journal of Cell Biology, 220(11): e202103122, November 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/220/11/e202103122/212670/The-
modified-mitochondrial-outer-membrane-carrier.

Larsson:2023:DFA

[LMM^{+}23] Elin Larsson, Björn Morén, Kerrie-Ann McMahon, Robert G. Parton, and Richard Lundmark. Dynamin2 functions as an accessory protein to reduce the rate of caveola internalization. Journal of Cell Biology, 222(4):e202205122, April 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/4/
e202205122/213853/Dynamin2-functions-as-an-accessory-
protein-to.

Larios:2020:AEI

[LMRG20] Jorge Larios, Vincent Mercier, Aurélien Roux, and Jean Gruenberg. ALIX- and ESCRT-III-dependent sorting of tetraspanins to exosomes. Journal of Cell Biology, 219(3):e201904113, March 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Larocque:2021:INM

[LMS^{+}21] Gabrielle Larocque, Daniel J. Moore, Méghane Sittewelle, Cansu Kuey, Joseph H. R. Hetmanski, Penelope J. La-Borde, Beverley J. Wilson, Nicholas I. Clarke, Patrick T. Caswell, and Stephen J. Royle. Intracellular nanovesicles mediate 51 integrin trafficking during cell migration. Journal of Cell Biology, 220(10):e202009028, October 4, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/10/
e202009028/212493/Intracellular-nanovesicles-mediate-
5-1-integrin.

Lee:2022:CPD

[LNY^{+}22] Moonsup Lee, Kunio Nagashima, Jaeho Yoon, Jian Sun, Ziqiu Wang, Christina Carpenter, Hyun-Kyung Lee, Yoo-Seok Hwang, Christopher J. Westlake, and Ira O. Daar. CEP97 phosphorylation by Dyrk1a is critical for centriole separation during multiciliogenesis. Journal of Cell Biology, 221(1):e202102110, January 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/1/
e202102110/212841/CEP97-phosphorylation-by-Dyrk1a-is-
critical-for.

Lowe:2021:GKG

[Low21] Martin Lowe. GOLPH3 keeps the Golgi residents at home. Journal of Cell Biology, 220(10):e202108147, October 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/10/
e202108147/212619/GOLPH3-keeps-the-Golgi-residents-at-
homeGOLPH3-and.

Li:2022:SAP

[LPMA^{+}22] Yihang Li, Maria F. Pazyra-Murphy, Daina Avizonis, Mariana de Sá Tavares Russo, Sophia Tang, Chiung-Ya Chen, Yi-Ping Hsueh, Johann S. Bergholz, Tao Jiang, Jean J. Zhao, Jian Zhu, Kwang Woo Ko, Jeffrey Milbrandt, Aaron DiAntonio, and Rosalind A. Segal. Sarm1 activation produces cADPR to increase intra-axonal Ca ++ and promote axon degeneration in PIPN. Journal of Cell Biology, 221(2):e202106080, February 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/2/
e202106080/212935/Sarm1-activation-produces-cADPR-to-
increase-intra.

Loi:2023:SAF

[LQS23] Jonathan Loi, Xiaofei Qu, and Aussie Suzuki. Semi-automated 3D fluorescence speckle analyzer (3D-Speckler) for microscope calibration and nanoscale measurement. Journal of Cell Biology, 222(4):e202202078, April 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/222/4/e202202078/213839/Semi-
automated-3D-fluorescence-speckle-analyzer-3D.

Lepeta:2022:EKT

[LRB^{+}22] Katarzyna Lepeta, Chantal Roubinet, Milena Bauer, M. Alessandra Vigano, Gustavo Aguilar, Oguz Kanca, Amanda Ochoa-Espinosa, Dimitri Bieli, Clemens Cabernard, Emmanuel Caussinus, and Markus Affolter. Engineered kinases as a tool for phosphorylation of selected targets in vivo. Journal of Cell Biology, 221(10):e202106179, October 3, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/10/
e202106179/213463/Engineered-kinases-as-a-tool-for-phosphorylation.

Liu:2020:MCB

[LRL^{+}20] Jie Liu, Manuel A. Riquelme, Zhen Li, Yuting Li, Yuxin Tong, Yumeng Quan, Cheng Pei, Sumin Gu, and Jean X. Jiang. Mechanosensitive collaboration between integrins and connexins allows nutrient and antioxidant transport into the lens. Journal of Cell Biology, 219(12):e202002154, December 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/
12/e202002154/211530/Mechanosensitive-collaboration-
between-integrins.

Levic:2020:DRL

[LRM^{+}20] Daniel S. Levic, Sean Ryan, Lindsay Marjoram, Jamie Honeycutt, Jennifer Bagwell, and Michel Bagnat. Distinct roles for luminal acidification in apical protein sorting and trafficking in zebrafish. Journal of Cell Biology, 219(4):e201908225, April 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/4/
e201908225/133852/Distinct-roles-for-luminal-acidification-
in-apical.

Lee:2020:FPN

[LSD^{+}20a] I-Ju Lee, Ema Stokasimov, Nathaniel Dempsey, Joseph M. Varberg, Etai Jacob, Sue L. Jaspersen, and David Pellman. Factors promoting nuclear envelope assembly independent of the canonical ESCRT pathway. Journal of Cell Biology, 219(6):e201908232, June 1, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/6/e201908232/151616/Factors-promoting-nuclear-envelope-
assembly.

Long:2020:IKF

[LSD20b] Alexandra F. Long, Pooja Suresh, and Sophie Dumont. Individual kinetochore-fibers locally dissipate force to maintain robust mammalian spindle structure. Journal of Cell Biology, 219(8): e201911090, August 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/219/8/e201911090/151795/Individual-kinetochore-
fibers-locally-dissipate.

Lv:2021:EUL

[LSD^{+}21] Bo Lv, Michael W. Stuck, Paurav B. Desai, Oscar A. Cabrera, and Gregory J. Pazour. E3 ubiquitin ligase Wwp1 regulates ciliary dynamics of the Hedgehog receptor Smoothened. Journal of Cell Biology, 220(9):e202010177, September 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/9/
e202010177/212435/E3-ubiquitin-ligase-Wwp1-regulates-
ciliary.

Lolicato:2022:CPC

[LSG^{+}22] Fabio Lolicato, Roberto Saleppico, Alessandra Griffo, Annalena Meyer, Federica Scollo, Bianca Pokrandt, Hans-Michael Müller, Helge Ewers, Hendrik Hähl, Jean-Baptiste Fleury, Ralf Seemann, Martin Hof, Britta Brügger, Karin Jacobs, Ilpo Vattulainen, and Walter Nickel. Cholesterol promotes clustering of PI(4,5)P 2 driving unconventional secretion of FGF2. Journal of Cell Biology, 221(11):e202106123, November 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/11/
e202106123/213511/Cholesterol-promotes-clustering-of-
PI-4-5-P2.

Lopes:2023:TTD

[LSOM23] Danilo Lopes, Alexandre L. Seabra, Bernardo Orr, and Helder Maiato. α-tubulin detyrosination links the suppression of MCAK activity with taxol cytotoxicity. Journal of Cell Biology, 222(2):e202205092, February 6, 2023. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/222/2/
e202205092/213730/Tubulin-detyrosination-links-the-suppression-
of.

Lleti:2023:CSS

[LSS^{+}23] José M. Serra Lleti, Anna M. Steyer, Nicole L. Schieber, Beate Neumann, Christian Tischer, Volker Hilsenstein, Mike Holtstrom, David Unrau, Robert Kirmse, John M. Lucocq, Rainer Pepperkok, and Yannick Schwab. CLEM site, a software for automated phenotypic screens using light microscopy and FIB-SEM. Journal of Cell Biology, 222(3):e202209127, March 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/3/
e202209127/213779/CLEMSite-a-software-for-automated-
phenotypic.

Liu:2022:AMB

[LSX^{+}22] Yan-Xia Liu, Wei-Yue Sun, Bin Xue, Rui-Kai Zhang, Wen-Juan Li, Xixian Xie, and Zhen-Chuan Fan. ARL3 mediates BBSome ciliary turnover by promoting its outward movement across the transition zone. Journal of Cell Biology, 221(10):e202111076, October 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/221/10/e202111076/213491/ARL3-mediates-BBSome-
ciliary-turnover-by-promoting.

Li:2020:GGC

[LTL^{+}20] Shuangxi Li, Aiguo Tian, Shuang Li, Yuhong Han, Bing Wang, and Jin Jiang. Gilgamesh (Gish)/CK1γ regulates tissue homeostasis and aging in adult Drosophila midgut. Journal of Cell Biology, 219(4):e.201909103, April 6, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/4/e.
201909103/133831/Gilgamesh-Gish-CK1-regulates-tissue-
homeostasis.

Lord:2020:SCR

[LVMFL20] Samuel J. Lord, Katrina B. Velle, R. Dyche Mullins, and Lillian K. Fritz-Laylin. SuperPlots: Communicating reproducibility and variability in cell biology. Journal of Cell Biology, 219(6): e202001064, June 1, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/219/6/e202001064/151717/SuperPlots-Communicating-
reproducibility-and.

Li:2020:RHS

[LW20a] Dan Li and Jianlong Wang. Ribosome heterogeneity in stem cells and development. Journal of Cell Biology, 219(6):e202001108, June 1, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/6/e202001108/151702/Ribosome-heterogeneity-in-stem-
cells-and.

Lord:2020:NEV

[LW20b] Christopher L. Lord and Susan R. Wente. Nuclear envelope– vacuole contacts mitigate nuclear pore complex assembly stress. Journal of Cell Biology, 219(12):e202001165, December 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/12/
e202001165/211463/Nuclear-envelope-vacuole-contacts-
mitigate-nuclear.

Li:2021:TVS

[LWD^{+}21] Yang Emma Li, Yichang Wang, Ximing Du, Tizhong Zhang, Hoi Yin Mak, Sarah E. Hancock, Holly McEwen, Elvis Pandzic, Renee M. Whan, Yvette Celine Aw, Ivan E. Lukmantara, Yiqiong Yuan, Xiuju Dong, Anthony Don, Nigel Turner, Shiqian Qi, and Hongyuan Yang. TMEM41B and VMP1 are scramblases and regulate the distribution of cholesterol and phosphatidylserine. Journal of Cell Biology, 220(6):e202103105, June 7, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
6/e202103105/212020/TMEM41B-and-VMP1-are-scramblases-
and-regulate-the.

Lim:2022:STB

[LWG^{+}22] Hui Jing Lim, Jacinta M. Wubben, Cristian Pinero Garcia, Sebastian Cruz-Gomez, Jieru Deng, Jeffrey Y. W. Mak, Abderrahman Hachani, Regan J. Anderson, Gavin F. Painter, Jesse Goyette, Shanika L. Amarasinghe, Matthew E. Ritchie, Antoine Roquilly, David P. Fairlie, Katharina Gaus, Jamie Rossjohn, Jose A. Villadangos, and Hamish E. G. McWilliam. A specialized tyrosine-based endocytosis signal in MR1 controls antigen presentation to MAIT cells. Journal of Cell Biology, 221(12):e202110125, December 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/221/12/e202110125/213489/A-
specialized-tyrosine-based-endocytosis-signal-in.

Liu:2023:EDA

[LWL^{+}23] Hui Liu, Jia-Yi Wei, Yuan Li, Meng Ban, Qi Sun, Hui-Jie Wang, Dan Zhao, Pai-Ge Tong, Li Wang, Kang-Ji Wang, Jin-Li Yue, Hong-Yan Zhang, Wen-Gang Fang, Dong-Xin Liu, De-Shu Shang, Bo Li, Ya-Ping Jin, Liu Cao, Wei-Dong Zhao, and Yu-Hua Chen. Endothelial depletion of atg7 triggers astrocyte–microvascular disassociation at blood–brain barrier. Journal of Cell Biology, 222(5): e202103098, May 1, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/222/5/e202103098/213981/Endothelial-
depletion-of-Atg7-triggers-astrocyte.

Liu:2023:APS

[LWZ^{+}23] Yi-Shi Liu, Yicheng Wang, Xiaoman Zhou, Linpei Zhang, Ganglong Yang, Xiao-Dong Gao, Yoshiko Murakami, Morihisa Fujita, and Taroh Kinoshita. Accumulated precursors of specific GPI-anchored proteins upregulate GPI biosynthesis with ARV1. Journal of Cell Biology, 222(5):e202208159, May 1, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/5/
e202208159/213904/Accumulated-precursors-of-specific-
GPI-anchored.

Li:2023:HIC

[LXJ^{+}23] Yanan Li, Baijie Xu, Mengmeng Jin, Hui Zhang, Ningxin Ren, Jinhui Hu, and Jie He. Homophilic interaction of cell adhesion molecule 3 coordinates retina neuroepithelial cell proliferation. Journal of Cell Biology, 222(6):e202204098, June 5, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/6/
e202204098/214005/Homophilic-interaction-of-cell-adhesion-
molecule-3.

Liu:2022:DHT

[LYL^{+}22] Liqing Liu, Shuxin Yang, Yang Liu, Xixia Li, Junjie Hu, Li Xiao, and Tao Xu. DeepContact: High-throughput quantification of membrane contact sites based on electron microscopy imaging. Journal of Cell Biology, 221(9):e202106190, September 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/9/e202106190/213379/DeepContact-High-throughput-
quantification-of.

Li:2023:GFR

[LYL^{+}23] Dong Li, Yihong Yang, Chenglin Lv, Yingjie Wang, Xiaoting Chao, Jiafeng Huang, Shashi P. Singh, Ye Yuan, Chengyu Zhang, Jizhong Lou, Pu Gao, Shanjin Huang, Bo Li, and Huaqing Cai. GxcM-Fbp17/RacC-WASP signaling regulates polarized cortex assembly in migrating cells via Arp2/3. Journal of Cell Biology, 222(6):e202208151, June 5, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/222/6/e202208151/213991/GxcM-
Fbp17-RacC-WASP-signaling-regulates-polarized.

Le:2021:CLI

[LYP^{+}21] Anh Hoang Le, Tamas Yelland, Nikki R. Paul, Loic Fort, Savvas Nikolaou, Shehab Ismail, and Laura M. Machesky. CYRI-A limits invasive migration through macropinosome formation and integrin uptake regulation. Journal of Cell Biology, 220(9):e202012114, September 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/220/9/e202012114/212437/CYRI-
A-limits-invasive-migration-through.

Liu:2020:RTD

[LYS^{+}20] Bing Liu, Muhammad Younus, Suhua Sun, Yiman Li, Yuan Wang, Xi Wu, Xiaoxuan Sun, Shujiang Shang, Changhe Wang, Michael X. Zhu, and Zhuan Zhou. Reply to “TRPA1-dependent calcium transients and CGRP release in DRG neurons require extracellular calcium”. Journal of Cell Biology, 219(6):e202004017, June 1, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/6/e202004017/151797/Reply-to-TRPA1-dependent-calcium-
transients-and. See [GKFR20].

Liang:2020:CLA

[LZC^{+}20] Cai Liang, Zhenlei Zhang, Qinfu Chen, Haiyan Yan, Miao Zhang, Linli Zhou, Junfen Xu, Weiguo Lu, and Fangwei Wang. Centromere-localized Aurora B kinase is required for the fidelity of chromosome segregation. Journal of Cell Biology, 219(2):e201907092, February 3, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/2/
e201907092/133535/Centromere-localized-Aurora-B-kinase-
is-required.

Li:2023:IAT

[LZT^{+}23] Pai Li, Ze Zhang, Yiying Tong, Bardees M. Foda, and Brad Day. ILEE: Algorithms and toolbox for unguided and accurate quantitative analysis of cytoskeletal images. Journal of Cell Biology, 222(2):e202203024, February 6, 2023. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/222/2/e202203024/213770/ILEE-
Algorithms-and-toolbox-for-unguided-and.

Liu:2021:ART

[LZZ^{+}21] Nan Liu, Hongyu Zhao, Yan G. Zhao, Junjie Hu, and Hong Zhang. Atlastin 2/3 regulate ER targeting of the ULK1 complex to initiate autophagy. Journal of Cell Biology, 220(7): e202012091, July 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/7/e202012091/212100/Atlastin-2-3-
regulate-ER-targeting-of-the-ULK1.

Meiring:2020:MKL

[MA20] Joyce C. M. Meiring and Anna Akhmanova. Microtubules keep large cells in shape. Journal of Cell Biology, 219(6): e202004031, June 1, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/6/e202004031/151751/Microtubules-
keep-large-cells-in-shapeMicrotubules.

Marston:2021:SWA

[Mar21] Adele L. Marston. A SUMOylation wave to anchor the genome. Journal of Cell Biology, 220(12):e202110031, December 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/220/12/e202110031/212822/A-SUMOylation-wave-
to-anchor-the-genomeA.

Maxson:2022:DQV

[MAW^{+}22] Michelle E. Maxson, Yazan M. Abbas, Jing Ze Wu, Jonathan D. Plumb, Sergio Grinstein, and John L. Rubinstein. Detection and quantification of the vacuolar H + ATPase using the Legionella effector protein SidK. Journal of Cell Biology, 221(3):e202107174, March 7, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/3/
e202107174/212963/Detection-and-quantification-of-the-
vacuolar-H.

Malin:2022:SDR

[MBA^{+}22] Jacob Malin, Christian Rosa Birriel, Sergio Astigarraga, Jessica E. Treisman, and Victor Hatini. Sidekick dynamically rebalances contractile and protrusive forces to control tissue morphogenesis. Journal of Cell Biology, 221(5):e202107035, May 2, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/
5/e202107035/213052/Sidekick-dynamically-rebalances-
contractile-and.

Moore:2023:CCC

[MBG^{+}23] Jessica L. Moore, Dhananjay Bhaskar, Feng Gao, Catherine Matte-Martone, Shuangshuang Du, Elizabeth Lathrop, Smirthy Ganesan, Lin Shao, Rachael Norris, Nil Campamà Sanz, Karl Annusver, Maria Kasper, Andy Cox, Caroline Hendry, Bastian Rieck, Smita Krishnaswamy, and Valentina Greco. Cell cycle controls long-range calcium signaling in the regenerating epidermis. Journal of Cell Biology, 222(7):e202302095, July 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/222/7/e202302095/214066/Cell-cycle-controls-
long-range-calcium-signaling.

Mesgarzadeh:2022:SRR

[MBW22] Jaleh S. Mesgarzadeh, Joel N. Buxbaum, and R. Luke Wiseman. Stress-responsive regulation of extracellular proteostasis. Journal of Cell Biology, 221(4):e202112104, April 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/4/e202112104/213026/Stress-responsive-regulation-
of-extracellular.

Milosevic:2021:FTA

[MC21] Ira Milosevic and Michael A. Cousin. Fine-tuning activitydependent bulk endocytosis via kinases and phosphatases. Journal of Cell Biology, 220(12):e202111056, December 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
12/e202111056/212866/Fine-tuning-activity-dependent-
bulk-endocytosis.

McClatchey:2021:ECC

[McC21] Andrea I. McClatchey. EPHecting cell contact by increasing cortical tension. Journal of Cell Biology, 220(6):e202105015, June 7, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/6/

e202105015/212115/EPHecting-cell-contact-by-increasing-
cortical.

McWilliams:2023:FTT

[McW23] Thomas G. McWilliams. Fine-tune TMEM11 to unleash basal mitophagy. Journal of Cell Biology, 222(4):e202302118, April 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/4/
e202302118/213947/Fine-tune-TMEM11-to-unleash-basal-
mitophagyTMEM11.

Martinez:2020:TMM

[MDB^{+}20] Pablo Martinez, Ram Dixit, Rachappa S. Balkunde, Antonia Zhang, Seán E. O’Leary, Kenneth A. Brakke, and Carolyn G. Rasmussen. TANGLED1 mediates microtubule interactions that may promote division plane positioning in maize. Journal of Cell Biology, 219(8):e201907184, August 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/8/
e201907184/151878/TANGLED1-mediates-microtubule-interactions-
that.

Milas:2023:FCC

[MdCT23] Ana Milas, Jorge de Carvalho, and Ivo A. Telley. Follicle cell contact maintains main body axis polarity in the Drosophila melanogaster oocyte. Journal of Cell Biology, 222(2): e202209052, February 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/222/2/e202209052/213703/Follicle-cell-
contact-maintains-main-body-axis.

Monster:2021:AJM

[MDV^{+}21] Jooske L. Monster, Lisa Donker, Marjolein J. Vliem, Zaw Win, Helen K. Matthews, Joleen S. Cheah, Soichiro Yamada, Johan de Rooij, Buzz Baum, and Martijn Gloerich. An asymmetric junctional mechanoresponse coordinates mitotic rounding with epithelial integrity. Journal of Cell Biology, 220(5):e202001042, May 3, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/220/5/e202001042/211872/An-
asymmetric-junctional-mechanoresponse.

Morao:2021:HST

[ME21] Ana Karina Morao and Sevinc Ercan. Hatched and starved: Two chromatin compaction mechanisms join forces to silence germ cell genome. Journal of Cell Biology, 220(9): e202107026, September 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/9/e202107026/212550/Hatched-and-
starved-Two-chromatin-compaction.

Merdes:2021:MNR

[Mer21] Andreas Merdes. Microtubule nucleation without a ring? Journal of Cell Biology, 220(3):e202101015, March 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/3/
e202101015/211792/Microtubule-nucleation-without-a-ring-
Microtubule.

Molinuevo:2020:DDR

[MFC^{+}20] Rut Molinuevo, Ana Freije, Lizbeth Contreras, Juan R. Sanz, and Alberto Gandarillas. The DNA damage response links human squamous proliferation with differentiation. Journal of Cell Biology, 219(11):e202001063, November 2, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/11/
e202001063/152154/The-DNA-damage-response-links-human-
squamous.

Miao:2022:BCP

[MGM22] Guangxia Miao, Li Guo, and Denise J. Montell. Border cell polarity and collective migration require the spliceosome component cactin. Journal of Cell Biology, 221(7):e202202146, July 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/221/7/e202202146/213245/Border-cell-polarity-
and-collective-migration.

Marsan:2022:ERD

[MH22] Elise Marsan and Eric J. Huang. Endosomal recycling defects link Huntington’s disease with McLeod syndrome. Journal of Cell Biology, 221(10):e202208164, October 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/

221/10/e202208164/213470/Endosomal-recycling-defects-
link-Huntington-s.

Marivin:2022:DOA

[MHGM22] Arthur Marivin, Rachel Xi-Yeen Ho, and Mikel Garcia-Marcos. DAPLE orchestrates apical actomyosin assembly from junctional polarity complexes. Journal of Cell Biology, 221(5):e202111002, May 2, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/5/e202111002/213115/DAPLE-orchestrates-apical-actomyosin-
assembly-from.

Maniscalco:2020:ICR

[MHN20] Chelsea Maniscalco, Allison E. Hall, and Jeremy Nance. An interphase contractile ring reshapes primordial germ cells to allow bulk cytoplasmic remodeling. Journal of Cell Biology, 219(2):e201906185, February 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/219/2/e201906185/132628/An-
interphase-contractile-ring-reshapes-primordial.

Malt:2020:WRP

[MHS^{+}20] Andre Landin Malt, Arielle K. Hogan, Connor D. Smith, Maxwell S. Madani, and Xiaowei Lu. Wnts regulate planar cell polarity via heterotrimeric G protein and PI3K signaling. Journal of Cell Biology, 219(10):e201912071, October 5, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/10/
e201912071/152025/Wnts-regulate-planar-cell-polarity-
via.

May:2021:TRP

[MKD^{+}21] Elena A. May, Marian Kalocsay, Inès Galtier D’Auriac, Patrick S. Schuster, Steven P. Gygi, Maxence V. Nachury, and David U. Mick. Time-resolved proteomics profiling of the ciliary hedgehog response. Journal of Cell Biology, 220 (5):e202007207, May 3, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/220/5/e202007207/211991/Time-
resolved-proteomics-profiling-of-the-ciliary.

Meitinger:2021:TPF

[MKO^{+}21] Franz Meitinger, Dong Kong, Midori Ohta, Arshad Desai, Karen Oegema, and Jadranka Loncarek. TRIM37 prevents formation of condensate-organized ectopic spindle poles to ensure mitotic fidelity. Journal of Cell Biology, 220(7):e202010180, July 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/7/e202010180/212098/TRIM37-prevents-formation-of-
condensate-organized.

Mannino:2022:QCM

[ML22] Philip J. Mannino and C. Patrick Lusk. Quality control mechanisms that protect nuclear envelope identity and function. Journal of Cell Biology, 221(9):e202205123, September 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/
9/e202205123/213424/Quality-control-mechanisms-that-
protect-nuclear.

Moulay:2020:ASC

[MLL^{+}20] Gilles Moulay, Jeanne Lainé, Mégane Lemaître, Masayuki Nakamori, Ichizo Nishino, Ghislaine Caillol, Kamel Mamchaoui, Laura Julien, Florent Dingli, Damarys Loew, Marc Bitoun, Christophe Leterrier, Denis Furling, and Stéphane Vassilopoulos. Alternative splicing of clathrin heavy chain contributes to the switch from coated pits to plaques. Journal of Cell Biology, 219(9):e201912061, September 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/
9/e201912061/151930/Alternative-splicing-of-clathrin-
heavy-chain.

Mao:2021:PSM

[MLQ^{+}21] Lejiao Mao, Chenyi Liao, Jiao Qin, Yanqiu Gong, Yifei Zhou, Shasha Li, Zhe Liu, Huaqing Deng, Wankun Deng, Qingxiang Sun, Xianming Mo, Yu Xue, Daniel D. Billadeau, Lunzhi Dai, Guohui Li, and Da Jia. Phosphorylation of SNX27 by MAPK11/14 links cellular stress–signaling pathways with endocytic recycling. Journal of Cell Biology, 220(4):??, April 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/4/
e202010048/211812/Phosphorylation-of-SNX27-by-MAPK11-
14-links.

Melia:2020:ABM

[MLS20] Thomas J. Melia, Alf H. Lystad, and Anne Simonsen. Autophagosome biogenesis: From membrane growth to closure. Journal of Cell Biology, 219(6):e202002085, June 1, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/6/
e202002085/151729/Autophagosome-biogenesis-From-membrane-
growth-to.

Michaud:2022:VCP

[MLS^{+}22] Ani Michaud, Marcin Leda, Zachary T. Swider, Songeun Kim, Jiaye He, Jennifer Landino, Jenna R. Valley, Jan Huisken, Andrew B. Goryachev, George von Dassow, and William M. Bement. A versatile cortical pattern-forming circuit based on Rho, F-actin, Ect2, and RGA-3/4. Journal of Cell Biology, 221(8): e202203017, August 1, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/221/8/e202203017/213290/A-versatile-
cortical-pattern-forming-circuit-based.

Marenda:2021:PFM

[MLvdL^{+}21] Mattia Marenda, Elena Lazarova, Sebastian van de Linde, Nick Gilbert, and Davide Michieletto. Parameter-free molecular super-structures quantification in single-molecule localization microscopy. Journal of Cell Biology, 220(5):e202010003, May 3, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/5/
e202010003/211893/Parameter-free-molecular-super-structures.

Mouery:2020:PPS

[MMC20] Brandon L. Mouery, Liu Mei, and Jeanette Gowen Cook. Programming pluripotent stem cells: Can’t teach an old cell new DNA replication tricks. Journal of Cell Biology, 219(9): e202008014, September 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/9/e202008014/152054/Programming-
pluripotent-stem-cells-Can-t-teach-an.

Mukherjee:2022:CLI

[MMDK^{+}22] Abhishek Mukherjee, Shay Melamed, Hana Damouny-Khoury, Malak Amer, Lea Feld, Elisabeth Nadjar-Boger, Michael P.

Sheetz, and Haguy Wolfenson. α-Catenin links integrin adhesions to F -actin to regulate ECM mechanosensing and rigidity dependence. Journal of Cell Biology, 221(8):e202102121, August 1, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/
8/e202102121/213257/Catenin-links-integrin-adhesions-
to-F-actin-to.

Muriel:2021:UPM

[MMKM21] Olivia Muriel, Laetitia Michon, Wanda Kukulski, and Sophie G. Martin. Ultrastructural plasma membrane asymmetries in tension and curvature promote yeast cell fusion. Journal of Cell Biology, 220(10):e202103142, October 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
10/e202103142/212551/Ultrastructural-plasma-membrane-
asymmetries-in.

Moon:2020:CTQ

[MMSP20] Stephanie L. Moon, Tatsuya Morisaki, Timothy J. Stasevich, and Roy Parker. Coupling of translation quality control and mRNA targeting to stress granules. Journal of Cell Biology, 219 (8):e202004120, August 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/8/e202004120/151851/Coupling-of-
translation-quality-control-and-mRNA.

Matikainen:2020:IEM

[MNC20] Sampsa Matikainen, Tuula A. Nyman, and Wojciech Cypryk. Inflammasomes: Exosomal miRNAs loaded for action. Journal of Cell Biology, 219(10):e202008130, October 5, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/10/
e202008130/152115/Inflammasomes-Exosomal-miRNAs-loaded-
for.

Mateska:2020:RSS

[MND^{+}20] Ivona Mateska, Kareena Nanda, Natalie A. Dye, Vasileia Ismini Alexaki, and Suzanne Eaton. Range of SHH signaling in adrenal gland is limited by membrane contact to cells with primary cilia. Journal of Cell Biology, 219(12):e201910087, December 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/12/

e201910087/211483/Range-of-SHH-signaling-in-adrenal-
gland-is-limited.

Mast:2020:CLS

[MNvdS^{+}20] Fred D. Mast, Arti T. Navare, Almer M. van der Sloot, Jasmin Coulombe-Huntington, Michael P. Rout, Nitin S. Baliga, Alexis Kaushansky, Brian T. Chait, Alan Aderem, Charles M. Rice, Andrej Sali, Mike Tyers, and John D. Aitchison. Crippling life support for SARS-CoV-2 and other viruses through synthetic lethality. Journal of Cell Biology, 219(10):e202006159, October 5, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/
10/e202006159/152015/Crippling-life-support-for-SARS-
CoV-2-and-other.

Mochida:2022:ALD

[MOK^{+}22] Keisuke Mochida, Toshifumi Otani, Yuto Katsumata, Hiromi Kirisako, Chika Kakuta, Tetsuya Kotani, and Hitoshi Nakatogawa. Atg39 links and deforms the outer and inner nuclear membranes in selective autophagy of the nucleus. Journal of Cell Biology, 221(2):e202103178, February 7, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/2/
e202103178/212974/Atg39-links-and-deforms-the-outer-
and-inner.

Mis:2020:IMT

[MOS^{+}20] Monika Mis, Siobhan O’Brien, Zachary Steinhart, Sichun Lin, Traver Hart, Jason Moffat, and Stephane Angers. IPO11 mediates β catenin nuclear import in a subset of colorectal cancers. Journal of Cell Biology, 219(2):e201903017, February 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/2/
e201903017/133547/IPO11-mediates-catenin-nuclear-import-
in-a-subset.

Matsumoto:2022:GPM

[MOS^{+}22] Shunsuke Matsumoto, Suzuka Ono, Saori Shinoda, Chika Kakuta, Satoshi Okada, Takashi Ito, Tomoyuki Numata, and Toshiya Endo. GET pathway mediates transfer of mislocalized tail-anchored proteins from mitochondria to the ER. Journal of Cell Biology, 221(6):e202104076, June 6, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/
6/e202104076/213171/GET-pathway-mediates-transfer-of-
mislocalized-tail.

Morgado-Palacin:2021:GPI

[MP21a]  Lucia Morgado-Palacin.  Gaia Pigino: Inside the cell.

Journal of Cell Biology, 220(9):e202108043, September 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
9/e202108043/212596/Gaia-Pigino-Inside-the-cellGaia-
Pigino-Inside-the.

Morgado-Palacin:2021:NYI

[MP21b] Lucia Morgado-Palacin. Nan Yan: Innate immune signaling goes beyond viral. Journal of Cell Biology, 220(11): e202109118, November 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/220/11/e202109118/212703/Nan-
Yan-Innate-immune-signaling-goes-beyond.

Morgado-Palacin:2021:VAA

[MP21c] Lucia Morgado-Palacin. Vaishnavi Ananthanarayanan: Advocating for women’s representation in science. Journal of Cell Biology, 220(12):e202111029, December 6, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/12/
e202111029/212817/Vaishnavi-Ananthanarayanan-Advocating-
for-women-s.

Morgado-Palacin:2021:YTS

[MP21d] Lucia Morgado-Palacin. Ye Tian: Surveilling stress to live longer. Journal of Cell Biology, 220(10):e202108098, October 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/10/e202108098/212642/Ye-Tian-Surveilling-stress-
to-live-longerYe-Tian.

Morgado-Palacin:2022:BZC

[MP22a] Lucia Morgado-Palacin. Bo Zhong: Captive by the viral immune escape. Journal of Cell Biology, 221(3):e202202057, March 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/221/3/e202202057/213029/Bo-Zhong-Captive-by-
the-viral-immune-escapeBo.

Morgado-Palacin:2022:DSS

[MP22b] Lucia Morgado-Palacin. Dorothy Schafer: Sculpting the next generation of microglia researchers. Journal of Cell Biology, 221(9):e202208061, September 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/221/9/e202208061/213412/Dorothy-
Schafer-Sculpting-the-next-generation-of.

Morgado-Palacin:2022:EGS

[MP22c] Lucia Morgado-Palacin. Elda Grabocka: Stress-buffering comes in granules. Journal of Cell Biology, 221(2):e202201081, February 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/221/2/e202201081/212981/Elda-Grabocka-Stress-
buffering-comes-in.

Morgado-Palacin:2022:EBL

[MP22d] Lucia Morgado-Palacin. Elvan Böke: Long live the oocyte. Journal of Cell Biology, 221(11):e202210049, November 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/11/
e202210049/213560/Elvan-Boke-Long-live-the-oocyteElvan-
Boke-Long.

Morgado-Palacin:2022:OAM

[MP22e] Lucia Morgado-Palacin. Ori Avinoam: Mind, body, and membranes in shape. Journal of Cell Biology, 221(4): e202203049, April 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/221/4/e202203049/213085/Ori-Avinoam-Mind-
body-and-membranes-in-shapeOri.

Morgado-Palacin:2022:RMP

[MP22f] Lucia Morgado-Palacin. Rushika M. Perera: Lysosomes unveil what cancer cells hide. Journal of Cell Biology, 221(1): e202112049, January 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/221/1/e202112049/212920/Rushika-M-Perera-
Lysosomes-unveil-what-cancer.

Morgado-Palacin:2022:SMF

[MP22g] Lucia Morgado-Palacin. Sachihiro Matsunaga: FISHing the nuclear architecture of plant cells. Journal of Cell Biology, 221(12): e202211043, December 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/221/12/e202211043/213701/Sachihiro-Matsunaga-
FISHing-the-nuclear.

Morgado-Palacin:2022:SCH

[MP22h] Lucia Morgado-Palacin. Sara Cuylen-Haering: Cellular soaps to keep neat chromosomes. Journal of Cell Biology, 221(8): e202206048, August 1, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/221/8/e202206048/213304/Sara-Cuylen-
Haering-Cellular-soaps-to-keep-neat.

Morgado-Palacin:2022:SWF

[MP22i] Lucia Morgado-Palacin. Sara Wickström: The forces controlling our cell fate. Journal of Cell Biology, 221(6): e202205077, June 6, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/221/6/e202205077/213238/Sara-
Wickstrom-The-forces-controlling-our-cell.

Morgado-Palacin:2023:CJC

[MP23a] Lucia Morgado-Palacin. Chii Jou Chan: The positives of being under “pressure”. Journal of Cell Biology, 222(2): e202211075, February 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/222/2/e202211075/213830/Chii-
Jou-Chan-The-positives-of-being-under.

Morgado-Palacin:2023:JAB

[MP23b] Lucia Morgado-Palacin. Judith Agudo: Beware of your inner self-immune attack. Journal of Cell Biology, 222(3): e202302027, March 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/222/3/e202302027/213881/Judith-Agudo-
Beware-of-your-inner-self-immune.

Morgado-Palacin:2023:YKD

[MP23c] Lucia Morgado-Palacin. Yogesh Kulathu: Decoding complex intracellular messages. Journal of Cell Biology, 222(4):e202303042, April 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/

222/4/e202303042/213967/Yogesh-Kulathu-Decoding-complex-
intracellular.

Marshall-Phelps:2020:NAD

[MPKB^{+}20] Katy L. H. Marshall-Phelps, Linde Kegel, Marion Baraban, Torben Ruhwedel, Rafael G. Almeida, Maria Rubio-Brotons, Anna Klingseisen, Silvia K. Benito-Kwiecinski, Jason J. Early, Jenea M. Bin, Daumante Suminaite, Matthew R. Livesey, Wiebke Möbius, Richard J. Poole, and David A. Lyons. Neuronal activity disrupts myelinated axon integrity in the absence of NKCC1b. Journal of Cell Biology, 219(7):e201909022, July 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/7/
e201909022/151733/Neuronal-activity-disrupts-myelinated-
axon.

Murillo-Pineda:2021:ISH

[MPVD^{+}21] Marina Murillo-Pineda, Luis P. Valente, Marie Dumont, João F. Mata, Daniele Fachinetti, and Lars E. T. Jansen. Induction of spontaneous human neocentromere formation and long-term maturation. Journal of Cell Biology, 220(3): e202007210, March 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/3/e202007210/211684/Induction-of-
spontaneous-human-neocentromere.

Mast:2020:PPE

[MRA20] Fred D. Mast, Richard A. Rachubinski, and John D. Aitchison. Peroxisome prognostications: Exploring the birth, life, and death of an organelle. Journal of Cell Biology, 219(3):e201912100, March 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/3/e201912100/133827/Peroxisome-prognostications-
Exploring-the-birth.

McQuown:2021:TPD

[MRD21] Alexander J. McQuown, Dvir Reif, and Vladimir Denic. A TR-Cky TA protein delivery service snubs the UPS. Journal of Cell Biology, 220(5):e202103196, May 3, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/220/5/e202103196/212007/A-
TRCky-TA-protein-delivery-service-snubs-the-UPSA.

Murari:2020:ATT

[MRG^{+}20] Anjaneyulu Murari, Shauna-Kay Rhooms, Naga Sri Goparaju, Maximino Villanueva, and Edward Owusu-Ansah. An antibody toolbox to track complex I assembly defines AIF’s mitochondrial function. Journal of Cell Biology, 219(10): e202001071, October 5, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/10/e202001071/152090/An-antibody-
toolbox-to-track-complex-I-assembly.

Miller:2023:ARE

[MRH^{+}23] Brittany K. Miller, Guendalina Rossi, Sara Hudson, David Cully, Richard W. Baker, and Patrick Brennwald. Allosteric regulation of exocyst: Discrete activation of tethering by two spatial signals. Journal of Cell Biology, 222(3):e202206108, March 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/
3/e202206108/213852/Allosteric-regulation-of-exocyst-
Discrete.

Ma:2021:DDP

[MRL^{+}21] Shuyun Ma, Zeming Rong, Chen Liu, Xiaobing Qin, Xiaoyan Zhang, and Qiang Chen. DNA damage promotes microtubule dynamics through a DNA-PK-AKT axis for enhanced repair. Journal of Cell Biology, 220(2):e201911025, February 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
2/e201911025/211656/DNA-damage-promotes-microtubule-
dynamics-through-a.

Montano-Rendon:2022:PPL

[MRWK^{+}22] Fernando Montaño-Rendón, Glenn F. W. Walpole, Matthias Krause, Gerald R. V. Hammond, Sergio Grinstein, and Gregory D. Fairn. PtdIns(3,4)P 2, lamellipodin, and VASP coordinate actin dynamics during phagocytosis in macrophages. Journal of Cell Biology, 221(11):e202207042, November 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/
11/e202207042/213505/PtdIns-3-4-P2-Lamellipodin-and-
VASP-coordinate.

Ma:2020:MFD

[MS20] Linda Ma and Daniel A. Starr. Membrane fusion drives pronuclear meeting in the one-cell embryo. Journal of Cell Biology, 219(2):e202001048, February 3, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/2/
e202001048/133643/Membrane-fusion-drives-pronuclear-
meeting-in-the.

Maiato:2023:DCC

[MS23] Helder Maiato and Sónia Silva. Double-checking chromosome segregation. Journal of Cell Biology, 222(5):e202301106, May 1, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/5/
e202301106/214000/Double-checking-chromosome-segregationDouble.

Mangon:2021:ICC

[MSB^{+}21] Aurélie Mangon, Danièle Salaün, Mohamed Lala Bouali, Mira Kuzmić, Sabine Quitard, Sylvie Thuault, Daniel Isnardon, Stéphane Audebert, Pierre-Henri Puech, Pascal Verdier-Pinard, and Ali Badache. iASPP contributes to cell cortex rigidity, mitotic cell rounding, and spindle positioning. Journal of Cell Biology, 220(12):e202012002, December 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/12/
e202012002/212730/iASPP-contributes-to-cell-cortex-rigidity-
mitotic.

Magliozzi:2020:FYP

[MSC^{+}20] Joseph O. Magliozzi, Jack Sears, Lauren Cressey, Marielle Brady, Hannah E. Opalko, Arminja N. Kettenbach, and James B. Moseley. Fission yeast Pak1 phosphorylates anillinlike Mid1 for spatial control of cytokinesis. Journal of Cell Biology, 219(8):e201908017, August 3, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/8/
e201908017/151784/Fission-yeast-Pak1-phosphorylates-
anillin-like.

Maranon:2020:NPR

[MSH^{+}20] David G. Maranon, Neelam Sharma, Yuxin Huang, Platon Selemenakis, Meiling Wang, Noelia Altina, Weixing Zhao, and Claudia Wiese. NUCKS1 promotes RAD54 activity in homologous recombination DNA repair. Journal of Cell Biology, 219(10): e201911049, October 5, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/219/10/e201911049/152064/NUCKS1-promotes-
RAD54-activity-in-homologous.

Mitra:2020:SIC

[MSJ20] Sreyoshi Mitra, Bharath Srinivasan, and Lars E. T. Jansen. Stable inheritance of CENP-a chromatin: Inner strength versus dynamic control. Journal of Cell Biology, 219(10):e202005099, October 5, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/10/e202005099/152089/Stable-inheritance-of-CENP-
A-chromatin-Inner.

McLamarrah:2020:MMP

[MSR^{+}20] Tiffany A. McLamarrah, Sarah K. Speed, John M. Ryniawec, Daniel W. Buster, Carey J. Fagerstrom, Brian J. Galletta, Nasser M. Rusan, and Gregory C. Rogers. A molecular mechanism for the procentriole recruitment of Ana2. Journal of Cell Biology, 219(2):e201905172, February 3, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/2/
e201905172/132764/A-molecular-mechanism-for-the-procentriole.

Muller:2021:FSR

[MSX^{+}21] Andreas Müller, Deborah Schmidt, C. Shan Xu, Song Pang, Joyson Verner D’Costa, Susanne Kretschmar, Carla Münster, Thomas Kurth, Florian Jug, Martin Weigert, Harald F. Hess, and Michele Solimena. 3D FIB-SEM reconstruction of microtubule–organelle interaction in whole primary mouse β cells. Journal of Cell Biology, 220(2):e202010039, February 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/2/e202010039/211599/3D-FIB-SEM-reconstruction-of-
microtubule-organelle.

Martins:2023:HSO

[MTCL^{+}23] Carla Silva Martins, Cyntia Taveneau, Gerard Castro-Linares, Mikhail Baibakov, Nicolas Buzhinsky, Mar Eroles, Violeta Milanović, Shizue Omi, Jean-Denis Pedelacq, Francois Iv, Léa Bouillard, Alexander Llewellyn, Maxime Gomes, Mayssa Belhabib, Mira Kuzmić, Pascal Verdier-Pinard, Stacey Lee, Ali Badache, Sanjay Kumar, Cristel Chandre, Sophie Brasselet, Felix Rico, Olivier Rossier, Gijsje H. Koenderink, Jerome Wenger, Stéphanie Cabantous, and Manos Mavrakis. Human septins organize as octamer-based filaments and mediate actin-membrane anchoring in cells. Journal of Cell Biology, 222(3):e202203016, March 6, 2023. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/222/3/
e202203016/213778/Human-septins-organize-as-octamer-
based-filaments.

Martinez-Terroba:2020:LNR

[MTD20] Elena Martínez-Terroba and Nadya Dimitrova. Long noncoding RNA amplified in lung cancer rewires cancer pathways. Journal of Cell Biology, 219(9):e202007098, September 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/9/e202007098/152053/Long-noncoding-RNA-amplified-
in-lung-cancer.

Mittasch:2020:RCM

[MTR^{+}20] Matthäus Mittasch, Vanna M. Tran, Manolo U. Rios, Anatol W. Fritsch, Stephen J. Enos, Beatriz Ferreira Gomes, Alec Bond, Moritz Kreysing, and Jeffrey B. Woodruff. Regulated changes in material properties underlie centrosome disassembly during mitotic exit. Journal of Cell Biology, 219(4):e201912036, April 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Mund:2023:CCP

[MTW^{+}23] Markus Mund, Aline Tschanz, Yu-Le Wu, Felix Frey, Johanna L. Mehl, Marko Kaksonen, Ori Avinoam, Ulrich S. Schwarz, and Jonas Ries. Clathrin coats partially preassemble and subsequently bend during endocytosis. Journal of Cell Biology, 222(3):e202206038, March 6, 2023. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/222/3/
e202206038/213855/Clathrin-coats-partially-preassemble-
and.

Meroni:2020:IFT

[MV20] Alice Meroni and Alessandro Vindigni. ISG15 fast-tracks DNA replication. Journal of Cell Biology, 219(8):e202007028, August 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/8/
e202007028/151958/ISG15-fast-tracks-DNA-replicationISG15-
fast-tracks.

Marek:2020:SBR

[MVM20] Magdalena Marek, Vincent Vincenzetti, and Sophie G. Martin. Sterol biosensor reveals LAM-family ltc1-dependent sterol flow to endosomes upon Arp2/3 inhibition. Journal of Cell Biology, 219(6):e202001147, June 1, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/6/
e202001147/151694/Sterol-biosensor-reveals-LAM-family-
Ltc1-dependent.

Moore:2021:STT

[MW21] Jeffrey K. Moore and Linnea Wethekam. Specialist -tubulins for pluralist microtubules. Journal of Cell Biology, 220(12): e202110038, December 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/12/e202110038/212789/Specialist-
tubulins-for-pluralist.

Mascarau:2023:PHI

[MWF^{+}23] Rémi Mascarau, Marie Woottum, Léa Fromont, Rémi Gence, Vincent Cantaloube-Ferrieu, Zoï Vahlas, Kevin Lévêque, Florent Bertrand, Thomas Beunon, Arnaud Métais, Hicham El Costa, Nabila Jabrane-Ferrat, Yohan Gallois, Nicolas Guibert, Jean-Luc Davignon, Gilles Favre, Isabelle Maridonneau-Parini, Renaud Poincloux, Bernard Lagane, Serge Bénichou, Brigitte Raynaud-Messina, and Christel Vérollet. Productive HIV-1 infection of tissue macrophages by fusion with infected CD4 + T cells. Journal of Cell Biology, 222(5): e202205103, May 1, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/222/5/e202205103/213978/Productive-HIV-1-
infection-of-tissue-macrophages.

Meng:2023:TGM

[MWSX23] Xinan Meng, Chandra Sugiarto Wijaya, Qingfang Shao, and Suhong Xu. Triggered Golgi membrane enrichment promotes PtdIns(4,5)P2 generation for plasma membrane repair. Journal of Cell Biology, 222(8):e202303017, August 7, 2023. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/222/8/
e202303017/214098/Triggered-Golgi-membrane-enrichment-
promotes.

Moore:2023:RPZ

[MYC^{+}23] Chandler E. Moore, Selin E. Yalcindag, Hanna Czeladko, Ramya Ravindranathan, Yodhara Wijesekara Hanthi, Juliana C. Levy, Vincenzo Sannino, Detlev Schindler, Alberto Ciccia, Vincenzo Costanzo, and Andrew E. H. Elia. RFWD3 promotes ZRANB3 recruitment to regulate the remodeling of stalled replication forks. Journal of Cell Biology, 222(5):e202106022, May 1, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/5/
e202106022/214012/RFWD3-promotes-ZRANB3-recruitment-
to-regulate-the.

Ma:2020:EED

[MYK^{+}20] Cheng-I J. Ma, Yitong Yang, Taeah Kim, Chang Hua Chen, Gordon Polevoy, Miluska Vissa, Jason Burgess, and Julie A. Brill. An early endosome-derived retrograde trafficking pathway promotes secretory granule maturation. Journal of Cell Biology, 219(3):e201808017, March 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/219/3/e201808017/133712/An-
early-endosome-derived-retrograde-trafficking. See cor-

rection [MYK^{+}21].

Ma:2021:CEE

[MYK^{+}21] Cheng-I J. Ma, Yitong Yang, Taeah Kim, Chang Hua Chen, Gordon Polevoy, Miluska Vissa, Jason Burgess, and Julie A. Brill. Correction: An early endosome-derived retrograde trafficking pathway promotes secretory granule maturation. Journal of Cell Biology, 220(3):??, March 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/3/jcb.20180801701192021c/211708/
Correction-An-early-endosome-derived-retrograde. See

[MYK^{+}20].

Ma:2022:CEE

[MYK^{+}22] Cheng-I J. Ma, Yitong Yang, Taeah Kim, Chang Hua Chen, Gordon Polevoy, Miluska Vissa, Jason Burgess, and Julie A. Brill. Correction: An early endosome–derived retrograde trafficking pathway promotes secretory granule maturation. Journal of Cell Biology, 221(3):e202107174, March 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/
3/jcb.20180801712202021c/213010/Correction-An-early-
endosome-derived-retrograde.

Mori:2021:RAC

[MYM^{+}21] Masashi Mori, Tatsuma Yao, Tappei Mishina, Hiromi Endoh, Masahito Tanaka, Nao Yonezawa, Yuta Shimamoto, Shigenobu Yonemura, Kazuo Yamagata, Tomoya S. Kitajima, and Masahito Ikawa. RanGTP and the actin cytoskeleton keep paternal and maternal chromosomes apart during fertilization. Journal of Cell Biology, 220(10):e202012001, October 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/10/
e202012001/212591/RanGTP-and-the-actin-cytoskeleton-
keep-paternal.

Molenaar:2021:REF

[MYT^{+}21] Martijn R. Molenaar, Kamlesh K. Yadav, Alexandre Toulmay, Tsjerk A. Wassenaar, Muriel C. Mari, Lucie Caillon, Aymeric Chorlay, Ivan E. Lukmantara, Maya W. Haaker, Richard W. Wubbolts, Martin Houweling, Arie Bas Vaandrager, Xavier Prieur, Fulvio Reggiori, Vineet Choudhary, Hongyuan Yang, Roger Schneiter, Abdou Rachid Thiam, William A. Prinz, and J. Bernd Helms. Retinyl esters form lipid droplets independently of triacylglycerol and seipin. Journal of Cell Biology, 220(10): e202011071, October 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/220/10/e202011071/212517/Retinyl-esters-
form-lipid-droplets-independently.

Naghavi:2023:VUM

[Nag23] Mojgan H. Naghavi. Virus update for the M2 “mac-in-touch”. Journal of Cell Biology, 222(5):e202303016, May 1, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
222/5/e202303016/214021/Virus-update-for-the-M2-mac-
in-touch-Virus-update.

Nsamba:2021:TIO

[NBC^{+}21] Emmanuel T. Nsamba, Abesh Bera, Michael Costanzo, Charles Boone, and Mohan L. Gupta, Jr. Tubulin isotypes optimize distinct spindle positioning mechanisms during yeast mitosis. Journal of Cell Biology, 220(12):e202010155, December 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/12/
e202010155/212745/Tubulin-isotypes-optimize-distinct-
spindle.

Nedozralova:2022:SCE

[NBI^{+}22] Hana Nedozralova, Nirakar Basnet, Iosune Ibiricu, Satish Bodakuntla, Christian Biertümpfel, and Naoko Mizuno. In situ cryo-electron tomography reveals local cellular machineries for axon branch development. Journal of Cell Biology, 221(4): e202106086, April 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/221/4/e202106086/213057/In-situ-cryo-
electron-tomography-reveals-local.

Narendra:2020:CAL

[NGG^{+}20] Derek P. Narendra, Christelle Guillermier, Frank Gyngard, Xiaoping Huang, Michael E. Ward, and Matthew L. Steinhauser. Coupling APEX labeling to imaging mass spectrometry of single organelles reveals heterogeneity in lysosomal protein turnover. Journal of Cell Biology, 219(1):e201901097, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Nong:2021:PSA

[NKS^{+}21] Junxiu Nong, Kexin Kang, Qiaoni Shi, Xuechen Zhu, Qinghua Tao, and Ye-Guang Chen. Phase separation of Axin organizes the β-catenin destruction complex. Journal of Cell Biology, 220(4):??, April 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/220/4/e202012112/211840/Phase-separation-
of-Axin-organizes-the-catenin.

Navare:2022:VPE

[NMO^{+}22] Arti T. Navare, Fred D. Mast, Jean Paul Olivier, Thierry Bertomeu, Maxwell L. Neal, Lindsay N. Carpp, Alexis Kaushansky, Jasmin Coulombe-Huntington, Mike Tyers, and John D. Aitchison. Viral protein engagement of GBF1 induces host cell vulnerability through synthetic lethality. Journal of Cell Biology, 221(11):e202011050, November 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/11/e202011050/213618/Viral-protein-engagement-of-
GBF1-induces-host-cell.

Nabais:2021:PTA

[NPdC^{+}21] Catarina Nabais, Delphine Pessoa, Jorge de Carvalho, Thomas van Zanten, Paulo Duarte, Satyajit Mayor, Jorge Carneiro, Ivo A. Telley, and Mónica Bettencourt-Dias. Plk4 triggers autonomous de novo centriole biogenesis and maturation. Journal of Cell Biology, 220(5):e202008090, May 3, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
5/e202008090/211915/Plk4-triggers-autonomous-de-novo-
centriole.

Nguyen:2022:LDA

[NR22] Ly T. S. Nguyen and Douglas N. Robinson. The lectin Discoidin I acts in the cytoplasm to help assemble the contractile machinery. Journal of Cell Biology, 221(11):e202202063, November 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/11/e202202063/213504/The-lectin-Discoidin-I-acts-
in-the-cytoplasm-to.

Nguyen:2020:MIM

[NS20] Kim Bich Nguyen and Stefani Spranger. Modulation of the immune microenvironment by tumor-intrinsic oncogenic signaling. Journal of Cell Biology, 219(1):e201908224, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Nthiga:2021:RGT

[NSB^{+}21] Thaddaeus Mutugi Nthiga, Birendra Kumar Shrestha, Jack-Ansgar Bruun, Kenneth Bowitz Larsen, Trond Lamark, and Terje Johansen. Regulation of Golgi turnover by CALCOCO1mediated selective autophagy. Journal of Cell Biology, 220(6): e202006128, June 7, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/220/6/e202006128/212004/Regulation-of-Golgi-
turnover-by-CALCOCO1-mediated.

Nanba:2021:EME

[NTA^{+}21] Daisuke Nanba, Fujio Toki, Kyosuke Asakawa, Hiroyuki Matsumura, Ken Shiraishi, Koji Sayama, Kyoichi Matsuzaki, Hiroshi Toki, and Emi K. Nishimura. EGFR-mediated epidermal stem cell motility drives skin regeneration through COL17A1 proteolysis. Journal of Cell Biology, 220(11):e202012073, November 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
11/e202012073/212656/EGFR-mediated-epidermal-stem-cell-
motility-drives.

Nijenhuis:2020:OTO

[NvGK20] Wilco Nijenhuis, Mariëlle M. P. van Grinsven, and Lukas C. Kapitein. An optimized toolbox for the optogenetic control of intracellular transport. Journal of Cell Biology, 219(4): e201907149, April 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/4/e201907149/133834/An-optimized-
toolbox-for-the-optogenetic-control.

Nakamura:2020:KLP

[NVPP20] Mitsutoshi Nakamura, Jeffrey M. Verboon, Clara L. Prentiss, and Susan M. Parkhurst. The kinesin-like protein Pavarotti functions noncanonically to regulate actin dynamics. Journal of Cell Biology, 219(9):e201912117, September 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/9/
e201912117/151940/The-kinesin-like-protein-Pavarotti-
functions.

Noda:2020:LLP

[NWZ20] Nobuo N. Noda, Zheng Wang, and Hong Zhang. Liquid–liquid phase separation in autophagy. Journal of Cell Biology, 219(8): e202004062, August 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/219/8/e202004062/151909/Liquid-liquid-phase-
separation-in-autophagyPhase.

Nakayama:2021:PCP

[NYN^{+}21] Shogo Nakayama, Tomoki Yano, Toshinori Namba, Satoshi Konishi, Maki Takagishi, Elisa Herawati, Tomoki Nishida, Yasuo Imoto, Shuji Ishihara, Masahide Takahashi, Ken’ya Furuta, Kazuhiro Oiwa, Atsushi Tamura, and Sachiko Tsukita. Planar cell polarity induces local microtubule bundling for coordinated ciliary beating. Journal of Cell Biology, 220(7):e202010034, July 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/7/
e202010034/212042/Planar-cell-polarity-induces-local-
microtubule.

Oshima:2021:PIM

[OCB^{+}21] Yumiko Oshima, Etienne Cartier, Liron Boyman, Nicolas Verhoeven, Brian M. Polster, Weiliang Huang, Maureen Kane, W. Jonathan Lederer, and Mariusz Karbowski. Parkin-independent mitophagy via drp1-mediated outer membrane severing and inner membrane ubiquitination. Journal of Cell Biology, 220(6):e202006043, June 7, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/6/
e202006043/211984/Parkin-independent-mitophagy-via-Drp1-
mediated.

Otto:2021:PCC

[OCLB21] George Maxwell Otto, Tia Cheunkarndee, Jessica Mae Leslie, and Gloria Ann Brar. Programmed cortical ER collapse drives selective ER degradation and inheritance in yeast meiosis. Journal of Cell Biology, 220(12):e202108105, December 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
12/e202108105/212710/Programmed-cortical-ER-collapse-
drives-selective.

ODonnell:2020:SGP

[O’D20a] Marie Anne O’Donnell. Susana Godinho: Placing cell biology at the center of cancer research. Journal of Cell Biology, 219(2): e202001060, February 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/2/e202001060/133642/Susana-Godinho-
Placing-cell-biology-at-the-center.

ODonnell:2020:YSH

[O’D20b] Marie Anne O’Donnell. Yan Song: How time flies. Journal of Cell Biology, 219(1):e201912056, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

ODonnell:2022:SCY

[O’D22] Allyson F. O’Donnell. A second chance at yeast early endosomes. Journal of Cell Biology, 221(11):e202210014, November 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/
11/e202210014/213557/A-second-chance-at-yeast-early-
endosomesIst1.

OShaughnessy:2023:TEP

[OHHR23] William J. O’Shaughnessy, Xiaoyu Hu, Sarah Ana Henriquez, and Michael L. Reese. Toxoplasma ERK7 protects the apical complex from premature degradation. Journal of Cell Biology, 222(6):e202209098, June 5, 2023. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/222/6/
e202209098/214008/Toxoplasma-ERK7-protects-the-apical-
complex-from.

Ozawa:2020:AJR

[OHY^{+}20] Masayuki Ozawa, Sylvain Hiver, Takaki Yamamoto, Tatsuo Shibata, Srigokul Upadhyayula, Yuko Mimori-Kiyosue, and Masatoshi Takeichi. Adherens junction regulates cryptic lamellipodia formation for epithelial cell migration. Journal of Cell Biology, 219(10):e202006196, October 5, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/10/
e202006196/152072/Adherens-junction-regulates-cryptic-
lamellipodia.

Odierna:2020:DSS

[OKH^{+}20] G. Lorenzo Odierna, Sarah K. Kerwin, Lucy E. Harris, Grace Ji eun Shin, Nickolas A. Lavidis, Peter G. Noakes, and S. Sean Millard. Dscam2 suppresses synaptic strength through a PI3K-dependent endosomal pathway. Journal of Cell Biology, 219(6):e201909143, June 1, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/6/
e201909143/151621/Dscam2-suppresses-synaptic-strength-
through-a-PI3K.

Ono:2022:MSC

[OMI22] Yumiko Ono, Kenji Matsuzawa, and Junichi Ikenouchi. mTORC2 suppresses cell death induced by hypo-osmotic stress by promoting sphingomyelin transport. Journal of Cell Biology, 221(4):e202106160, April 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/221/4/e202106160/213090/mTORC2-suppresses-
cell-death-induced-by-hypo.

Opalko:2022:AAC

[OMK^{+}22] Hannah E. Opalko, Kristi E. Miller, Hyun-Soo Kim, Cesar Augusto Vargas-Garcia, Abhyudai Singh, Michael-Christopher Keogh, and James B. Moseley. Arf6 anchors Cdr2 nodes at the cell cortex to control cell size at division. Journal of Cell Biology, 221(2):e202109152, February 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/221/2/e202109152/212942/Arf6-
anchors-Cdr2-nodes-at-the-cell-cortex-to.

Osseni:2020:HRM

[ORCT^{+}20] Alexis Osseni, Aymeric Ravel-Chapuis, Jean-Luc Thomas, Vincent Gache, Laurent Schaeffer, and Bernard J. Jasmin. HDAC6 regulates microtubule stability and clustering of AChRs at neuromuscular junctions. Journal of Cell Biology, 219(8): e201901099, August 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/219/8/e201901099/151966/HDAC6-regulates-
microtubule-stability-and.

Orii:2021:TPT

[OTOF21] Minami Orii, Takuma Tsuji, Yuta Ogasawara, and Toyoshi Fujimoto. Transmembrane phospholipid translocation mediated by Atg9 is involved in autophagosome formation. Journal of Cell Biology, 220(3):e202009194, March 1, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/3/
e202009194/211678/Transmembrane-phospholipid-translocation-
mediated.

Overholtzer:2021:CDL

[Ove21] Michael Overholtzer. Cell death leaves a new TRAIL. Journal of Cell Biology, 220(11):e202109018, November 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/11/
e202109018/212699/Cell-death-leaves-a-new-TRAILCell-
death-leaves-a.

Olivas:2023:AVC

[OWY^{+}23] Taryn J. Olivas, Yumei Wu, Shenliang Yu, Lin Luan, Peter Choi, Emily D. Guinn, Shanta Nag, Pietro V. De Camilli, Kallol Gupta, and Thomas J. Melia. ATG9 vesicles comprise the seed membrane of mammalian autophagosomes. Journal of Cell Biology, 222(7):e202208088, July 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/222/7/e202208088/214071/ATG9-
vesicles-comprise-the-seed-membrane-of.

OBrien:2023:GCP

[OYJJ23] Caitlin E. O’Brien, Susan H. Younger, Lily Yeh Jan, and Yuh Nung Jan. The GARP complex prevents sterol accumulation at the trans-Golgi network during dendrite remodeling. Journal of Cell Biology, 222(1):e202112108, January 2, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/
1/e202112108/213548/The-GARP-complex-prevents-sterol-
accumulation-at.

Ohshima:2022:NDF

[OYS^{+}22] Tomoko Ohshima, Hayashi Yamamoto, Yuriko Sakamaki, Chieko Saito, and Noboru Mizushima. NCOA4 drives ferritin phase separation to facilitate macroferritinophagy and microferritinophagy. Journal of Cell Biology, 221(10):e202203102, October 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/10/e202203102/213442/NCOA4-drives-ferritin-phase-
separation-to.

Ohta:2021:PLK

[OZW^{+}21] Midori Ohta, Zhiling Zhao, Di Wu, Shaohe Wang, Jennifer L. Harrison, J. Sebastián Gómez-Cavazos, Arshad Desai, and Karen F. Oegema. Polo-like kinase 1 independently controls microtubule-nucleating capacity and size of the centrosome. Journal of Cell Biology, 220(2):e202009083, February 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
2/e202009083/211652/Polo-like-kinase-1-independently-
controls.

Prechova:2022:PMC

[PAS^{+}22] Magdalena Prechova, Zuzana Adamova, Anna-Lena Schweizer, Miloslava Maninova, Andreas Bauer, Delf Kah, Samuel M. Meier-Menches, Gerhard Wiche, Ben Fabry, and Martin Gregor. Plectin-mediated cytoskeletal crosstalk controls cell tension and cohesion in epithelial sheets. Journal of Cell Biology, 221(3): e202105146, March 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/221/3/e202105146/212995/Plectin-mediated-
cytoskeletal-crosstalk-controls.

Pennauer:2022:SSF

[PBPBS22] Mirjam Pennauer, Katarzyna Buczak, Cristina Prescianotto-Baschong, and Martin Spiess. Shared and specific functions of Arfs 1–5 at the Golgi revealed by systematic knockouts. Journal of Cell Biology, 221(1):e202106100, January 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/
1/e202106100/212750/Shared-and-specific-functions-of-
Arfs-1-5-at-the.

Poser:2020:APC

[PCGB20] Elena Poser, Renaud Caous, Ulrike Gruneberg, and Francis A. Barr. Aurora A promotes chromosome congression by activating the condensin-dependent pool of KIF4A. Journal of Cell Biology, 219(2):e201905194, February 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/2/e201905194/133548/Aurora-A-promotes-chromosome-
congression-by.

Pacheco:2023:PPD

[PCZ^{+}23] Jonathan Pacheco, Anna C. Cassidy, James P. Zewe, Rachel C. Wills, and Gerald R. V. Hammond. PI(4,5)P 2 diffuses freely in the plasma membrane even within high-density effector protein complexes. Journal of Cell Biology, 222(2):e202204099, February 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/
2/e202204099/213711/PI-4-5-P2-diffuses-freely-in-the-
plasma-membrane.

Paniza:2020:PSC

[PDW^{+}20] Theodore Paniza, Madhura Deshpande, Ning Wang, Ryan O’Neil, Michael V. Zuccaro, Morgan Elizabeth Smith, Advaitha Madireddy, Daylon James, Joseph Ecker, Zev Rosenwaks, Dieter Egli, and Jeannine Gerhardt. Pluripotent stem cells with low differentiation potential contain incompletely reprogrammed DNA replication. Journal of Cell Biology, 219(9):e201909163, September 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/219/9/e201909163/151954/Pluripotent-stem-
cells-with-low-differentiation.

Platta:2022:IME

[PE22] Harald W. Platta and Ralf Erdmann. The ides of MARCH5: The E3 ligase essential for peroxisome degradation by pexophagy. Journal of Cell Biology, 221(1):e202111008, January 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/
1/e202111008/212897/The-ides-of-MARCH5-The-E3-ligase-
essential-for.

Pederson:2022:SPV

[Ped22] Thoru Pederson. Sheldon Penman: Visionary of cell form and function. Journal of Cell Biology, 221(7):e202205033, July 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/
7/e202205033/213253/Sheldon-Penman-Visionary-of-cell-
form-and.

Phuyal:2021:WLW

[PF21] Santosh Phuyal and Hesso Farhan. Want to leave the ER? We offer vesicles, tubules, and tunnels. Journal of Cell Biology, 220(6):e202104062, June 7, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/220/6/e202104062/212113/Want-
to-leave-the-ER-We-offer-vesicles-tubules-and.

Pacheco-Fernandez:2020:NRE

[PFPB^{+}20] Natalia Pacheco-Fernandez, Mehrshad Pakdel, Birgit Blank, Ismael Sanchez-Gonzalez, Kathrin Weber, Mai Ly Tran, Tobias Karl-Heinz Hecht, Renate Gautsch, Gisela Beck, Franck Perez, Angelika Hausser, Stefan Linder, and Julia von Blume.

Nucleobindin-1 regulates ECM degradation by promoting intra-Golgi trafficking of MMPs. Journal of Cell Biology, 219(8): e201907058, August 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/8/e201907058/151825/Nucleobindin-
1-regulates-ECM-degradation-by.

Paramasivam:2022:EED

[PFS^{+}22] Prasath Paramasivam, Christian Franke, Martin Stöter, Andreas Höijer, Stefano Bartesaghi, Alan Sabirsh, Lennart Lindfors, Marianna Yanez Arteta, Anders Dahlén, Annette Bak, Shalini Andersson, Yannis Kalaidzidis, Marc Bickle, and Marino Zerial. Endosomal escape of delivered mRNA from endosomal recycling tubules visualized at the nanoscale. Journal of Cell Biology, 221(2):e202110137, February 7, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/2/
e202110137/212896/Endosomal-escape-of-delivered-mRNA-
from-endosomal.

Petzoldt:2020:RBP

[PGD^{+}20] Astrid G. Petzoldt, Torsten W. B. Götz, Jan Heiner Driller, Janine Lützkendorf, Suneel Reddy-Alla, Tanja Matkovic-Rachid, Sunbin Liu, Elena Knoche, Sara Mertel, Vladimir Ugorets, Martin Lehmann, Niraja Ramesh, Christine Brigitte Beuschel, Benno Kuropka, Christian Freund, Ulrich Stelzl, Bernhard Loll, Fan Liu, Markus C. Wahl, and Stephan J. Sigrist. RIMbinding protein couples synaptic vesicle recruitment to release sites. Journal of Cell Biology, 219(7):e201902059, July 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/7/e201902059/151735/RIM-binding-protein-couples-
synaptic-vesicle.

Panicker:2021:CBP

[PGDD21] Nikhil Panicker, Preston Ge, Valina L. Dawson, and Ted M. Dawson. The cell biology of Parkinson’s disease. Journal of Cell Biology, 220(4):??, April 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:/

/rupress.org/jcb/article/220/4/e202012095/211906/The-
cell-biology-of-Parkinson-s-diseaseThe-cell.

Parmar:2023:RMD

[PGH^{+}23] Sneha Parmar, Samuel J. Gonzalez, Julia M. Heckel, Soumya Mukherjee, Mark McClellan, Duncan J. Clarke, Marnie Johansson, Damien Tank, Athena Geisness, David K. Wood, and Melissa K. Gardner. Robust microtubule dynamics facilitate low-tension kinetochore detachment in metaphase. Journal of Cell Biology, 222(8):e202202085, August 7, 2023. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/222/8/
e202202085/214105/Robust-microtubule-dynamics-facilitate-
low-tension.

Peng:2021:MKC

[PGW^{+}21] Yi-Jheng Peng, Junhua Geng, Ying Wu, Cristian Pinales, Jennifer Langen, Yen-Ching Chang, Christopher Buser, and Karen T. Chang. Minibrain kinase and calcineurin coordinate activity-dependent bulk endocytosis through synaptojanin. Journal of Cell Biology, 220(12):e202011028, December 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/12/
e202011028/212674/Minibrain-kinase-and-calcineurin-coordinate.

Prinz:2020:FP

[PH20] William A. Prinz and James H. Hurley. A firehose for phospholipids. Journal of Cell Biology, 219(5):e202003132, May 4, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/5/
e202003132/151705/A-firehose-for-phospholipidsA-firehose-
for.

Pineiro-Hermida:2020:TTP

[PHAM^{+}20] Sergio Piñeiro-Hermida, Chiara Autilio, Paula Martínez, Fátima Bosch, Jesús Pérez-Gil, and Maria A. Blasco. Telomerase treatment prevents lung profibrotic pathologies associated with physiological aging. Journal of Cell Biology, 219(10): e202002120, October 5, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/10/e202002120/152010/Telomerase-
treatment-prevents-lung-profibrotic.

Pedersen:2020:SRC

[PHMD20] Ross T. A. Pedersen, Julian E. Hassinger, Paul Marchando, and David G. Drubin. Spatial regulation of clathrinmediated endocytosis through position-dependent site maturation. Journal of Cell Biology, 219(11):e202002160, November 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/
11/e202002160/211446/Spatial-regulation-of-clathrin-
mediated.

Pleiner:2023:SFM

[PHT^{+}23] Tino Pleiner, Masami Hazu, Giovani Pinton Tomaleri, Vy N. Nguyen, Kurt Januszyk, and Rebecca M. Voorhees. A selectivity filter in the ER membrane protein complex limits protein misinsertion at the ER. Journal of Cell Biology, 222(8):e202212007, August 7, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/222/8/e202212007/214118/A-
selectivity-filter-in-the-ER-membrane-protein.

Pierchala:2020:NSY

[Pie20] Brian A. Pierchala. Necroptosis is SARMful to your health. Journal of Cell Biology, 219(8):e202006090, August 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/8/
e202006090/151962/Necroptosis-is-SARMful-to-your-healthNecroptosis.

Phalora:2022:MUT

[PK22] Prabhjeet Phalora and Paul Klenerman. MR1: an unconventional twist in the tail. Journal of Cell Biology, 221(12): e202211016, December 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/221/12/e202211016/213684/MR1-
An-unconventional-twist-in-the-tailMR1-An.

Parker:2023:QCE

[PK23] Melissa D. Parker and Katrin Karbstein. Quality control ensures fidelity in ribosome assembly and cellular health. Journal of Cell Biology, 222(4):e202209115, April 3, 2023. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/222/4/
e202209115/213871/Quality-control-ensures-fidelity-in-
ribosome.

Parton:2020:CLS

[PKA20] Robert G. Parton, Michael M. Kozlov, and Nicholas Ariotti. Caveolae and lipid sorting: Shaping the cellular response to stress. Journal of Cell Biology, 219(4):e201905071, April 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/4/
e201905071/133844/Caveolae-and-lipid-sorting-Shaping-
the-cellular.

Park:2022:VIR

[PKC^{+}22] Hyun Gwan Park, Yeongjin David Kim, Eunsang Cho, Ting-Yi Lu, Chi-Kuang Yao, Jihye Lee, and Seungbok Lee. Vav independently regulates synaptic growth and plasticity through distinct actin-based processes. Journal of Cell Biology, 221 (10):e202203048, October 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/221/10/e202203048/213401/Vav-
independently-regulates-synaptic-growth-and.

Panda:2020:TSR

[PKD^{+}20] Pallavi Panda, Levente Kovacs, Nikola Dzhindzhev, Agnieszka Fatalska, Veronica Persico, Marco Geymonat, Maria Giovanna Riparbelli, Giuliano Callaini, and David M. Glover. Tissue specific requirement of Drosophila Rcd4 for centriole duplication and ciliogenesis. Journal of Cell Biology, 219(8):e201912154, August 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/8/e201912154/151861/Tissue-specific-requirement-
of-Drosophila-Rcd4-for.

Pemberton:2020:DSD

[PKH^{+}20] Joshua G. Pemberton, Yeun Ju Kim, Jana Humpolickova, Andrea Eisenreichova, Nivedita Sengupta, Daniel J. Toth, Evzen Boura, and Tamas Balla. Defining the subcellular distribution and metabolic channeling of phosphatidylinositol. Journal of Cell Biology, 219(3):e201906130, March 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Pandey:2020:TIS

[PKY^{+}20] Nootan Pandey, Daniel Keifenheim, Makoto Michael Yoshida, Victoria A. Hassebroek, Caitlin Soroka, Yoshiaki Azuma, and Duncan J. Clarke. Topoisomerase II SUMOylation activates a metaphase checkpoint via Haspin and Aurora B kinases. Journal of Cell Biology, 219(1):e201807189, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Park:2022:ITO

[PL22] Kwangjin Park and Michel R. Leroux. IFT trains overcome an NPHP module barrier at the transition zone. Journal of Cell Biology, 221(1):e202112015, January 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:/

/rupress.org/jcb/article/221/1/e202112015/212898/IFT-
trains-overcome-an-NPHP-module-barrier-at-the.

Parker:2023:EMM

[PLG^{+}23] Sara S. Parker, Kenneth Tran Ly, Adam D. Grant, Jillian Sweetland, Ashley M. Wang, James D. Parker, Mackenzie R. Roman, Kathylynn Saboda, Denise J. Roe, Megha Padi, Charles W. Wolgemuth, Paul Langlais, and Ghassan Mouneimne. EVL and MIM/MTSS1 regulate actin cytoskeletal remodeling to promote dendritic filopodia in neurons. Journal of Cell Biology, 222(5):e202106081, May 1, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:/

/rupress.org/jcb/article/222/5/e202106081/213902/EVL-
and-MIM-MTSS1-regulate-actin-cytoskeletal.

Pfisterer:2020:FRD

[PLL^{+}20] Karin Pfisterer, James Levitt, Campbell D. Lawson, Richard J. Marsh, John M. Heddleston, Eric Wait, Simon Morris Ameer-Beg, Susan Cox, and Maddy Parsons. FMNL2 regulates dynamics of fascin in filopodia. Journal of Cell Biology, 219(5): e201906111, May 4, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/219/5/e201906111/151666/FMNL2-regulates-
dynamics-of-fascin-in.

Paul:2020:SCE

[PMB^{+}20] Danielle M. Paul, Judith Mantell, Ufuk Borucu, Jennifer Coombs, Katherine J. Surridge, John M. Squire, Paul Verkade, and Mark P. Dodding. In situ cryo-electron tomography reveals filamentous actin within the microtubule lumen. Journal of Cell Biology, 219(9):e201911154, September 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/
9/e201911154/151828/In-situ-cryo-electron-tomography-
reveals.

Pande:2022:FOB

[PMB^{+}22] Vani Pande, Nivedita Mitra, Saket Rahul Bagde, Ramanujam Srinivasan, and Pananghat Gayathri. Filament organization of the bacterial actin MreB is dependent on the nucleotide state. Journal of Cell Biology, 221(5):e202106092, May 2, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/5/e202106092/213108/Filament-organization-of-the-
bacterial-actin-MreB.

Perez-Moreno:2023:DSO

[PMSO^{+}23] Juan José Pérez-Moreno, Rebecca C. Smith, Megan K. Oliva, Filomena Gallo, Shainy Ojha, Karin H. Müller, and Cahir J. O’Kane. Drosophila SPG12 ortholog, reticulon-like 1, governs presynaptic ER organization and Ca^{2+} dynamics. Journal of Cell Biology, 222(6):e202112101, June 5, 2023. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/222/6/
e202112101/213965/Drosophila-SPG12-ortholog-reticulon-
like-1-governs.

Phatarpekar:2020:SCR

[POL^{+}20] Prasad V. Phatarpekar, Brittany L. Overlee, Alexander Leehan, Katelynn M. Wilton, Hyoungjun Ham, and Daniel D. Billadeau. The septin cytoskeleton regulates natural killer cell lytic granule release. Journal of Cell Biology, 219(11): e202002145, November 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/219/11/e202002145/152040/The-
septin-cytoskeleton-regulates-natural-killer.

Puri:2021:WSE

[PPB^{+}21] Dharmendra Puri, Keerthana Ponniah, Kasturi Biswas, Atrayee Basu, Swagata Dey, Erik A. Lundquist, and Anindya Ghosh-Roy. Wnt signaling establishes the microtubule polarity in neurons through regulation of Kinesin-13. Journal of Cell Biology, 220(9):e202005080, September 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:/

/rupress.org/jcb/article/220/9/e202005080/212396/Wnt-
signaling-establishes-the-microtubule-polarity.

Petit-Pedrol:2021:RMN

[PPG21] Mar Petit-Pedrol and Laurent Groc. Regulation of membrane NMDA receptors by dynamics and protein interactions. Journal of Cell Biology, 220(1):e202006101, January 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/1/e202006101/211609/Regulation-of-membrane-NMDA-
receptors-by-dynamics.

Pejskova:2020:KCC

[PRB^{+}20] Petra Pejskova, Madeline Louise Reilly, Lucia Bino, Ondrej Bernatik, Linda Dolanska, Ranjani Sri Ganji, Zbynek Zdrahal, Alexandre Benmerah, and Lukas Cajanek. KIF14 controls ciliogenesis via regulation of Aurora A and is important for Hedgehog signaling. Journal of Cell Biology, 219(6):e201904107, June 1, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/
6/e201904107/151721/KIF14-controls-ciliogenesis-via-
regulation-of.

Ponce:2023:STS

[PRMF^{+}23] Maria Laura Sosa Ponce, Mayrene Horta Remedios, Sarah Moradi-Fard, Jennifer A. Cobb, and Vanina Zaremberg. SIR telomere silencing depends on nuclear envelope lipids and modulates sensitivity to a lysolipid. Journal of Cell Biology, 222 (7):e202206061, July 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/222/7/e202206061/214023/SIR-telomere-
silencing-depends-on-nuclear-envelope.

Prokop:2020:COA

[Pro20] Andreas Prokop. Cytoskeletal organization of axons in vertebrates and invertebrates. Journal of Cell Biology, 219(7): e201912081, July 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/7/e201912081/151734/Cytoskeletal-
organization-of-axons-in-vertebrates.

Pereira:2023:ECE

[PSA^{+}23] Conceição Pereira, Danièle Stalder, Georgina S. F. Anderson, Amber S. Shun-Shion, Jack Houghton, Robin Antrobus, Michael A. Chapman, Daniel J. Fazakerley, and David C. Gershlick. The exocyst complex is an essential component of the mammalian constitutive secretory pathway. Journal of Cell Biology, 222(5):e202205137, May 1, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:/

/rupress.org/jcb/article/222/5/e202205137/213943/The-
exocyst-complex-is-an-essential-component-of.

Park:2020:WRS

[PSC^{+}20] Elizabeth M. Park, Phillip M. Scott, Kevin Clutario, Katelyn B. Cassidy, Kevin Zhan, Scott A. Gerber, and Andrew J. Holland. WBP11 is required for splicing the TUBGCP6 pre-mRNA to promote centriole duplication. Journal of Cell Biology, 219(1): e201904203, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Ptak:2021:PDM

[PSP^{+}21] Christopher Ptak, Natasha O. Saik, Ashwini Premashankar, Diego L. Lapetina, John D. Aitchison, Ben Montpetit, and Richard W. Wozniak. Phosphorylation-dependent mitotic SUMOylation drives nuclear envelope–chromatin interactions. Journal of Cell Biology, 220(12):e202103036, December 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/12/
e202103036/212843/Phosphorylation-dependent-mitotic-
SUMOylation.

Penfield:2020:RLS

[PSS^{+}20] Lauren Penfield, Raakhee Shankar, Erik Szentgyörgyi, Alyssa Laffitte, Michael Sean Mauro, Anjon Audhya, Thomas Müller-Reichert, and Shirin Bahmanyar. Regulated lipid synthesis and LEM2/CHMP7 jointly control nuclear envelope closure. Journal of Cell Biology, 219(5):e201908179, May 4, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/5/
e201908179/151636/Regulated-lipid-synthesis-and-LEM2-
CHMP7-jointly.

Parchure:2022:LLP

[PTS^{+}22] Anup Parchure, Meng Tian, Danièle Stalder, Cierra K. Boyer, Shelby C. Bearrows, Kristen E. Rohli, Jianchao Zhang, Felix Rivera-Molina, Bulat R. Ramazanov, Sushil K. Mahata, Yanzhuang Wang, Samuel B. Stephens, David C. Gershlick, and Julia von Blume. Liquid–liquid phase separation facilitates the biogenesis of secretory storage granules. Journal of Cell Biology, 221(12):e202206132, December 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/
12/e202206132/213510/Liquid-liquid-phase-separation-
facilitates-the.

Perez-Vale:2021:MIM

[PVYJ^{+}21] Kia Z. Perez-Vale, Kristi D. Yow, Ruth I. Johnson, Amy E. Byrnes, Tara M. Finegan, Kevin C. Slep, and Mark Peifer. Multivalent interactions make adherens junction–cytoskeletal linkage robust during morphogenesis. Journal of Cell Biology, 220(12): e202104087, December 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/12/e202104087/212790/Multivalent-
interactions-make-adherens-junction.

Pedersen:2020:PME

[PWW^{+}20] Nina Marie Pedersen, Eva Maria Wenzel, Ling Wang, Sandra Antoine, Philippe Chavrier, Harald Stenmark, and Camilla Raiborg. Protrudin-mediated ER–endosome contact sites promote MT1-MMP exocytosis and cell invasion. Journal of Cell Biology, 219(8):e202003063, August 3, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/8/
e202003063/151827/Protrudin-mediated-ER-endosome-contact-
sites.

Petsalaki:2021:ACI

[PZ21] Eleni Petsalaki and George Zachos. An ATM–chk2–INCENP pathway activates the abscission checkpoint. Journal of Cell Biology, 220(2):e202008029, February 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/220/2/e202008029/211635/An-
ATM-Chk2-INCENP-pathway-activates-the.

Pal:2021:UMB

[PZWW21] Kaushik Pal, Yuanchang Zhao, Yongliang Wang, and Xuefeng Wang. Ubiquitous membrane-bound DNase activity in podosomes and invadopodia. Journal of Cell Biology, 220(7): e202008079, July 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/220/7/e202008079/212028/Ubiquitous-membrane-
bound-DNase-activity-in.

Qi:2020:PDA

[QLC^{+}20] Yanmei Qi, Jie Liu, Joshua Chao, Peter A. Greer, and Shaohua Li. PTEN dephosphorylates Abi1 to promote epithelial morphogenesis. Journal of Cell Biology, 219(9): e201910041, September 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/219/9/e201910041/151941/PTEN-
dephosphorylates-Abi1-to-promote-epithelial.

Qiu:2023:KAF

[QZX23] Rongde Qiu, Jun Zhang, and Xin Xiang. Kinesin-1 autoinhibition facilitates the initiation of dynein cargo transport. Journal of Cell Biology, 222(3):e202205136, March 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/3/
e202205136/213766/Kinesin-1-autoinhibition-facilitates-
the.

Rios-Barrera:2022:EAA

[RBL22] Luis Daniel Ríos-Barrera and Maria Leptin. An endosomeassociated actin network involved in directed apical plasma membrane growth. Journal of Cell Biology, 221(3):e202106124, March 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/3/e202106124/212975/An-endosome-associated-actin-
network-involved-in.

Robertson:2021:LES

[RCA^{+}21] Tanner F. Robertson, Pragati Chengappa, Daniela Gomez Atria, Christine F. Wu, Lyndsay Avery, Nathan H. Roy, Ivan Maillard, Ryan J. Petrie, and Janis K. Burkhardt. Lymphocyte egress signal sphingosine-1-phosphate promotes ERM-guided, bleb-based migration. Journal of Cell Biology, 220(6):e202007182, June 7, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/6/
e202007182/211919/Lymphocyte-egress-signal-sphingosine-
1-phosphate.

Rahi:2023:NCS

[RCA^{+}23] Amit Rahi, Manas Chakraborty, Shivangi Agarwal, Kristen M. Vosberg, Shivani Agarwal, Annie Y. Wang, Richard J. McKenney, and Dileep Varma. The Ndc80-Cdt1-Ska1 complex is a central processive kinetochore–microtubule coupling unit. Journal of Cell Biology, 222(8):e202208018, August 7, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/
8/e202208018/214139/The-Ndc80-Cdt1-Ska1-complex-is-a-
central.

Reina-Campos:2020:CSG

[RCDMM20] Miguel Reina-Campos, Maria T. Diaz-Meco, and Jorge Moscat. The complexity of the serine glycine one-carbon pathway in cancer. Journal of Cell Biology, 219(1):e201907022, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Renne:2022:SCD

[RCF^{+}22] Mike F. Renne, Robin A. Corey, Joana Veríssimo Ferreira, Phillip J. Stansfeld, and Pedro Carvalho. Seipin concentrates distinct neutral lipids via interactions with their acyl chain carboxyl esters. Journal of Cell Biology, 221(9):e202112068, September 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/9/e202112068/213387/Seipin-concentrates-distinct-
neutral-lipids-via.

Rahman:2020:CEP

[RCH^{+}20] Mohammad Rahman, Irene Y. Chang, Adam Harned, Richa Maheshwari, Kwabena Amoateng, Kedar Narayan, and Orna Cohen-Fix. C. elegans pronuclei fuse after fertilization through a novel membrane structure. Journal of Cell Biology, 219(2):e201909137, February 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/219/2/e201909137/132753/C-
elegans-pronuclei-fuse-after-fertilization.

Ralhan:2023:AEL

[RCM^{+}23a] Isha Ralhan, Jinlan Chang, Matthew J. Moulton, Lindsey D. Goodman, Nathanael Y. J. Lee, Greg Plummer, H. Amalia Pasolli, Doreen Matthies, Hugo J. Bellen, and Maria S. Ioannou. Autolysosomal exocytosis of lipids protect neurons from ferroptosis. Journal of Cell Biology, 222(6):e202207130, June 5, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/6/
e202207130/214014/Autolysosomal-exocytosis-of-lipids-
protect-neurons.

Rawat:2023:RBA

[RCM^{+}23b] Shalini Rawat, Dhruba Chatterjee, Rituraj Marwaha, Gitanjali Charak, Gaurav Kumar, Shrestha Shaw, Divya Khatter, Sheetal Sharma, Cecilia de Heus, Nalan Liv, Judith Klumperman, Amit Tuli, and Mahak Sharma. RUFY1 binds Arl8b and mediates endosome-to-TGN CI-M6PR retrieval for cargo sorting to lysosomes. Journal of Cell Biology, 222(1): e202108001, January 2, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/222/1/e202108001/213572/RUFY1-binds-
Arl8b-and-mediates-endosome-to-TGN-CI.

Roney:2022:NET

[RCS22] Joseph C. Roney, Xiu-Tang Cheng, and Zu-Hang Sheng. Neuronal endolysosomal transport and lysosomal functionality in maintaining axonostasis. Journal of Cell Biology, 221(3): e202111077, March 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/221/3/e202111077/213000/Neuronal-endolysosomal-
transport-and-lysosomal.

Reichmann:2020:TIE

[RDL^{+}20] Judith Reichmann, Karen Dobie, Lisa M. Lister, James H. Crichton, Diana Best, Marie MacLennan, David Read, Eleanor S. Raymond, Chao-Chun Hung, Shelagh Boyle, Katsuhiko Shirahige, Howard J. Cooke, Mary Herbert, and Ian R. Adams. Tex19.1 inhibits the N-end rule pathway and maintains acetylated SMC3 cohesin and sister chromatid cohesion in oocytes. Journal of Cell Biology, 219(5):e201702123, May 4, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/
5/e201702123/151600/Tex19-1-inhibits-the-N-end-rule-
pathway-and.

Raso:2020:ISG

[RDW^{+}20] Maria Chiara Raso, Nikola Djoric, Franziska Walser, Sandra Hess, Fabian Marc Schmid, Sibylle Burger, Klaus-Peter Knobeloch, and Lorenza Penengo. Interferon-stimulated gene 15 accelerates replication fork progression inducing chromosomal breakage. Journal of Cell Biology, 219(8):e202002175, August 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/8/
e202002175/151903/Interferon-stimulated-gene-15-accelerates.

Renne:2020:ORP

[RE20] Mike F. Renne and Brooke M. Emerling. ORP5 regulates PI(4)P on the lipid droplet: Novel players on the monolayer. Journal of Cell Biology, 219(1):e201912010, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Ryder:2020:CER

[RFL20] Pearl V. Ryder, Junnan Fang, and Dorothy A. Lerit. c entrocortin RNA localization to centrosomes is regulated by FMRP and facilitates error-free mitosis. Journal of Cell Biology, 219(12):e202004101, December 7, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/12/
e202004101/211538/centrocortin-RNA-localization-to-centrosomes-
is.

Rothlin:2023:WAG

[RG23] Carla V. Rothlin and Sourav Ghosh. When aging gets on the way of disposal: Senescent cell suppression of efferocytosis. Journal of Cell Biology, 222(2):e202212023, February 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/222/2/e202212023/213792/When-aging-gets-
on-the-way-of-disposal-Senescent.

Rogers:2022:TLT

[RGK^{+}22] Sean Rogers, Long Gui, Anastasiia Kovalenko, Valeria Zoni, Maxime Carpentier, Kamran Ramji, Kalthoum Ben Mbarek, Amelie Bacle, Patrick Fuchs, Pablo Campomanes, Evan Reetz, Natalie Ortiz Speer, Emma Reynolds, Abdou Rachid Thiam, Stefano Vanni, Daniela Nicastro, and W. Mike Henne. Triglyceride lipolysis triggers liquid crystalline phases in lipid droplets and alters the LD proteome. Journal of Cell Biology, 221(11):e202205053, November 7, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/11/

e202205053/213472/Triglyceride-lipolysis-triggers-liquid-
crystalline.

Roy:2022:NCA

[RGP^{+}22] Abhijit Deb Roy, Evan G. Gross, Gayatri S. Pillai, Shailaja Seetharaman, Sandrine Etienne-Manneville, and Takanari Inoue. Non-catalytic allostery in α-TAT1 by a phospho-switch drives dynamic microtubule acetylation. Journal of Cell Biology, 221(11):e202202100, November 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/221/11/e202202100/213540/Non-
catalytic-allostery-in-TAT1-by-a-phospho.

Ramsey:2023:PRB

[RH23] Arren Ramsey and Eric J. Huang. Plastin 3 rescues BDNF-TrkB signaling in spinal muscular atrophy. Journal of Cell Biology, 222(3):e202301036, March 6, 2023. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/222/3/
e202301036/213868/Plastin-3-rescues-BDNF-TrkB-signaling-
in-spinal.

Rasmussen:2022:NAS

[RKLJ22] Nikoline Lander Rasmussen, Athanasios Kournoutis, Trond Lamark, and Terje Johansen. NBR1: The archetypal selective autophagy receptor. Journal of Cell Biology, 221(11): e202208092, November 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/221/11/e202208092/213552/NBR1-The-archetypal-
selective-autophagy.

Rossi:2020:ESC

[RLK^{+}20] Guendalina Rossi, Dante Lepore, Lillian Kenner, Alexander B. Czuchra, Melissa Plooster, Adam Frost, Mary Munson, and Patrick Brennwald. Exocyst structural changes associated with activation of tethering downstream of Rho/Cdc42 GT-Pases. Journal of Cell Biology, 219(2):e201904161, February 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/2/
e201904161/133563/Exocyst-structural-changes-associated-
with.

Rondelet:2020:CAI

[RLS^{+}20] Arnaud Rondelet, Yu-Chih Lin, Divya Singh, Arthur T. Porfetye, Harish C. Thakur, Andreas Hecker, Pia Brinkert, Nadine Schmidt, Shweta Bendre, Franziska Müller, Lisa Mazul, Per O. Widlund, Tanja Bange, Michael Hiller, Ingrid R. Vetter, and Alexander W. Bird. Clathrin’s adaptor interaction sites are repurposed to stabilize microtubules during mitosis. Journal of Cell Biology, 219(2):e201907083, February 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/2/
e201907083/133599/Clathrin-s-adaptor-interaction-sites-
are.

Rampello:2020:TAD

[RLV^{+}20] Anthony J. Rampello, Ethan Laudermilch, Nidhi Vishnoi, Sarah M. Prophet, Lin Shao, Chenguang Zhao, C. Patrick Lusk, and Christian Schlieker. Torsin ATPase deficiency leads to defects in nuclear pore biogenesis and sequestration of MLF2. Journal of Cell Biology, 219(6):e201910185, June 1, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/6/
e201910185/151708/Torsin-ATPase-deficiency-leads-to-
defects-in.

Rice:2021:MFP

[RMA21] Luke M. Rice, Michelle Moritz, and David A. Agard. Microtubules form by progressively faster tubulin accretion, not by nucleation–elongation. Journal of Cell Biology, 220(5): e202012079, May 3, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/5/e202012079/211894/Microtubules-
form-by-progressively-faster-tubulin.

Ronchi:2021:HPT

[RMM^{+}21] Paolo Ronchi, Giulia Mizzon, Pedro Machado, Edoardo D’Imprima, Benedikt T. Best, Lucia Cassella, Sebastian Schnorrenberg, Marta G. Montero, Martin Jechlinger, Anne Ephrussi, Maria Leptin, Julia Mahamid, and Yannick Schwab. Highprecision targeting workflow for volume electron microscopy. Journal of Cell Biology, 220(9):e202104069, September 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/9/

e202104069/212433/High-precision-targeting-workflow-
for-volume.

Reinisch:2021:MNL

[RP21] Karin M. Reinisch and William A. Prinz. Mechanisms of nonvesicular lipid transport. Journal of Cell Biology, 220(3): e202012058, March 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/220/3/e202012058/211813/Mechanisms-of-nonvesicular-
lipid.

Rabas:2021:PDP

[RPM^{+}21] Nicolas Rabas, Sarah Palmer, Louise Mitchell, Shehab Ismail, Andrea Gohlke, Joel S. Riley, Stephen W. G. Tait, Payam Gammage, Leandro Lemgruber Soares, Iain R. Macpherson, and Jim C. Norman. PINK1 drives production of mtDNA-containing extracellular vesicles to promote invasiveness. Journal of Cell Biology, 220(12):e202006049, December 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/12/e202006049/212691/PINK1-drives-production-of-
mtDNA-containing.

Rabanal-Ruiz:2021:MAS

[RRBW^{+}21] Yoana Rabanal-Ruiz, Adam Byron, Alexander Wirth, Ralitsa Madsen, Lucia Sedlackova, Graeme Hewitt, Glyn Nelson, Julian Stingele, Jimi C. Wills, Tong Zhang, André Zeug, Reinhard Fässler, Bart Vanhaesebroeck, Oliver D. K. Maddocks, Evgeni Ponimaskin, Bernadette Carroll, and Viktor I. Korolchuk. mTORC1 activity is supported by spatial association with focal adhesions. Journal of Cell Biology, 220(5):e202004010, May 3, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/5/e202004010/211831/mTORC1-activity-is-supported-
by-spatial.

Ramirez-Rios:2023:VSV

[RRCS^{+}23] Sacnicte Ramirez-Rios, Sung Ryul Choi, Chadni Sanyal, Thorsten B. Blum, Christophe Bosc, Fatma Krichen, Eric Denarier, Jean-Marc Soleilhac, Béatrice Blot, Carsten Janke, Virginie Stoppin-Mellet, Maria M. Magiera, Isabelle Arnal, Michel O. Steinmetz, and Marie-Jo Moutin. VASH1–SVBP and VASH2–SVBP generate different detyrosination profiles on microtubules. Journal of Cell Biology, 222(2):e202205096, February 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/2/
e202205096/213744/VASH1-SVBP-and-VASH2-SVBP-generate-
different.

Rottner:2022:WSM

[RS22] Klemens Rottner and Theresia E. B. Stradal. WASP stings into matrix to lead immune cell migration. Journal of Cell Biology, 221(2):e202112087, February 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/221/2/e202112087/212973/WASP-
stings-into-matrix-to-lead-immune-cell.

Rocha:2023:NEP

[RSB^{+}23] Helder Rocha, Patrícia A. Simões, Jacqueline Budrewicz, Pablo Lara-Gonzalez, Ana Xavier Carvalho, Julien Dumont, Arshad Desai, and Reto Gassmann. Nuclear-enriched protein phosphatase 4 ensures outer kinetochore assembly prior to nuclear dissolution. Journal of Cell Biology, 222(3): e202208154, March 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/222/3/e202208154/213846/Nuclear-enriched-
protein-phosphatase-4-ensures.

Ruehle:2020:SLB

[RSWP20] Marisa D. Ruehle, Alexander J. Stemm-Wolf, and Chad G. Pearson. Sas4 links basal bodies to cell division via Hippo signaling. Journal of Cell Biology, 219(8):e201906183, August 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/
8/e201906183/151794/Sas4-links-basal-bodies-to-cell-
division-via-Hippo.

Ruthnick:2021:TSY

[RVNS21] Diana Rüthnick, Jlenia Vitale, Annett Neuner, and Elmar Schiebel. The N-terminus of Sfi1 and yeast centrin Cdc31 provide the assembly site for a new spindle pole body. Journal of Cell Biology, 220(3):e202004196, March 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
3/e202004196/211743/The-N-terminus-of-Sfi1-and-yeast-
centrin-Cdc31.

Romero-Wolf:2020:NCN

[RWSZ^{+}20] Maile Romero-Wolf, Boyoung Shin, Wen Zhou, Maria Koizumi, Ellen V. Rothenberg, and Hiroyuki Hosokawa. Notch2 complements Notch1 to mediate inductive signaling that initiates early T cell development. Journal of Cell Biology, 219(10):e202005093, October 5, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/219/10/e202005093/152003/Notch2-complements-
Notch1-to-mediate-inductive.

Rong:2022:SCE

[RZN^{+}22] Yueguang Rong, Shen Zhang, Nilay Nandi, Zhe Wu, Linsen Li, Yang Liu, Yuehan Wei, Yuan Zhao, Weigang Yuan, Chuchu Zhou, Guanghua Xiao, Beth Levine, Nan Yan, Shan Mou, Liufu Deng, Zaiming Tang, Xiaoxia Liu, Helmut Kramer, and Qing Zhong. STING controls energy stressinduced autophagy and energy metabolism via STX17. Journal of Cell Biology, 221(7):e202202060, July 4, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/7/
e202202060/213198/STING-controls-energy-stress-induced-
autophagy-and.

Stritto:2021:DTR

[SBBJ21] Maria Rosaria Dello Stritto, Bernd Bauer, Pierre Barraud, and Verena Jantsch. DNA topoisomerase 3 is required for efficient germ cell quality control. Journal of Cell Biology, 220(6):e202012057, June 7, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:/

/rupress.org/jcb/article/220/6/e202012057/211935/DNA-
topoisomerase-3-is-required-for-efficient-germ.

Serena:2020:MBM

[SBEB20] Michela Serena, Ricardo Nunes Bastos, Paul R. Elliott, and Francis A. Barr. Molecular basis of MKLP2-dependent Aurora B transport from chromatin to the anaphase central spindle. Journal of Cell Biology, 219(7):e201910059, July 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/7/
e201910059/151730/Molecular-basis-of-MKLP2-dependent-
Aurora-B.

Stevenson:2021:GRI

[SBL^{+}21] Nicola L. Stevenson, Dylan J. M. Bergen, Yinhui Lu, M. Esther Prada-Sanchez, Karl E. Kadler, Chrissy L. Hammond, and David J. Stephens. Giantin is required for intracellular N-terminal processing of type I procollagen. Journal of Cell Biology, 220(6):e202005166, June 7, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/6/
e202005166/212045/Giantin-is-required-for-intracellular-
N-terminal.

Spriggs:2020:GAB

[SBV^{+}20] Chelsey C. Spriggs, Somayesadat Badieyan, Kristen J. Verhey, Michael A. Cianfrocco, and Billy Tsai. Golgi-associated BICD adaptors couple ER membrane penetration and disassembly of a viral cargo. Journal of Cell Biology, 219(5): e201908099, May 4, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/219/5/e201908099/151622/Golgi-associated-
BICD-adaptors-couple-ER-membrane.

Smith:2020:FSI

[SCB^{+}20] Jean A. Smith, Ennessa G. Curry, R. Eric Blue, Christine Roden, Samantha E. R. Dundon, Anthony Rodríguez-Vargas, Danielle C. Jordan, Xiaomin Chen, Shawn M. Lyons, John Crutchley, Paul Anderson, Marko E. Horb, Amy S. Gladfelter, and Jimena Giudice. FXR1 splicing is important for muscle development and biomolecular condensates in muscle cells. Journal of Cell Biology, 219(4): e201911129, April 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/219/4/e201911129/133869/FXR1-splicing-is-
important-for-muscle-development.

Skouloudaki:2019:YCT

[SCK^{+}19] Kassiani Skouloudaki, Ioannis Christodoulou, Dilan Khalili, Vasilios Tsarouhas, Christos Samakovlis, Pavel Tomancak, Elisabeth Knust, and Dimitrios K. Papadopoulos. Yorkie controls tube length and apical barrier integrity during airway development. Journal of Cell Biology, 218(8):2762–??, August 2019. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/218/8/2762. See cor-

rection [SCK^{+}23].

Sicari:2020:CRE

[SCK^{+}20a] Daria Sicari, Aristotelis Chatziioannou, Theodoros Koutsandreas, Roberto Sitia, and Eric Chevet. Correction: Role of the early secretory pathway in SARS-CoV-2 infection. Journal of Cell Biology, 219(9):e20200600508132020c, September 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/
9/e20200600508132020c/152042/Correction-Role-of-the-
early-secretory-pathway-in. See [SCK^{+}20b].

Sicari:2020:RES

[SCK^{+}20b] Daria Sicari, Aristotelis Chatziioannou, Theodoros Koutsandreas, Roberto Sitia, and Eric Chevet. Role of the early secretory pathway in SARS-CoV-2 infection. Journal of Cell Biology, 219(9):e202006005, September 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/219/9/e202006005/151984/Role-
of-the-early-secretory-pathway-in-SARS-CoV-2. See cor-

rection [SCK^{+}20a].

Skouloudaki:2023:CYC

[SCK^{+}23] Kassiani Skouloudaki, Ioannis Christodoulou, Dilan Khalili, Vasilios Tsarouhas, Christos Samakovlis, Pavel Tomancak, Elisabeth Knust, and Dimitrios K. Papadopoulos. Correction: Yorkie controls tube length and apical barrier integrity during airway development. Journal of Cell Biology, 222 (5):e20180912104072023c, May 1, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/222/5/e20180912104072023c/
214029/Correction-Yorkie-controls-tube-length-and-apical.

See [SCK^{+}19].

Sun:2021:KPL

[SCL^{+}21] Landi Sun, Lihong Cui, Zhen Liu, Qixuan Wang, Zhaoyu Xue, Menghua Wu, Tianhui Sun, Decai Mao, Jianquan Ni, José Carlos Pastor-Pareja, and Xin Liang. Katanin p60-like 1 sculpts the cytoskeleton in mechanosensory cilia. Journal of Cell Biology, 220 (1):e202004184, January 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/1/e202004184/211570/Katanin-p60-
like-1-sculpts-the-cytoskeleton-in.

Silva:2023:HSS

[SCN^{+}23] Ana Marta Silva, Fung-Yi Chan, Michael J. Norman, Ana Filipa Sobral, Esther Zanin, Reto Gassmann, Julio Monti Belmonte, and Ana Xavier Carvalho. β-heavy-spectrin stabilizes the constricting contractile ring during cytokinesis. Journal of Cell Biology, 222(1):e202202024, January 2, 2023. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/222/1/
e202202024/213538/heavy-spectrin-stabilizes-the-constricting.

Shkarina:2022:OAA

[SdCS^{+}22] Kateryna Shkarina, Eva Hasel de Carvalho, José Carlos Santos, Saray Ramos, Maria Leptin, and Petr Broz. Optogenetic activators of apoptosis, necroptosis, and pyroptosis. Journal of Cell Biology, 221(6):e202109038, June 6, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/6/
e202109038/213145/Optogenetic-activators-of-apoptosis-
necroptosis.

Martin:2022:CCA

[SDD^{+}22] Rebeca San Martin, Priyojit Das, Renata Dos Reis Marques, Yang Xu, Justin M. Roberts, Jacob T. Sanders, Rosela Golloshi, and Rachel Patton McCord. Chromosome compartmentalization alterations in prostate cancer cell lines model disease progression. Journal of Cell Biology, 221(2):e202104108, February 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/
2/e202104108/212899/Chromosome-compartmentalization-
alterations-in.

Schneider:2021:DSA

[SdRVH^{+}21] Isabell Schneider, Marta de Ruijter-Villani, M. Julius Hossain, Tom A. E. Stout, and Jan Ellenberg. Dual spindles assemble in bovine zygotes despite the presence of paternal centrosomes. Journal of Cell Biology, 220(11):e202010106, November 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/11/
e202010106/212655/Dual-spindles-assemble-in-bovine-zygotes-
despite.

Seaman:2021:DSE

[Sea21] Matthew N. J. Seaman. A dimmer switch for endosome-to-cell surface recycling. Journal of Cell Biology, 220(4):??, April 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/4/e202102130/211878/A-dimmer-switch-for-endosome-
to-cell-surface.

SethGHaddix:2021:LIU

[Set21] Matthew N. Rasband Seth G.Haddix . Lose it to use it. Journal of Cell Biology, 220(4):??, April 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/220/4/e202102030/211898/Lose-
it-to-use-itMyelination-of-terminal-motor.

Sandilands:2023:SGA

[SFC^{+}23] Emma Sandilands, Eva C. Freckmann, Erin M. Cumming, Alvaro Román-Fernández, Lynn McGarry, Jayanthi Anand, Laura Galbraith, Susan Mason, Rachana Patel, Colin Nixon, Jared Cartwright, Hing Y. Leung, Karen Blyth, and David M. Bryant. The small GTPase ARF3 controls invasion modality and metastasis by regulating N-cadherin levels. Journal of Cell Biology, 222(4):e202206115, April 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:/

/rupress.org/jcb/article/222/4/e202206115/213909/The-
small-GTPase-ARF3-controls-invasion-modality.

Sugawara:2021:AST

[SFO^{+}21] Taichi Sugawara, Kyoko Furuse, Tetsuhisa Otani, Tomohiko Wakayama, and Mikio Furuse. Angulin-1 seals tricellular contacts independently of tricellulin and claudins. Journal of Cell Biology, 220(9):e202005062, September 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/9/e202005062/212477/Angulin-1-seals-tricellular-
contacts-independently.

Shen:2021:VFP

[SFWB21] James L. Shen, Tina M. Fortier, Ruoxi Wang, and Eric H. Baehrecke. Vps13D functions in a Pink1-dependent and Parkinindependent mitophagy pathway. Journal of Cell Biology, 220(11):e202104073, November 1, 2021. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/220/11/e202104073/212607/Vps13D-
functions-in-a-Pink1-dependent-and-Parkin.

Sheppard:2023:TCM

[SGL^{+}23] Luka Sheppard, David G. Green, Gerald Lerchbaumer, Katheryn E. Rothenberg, Rodrigo Fernandez-Gonzalez, and Ulrich Tepass. The α-Catenin mechanosensing M region is required for cell adhesion during tissue morphogenesis. Journal of Cell Biology, 222(2):e202108091, February 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:/

/rupress.org/jcb/article/222/2/e202108091/213759/The-
Catenin-mechanosensing-M-region-is-required.

Szikora:2020:NRL

[SGN^{+}20] Szilárd Szikora, Tamás Gajdos, Tibor Novák, Dávid Farkas, István Földi, Peter Lenart, Miklós Erdélyi, and József Mihály. Nanoscopy reveals the layered organization of the sarcomeric Hzone and I-band complexes. Journal of Cell Biology, 219(1): e201907026, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Subramanian:2020:OMN

[SGW^{+}20] Goutham Narayanan Subramanian, Jessica Greaney, Zhe Wei, Olivier Becherel, Martin Lavin, and Hayden Anthony Homer. Oocytes mount a noncanonical DNA damage response involving APC-Cdh1–mediated proteolysis. Journal of Cell Biology, 219(4):e201907213, April 6, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/4/
e201907213/151594/Oocytes-mount-a-noncanonical-DNA-damage-
response.

Sixt:2020:ZWC

[SH20] Michael Sixt and Anna Huttenlocher. Zena Werb (1945– 2020): Cell biology in context. Journal of Cell Biology, 219 (8):e202007029, August 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/219/8/e202007029/151969/Zena-
Werb-1945-2020-Cell-biology-in-contextZena.

Schormann:2020:RLA

[SHA20] Wiebke Schormann, Santosh Hariharan, and David W. Andrews. A reference library for assigning protein subcellular localizations by image-based machine learning. Journal of Cell Biology, 219 (3):e201904090, March 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/3/e201904090/133635/A-reference-
library-for-assigning-protein.

Sanchez-Huertas:2020:TNA

[SHBF^{+}20] Carlos Sánchez-Huertas, Marion Bonhomme, Amandine Falco, Christine Fagotto-Kaufmann, Jeffrey van Haren, Freddy Jeanneteau, Niels Galjart, Anne Debant, and Jérôme Boudeau. The +TIP Navigator-1 is an actin–microtubule crosslinker that regulates axonal growth cone motility. Journal of Cell Biology, 219(9):e201905199, September 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:/

/rupress.org/jcb/article/219/9/e201905199/151835/The-
TIP-Navigator-1-is-an-actin-microtubule.

Schmidt:2021:SRI

[SHD^{+}21] Vanessa Schmidt, Carla Horváth, Hua Dong, Matthias Blüher, Per Qvist, Christian Wolfrum, and Thomas E. Willnow. SORLA is required for insulin-induced expansion of the adipocyte precursor pool in visceral fat. Journal of Cell Biology, 220(12): e202006058, December 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/12/e202006058/212816/SORLA-is-required-
for-insulin-induced-expansion-of.

Shekhar:2021:TBA

[SHGG21] Shashank Shekhar, Gregory J. Hoeprich, Jeff Gelles, and Bruce L. Goode. Twinfilin bypasses assembly conditions and actin filament aging to drive barbed end depolymerization. Journal of Cell Biology, 220(1):e202006022, January 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/1/e202006022/211555/Twinfilin-bypasses-assembly-
conditions-and-actin.

Sando:2022:EST

[SHLS22] Richard Sando, Milan Lyndie Ho, Xinran Liu, and Thomas C. Südhof. Engineered synaptic tools reveal localized cAMP signaling in synapse assembly. Journal of Cell Biology, 221(2): e202109111, February 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/221/2/e202109111/212916/Engineered-synaptic-
tools-reveal-localized-cAMP.

Sanchez:2023:CPC

[SIP^{+}23] Gonzalo Manuel Sanchez, Tugce Ceren Incedal, Juan Prada, Paul O’Callaghan, Oleg Dyachok, Santiago Echeverry, Özge Dumral, Phuoc My Nguyen, Beichen Xie, Sebastian Barg, Johan Kreuger, Thomas Dandekar, and Olof Idevall-Hagren. The β-cell primary cilium is an autonomous Ca^{2+} compartment for paracrine GABA signaling. Journal of Cell Biology, 222(1): e202108101, January 2, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/222/1/e202108101/213674/The-cell-primary-
cilium-is-an-autonomous-Ca2.

Sirotkin:2023:CFF

[Sir23] Vladimir Sirotkin. Cappin’ or formin’: Formin and capping protein competition for filament ends shapes actin networks. Journal of Cell Biology, 222(4):e202302009, April 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
222/4/e202302009/213944/Cappin-or-formin-Formin-and-
capping-protein.

Shi:2022:FFP

[SJL^{+}22] Leiling Shi, Youli Jian, Meijiao Li, Tianchao Hao, Chonglin Yang, and Xiaochen Wang. Filamin FLN-2 promotes MVB biogenesis by mediating vesicle docking on the actin cytoskeleton. Journal of Cell Biology, 221(7):e202201020, July 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/7/
e202201020/213219/Filamin-FLN-2-promotes-MVB-biogenesis-
by-mediating.

Sell:2023:OSC

[SKF^{+}23] Thomas Sell, Christian Klotz, Matthias M. Fischer, Rosario Astaburuaga-García, Susanne Krug, Jarno Drost, Hans Clevers, Christine Sers, Markus Morkel, and Nils Blüthgen. Oncogenic signaling is coupled to colorectal cancer cell differentiation state. Journal of Cell Biology, 222(6):e202204001, June 5, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/6/
e202204001/213999/Oncogenic-signaling-is-coupled-to-
colorectal.

Sarangapani:2021:KBM

[SKN^{+}21] Krishna K. Sarangapani, Lori B. Koch, Christian R. Nelson, Charles L. Asbury, and Sue Biggins. Kinetochorebound Mps1 regulates kinetochore–microtubule attachments via Ndc80 phosphorylation. Journal of Cell Biology, 220(12): e202106130, December 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/12/e202106130/212702/Kinetochore-
bound-Mps1-regulates-kinetochore.

Sobajima:2023:PRA

[SKS^{+}23] Tomoaki Sobajima, Katarzyna M. Kowalczyk, Stefanos Skylakakis, Daniel Hayward, Luke J. Fulcher, Colette Neary, Caleb Batley, Samvid Kurlekar, Emile Roberts, Ulrike Gruneberg, and Francis A. Barr. PP6 regulation of Aurora A–TPX2 limits NDC80 phosphorylation and mitotic spindle size. Journal of Cell Biology, 222(5):e202205117, May 1, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:/

/rupress.org/jcb/article/222/5/e202205117/213930/PP6-
regulation-of-Aurora-A-TPX2-limits-NDC80.

Safavian:2023:SMR

[SKX^{+}23] Darya Safavian, Moshe S. Kim, Hong Xie, Maha El-Zeiry, Oliva Palander, Lu Dai, Richard F. Collins, Carol Froese, Rachel Shannon, Koh ichi Nagata, and William S. Trimble. Septinmediated RhoA activation engages the exocyst complex to recruit the cilium transition zone. Journal of Cell Biology, 222(4): e201911062, April 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/222/4/e201911062/213933/Septin-mediated-
RhoA-activation-engages-the.

Shi:2021:LRE

[SLD^{+}21] Xiaoyu Shi, Qi Li, Zhipeng Dai, Arthur A. Tran, Siyu Feng, Alejandro D. Ramirez, Zixi Lin, Xiaomeng Wang, Tracy T. Chow, Jiapei Chen, Dhivya Kumar, Andrew R. McColloch, Jeremy F. Reiter, Eric J. Huang, Ian B. Seiple, and Bo Huang. Label-retention expansion microscopy. Journal of Cell Biology, 220(9):e202105067, September 6, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/9/
e202105067/212454/Label-retention-expansion-microscopyLabel.

Shard:2020:TWC

[SLES20] Chloé Shard, Juan Luna-Escalante, and François Schweisguth. Tissue-wide coordination of epithelium-to-neural stem cell transition in the Drosophila optic lobe requires Neuralized. Journal of Cell Biology, 219(11):e202005035, November 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/11/e202005035/152101/Tissue-wide-coordination-of-
epithelium-to-neural.

Seoane:2020:NIS

[SLH^{+}20a] Paula I. Seoane, Bali Lee, Christopher Hoyle, Shi Yu, Gloria Lopez-Castejon, Martin Lowe, and David Brough. The NLRP3–inflammasome as a sensor of organelle dysfunction. Journal of Cell Biology, 219(12):e202006194, December 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/12/
e202006194/191204/The-NLRP3-inflammasome-as-a-sensor-
of-organelle.

Stancheva:2020:CMI

[SLH^{+}20b] Viktoriya G. Stancheva, Xiao-Han Li, Joshua Hutchings, Natalia Gomez-Navarro, Balaji Santhanam, M. Madan Babu, Giulia Zanetti, and Elizabeth A. Miller. Combinatorial multivalent interactions drive cooperative assembly of the COPII coat. Journal of Cell Biology, 219(11):e202007135, November 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/11/
e202007135/152133/Combinatorial-multivalent-interactions-
drive.

Sato:2021:UCA

[SLL^{+}21] Mai Sato, Andrew Walter Liebau, Zhaoqi Liu, Lizhi Liu, Raul Rabadan, and Jean Gautier. The UVSSA complex alleviates MYC-driven transcription stress. Journal of Cell Biology, 220(2):e201807163, February 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:/

/rupress.org/jcb/article/220/2/e201807163/211657/The-
UVSSA-complex-alleviates-MYC-driven. See correction [SLL^{+}23].

Sato:2023:CUC

[SLL^{+}23] Mai Sato, Rowyn C. Liebau, Zhaoqi Liu, Lizhi Liu, Raul Rabadan, and Jean Gautier. Correction: The UVSSA complex alleviates MYC-driven transcription stress. Journal of Cell Biology, 222(3):e20180716302092023c, March 6, 2023. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/222/3/
e20180716302092023c/213897/Correction-The-UVSSA-complex-
alleviates-MYC-driven. See [SLL^{+}21].

Singhal:2020:NGG

[SLM20] Neel S. Singhal, Evan M. Lee, and Dengke K. Ma. Neuronal GDPGP1 and glycogen metabolism: friend or foe? Journal of Cell Biology, 219(2):e202001006, February 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/2/e202001006/133641/Neuronal-GDPGP1-and-glycogen-
metabolism-friend-or.

Sun:2023:SSE

[SLM23] Sha Sun, Xia Li, and Malaiyalam Mariappan. Signal sequences encode information for protein folding in the endoplasmic reticulum. Journal of Cell Biology, 222(1):e202203070, January 2, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/222/1/e202203070/213733/Signal-sequences-
encode-information-for-protein.

Simoes:2022:CCD

[SLP^{+}22] Sérgio Simões, Gerald Lerchbaumer, Milena Pellikka, Paraskevi Giannatou, Thomas Lam, Dohyun Kim, Jessica Yu, David ter Stal, Kenana Al Kakouni, Rodrigo Fernandez-Gonzalez, and Ulrich Tepass. Crumbs complex–directed apical membrane dynamics in epithelial cell ingression. Journal of Cell Biology, 221(7): e202108076, July 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/221/7/e202108076/213229/Crumbs-complex-directed-
apical-membrane-dynamics.

Schloesser:2023:SCS

[SLS^{+}23] Daniela Schloesser, Laura Lindenthal, Julia Sauer, Kyoung-Jin Chung, Triantafyllos Chavakis, Eva Griesser, Praveen Baskaran, Ulrike Maier-Habelsberger, Katrin Fundel-Clemens, Ines Schlotthauer, Carolin Kirsten Watson, Lee Kim Swee, Frederik Igney, John Edward Park, Markus S. Huber-Lang, Matthew-James Thomas, Karim Christian El Kasmi, and Peter J. Murray. Senescent cells suppress macrophagemediated corpse removal via upregulation of the CD47-QPCT/L axis. Journal of Cell Biology, 222(2):e202207097, February 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/222/2/e202207097/213731/Senescent-cells-
suppress-macrophage-mediated.

Subramanian:2020:LBA

[SMC^{+}20] Bhagawat C. Subramanian, Nicolas Melis, Desu Chen, Weiye Wang, Devorah Gallardo, Roberto Weigert, and Carole A. Parent. The LTB_{4}–BLT1 axis regulates actomyosin and β_{2}-integrin dynamics during neutrophil extravasation. Journal of Cell Biology, 219(10):e201910215, October 5, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/219/10/e201910215/152048/The-
LTB4-BLT1-axis-regulates-actomyosin-and-2.

Sepaniac:2021:MKM

[SMD^{+}21] Leslie A. Sepaniac, Whitney Martin, Louise A. Dionne, Timothy M. Stearns, Laura G. Reinholdt, and Jason Stumpff. Micronuclei in Kif18a mutant mice form stable micronuclear envelopes and do not promote tumorigenesis. Journal of Cell Biology, 220(11):e202101165, November 1, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/11/
e202101165/212637/Micronuclei-in-Kif18a-mutant-mice-
form-stable.

Segal:2022:VPS

[SMFC^{+}22] Dagan Segal, Hanieh Mazloom-Farsibaf, Bo-Jui Chang, Philippe Roudot, Divya Rajendran, Stephan Daetwyler, Reto Fiolka, Mikako Warren, James F. Amatruda, and Gaudenz Danuser. In vivo 3D profiling of site-specific human cancer cell morphotypes in zebrafish. Journal of Cell Biology, 221(11): e202109100, November 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/221/11/e202109100/213501/In-vivo-3D-
profiling-of-site-specific-human-cancer.

Serra-Marques:2020:MPN

[SMHH^{+}20] Andrea Serra-Marques, Ronja Houtekamer, Dorine Hintzen, John T. Canty, Ahmet Yildiz, and Sophie Dumont. The mitotic protein NuMA plays a spindle-independent role in nuclear formation and mechanics. Journal of Cell Biology, 219(12): e202004202, December 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/12/e202004202/211454/The-mitotic-
protein-NuMA-plays-a-spindle.

Sanchez-Martin:2020:HSC

[SMK20] Pablo Sánchez-Martín and Masaaki Komatsu. Heparan sulfate and clusterin: Cleaning squad for extracellular protein degradation. Journal of Cell Biology, 219(3):e202001159, March 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/219/3/e202001159/133822/Heparan-sulfate-and-
clusterin-Cleaning-squad-for.

Sengupta:2021:SPC

[SMM^{+}21] Nayanika Sengupta, Anish Kumar Mondal, Suman Mishra, Kausik Chattopadhyay, and Somnath Dutta. Single-particle cryo-EM reveals conformational variability of the oligomeric VCC β-barrel pore in a lipid bilayer. Journal of Cell Biology, 220(12):e202102035, December 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/220/12/e202102035/212683/Single-
particle-cryo-EM-reveals-conformational.

Sakakibara:2020:ARA

[SMS^{+}20] Shotaro Sakakibara, Kiyohito Mizutani, Ayumu Sugiura, Ayuko Sakane, Takuya Sasaki, Shigenobu Yonemura, and Yoshimi Takai. Afadin regulates actomyosin organization through αecatenin at adherens junctions. Journal of Cell Biology, 219(5): e201907079, May 4, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/219/5/e201907079/151595/Afadin-regulates-
actomyosin-organization-through-E.

Smythe:2022:CCP

[Smy22] Elizabeth Smythe. Clathrin coated pits as signaling platforms for Akt signaling. Journal of Cell Biology, 221(4):e202203026, April 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/4/
e202203026/213084/Clathrin-coated-pits-as-signaling-
platforms-for.

Solvik:2022:SAM

[SNL^{+}22] Tina A. Solvik, Tan A. Nguyen, Yu-Hsiu Tony Lin, Timothy Marsh, Eric J. Huang, Arun P. Wiita, Jayanta Debnath, and Andrew M. Leidal. Secretory autophagy maintains proteostasis upon lysosome inhibition. Journal of Cell Biology, 221(6): e202110151, June 6, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/221/6/e202110151/213172/Secretory-autophagy-
maintains-proteostasis-upon.

Shinde:2020:UCE

[SNN20] Swapnil Rohidas Shinde, Andrew R. Nager, and Maxence V. Nachury. Ubiquitin chains earmark GPCRs for BBSomemediated removal from cilia. Journal of Cell Biology, 219(12): e202003020, December 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/12/e202003020/211536/Ubiquitin-chains-
earmark-GPCRs-for-BBSome-mediated.

Sosicka:2022:OCG

[SNP^{+}22] Paulina Sosicka, Bobby G. Ng, Lauren E. Pepi, Asif Shajahan, Maurice Wong, David A. Scott, Kenjiroo Matsumoto, Zhi-Jie Xia, Carlito B. Lebrilla, Robert S. Haltiwanger, Parastoo Azadi, and Hudson H. Freeze. Origin of cytoplasmic GDPfucose determines its contribution to glycosylation reactions. Journal of Cell Biology, 221(10):e202205038, October 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/10/
e202205038/213437/Origin-of-cytoplasmic-GDP-fucose-determines-
its.

Shomron:2021:CCD

[SNYA^{+}21] Olga Shomron, Inbar Nevo-Yassaf, Tamar Aviad, Yakey Yaffe, Eitan Erez Zahavi, Anna Dukhovny, Eran Perlson, Ilya Brodsky, Adva Yeheskel, Metsada Pasmanik-Chor, Anna Mironov, Galina V. Beznoussenko, Alexander A. Mironov, Ella H. Sklan, George H. Patterson, Yoji Yonemura, Mara Sannai, Christoph Kaether, and Koret Hirschberg. COPII collar defines the boundary between ER and ER exit site and does not coat cargo containers. Journal of Cell Biology, 220(6):e201907224, June 7, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/6/
e201907224/211990/COPII-collar-defines-the-boundary-
between-ER-and.

Soltysik:2021:NLD

[SOT^{+}21] Kamil Sołtysik, Yuki Ohsaki, Tsuyako Tatematsu, Jinglei Cheng, Asami Maeda, Shin ya Morita, and Toyoshi Fujimoto. Nuclear lipid droplets form in the inner nuclear membrane in a seipin-independent manner. Journal of Cell Biology, 220(1):e202005026, January 4, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/1/
e202005026/211592/Nuclear-lipid-droplets-form-in-the-
inner-nuclear.

Simoes:2022:IAI

[SPKP22] Davina Camargo Madeira Simoes, Nikolaos Paschalidis, Evangelia Kourepini, and Vily Panoutsakopoulou. An integrin axis induces IFN-β production in plasmacytoid dendritic cells. Journal of Cell Biology, 221(9):e202102055, September 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/9/e202102055/213363/An-integrin-axis-induces-IFN-
production-in.

Silva:2020:WRS

[SPL^{+}20] Mariana C. C. Silva, Sean Powell, Sabrina Ladstätter, Johanna Gassler, Roman Stocsits, Antonio Tedeschi, Jan-Michael Peters, and Kikuë Tachibana. Wapl releases Scc1cohesin and regulates chromosome structure and segregation in mouse oocytes. Journal of Cell Biology, 219(4): e201906100, April 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/219/4/e201906100/151562/Wapl-
releases-Scc1-cohesin-and-regulates.

Sharp:2020:CDR

[SPRWB20] Judith A. Sharp, Carlos Perea-Resa, Wei Wang, and Michael D. Blower. Cell division requires RNA eviction from condensing chromosomes. Journal of Cell Biology, 219(11):e201910148, November 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/219/11/e201910148/211450/Cell-division-requires-
RNA-eviction-from.

Sharma:2020:SAR

[SPS^{+}20] Priyanka Sharma, Sameena Parveen, Lekha V. Shah, Madhumita Mukherjee, Yannis Kalaidzidis, Anthony J. Kozielski, Roberto Rosato, Jenny C. Chang, and Sunando Datta. SNX27 retromer assembly recycles MT1-MMP to invadopodia and promotes breast cancer metastasis. Journal of Cell Biology, 219(1): e201812098, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Schuck:2009:MEA

[SPT^{+}09] Sebastian Schuck, William A. Prinz, Kurt S. Thorn, Christiane Voss, and Peter Walter. Membrane expansion alleviates endoplasmic reticulum stress independently of the unfolded protein response. Journal of Cell Biology, 187(4):525–??, November 2009. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/187/4/

525. See correction [SPT^{+}21].

Schuck:2021:CME

[SPT^{+}21] Sebastian Schuck, William A. Prinz, Kurt S. Thorn, Christiane Voss, and Peter Walter. Correction: Membrane expansion alleviates endoplasmic reticulum stress independently of the unfolded protein response. Journal of Cell Biology, 220(4):??, April 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
4/jcb.20090707402092021c/211800/Correction-Membrane-
expansion-alleviates. See [SPT^{+}09].

Sana:2022:MCE

[SRK22] Shrividya Sana, Ashwathi Rajeevan, and Sachin Kotak. Membrane compartmentalization of Ect2/Cyk4/Mklp1 and NuMA/ dynein regulates cleavage furrow formation. Journal of Cell Biology, 221(12):e202203127, December 5, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/12/
e202203127/213522/Membrane-compartmentalization-of-Ect2-
Cyk4-Mklp1.

Sanchez-Ramirez:2022:CMT

[SRUdC^{+}22] Edgar Sánchez-Ramírez, Thi Phuong Lien Ung, Alejandro Alarcón del Carmen, Ximena del Toro-Ríos, Guadalupe R. Fajardo-Orduña, Lilia G. Noriega, Victor A. Cortés-Morales, Armando R. Tovar, Juan José Montesinos, Ricardo Orozco-Solís, Chiara Stringari, and Lorena Aguilar-Arnal. Coordinated metabolic transitions and gene expression by NAD+ during adipogenesis. Journal of Cell Biology, 221(12):e202111137, December 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/221/12/e202111137/213521/Coordinated-
metabolic-transitions-and-gene.

Shrestha:2021:COP

[SRW^{+}21] Roshan L. Shrestha, Austin Rossi, Darawalee Wangsa, Ann K. Hogan, Kimberly S. Zaldana, Evelyn Suva, Yang Jo Chung, Chelsea L. Sanders, Simone Difilippantonio, Tatiana S. Karpova, Baktiar Karim, Daniel R. Foltz, Daniele Fachinetti, Peter D. Aplan, Thomas Ried, and Munira A. Basrai. CENP

— a overexpression promotes aneuploidy with karyotypic heterogeneity. Journal of Cell Biology, 220(4):??, April 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/4/
e202007195/211820/CENP-A-overexpression-promotes-aneuploidy-
with.

Stopp:2022:PYT

[SS22] Julian Stopp and Michael Sixt. Plan your trip before you leave: The neutrophils’ search-and-run journey. Journal of Cell Biology, 221(8):e202206127, August 1, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/221/8/e202206127/213358/Plan-
your-trip-before-you-leave-The-neutrophils.

Stow:2020:LPM

[SSB20] Jennifer L. Stow, Jaakko Saraste, and William J. Brown. A life in pictures — Marilyn Gist Farquhar. Journal of Cell Biology, 219(2):e202001010, February 3, 2020. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/219/2/e202001010/133633/A-
life-in-pictures-Marilyn-Gist-FarquharA-life-in.

Stormo:2022:ELT

[SSF^{+}22] Adrienne E. D. Stormo, Farbod Shavarebi, Molly FitzGibbon, Elizabeth M. Earley, Hannah Ahrendt, Lotus S. Lum, Erik Verschueren, Danielle L. Swaney, Gaia Skibinski, Abinaya Ravisankar, Jeffrey van Haren, Emily J. Davis, Jeffrey R. Johnson, John Von Dollen, Carson Balen, Jacob Porath, Claudia Crosio, Christian Mirescu, Ciro Iaccarino, William T. Dauer, R. Jeremy Nichols, Torsten Wittmann, Timothy C. Cox, Steve Finkbeiner, Nevan J. Krogan, Scott A. Oakes, and Annie Hiniker. The E3 ligase TRIM1 ubiquitinates LRRK2 and controls its localization, degradation, and toxicity. Journal of Cell Biology, 221(4):e202010065, April 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:/

/rupress.org/jcb/article/221/4/e202010065/213061/The-
E3-ligase-TRIM1-ubiquitinates-LRRK2-and.

Smolka:2021:RRS

[SSHC21] John A. Smolka, Lionel A. Sanz, Stella R. Hartono, and Frédéric Chédin. Recognition of RNA by the S9.6 antibody creates pervasive artifacts when imaging RNA:DNA hybrids. Journal of Cell Biology, 220(6):e202004079, June 7, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/6/e202004079/211957/Recognition-of-RNA-by-the-S9-
6-antibody-creates.

Schulz:2020:SRG

[SSO^{+}20] Alexander Schulz, Yuichi Sekine, Motunrayo J. Oyeyemi, Alexander J. Abrams, Manasa Basavaraju, Sung Min Han, Marco Groth, Helen Morrison, Stephen M. Strittmatter, and Marc Hammarlund. The stress-responsive gene GDPGP1/ mcp-1 regulates neuronal glycogen metabolism and survival. Journal of Cell Biology, 219(2):e201807127, February 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/2/
e201807127/133634/The-stress-responsive-gene-GDPGP1-
mcp-1-regulates.

Shrivastava:2022:CSA

[SSR^{+}22] Aseem Shrivastava, Carl Alexander Sandhof, Kevin Reinle, Areeb Jawed, Carmen Ruger-Herreros, Dominic Schwarz, Declan Creamer, Carmen Nussbaum-Krammer, Axel Mogk, and Bernd Bukau. The cytoprotective sequestration activity of small heat shock proteins is evolutionarily conserved. Journal of Cell Biology, 221(10):e202202149, October 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/221/10/e202202149/213447/The-
cytoprotective-sequestration-activity-of-small.

Siu:2021:CBF

[SSZL21] Karen K. Siu, Vitor Hugo B. Serrão, Ahmed Ziyyat, and Jeffrey E. Lee. The cell biology of fertilization: Gamete attachment and fusion. Journal of Cell Biology, 220(10):e202102146, October 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/220/10/e202102146/212606/The-cell-biology-of-
fertilization-Gamete.

Song:2021:EVN

[STS21] Lu Song, Xinran Tian, and Randy Schekman. Extracellular vesicles from neurons promote neural induction of stem cells through cyclin D1. Journal of Cell Biology, 220(9):e202101075, September 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/9/e202101075/212508/Extracellular-vesicles-from-
neurons-promote-neural.

Simon:2023:ECT

[STvT23] François Simon, Jean-Yves Tinevez, and Sven van Teeffelen. ExTrack characterizes transition kinetics and diffusion in noisy single-particle tracks. Journal of Cell Biology, 222(5): e202208059, May 1, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/222/5/e202208059/213911/ExTrack-characterizes-
transition-kinetics-and.

Sofroni:2020:CDA

[STY^{+}20] Kostika Sofroni, Hirotomo Takatsuka, Chao Yang, Nico Dissmeyer, Shinichiro Komaki, Yuki Hamamura, Lev Böttger, Masaaki Umeda, and Arp Schnittger. CDKD-dependent activation of CDKA;1 controls microtubule dynamics and cytokinesis during meiosis. Journal of Cell Biology, 219(8):e201907016, August 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/8/e201907016/151917/CDKD-dependent-activation-of-
CDKA-1-controls.

Striepen:2022:CRE

[SV22] Jonathan F. Striepen and Gia K. Voeltz. Coronin 1C restricts endosomal branched actin to organize ER contact and endosome fission. Journal of Cell Biology, 221(8):e202110089, August 1, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/221/8/e202110089/213342/Coronin-1C-restricts-
endosomal-branched-actin-to.

Soh:2020:CFR

[SvDSW^{+}20] Adam W. J. Soh, Teunis J. P. van Dam, Alexander J. Stemm-Wolf, Andrew T. Pham, Garry P. Morgan, Eileen T. O’Toole, and Chad G. Pearson. Ciliary force-responsive striated fibers promote basal body connections and cortical interactions. Journal of Cell Biology, 219(1):e201904091, January 6, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Sassano:2023:PRS

[SvVV^{+}23] Maria Livia Sassano, Alexander R. van Vliet, Ellen Vervoort, Sofie Van Eygen, Chris Van den Haute, Benjamin Pavie, Joris Roels, Johannes V. Swinnen, Marco Spinazzi, Leen Moens, Kristina Casteels, Isabelle Meyts, Paolo Pinton, Saverio Marchi, Leila Rochin, Francesca Giordano, Blanca Felipe-Abrio, and Patrizia Agostinis. PERK recruits E-Syt1 at ER-mitochondria contacts for mitochondrial lipid transport and respiration. Journal of Cell Biology, 222(3):e202206008, March 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/222/3/e202206008/213891/PERK-recruits-E-Syt1-
at-ER-mitochondria-contacts.

Steinacker:2022:CGL

[SWN^{+}22] Thomas L. Steinacker, Siu-Shing Wong, Zsofia A. Novak, Saroj Saurya, Lisa Gartenmann, Eline J. H. van Houtum, Judith R. Sayers, B. Christoffer Lagerholm, and Jordan W. Raff. Centriole growth is limited by the Cdk/Cyclin-dependent phosphorylation of Ana2/STIL. Journal of Cell Biology, 221(9):e202205058, September 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/221/9/e202205058/213359/Centriole-growth-
is-limited-by-the-Cdk-Cyclin.

Su:2021:SSS

[SWS^{+}21a] Xue Bessie Su, Menglu Wang, Claudia Schaffner, Olga O. Nerusheva, Dean Clift, Christos Spanos, David A. Kelly, Michael Tatham, Andreas Wallek, Yehui Wu, Juri Rappsilber, A. Arockia Jeyaprakash, Zuzana Storchova, Ronald T. Hay, and Adèle L. Marston. SUMOylation stabilizes sister kinetochore biorientation to allow timely anaphase. Journal of Cell Biology, 220 (7):e202005130, July 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/7/e202005130/212041/SUMOylation-
stabilizes-sister-kinetochore.

Surve:2021:ERM

[SWS21b] Sachin Surve, Simon C. Watkins, and Alexander Sorkin. EGFR-RAS-MAPK signaling is confined to the plasma membrane and associated endorecycling protrusions. Journal of Cell Biology, 220(11):e202107103, November 1, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/11/
e202107103/212639/EGFR-RAS-MAPK-signaling-is-confined-
to-the-plasma.

Sofi:2022:PLD

[SWT^{+}22] Sajad Sofi, Louisa Williamson, Gabrielle L. Turvey, Charlotte Scoynes, Claire Hirst, Jonathan Godwin, Neil Brockdorff, Justin Ainscough, and Dawn Coverley. Prion-like domains drive CIZ1 assembly formation at the inactive X chromosome. Journal of Cell Biology, 221(4):e202103185, April 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/4/
e202103185/213067/Prion-like-domains-drive-CIZ1-assembly-
formation.

Shen:2022:LRC

[SYQ^{+}22] Xiao-Lin Shen, Jin-Feng Yuan, Xuan-He Qin, Guang-Ping Song, Huai-Bin Hu, Hai-Qing Tu, Zeng-Qing Song, Pei-Yao Li, Yu-Ling Xu, Sen Li, Xiao-Xiao Jian, Jia-Ning Li, Chun-Yu He, Xi-Ping Yu, Li-Yun Liang, Min Wu, Qiu-Ying Han, Kai Wang, Ai-Ling Li, Tao Zhou, Yu-Cheng Zhang, Na Wang, and Hui-Yan Li. LUBAC regulates ciliogenesis by promoting CP110 removal from the mother centriole. Journal of Cell Biology, 221(1): e202105092, January 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/221/1/e202105092/212875/LUBAC-regulates-
ciliogenesis-by-promoting-CP110.

Shi:2020:UCO

[SYW^{+}20] Xiaoshan Shi, Adam L. Yokom, Chunxin Wang, Lindsey N. Young, Richard J. Youle, and James H. Hurley. ULK complex organization in autophagy by a C-shaped FIP200 N-terminal domain dimer. Journal of Cell Biology, 219(7):e201911047, July 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/7/e201911047/151844/ULK-complex-organization-in-
autophagy-by-a-C.

Tait:2022:KCU

[Tai22] Stephen W. G. Tait. Killing cells using light (activated) sabers. Journal of Cell Biology, 221(6):e202205018, June 6, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/221/6/e202205018/213218/Killing-cells-using-
light-activated-sabersAn.

Tanner:2023:VPP

[Tan23] Kandice Tanner. An in vivo phosphoregulation paradox for focal adhesions. Journal of Cell Biology, 222(3):e202301060, March 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
222/3/e202301060/213876/An-in-vivo-phosphoregulation-
paradox-for-focal.

Tarabykin:2021:SAS

[Tar21] Victor Tarabykin. SNAP to attention: a SNARE complex regulates neuronal progenitor polarity. Journal of Cell Biology, 220(1):e202011052, January 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/220/1/e202011052/211604/SNAP-
to-attention-A-SNARE-complex-regulates.

Tei:2020:SCP

[TB20a] Reika Tei and Jeremy M. Baskin. Spatiotemporal control of phosphatidic acid signaling with optogenetic, engineered phospholipase Ds. Journal of Cell Biology, 219(3):e201907013, March 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/3/
e201907013/133675/Spatiotemporal-control-of-phosphatidic-
acid.

Traub:2020:EJU

[TB20b] Linton M. Traub and Frances M. Brodsky. Ernst Joachim Ungewickell: 1950–2020. Journal of Cell Biology, 219(11): e202009044, November 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/11/e202009044/182191/Ernst-Joachim-
Ungewickell-1950-2020Ernst-Joachim.

Theisen:2020:MMP

[TEH^{+}20] Ulrike Theisen, Alexander U. Ernst, Ronja L. S. Heyne, Tobias P. Ring, Oliver Thorn-Seshold, and Reinhard W. Köster. Microtubules and motor proteins support zebrafish neuronal migration by directing cargo. Journal of Cell Biology, 219(10):e201908040, October 5, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/219/10/e201908040/151951/Microtubules-and-motor-
proteins-support-zebrafish.

Teves:2020:FTA

[Tev20] Sheila S. Teves. Function through absence: Active RNA exclusion from chromosomes leads to proper cell division. Journal of Cell Biology, 219(11):e202009193, November 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/
11/e202009193/211462/Function-through-absence-Active-
RNA-exclusion-from.

Tan:2020:MIS

[TF20] Jay X. Tan and Toren Finkel. Mitochondria as intracellular signaling platforms in health and disease. Journal of Cell Biology, 219(5):e202002179, May 4, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/5/e202002179/151695/Mitochondria-
as-intracellular-signaling-platforms.

Tsuchiya:2021:MAP

[TG21] Kenta Tsuchiya and Gohta Goshima. Microtubule-associated proteins promote microtubule generation in the absence of - tubulin in human colon cancer cells. Journal of Cell Biology, 220(12):e202104114, December 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/220/12/e202104114/212818/Microtubule-
associated-proteins-promote.

Trillet:2021:GGG

[TJAG^{+}21] Kilian Trillet, Kathryn A. Jacobs, Gwennan André-Grégoire, An Thys, Clément Maghe, Jonathan Cruard, Stéphane Minvielle, Sara Gonzalez Diest, Guillaume Montagnac, Nicolas Bidère, and Julie Gavard. The glycoprotein GP130 governs the surface presentation of the G protein–coupled receptor APLNR. Journal of Cell Biology, 220(9):e202004114, September 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/9/
e202004114/212489/The-glycoprotein-GP130-governs-the-
surface.

Tomioka:2020:TIS

[TKK^{+}20] Yui Tomioka, Tetsuya Kotani, Hiromi Kirisako, Yu Oikawa, Yayoi Kimura, Hisashi Hirano, Yoshinori Ohsumi, and Hitoshi Nakatogawa. TORC1 inactivation stimulates autophagy of nucleoporin and nuclear pore complexes. Journal of Cell Biology, 219(7):e201910063, July 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/7/e201910063/151819/TORC1-inactivation-
stimulates-autophagy-of.

Tordonato:2021:MCS

[TMG^{+}21] Chiara Tordonato, Matteo Jacopo Marzi, Giovanni Giangreco, Stefano Freddi, Paola Bonetti, Daniela Tosoni, Pier Paolo Di Fiore, and Francesco Nicassio. miR-146 connects stem cell identity with metabolism and pharmacological resistance in breast cancer. Journal of Cell Biology, 220(5): e202009053, May 3, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/220/5/e202009053/211945/miR-146-connects-
stem-cell-identity-with.

Tie:2022:VIG

[TML22] Hieng Chiong Tie, Divyanshu Mahajan, and Lei Lu. Visualizing intra-Golgi localization and transport by side-averaging Golgi ministacks. Journal of Cell Biology, 221(6):e202109114, June 6, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/6/
e202109114/213180/Visualizing-intra-Golgi-localization-
and-transport.

Taskinen:2020:MPC

[TNC^{+}20] Maria Emilia Taskinen, Elisa Närvä, James R. W. Conway, Laura Soto Hinojosa, Sergio Lilla, Anja Mai, Nicola De Franceschi, Laura L. Elo, Robert Grosse, Sara Zanivan, Jim C. Norman, and Johanna Ivaska. MASTL promotes cell contractility and motility through kinase-independent signaling. Journal of Cell Biology, 219(6):e201906204, June 1, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/6/
e201906204/151688/MASTL-promotes-cell-contractility-
and-motility.

Tirrell:2020:MSA

[TNLPF20] Parker S. Tirrell, Kailey N. Nguyen, Katherine Luby-Phelps, and Jonathan R. Friedman. MICOS subcomplexes assemble independently on the mitochondrial inner membrane in proximity to ER contact sites. Journal of Cell Biology, 219(11):e202003024, November 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/219/11/e202003024/211445/MICOS-subcomplexes-
assemble-independently-on-the.

Torii:2020:NPA

[TOL^{+}20] Tomohiro Torii, Yuki Ogawa, Cheng-Hsin Liu, Tammy Szu-Yu Ho, Hamdan Hamdan, Chih chuan Wang, Juan A. Oses-Prieto, Alma L. Burlingame, and Matthew N. Rasband. NuMA1 promotes axon initial segment assembly through inhibition of endocytosis. Journal of Cell Biology, 219(2):e201907048, February 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/2/e201907048/132495/NuMA1-promotes-axon-initial-
segment-assembly.

Taylor:2020:CSD

[TP20] Samuel J. P. Taylor and Federico Pelisch. Chromosome segregation during female meiosis in C. elegans: a tale of pushing and pulling. Journal of Cell Biology, 219(12): e202011035, December 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/12/e202011035/211548/Chromosome-
segregation-during-female-meiosis-in-C.

Tran:2021:APL

[TPM^{+}21] Joseph R. Tran, Danielle I. Paulson, James J. Moresco, Stephen A. Adam, John R. Yates III, Robert D. Goldman, and Yixian Zheng. An APEX2 proximity ligation method for mapping interactions with the nuclear lamina. Journal of Cell Biology, 220(1):e202002129, January 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/220/1/e202002129/211588/An-
APEX2-proximity-ligation-method-for-mapping.

Thines:2023:ISC

[TRHS23] Louise Thines, Francis J. Roushar, Andrew C. Hedman, and David B. Sacks. The IQGAP scaffolds: Critical nodes bridging receptor activation to cellular signaling. Journal of Cell Biology, 222(6):e202205062, June 5, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:/

/rupress.org/jcb/article/222/6/e202205062/214044/The-
IQGAP-scaffolds-Critical-nodes-bridging.

Titlow:2020:SHQ

[TRJ^{+}20] Joshua Titlow, Francesca Robertson, Aino Järvelin, David Ish-Horowicz, Carlas Smith, Enrico Gratton, and Ilan Davis. Syncrip/hnRNP Q is required for activity-induced Msp300/ Nesprin-1 expression and new synapse formation. Journal of Cell Biology, 219(3):e201903135, March 2, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/3/
e201903135/133707/Syncrip-hnRNP-Q-is-required-for-activity-
induced.

Tang:2020:UEL

[TSL^{+}20] Danming Tang, Wendy Sandoval, Cynthia Lam, Benjamin Haley, Peter Liu, Di Xue, Deepankar Roy, Tom Patapoff, Salina Louie, Brad Snedecor, and Shahram Misaghi. UBR E3 ligases and the PDIA3 protease control degradation of unfolded antibody heavy chain by ERAD. Journal of Cell Biology, 219(7):e201908087, July 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:/

/rupress.org/jcb/article/219/7/e201908087/151862/UBR-
E3-ligases-and-the-PDIA3-protease-control.

Tavernier:2021:AKA

[TSP21] Nicolas Tavernier, Frank Sicheri, and Lionel Pintard. Aurora A kinase activation: Different means to different ends. Journal of Cell Biology, 220(9):e202106128, September 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/9/
e202106128/212490/Aurora-A-kinase-activation-Different-
means-to.

Thaller:2021:DBE

[TTM^{+}21] David J. Thaller, Danqing Tong, Christopher J. Marklew, Nicholas R. Ader, Philip J. Mannino, Sapan Borah, Megan C. King, Barbara Ciani, and C. Patrick Lusk. Direct binding of ESCRT protein Chm7 to phosphatidic acid–rich membranes at nuclear envelope herniations. Journal of Cell Biology, 220(3):e202004222, March 1, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/3/
e202004222/211693/Direct-binding-of-ESCRT-protein-Chm7-
to.

Tian:2021:SCC

[TWH^{+}21] Yuan Tian, Chenxi Wei, Jianfeng He, Yuxuan Yan, Nan Pang, Xiaomin Fang, Xin Liang, and Jingyan Fu. Superresolution characterization of core centriole architecture. Journal of Cell Biology, 220(4):??, April 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/220/4/e202005103/211748/Superresolution-
characterization-of-core-centriole.

Towers:2020:ACM

[TWT20] Christina G. Towers, Darya Wodetzki, and Andrew Thorburn. Autophagy and cancer: Modulation of cell death pathways and cancer cell adaptations. Journal of Cell Biology, 219(1): e201909033, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Toulmay:2022:VLP

[TWY^{+}22] Alexandre Toulmay, Fawn B. Whittle, Jerry Yang, Xiaofei Bai, Jessica Diarra, Subhrajit Banerjee, Tim P. Levine, Andy Golden, and William A. Prinz. Vps13-like proteins provide phosphatidylethanolamine for GPI anchor synthesis in the ER. Journal of Cell Biology, 221(3):e202111095, March 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/3/
e202111095/212952/Vps13-like-proteins-provide.

Uchino:2022:LIT

[UIS^{+}22] Satoshi Uchino, Yuma Ito, Yuko Sato, Tetsuya Handa, Yasuyuki Ohkawa, Makio Tokunaga, and Hiroshi Kimura. Live imaging of transcription sites using an elongating RNA polymerase II–specific probe. Journal of Cell Biology, 221(2): e202104134, February 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/221/2/e202104134/212888/Live-imaging-
of-transcription-sites-using-an.

Umebayashi:2023:CLP

[UTR^{+}23] Miwa Umebayashi, Satoko Takemoto, Luc Reymond, Mayya Sundukova, Ruud Hovius, Annalisa Bucci, Paul A. Heppenstall, Hideo Yokota, Kai Johnsson, and Howard Riezman. A covalently linked probe to monitor local membrane properties surrounding plasma membrane proteins. Journal of Cell Biology, 222(3): e202206119, March 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/222/3/e202206119/213783/A-covalently-
linked-probe-to-monitor-local.

Udi:2023:GMQ

[UZS^{+}23] Yael Udi, Wenzhu Zhang, Milana E. Stein, Inna Ricardo-Lax, Hilda A. Pasolli, Brian T. Chait, and Michael P. Rout. A general method for quantitative fractionation of mammalian cells. Journal of Cell Biology, 222(6):e202209062, June 5, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/6/
e202209062/213941/A-general-method-for-quantitative-
fractionation-of.

Verweij:2022:MCS

[VBG^{+}22] Frederik J. Verweij, Maarten P. Bebelman, Anna E. George, Mickael Couty, Anaïs Bécot, Roberta Palmulli, Xavier Heiligenstein, Julia Sirés-Campos, Graça Raposo, Dirk Michiel Pegtel, and Guillaume van Niel. ER membrane contact sites support endosomal small GTPase conversion for exosome secretion. Journal of Cell Biology, 221(12):e202112032, December 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/12/
e202112032/213494/ER-membrane-contact-sites-support-
endosomal-small.

Varadarajan:2022:MCF

[VCS^{+}22] Saranyaraajan Varadarajan, Shahana A. Chumki, Rachel E. Stephenson, Eileen R. Misterovich, Jessica L. Wu, Claire E. Dudley, Ivan S. Erofeev, Andrew B. Goryachev, and Ann L. Miller. Mechanosensitive calcium flashes promote sustained RhoA activation during tight junction remodeling. Journal of Cell Biology, 221(4):e202105107, April 4, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/4/
e202105107/213049/Mechanosensitive-calcium-flashes-promote-
sustained.

vanderBeek:2022:QCM

[vdBdHLK22] Jan van der Beek, Cecilia de Heus, Nalan Liv, and Judith Klumperman. Quantitative correlative microscopy reveals the ultrastructural distribution of endogenous endosomal proteins. Journal of Cell Biology, 221(1):e202106044, January 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/221/1/e202106044/212877/Quantitative-
correlative-microscopy-reveals-the.

vandenBerg:2023:CSG

[vdBVS^{+}23] Cyntha M. van den Berg, Vladimir A. Volkov, Sebastian Schnorrenberg, Ziqiang Huang, Kelly E. Stecker, Ilya Grigoriev, Sania Gilani, Kari-Anne M. Frikstad, Sebastian Patzke, Timo Zimmermann, Marileen Dogterom, and Anna Akhmanova. CSPP1 stabilizes growing microtubule ends and damaged lattices from the luminal side. Journal of Cell Biology, 222(4): e202208062, April 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/222/4/e202208062/213861/CSPP1-stabilizes-
growing-microtubule-ends-and.

Vishnoi:2020:DTN

[VDC^{+}20] Nidhi Vishnoi, Karthigeyan Dhanasekeran, Madeleine Chalfant, Ivan Surovstev, Mustafa K. Khokha, and C. Patrick Lusk. Differential turnover of Nup188 controls its levels at centrosomes and role in centriole duplication. Journal of Cell Biology, 219 (3):e201906031, March 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/3/e201906031/133835/Differential-
turnover-of-Nup188-controls-its.

vandenGoor:2022:CGT

[vdGM22] Lotte van den Goor and Ann L. Miller. Closing the gap: Tricellulin/-catenin interaction maintains epithelial integrity at vertices. Journal of Cell Biology, 221(4):e202202009, April 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/4/e202202009/213024/Closing-the-gap-Tricellulin-
catenin-interaction.

Verlhac:2021:PAG

[Ver21] Marie-Hélène Verlhac. Preventing aneuploidy: The groom must wait until the bride is ready. Journal of Cell Biology, 220(10): e202108030, October 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/10/e202108030/212627/Preventing-
aneuploidy-The-groom-must-wait-until.

Velle:2020:CAM

[VFL20] Katrina B. Velle and Lillian K. Fritz-Laylin. Conserved actin machinery drives microtubule-independent motility and phagocytosis in Naegleria. Journal of Cell Biology, 219(11):e202007158, November 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/219/11/e202007158/152108/Conserved-actin-
machinery-drives-microtubule.

Venkatesan:2021:MMC

[VGK^{+}21] Arunkumar Venkatesan, Jie Geng, Malathi Kandarpa, Sanjeeva Joseph Wijeyesakere, Ashwini Bhide, Moshe Talpaz, Irina D. Pogozheva, and Malini Raghavan. Mechanism of mutant calreticulin-mediated activation of the thrombopoietin receptor in cancers. Journal of Cell Biology, 220(7):e202009179, July 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
7/e202009179/212031/Mechanism-of-mutant-calreticulin-
mediated.

Wende:2023:MRC

[VGO^{+}23] Helen M. Vander Wende, Mounika Gopi, Megan Onyundo, Claudia Medrano, Temiloluwa Adanlawo, and Gloria Ann Brar. Meiotic resetting of the cellular Sod1 pool is driven by protein aggregation, degradation, and transient LUTI-mediated repression. Journal of Cell Biology, 222(3):e202206058, March 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/3/
e202206058/213795/Meiotic-resetting-of-the-cellular-
Sod1-pool-is.

Viol:2020:NKD

[VHPP^{+}20] Linda Viol, Shoji Hata, Ana Pastor-Peidro, Annett Neuner, Florian Murke, Patrick Wuchter, Anthony D. Ho, Bernd Giebel, and Gislene Pereira. Nek2 kinase displaces distal appendages from the mother centriole prior to mitosis. Journal of Cell Biology, 219(3):e201907136, March 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/219/3/e201907136/133805/Nek2-
kinase-displaces-distal-appendages-from-the.

Vega-Lugo:2022:ACC

[VLdRADJ22] Jesus Vega-Lugo, Bruno da Rocha-Azevedo, Aparajita Dasgupta, and Khuloud Jaqaman. Analysis of conditional colocalization relationships and hierarchies in three-color microscopy images. Journal of Cell Biology, 221(7):e202106129, July 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/7/
e202106129/213216/Analysis-of-conditional-colocalization.

vanLoon:2020:CCD

[vLEM^{+}20] Aaron P. van Loon, Ivan S. Erofeev, Ivan V. Maryshev, Andrew B. Goryachev, and Alvaro Sagasti. Cortical contraction drives the 3d patterning of epithelial cell surfaces. Journal of Cell Biology, 219(3):e201904144, March 2, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/3/
e201904144/133677/Cortical-contraction-drives-the-3D-
patterning-of.

Visintin:2021:AAB

[VM21] Rosella Visintin and Adele L. Marston. Angelika Amon (1967– 2020): Breakthrough scientist, extraordinary mentor, and loyal friend. Journal of Cell Biology, 220(2):e202012031, February 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/2/
e202012031/211643/Angelika-Amon-1967-2020-Breakthrough-
scientist.

Vellino:2021:CTB

[VOR^{+}21] Sanela Vellino, Christiane Oddou, Paul Rivier, Cyril Boyault, Edwige Hiriart-Bryant, Alexandra Kraut, René Martin, Yohann Coute, Hans-Joachim Knölker, Miguel A. Valverde, Corinne Albigès-Rizo, and Olivier Destaing. Cross-talk between the calcium channel TRPV4 and reactive oxygen species interlocks adhesive and degradative functions of invadosomes. Journal of Cell Biology, 220(2):e201910079, February 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/2/
e201910079/211651/Cross-talk-between-the-calcium-channel-
TRPV4-and.

Varadarajan:2023:CER

[VRSN23] Saranyaraajan Varadarajan, Arturo Raya-Sandino, and Asma Nusrat. Clipping EpCAM to release Claudin-7 for the greater good of the epithelial barrier. Journal of Cell Biology, 222(1): e202211127, January 2, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/222/1/e202211127/213749/Clipping-EpCAM-to-
release-Claudin-7-for-the.

Valente:2020:PDT

[VTL^{+}20] Liz J. Valente, Amy Tarangelo, Albert Mao Li, Marwan Naciri, Nitin Raj, Anthony M. Boutelle, Yang Li, Stephano Spano Mello, Kathryn Bieging-Rolett, Ralph J. DeBerardinis, Jiangbin Ye, Scott J. Dixon, and Laura D. Attardi. p53 deficiency triggers dysregulation of diverse cellular processes in physiological oxygen. Journal of Cell Biology, 219(11):e201908212, November 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/219/11/e201908212/152074/p53-deficiency-
triggers-dysregulation-of-diverse.

Veldsink:2023:HUB

[VV23] Annemiek C. Veldsink and Liesbeth M. Veenhoff. How to unravel a basket: NPC reorganization during meiosis. Journal of Cell Biology, 222(2):e202301044, February 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:/

/rupress.org/jcb/article/222/2/e202301044/213821/How-
to-unravel-a-basket-NPC-reorganization-during.

Virant:2023:UKN

[VVW^{+}23] David Virant, Ilijana Vojnovic, Jannik Winkelmeier, Marc Endesfelder, Bartosz Turkowyd, David Lando, and Ulrike Endesfelder. Unraveling the kinetochore nanostructure in schizosaccharomyces pombe using multi-color SMLM imaging. Journal of Cell Biology, 222(4):e202209096, April 3, 2023. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/222/4/
e202209096/213836/Unraveling-the-kinetochore-nanostructure-
in.

Vessoni:2021:TEH

[VZQ^{+}21] Alexandre T. Vessoni, Tianpeng Zhang, Annabel Quinet, Ho-Chang Jeong, Michael Munroe, Matthew Wood, Enzo Tedone, Alessandro Vindigni, Jerry W. Shay, Roger A. Greenberg, and Luis F. Z. Batista. Telomere erosion in human pluripotent stem cells leads to ATR-mediated mitotic catastrophe. Journal of Cell Biology, 220(6):e202011014, June 7, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/220/6/e202011014/211982/Telomere-erosion-in-
human-pluripotent-stem-cells.

Wozniak:2020:RBP

[WAK^{+}20] Ann L. Wozniak, Abby Adams, Kayla E. King, Winston Dunn, Lane K. Christenson, Wei-Ting Hung, and Steven A. Weinman. The RNA binding protein FMR1 controls selective exosomal miRNA cargo loading during inflammation. Journal of Cell Biology, 219(10):e201912074, October 5, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/219/10/e201912074/152116/The-
RNA-binding-protein-FMR1-controls-selective.

Watson:2021:HES

[WAOS^{+}21] Joseph L. Watson, Samya Aich, Benjamí Oller-Salvia, Andrew A. Drabek, Stephen C. Blacklow, Jason Chin, and Emmanuel Derivery. High-efficacy subcellular micropatterning of proteins using fibrinogen anchors. Journal of Cell Biology, 220(2):e202009063, February 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/220/2/e202009063/211662/High-
efficacy-subcellular-micropatterning-of.

Wang:2020:CLA

[WB20] Yuou Wang and William M. Brieher. CD2AP links actin to PI3 kinase activity to extend epithelial cell height and constrain cell area. Journal of Cell Biology, 219(1):e201812087, January 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Work:2021:AUP

[WB21] Jeremy J. Work and Onn Brandman. Adaptability of the ubiquitin-proteasome system to proteolytic and folding stressors. Journal of Cell Biology, 220(3):e201912041, March 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/220/3/e201912041/211650/Adaptability-of-the-
ubiquitin-proteasome-system-to.

Wood:2021:CCG

[WBH^{+}21] Brent M. Wood, Valentina Baena, Hai Huang, Danielle M. Jorgens, Mark Terasaki, and Thomas B. Kornberg. Cytonemes with complex geometries and composition extend into invaginations of target cells. Journal of Cell Biology, 220(5):e202101116, May 3, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/5/
e202101116/211896/Cytonemes-with-complex-geometries-
and-composition.

Waithe:2020:ODN

[WBR^{+}20] Dominic Waithe, Jill M. Brown, Katharina Reglinski, Isabel Diez-Sevilla, David Roberts, and Christian Eggeling. Object detection networks and augmented reality for cellular detection in fluorescence microscopy. Journal of Cell Biology, 219(10): e201903166, October 5, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/219/10/e201903166/152049/Object-detection-
networks-and-augmented-reality.

Windham:2022:LDG

[WC22] Ian A. Windham and Sarah Cohen. Lipid droplets go through a (liquid crystalline) phase. Journal of Cell Biology, 221(11): e202210008, November 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/221/11/e202210008/213559/Lipid-droplets-
go-through-a-liquid-crystalline.

Williams:2023:RPF

[WCC^{+}23] Jeffrey M. Williams, Yu-Jie Chen, Woo Jung Cho, Andrew W. Tai, and Billy Tsai. Reticulons promote formation of ERderived double-membrane vesicles that facilitate SARS-CoV-2 replication. Journal of Cell Biology, 222(7):e202203060, July 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/7/
e202203060/214051/Reticulons-promote-formation-of-ER-
derived-double.

Wang:2022:MPP

[WCG^{+}22] Zheng Wang, Di Chen, Dongshi Guan, Xiaobo Liang, Jianfeng Xue, Hongyu Zhao, Guangtao Song, Jizhong Lou, Yan He, and Hong Zhang. Material properties of phase-separated TFEB condensates regulate the autophagy-lysosome pathway. Journal of Cell Biology, 221(5):e202112024, May 2, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/5/e202112024/213069/Material-properties-of-phase-
separated-TFEB.

Wu:2023:CSL

[WCL^{+}23] Shu-Zon Wu, Arielle M. Chaves, Rongrong Li, Alison W. Roberts, and Magdalena Bezanilla. Cellulose synthase-like D movement in the plasma membrane requires enzymatic activity. Journal of Cell Biology, 222(6):e202212117, June 5, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/6/

e202212117/214040/Cellulose-synthase-like-D-movement-
in-the-plasma.

Walsh:2021:OFR

[WDB^{+}21] Rylie B. Walsh, Erica C. Dresselhaus, Agata N. Becalska, Matthew J. Zunitch, Cassandra R. Blanchette, Amy L. Scalera, Tania Lemos, So Min Lee, Julia Apiki, ShiYu Wang, Berith Isaac, Anna Yeh, Kate Koles, and Avital A. Rodal. Opposing functions for retromer and Rab11 in extracellular vesicle traffic at presynaptic terminals. Journal of Cell Biology, 220(8):e202012034, August 2, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/8/
e202012034/212178/Opposing-functions-for-retromer-and-
Rab11-in.

Wang:2021:BDR

[WDJ^{+}21] Lin-Ing Wang, Tyler DeFosse, Janet K. Jang, Rachel A. Battaglia, Victoria F. Wagner, and Kim S. McKim. Borealin directs recruitment of the CPC to oocyte chromosomes and movement to the microtubules. Journal of Cell Biology, 220(6): e202006018, June 7, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/220/6/e202006018/211972/Borealin-directs-
recruitment-of-the-CPC-to-oocyte.

Wang:2020:DLC

[WDL^{+}20] Mei Wang, Lei Du, Aih Cheun Lee, Yan Li, Huiwen Qin, and Jie He. Different lineage contexts direct common proneural factors to specify distinct retinal cell subtypes. Journal of Cell Biology, 219(9):e202003026, September 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/9/
e202003026/151968/Different-lineage-contexts-direct-
common-pro.

Wong-Dilworth:2023:SIE

[WDRRF^{+}23] Luis Wong-Dilworth, Carmen Rodilla-Ramirez, Eleanor Fox, Steffen D. Restel, Alexander Stockhammer, Petia Adarska, and Francesca Bottanelli. STED imaging of endogenously tagged ARF GTPases reveals their distinct nanoscale localizations. Journal of Cell Biology, 222(7):e202205107, July 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
222/7/e202205107/214063/STED-imaging-of-endogenously-
tagged-ARF-GTPases.

Wenzel:2022:MRM

[WESR22] Eva Maria Wenzel, Liv Anker Elfmark, Harald Stenmark, and Camilla Raiborg. ER as master regulator of membrane trafficking and organelle function. Journal of Cell Biology, 221(10):e202205135, October 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:/

/rupress.org/jcb/article/221/10/e202205135/213468/ER-
as-master-regulator-of-membrane-trafficking-and.

Wang:2022:PPA

[WH22] Mitchell S. Wang and Morgan Huse. Phollow the phosphoinositol: Actin dynamics at the B cell immune synapse. Journal of Cell Biology, 221(9):e202208015, September 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/9/
e202208015/213394/Phollow-the-phosphoinositol-Actin-
dynamics-at-the.

Wust:2020:CFI

[WHA20] Rob C. I. Wüst, Riekelt H. Houtkooper, and Johan Auwerx. Confounding factors from inducible systems for spatiotemporal gene expression regulation. Journal of Cell Biology, 219 (7):e202003031, July 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/7/e202003031/151788/Confounding-
factors-from-inducible-systems-for.

Weier:2022:MCE

[WHE^{+}22] Ann-Kathrin Weier, Mirka Homrich, Stephanie Ebbinghaus, Pavel Juda, Eliška Miková, Robert Hauschild, Lili Zhang, Thomas Quast, Elvira Mass, Andreas Schlitzer, Waldemar Kolanus, Sven Burgdorf, Oliver J. Gruß, Miroslav Hons, Stefan Wieser, and Eva Kiermaier. Multiple centrosomes enhance migration and immune cell effector functions of mature dendritic cells. Journal of Cell Biology, 221(12):e202107134, December 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/

12/e202107134/213533/Multiple-centrosomes-enhance-migration-
and-immune.

Wakana:2021:CSS

[WHN^{+}21] Yuichi Wakana, Kaito Hayashi, Takumi Nemoto, Chiaki Watanabe, Masato Taoka, Jessica Angulo-Capel, Maria F. Garcia-Parajo, Hidetoshi Kumata, Tomonari Umemura, Hiroki Inoue, Kohei Arasaki, Felix Campelo, and Mitsuo Tagaya. The ER cholesterol sensor SCAP promotes CARTS biogenesis at ER–Golgi membrane contact sites. Journal of Cell Biology, 220(1):e202002150, January 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:/

/rupress.org/jcb/article/220/1/e202002150/211521/The-
ER-cholesterol-sensor-SCAP-promotes-CARTS.

Wilson:2022:FFN

[WI22] Darren Graham Samuel Wilson and Thomas Iskratsch. Factoring in the force: a novel role for eIF6. Journal of Cell Biology, 221(2): e202201002, February 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/221/2/e202201002/212979/Factoring-in-
the-force-A-novel-role-for.

Wang:2023:RCC

[WJL^{+}23] Xinyi Wang, Xiao Jiang, Boran Li, Jiahua Zheng, Jiansheng Guo, Lei Gao, Mengjie Du, Xialian Weng, Lin Li, She Chen, Jingzi Zhang, Lei Fang, Ting Liu, Liang Wang, Wei Liu, Dante Neculai, and Qiming Sun. A regulatory circuit comprising the CBP and SIRT7 regulates FAM134B-mediated ERphagy. Journal of Cell Biology, 222(5):e202201068, May 1, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/
5/e202201068/214020/A-regulatory-circuit-comprising-
the-CBP-and-SIRT7.

Wang:2022:PEG

[WJW^{+}22] Yichen Wang, Lin Jia, Cong Wang, Zhonghua Du, Shilin Zhang, Lei Zhou, Xue Wen, Hui Li, Huiling Chen, Yuanyuan Nie, Dan Li, Shanshan Liu, Daniela Salgado Figueroa, Ferhat Ay, Wei Xu, Songling Zhang, Wei Li, Jiuwei Cui, Andrew R. Hoffman, Hui Guo, and Ji-Fan Hu. Pluripotency exit is guided by the Peln1-mediated disruption of intrachromosomal architecture. Journal of Cell Biology, 221(4):e202009134, April 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/4/
e202009134/213009/Pluripotency-exit-is-guided-by-the-
Peln1-mediated.

Wong:2022:MFM

[WKC^{+}22] Yvette C. Wong, Soojin Kim, Jasmine Cisneros, Catherine G. Molakal, Pingping Song, Steven J. Lubbe, and Dimitri Krainc. Mid51/Fis1 mitochondrial oligomerization complex drives lysosomal untethering and network dynamics. Journal of Cell Biology, 221(10):e202206140, October 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/10/
e202206140/213434/Mid51-Fis1-mitochondrial-oligomerization-
complex.

Wang:2021:CNR

[WKX^{+}21] Mengzhe Wang, Tatjana Kleele, Yan Xiao, Gabriela Plucinska, Petros Avramopoulos, Stefan Engelhardt, Markus H. Schwab, Matthias Kneussel, Tim Czopka, Diane L. Sherman, Peter J. Brophy, Thomas Misgeld, and Monika S. Brill. Completion of neuronal remodeling prompts myelination along developing motor axon branches. Journal of Cell Biology, 220(4):??, April 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/4/
e201911114/211755/Completion-of-neuronal-remodeling-
prompts.

Warecki:2020:EIM

[WLBS20] Brandt Warecki, Xi Ling, Ian Bast, and William Sullivan. ESCRT-III–mediated membrane fusion drives chromosome fragments through nuclear envelope channels. Journal of Cell Biology, 219(3):e201905091, March 2, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/3/
e201905091/133702/ESCRT-III-mediated-membrane-fusion-
drives.

Wang:2020:RCA

[WLM^{+}20] Tong Wang, Wei Li, Sally Martin, Andreas Papadopulos, Merja Joensuu, Chunxia Liu, Anmin Jiang, Golnoosh Shamsollahi, Rumelo Amor, Vanessa Lanoue, Pranesh Padmanabhan, and Frédéric A. Meunier. Radial contractility of actomyosin rings facilitates axonal trafficking and structural stability. Journal of Cell Biology, 219(5):e201902001, May 4, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/5/
e201902001/151566/Radial-contractility-of-actomyosin-
rings.

Wu:2021:LCR

[WLM^{+}21] Tianyu Wu, Simon I. R. Lane, Stephanie L. Morgan, Feng Tang, and Keith T. Jones. Loss of centromeric RNA activates the spindle assembly checkpoint in mammalian female meiosis I. Journal of Cell Biology, 220(10):e202011153, October 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/10/
e202011153/212548/Loss-of-centromeric-RNA-activates-
the-spindle.

Wang:2022:SNA

[WLW^{+}22] Xianghong Wang, Xinxin Li, Junkai Wang, Jiabin Wang, Can Hu, Jia Zeng, Anbing Shi, and Long Lin. SMGL-1/NBAS acts as a RAB-8 GEF to regulate unconventional protein secretion. Journal of Cell Biology, 221(7):e202111125, July 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/7/
e202111125/213235/SMGL-1-NBAS-acts-as-a-RAB-8-GEF-to-
regulate.

Weavers:2020:CBI

[WM20] Helen Weavers and Paul Martin. The cell biology of inflammation: From common traits to remarkable immunological adaptations. Journal of Cell Biology, 219(7):e202004003, July 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/
7/e202004003/151857/The-cell-biology-of-inflammation-
From-common.

Wethekam:2023:TIR

[WM23] Linnea C. Wethekam and Jeffrey K. Moore. Tubulin isotype regulation maintains asymmetric requirement for αtubulin over β-tubulin. Journal of Cell Biology, 222(3): e202202102, March 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/222/3/e202202102/213847/Tubulin-isotype-
regulation-maintains-asymmetric.

Wildenberg:2023:ARF

[WMA^{+}23] Gregg Wildenberg, Deborah J. Mariner, Panos Z. Anastasiadis, Micheal A. Davis, Renee Ireton, Huapeng Yu, Andrew L. Smith, Sarah Kurley, Juliet M. Daniel, Agnes Roczniak-Ferguson, Xiaobo Xia, Meredith Brown, Manish Tripathi, Molly Seale, Robert Carnahan, Whitney Smalley-Freed, and Michael Dohn. Albert Reynolds (1956–2022): The father of p120. Journal of Cell Biology, 222(2):e202211100, February 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/222/2/e202211100/213832/Albert-Reynolds-1956-
2022-The-father-of-p120Albert.

Wong:2022:TOF

[WME22] Louise H. Wong, Andrea Martello, and Emily R. Eden. Thank ORP9 for FFAT: With endosomal ORP10, it’s fission accomplished! Journal of Cell Biology, 221(1):e202112057, January 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/1/
e202112057/212925/Thank-ORP9-for-FFAT-With-endosomal-
ORP10-it-s.

Wan:2023:KRN

[WMM^{+}23] Yuansong Wan, Momo Morikawa, Manatsu Morikawa, Suguru Iwata, Muhammad Imran Naseer, Adeel Gulzar Ahmed Chaudhary, Yosuke Tanaka, and Nobutaka Hirokawa. KIF4 regulates neuronal morphology and seizure susceptibility via the PARP1 signaling pathway. Journal of Cell Biology, 222 (2):e202208108, February 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/222/2/e202208108/213741/KIF4-
regulates-neuronal-morphology-and-seizure.

Watanabe:2020:CIM

[WMS^{+}20] Sadanori Watanabe, Franz Meitinger, Andrew K. Shiau, Karen Oegema, and Arshad Desai. Centriole-independent mitotic spindle assembly relies on the PCNT–CDK5RAP2 pericentriolar matrix. Journal of Cell Biology, 219(12):e202006010, December 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/
12/e202006010/211524/Centriole-independent-mitotic-spindle-
assembly.

Wouters:2021:ATS

[WMS^{+}21] Rosanne Wouters, Christine Michiels, Ragna Sannerud, Bertrand Kleizen, Katleen Dillen, Wendy Vermeire, Abril Escamilla Ayala, David Demedts, Randy Schekman, and Wim Annaert. Assembly of -secretase occurs through stable dimers after exit from the endoplasmic reticulum. Journal of Cell Biology, 220(9): e201911104, September 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/9/e201911104/212501/Assembly-of-
secretase-occurs-through-stable-dimers.

Welch:2021:GGB

[WPCB^{+}21] Lawrence G. Welch, Sew-Yeu Peak-Chew, Farida Begum, Tim J. Stevens, and Sean Munro. GOLPH3 and GOLPH3L are broad-spectrum COPI adaptors for sorting into intra-Golgi transport vesicles. Journal of Cell Biology, 220(10): e202106115, October 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/10/e202106115/212611/GOLPH3-and-
GOLPH3L-are-broad-spectrum-COPI.

Waghmare:2021:GCU

[WPM21] Indrayani Waghmare and Andrea Page-McCaw. Glypicans and cytonemes unite to distribute Wnt ligands. Journal of Cell Biology, 220(12):e202110033, December 6, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/12/
e202110033/212815/Glypicans-and-cytonemes-unite-to-distribute-
Wnt.

Wang:2022:GTM

[WPS22] Xin Wang, Chih-Yu Pai, and David E. Stone. Gradient tracking in mating yeast depends on bud1 inactivation and actin-independent vesicle delivery. Journal of Cell Biology, 221(12):e202203004, December 5, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/12/
e202203004/213500/Gradient-tracking-in-mating-yeast-
depends-on-Bud1.

Wenzel:2022:ECM

[WR22] Eva Maria Wenzel and Camilla Raiborg. ER–endosome contacts master the ins and outs of secretory endosomes.

Journal of Cell Biology, 221(12):e202210033, December 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/
12/e202210033/213676/ER-endosome-contacts-master-the-
ins-and-outs-of.

Wirshing:2023:ETB

[WRG23] Alison C. E. Wirshing, Sofia Gonzalez Rodriguez, and Bruce L. Goode. Evolutionary tuning of barbed end competition allows simultaneous construction of architecturally distinct actin structures. Journal of Cell Biology, 222(4):e202209105, April 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/4/
e202209105/213854/Evolutionary-tuning-of-barbed-end-
competition.

Wakatsuki:2021:SNV

[WTS^{+}21] Shuji Wakatsuki, Yoko Takahashi, Megumi Shibata, Naoki Adachi, Tadahiro Numakawa, Hiroshi Kunugi, and Toshiyuki Araki. Small noncoding vault RNA modulates synapse formation by amplifying MAPK signaling. Journal of Cell Biology, 220(2): e201911078, February 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/2/e201911078/211679/Small-noncoding-
vault-RNA-modulates-synapse.

Wieczorek:2021:BRR

[WTU^{+}21] Michal Wieczorek, Shih-Chieh Ti, Linas Urnavicius, Kelly R. Molloy, Amol Aher, Brian T. Chait, and Tarun M. Kapoor. Biochemical reconstitutions reveal principles of human γ-TuRC assembly and function. Journal of Cell Biology, 220(3):e202009146, March 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/3/e202009146/211719/Biochemical-reconstitutions-
reveal-principles-of.

Wang:2022:PLP

[WXM22] Mengdie Wang, Choua Xiong, and Arthur M. Mercurio. PD-LI promotes rear retraction during persistent cell migration by altering integrin β4 dynamics. Journal of Cell Biology, 221(5):e202108083, May 2, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:

//rupress.org/jcb/article/221/5/e202108083/213099/PD-
LI-promotes-rear-retraction-during-persistent.

Wang:2020:NRP

[WYG^{+}20] Yaxi Wang, Peihua Yuan, Aby Grabon, Ashutosh Tripathi, Dongju Lee, Martin Rodriguez, Max Lönnfors, Michal Eisenberg-Bord, Zehua Wang, Sin Man Lam, Maya Schuldiner, and Vytas A. Bankaitis. Noncanonical regulation of phosphatidylserine metabolism by a Sec14-like protein and a lipid kinase. Journal of Cell Biology, 219(5):e201907128, May 4, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/5/
e201907128/151686/Noncanonical-regulation-of-phosphatidylserine.

Wong:2021:IRL

[WYL21] Andrew King On Wong, Barry Paul Young, and Christopher J. R. Loewen. Ist2 recruits the lipid transporters Osh6/7 to ER–PM contacts to maintain phospholipid metabolism. Journal of Cell Biology, 220(9):e201910161, September 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/9/
e201910161/212468/Ist2-recruits-the-lipid-transporters-
Osh6-7-to-ER.

Wang:2022:NMT

[WZ22] Zheng Wang and Hong Zhang. NCOA4: More than a receptor for ferritinophagy. Journal of Cell Biology, 221(10): e202209004, October 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/221/10/e202209004/213471/NCOA4-More-than-
a-receptor-for-ferritinophagyNCOA4.

Wang:2022:CEF

[WZG22] Haodong Wang, Min Zhang, and Liang Ge. Cholesterol: Enhancing FGF2 translocation in unconventional secretion. Journal of Cell Biology, 221(11):e202210007, November 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/11/e202210007/213551/Cholesterol-Enhancing-FGF2-
translocation-in.

Wu:2023:CDI

[WZK^{+}23] Jing Ze Wu, Mariia Zeziulia, Whijin Kwon, Thomas J. Jentsch, Sergio Grinstein, and Spencer A. Freeman. ClC-7 drives intraphagosomal chloride accumulation to support hydrolase activity and phagosome resolution. Journal of Cell Biology, 222 (6):e202208155, June 5, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/222/6/e202208155/213990/ClC-7-drives-
intraphagosomal-chloride-accumulation.

Wang:2020:HMF

[WZtM^{+}20] Wei Wang, Alba Zuidema, Lisa te Molder, Leila Nahidiazar, Liesbeth Hoekman, Thomas Schmidt, Stefano Coppola, and Arnoud Sonnenberg. Hemidesmosomes modulate force generation via focal adhesions. Journal of Cell Biology, 219(2):e201904137, February 3, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/2/
e201904137/133567/Hemidesmosomes-modulate-force-generation-
via-focal.

Wei:2023:CGR

[WZZ^{+}23] Wenfan Wei, Biyu Zheng, Shengnan Zheng, Daqiang Wu, Yongkang Chu, Shenghao Zhang, Dongmei Wang, Xiaopeng Ma, Xing Liu, Xuebiao Yao, and Chuanhai Fu. The Cdc42 GAP Rga6 promotes monopolar outgrowth of spores. Journal of Cell Biology, 222(1):e202202064, January 2, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:/

/rupress.org/jcb/article/222/1/e202202064/213678/The-
Cdc42-GAP-Rga6-promotes-monopolar-outgrowth-of.

Xin:2022:UMP

[XDY^{+}22] Nan Xin, Jenni Durieux, Chunxia Yang, Suzanne Wolff, Hyun-Eui Kim, and Andrew Dillin. The UPR mt preserves mitochondrial import to extend lifespan. Journal of Cell Biology, 221(7):e202201071, July 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:/

/rupress.org/jcb/article/221/7/e202201071/213239/The-
UPRmt-preserves-mitochondrial-import-to-extend.

Xie:2023:IBT

[XGD^{+}23] Boyang Xie, Clara Guillem, Swapneeta S. Date, Cameron I. Cohen, Christian Jung, Amy K. Kendall, Jordan T. Best, Todd R. Graham, and Lauren P. Jackson. An interaction between β′ − COP andtheArfGAP , Glo3, maintainspost − Golgicargorecycling.JournalofCellBiology, 222(4) : e202008061, April3, 2023.CODENJCLBA3.ISSN0021−

9525(print), 1540−8140(electronic).URLhttpsrupressorgjcbarticle2224e202008061213887AninteractionbetweenCOPandtheArfGAPGlo3.

Xu:2020:PBR

[XHF^{+}20] Xiaozheng Xu, Bowen Hou, Amitkumar Fulzele, Takeya Masubuchi, Yunlong Zhao, Zijun Wu, Yanyan Hu, Yong Jiang, Yanzhe Ma, Haopeng Wang, Eric J. Bennett, Guo Fu, and Enfu Hui. PD-1 and BTLA regulate T cell signaling differentially and only partially through SHP1 and SHP2. Journal of Cell Biology, 219(6):e201905085, June 1, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/219/6/e201905085/151801/PD-1-
and-BTLA-regulate-T-cell-signaling.

Xue:2023:LPP

[XVW^{+}23] Qian Xue, Sophia R. S. Varady, Trinity Q. Alaka’i Waddell, Mackenzie R. Roman, James Carrington, and Minna Roh-Johnson. Lack of Paxillin phosphorylation promotes singlecell migration in vivo. Journal of Cell Biology, 222(3): e202206078, March 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/222/3/e202206078/213850/Lack-of-Paxillin-
phosphorylation-promotes-single.

Xiong:2023:AIR

[XYG^{+}23] Mengneng Xiong, Lisha Yin, Yiqian Gui, Chunyu Lv, Xixiang Ma, Shuangshuang Guo, Yanqing Wu, Shenglei Feng, Xv Fan, Shumin Zhou, Lingjuan Wang, Yujiao Wen, Xiaoli Wang, Qingzhen Xie, Satoshi H. Namekawa, and Shuiqiao Yuan. ADAD2 interacts with RNF17 in P-bodies to repress the ping-pong cycle in pachytene piRNA biogenesis. Journal of Cell Biology, 222(5):e202206067, May 1, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/
5/e202206067/213948/ADAD2-interacts-with-RNF17-in-P-
bodies-to-repress.

Xing:2021:REW

[XZJ^{+}21] Ruxiao Xing, Hejiang Zhou, Youli Jian, Lingling Li, Min Wang, Nan Liu, Qiuyuan Yin, Ziqi Liang, Weixiang Guo, and Chonglin Yang. The Rab7 effector WDR91 promotes autophagylysosome degradation in neurons by regulating lysosome fusion. Journal of Cell Biology, 220(8):e202007061, August 2, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/

8/e202007061/212180/The-Rab7-effector-WDR91-promotes-
autophagy.

Yamamoto:2021:GGE

[Yam21] Ai Yamamoto. Go for the Golgi: Eating selectively with calcoco1. Journal of Cell Biology, 220(6):e202105005, June 7, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/6/
e202105005/212114/Go-for-the-Golgi-Eating-selectively-
with.

Yang:2021:EAD

[YCC^{+}21] Yanrui Yang, Jiang Chen, Xue Chen, Di Li, Jianfeng He, Shen Wang, Shun Zhao, Xiaoyu Yang, Shikun Deng, Chunfang Tong, Dou Wang, Zhenzhen Guo, Dong Li, Cong Ma, Xin Liang, Yun S. Shi, and Jia-Jia Liu. Endophilin A1 drives acute structural plasticity of dendritic spines in response to Ca^{2+}/calmodulin. Journal of Cell Biology, 220(6):e202007172, June 7, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/220/6/e202007172/212102/Endophilin-A1-drives-
acute-structural-plasticity.

Yin:2020:CRL

[YJX^{+}20] Qiuyuan Yin, Youli Jian, Meng Xu, Xiahe Huang, Niya Wang, Zhifang Liu, Qian Li, Jinglin Li, Hejiang Zhou, Lin Xu, Yingchun Wang, and Chonglin Yang. CDK4/6 regulate lysosome biogenesis through TFEB/TFE3. Journal of Cell Biology, 219(8):e201911036, August 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/219/8/e201911036/151944/CDK4-
6-regulate-lysosome-biogenesis-through-TFEB.

Yamano:2020:CRM

[YKK^{+}20] Koji Yamano, Reika Kikuchi, Waka Kojima, Ryota Hayashida, Fumika Koyano, Junko Kawawaki, Takuji Shoda, Yosuke Demizu, Mikihiko Naito, Keiji Tanaka, and Noriyuki Matsuda. Critical role of mitochondrial ubiquitination and the OPTN–ATG9A axis in mitophagy. Journal of Cell Biology, 219(9):e201912144, September 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/219/9/e201912144/151863/Critical-role-of-
mitochondrial-ubiquitination-and.

Yu-Kemp:2022:MSS

[YKSC^{+}22] Hui-Chia Yu-Kemp, Rachel A. Szymanski, Daniel B. Cortes, Nicole C. Gadda, Madeline L. Lillich, Amy S. Maddox, and Mark Peifer. Micron-scale supramolecular myosin arrays help mediate cytoskeletal assembly at mature adherens junctions. Journal of Cell Biology, 221(1):e202103074, January 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/1/
e202103074/212872/Micron-scale-supramolecular-myosin-
arrays-help.

Yang:2021:LIP

[YLC^{+}21] Yihong Yang, Dong Li, Xiaoting Chao, Shashi P. Singh, Peter Thomason, Yonghong Yan, Mengqiu Dong, Lei Li, Robert H. Insall, and Huaqing Cai. Leep1 interacts with PIP 3 and the Scar/ WAVE complex to regulate cell migration and macropinocytosis. Journal of Cell Biology, 220(7):e202010096, July 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/7/
e202010096/212090/Leep1-interacts-with-PIP3-and-the-
Scar-WAVE.

Yan:2021:RRS

[YLH^{+}21] Yanling Yan, Shuai Liu, Can Hu, Chaoyi Xie, Linyue Zhao, Shimin Wang, Wenjuan Zhang, Zihang Cheng, Jinghu Gao, Xin Fu, Zhenrong Yang, Xianghong Wang, Jing Zhang, Long Lin, and Anbing Shi. RTKN-1/Rhotekin shields endosomeassociated F-actin from disassembly to ensure endocytic recycling. Journal of Cell Biology, 220(5):e202007149, May 3, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
5/e202007149/211976/RTKN-1-Rhotekin-shields-endosome-
associated-F.

Yang:2022:TSB

[YLH^{+}22] Tzu-Jing Yang, Tian-Neng Li, Rih-Sheng Huang, Max Yu-Chen Pan, Shu-Yu Lin, Steven Lin, Kuen-Phon Wu, Lily Hui-Ching Wang, and Shang-Te Danny Hsu. Tumor suppressor BAP1 nuclear import is governed by transportin-1. Journal of Cell Biology, 221(6):e202201094, June 6, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/6/

e202201094/213174/Tumor-suppressor-BAP1-nuclear-import-
is-governed.

Chapa-y-Lazo:2020:PRD

[yLHW^{+}20] Bernardo Chapa y Lazo, Motonari Hamanaka, Alexander Wray, Mohan K. Balasubramanian, and Masanori Mishima. Polar relaxation by dynein-mediated removal of cortical myosin II. Journal of Cell Biology, 219(8):e201903080, August 3, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/219/8/e201903080/151836/Polar-relaxation-by-
dynein-mediated-removal-of.

Yoo:2021:GMN

[YM21] Tae Yeon Yoo and Timothy J. Mitchison. O-GlcNAc modification of nuclear pore complexes accelerates bidirectional transport. Journal of Cell Biology, 220(7):e202010141, July 5, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
7/e202010141/212033/O-GlcNAc-modification-of-nuclear-
pore-complexes.

Yaniv:2020:DAR

[YMAS20] Shiri P. Yaniv, Hagar Meltzer, Idan Alyagor, and Oren Schuldiner. Developmental axon regrowth and primary neuron sprouting utilize distinct actin elongation factors. Journal of Cell Biology, 219(5):e201903181, May 4, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/5/
e201903181/151569/Developmental-axon-regrowth-and-primary-
neuron.

Yang:2020:TWV

[YMH^{+}20] Zhe Yang, Brendan C. Mattingly, David H. Hall, Brian D. Ackley, and Matthew Buechner. Terminal web and vesicle trafficking proteins mediate nematode single-cell tubulogenesis. Journal of Cell Biology, 219(11):e202003152, November 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/11/
e202003152/152057/Terminal-web-and-vesicle-trafficking-
proteins.

Yahya:2021:SGD

[YPM^{+}21] Galal Yahya, Alexis P. Pérez, Mònica B. Mendoza, Eva Parisi, David F. Moreno, Marta H. Artés, Carme Gallego, and Martí Aldea. Stress granules display bistable dynamics modulated by Cdk. Journal of Cell Biology, 220(3):e202005102, March 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
3/e202005102/211705/Stress-granules-display-bistable-
dynamics.

You:2002:WSP

[YSC^{+}02] Zongbing You, Daniel Saims, Shaoqiong Chen, Zhaocheng Zhang, Denis C. Guttridge, Kun liang Guan, Ormond A. Mac-Dougald, Anthony M. C. Brown, Gerard Evan, Jan Kitajewski, and Cun-Yu Wang. Wnt signaling promotes oncogenic transformation by inhibiting c–Myc–induced apoptosis. Journal of Cell Biology, 157(3):429–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/3/429. See correction [YSC^{+}21].

You:2021:CWS

[YSC^{+}21] Zongbing You, Daniel Saims, Shaoqiong Chen, Zhaocheng Zhang, Denis C. Guttridge, Kun liang Guan, Ormond A. MacDougald, Anthony M. C. Brown, Gerard Evan, Jan Kitajewski, and Cun-Yu Wang. Correction: Wnt signaling promotes oncogenic transformation by inhibiting c-Myc-induced apoptosis. Journal of Cell Biology, 220(5): ??, May 3, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/5/jcb.20020111004052021c/211961/
Correction-Wnt-signaling-promotes-oncogenic.

Yan:2021:HCI

[YSR^{+}21] Xiaowei Yan, Nico Stuurman, Susana A. Ribeiro, Marvin E. Tanenbaum, Max A. Horlbeck, Christina R. Liem, Marco Jost, Jonathan S. Weissman, and Ronald D. Vale. Highcontent imaging-based pooled CRISPR screens in mammalian cells. Journal of Cell Biology, 220(2):e202008158, February 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/2/
e202008158/211696/High-content-imaging-based-pooled-
CRISPR-screens.

Yamashita:2020:MPP

[YTH^{+}20] Koichiro Yamashita, Shigehiko Tamura, Masanori Honsho, Hiroto Yada, Yuichi Yagita, Hidetaka Kosako, and Yukio Fujiki. Mitotic phosphorylation of Pex14p regulates peroxisomal import machinery. Journal of Cell Biology, 219(10):e202001003, October 5, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
219/10/e202001003/152047/Mitotic-phosphorylation-of-
Pex14p-regulates.

Yang:2021:LBR

[YW21] Chonglin Yang and Xiaochen Wang. Lysosome biogenesis: Regulation and functions. Journal of Cell Biology, 220(6):e202102001, June 7, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/jcb/

article/220/6/e202102001/212053/Lysosome-biogenesis-
Regulation-and.

Yang:2020:TRV

[YZW^{+}20] Xi Yang, Weichao Zhang, Xin Wen, Patrick J. Bulinski, Dominic A. Chomchai, Felichi Mae Arines, Yun-Yu Liu, Simon Sprenger, David Teis, Daniel J. Klionsky, and Ming Li. TORC1 regulates vacuole membrane composition through ubiquitin- and ESCRT-dependent microautophagy. Journal of Cell Biology, 219(3):e201902127, March 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/
3/e201902127/133713/TORC1-regulates-vacuole-membrane-
composition.

Yu:2020:SIG

[YZY^{+}20] Caiting Yu, Jinghua Zhao, Liming Yan, Yuanbo Qi, Xiangyang Guo, Zhiyong Lou, Junjie Hu, and Zihe Rao. Structural insights into G domain dimerization and pathogenic mutation of OPA1. Journal of Cell Biology, 219(7):e201907098, July 6, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/219/7/
e201907098/151743/Structural-insights-into-G-domain-
dimerization-and.

Zuidema:2022:PYP

[ZAK^{+}22] Alba Zuidema, Paul Atherton, Maaike Kreft, Liesbeth Hoekman, Onno B. Bleijerveld, Nagarjuna Nagaraj, Nanpeng Chen, Reinhard Fässler, and Arnoud Sonnenberg. PEAK1 Y635 phosphorylation regulates cell migration through association with Tensin3 and integrins. Journal of Cell Biology, 221(8):e202108027, August 1, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/8/e202108027/213273/PEAK1-Y635-phosphorylation-regulates-
cell.

Zarnescu:2020:TGA

[Zar20] Daniela C. Zarnescu. Think globally, act locally: Centrosomelocalized mRNAs ensure mitotic fidelity. Journal of Cell Biology, 219(12):e202010172, December 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/219/12/e202010172/211553/Think-
globally-act-locally-Centrosome-localized.

Zhou:2021:CCA

[ZAR^{+}21] Yong Zhou, Nicholas Ariotti, James Rae, Hong Liang, Vikas Tillu, Shern Tee, Michele Bastiani, Adekunle T. Bademosi, Brett M. Collins, Frederic A. Meunier, John F. Hancock, and Robert G. Parton. Caveolin-1 and cavin1 act synergistically to generate a unique lipid environment in caveolae. Journal of Cell Biology, 220(3):e202005138, March 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/3/
e202005138/211716/Caveolin-1-and-cavin1-act-synergistically-
to.

Zhao:2022:PAM

[ZBM^{+}22] Jierui Zhao, Mai Thu Bui, Juncai Ma, Fabian Künzl, Lorenzo Picchianti, Juan Carlos De La Concepcion, Yixuan Chen, Sofia Petsangouraki, Azadeh Mohseni, Marta García-Leon, Marta Salas Gomez, Caterina Giannini, Dubois Gwennogan, Roksolana Kobylinska, Marion Clavel, Swen Schellmann, Yvon Jaillais, Jiri Friml, Byung-Ho Kang, and Yasin Dagdas. Plant autophagosomes mature into amphisomes prior to their delivery to the central vacuole. Journal of Cell Biology, 221(12):e202203139, December 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
221/12/e202203139/213556/Plant-autophagosomes-mature-
into-amphisomes-prior.

Zhang:2021:CPC

[ZBY^{+}21] Yu-Cheng Zhang, Yun-Feng Bai, Jin-Feng Yuan, Xiao-Lin Shen, Yu-Ling Xu, Xiao-Xiao Jian, Sen Li, Zeng-Qing Song, Huai-Bin Hu, Pei-Yao Li, Hai-Qing Tu, Qiu-Ying Han, Na Wang, Ai-Ling Li, Xue-Min Zhang, Min Wu, Tao Zhou, and Hui-Yan Li. CEP55 promotes cilia disassembly through stabilizing Aurora A kinase. Journal of Cell Biology, 220(2):e202003149, February 1, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
2/e202003149/211702/CEP55-promotes-cilia-disassembly-
through.

Zomot:2021:BRC

[ZCD^{+}21] Elia Zomot, Hadas Achildiev Cohen, Inbal Dagan, Ruslana Militsin, and Raz Palty. Bidirectional regulation of calcium release–activated calcium (CRAC) channel by SARAF. Journal of Cell Biology, 220(12):e202104007, December 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/12/e202104007/212731/Bidirectional-regulation-of-
calcium-release.

Zheng:2022:ULM

[ZCL^{+}22] Jun Zheng, Xi Chen, Qiang Liu, Guisheng Zhong, and Min Zhuang. Ubiquitin ligase MARCH5 localizes to peroxisomes to regulate pexophagy. Journal of Cell Biology, 221(1): e202103156, January 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/221/1/e202103156/212749/Ubiquitin-ligase-
MARCH5-localizes-to-peroxisomes.

Zhou:2022:PMS

[ZDGB^{+}22] Jun Zhou, Yasamin Dabiri, Rodrigo A. Gama-Brambila, Shahrouz Ghafoory, Mukaddes Altinbay, Arianeb Mehrabi, Mohammad Golriz, Biljana Blagojevic, Stefanie Reuter, Kang Han, Anna Seidel, Ivan ikić, Stefan Wölfl, and Xinlai Cheng. pVHL-mediated SMAD3 degradation suppresses TGF- signaling. Journal of Cell Biology, 221(1):e202012097, January 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/
1/e202012097/212891/pVHL-mediated-SMAD3-degradation-
suppresses-TGF.

Zouiouich:2022:MER

[ZDM^{+}22] Mehdi Zouiouich, Thomas Di Mattia, Arthur Martinet, Julie Eichler, Corinne Wendling, Nario Tomishige, Erwan Grandgirard, Nicolas Fuggetta, Catherine Fromental-Ramain, Giulia Mizzon, Calvin Dumesnil, Maxime Carpentier, Bernardo Reina-San-Martin, Carole Mathelin, Yannick Schwab, Abdou Rachid Thiam, Toshihide Kobayashi, Guillaume Drin, Catherine Tomasetto, and Fabien Alpy. MOSPD2 is an endoplasmic reticulum–lipid droplet tether functioning in LD homeostasis. Journal of Cell Biology, 221(6):e202110044, June 6, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/6/
e202110044/213116/MOSPD2-is-an-endoplasmic-reticulum-
lipid-droplet.

Zhou:2023:LLP

[ZFZ^{+}23] Sheng Zhou, Zhifei Fu, Ziwei Zhang, Xing Jia, Guangjun Xu, Long Sun, Fei Sun, Pu Gao, Pingyong Xu, and Hongyu Deng. Liquid–liquid phase separation mediates the formation of herpesvirus assembly compartments. Journal of Cell Biology, 222(1):e202201088, January 2, 2023. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/222/1/
e202201088/213550/Liquid-liquid-phase-separation-mediates-
the.

Zihni:2022:SCA

[ZGR^{+}22] Ceniz Zihni, Anastasios Georgiadis, Conor M. Ramsden, Elena Sanchez-Heras, Alexis J. Haas, Britta Nommiste, Olha Semenyuk, James W. B. Bainbridge, Peter J. Coffey, Alexander J. Smith, Robin R. Ali, Maria S. Balda, and Karl Matter. Spatiotemporal control of actomyosin contractility by MRCKβ signaling drives phagocytosis. Journal of Cell Biology, 221(11):e202012042, November 7, 2022. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/11/
e202012042/213476/Spatiotemporal-control-of-actomyosin-
contractility.

Zhang:2022:RMD

[ZHHJ22] Yaqian Zhang, Xing Hong, Shasha Hua, and Kai Jiang. Reconstitution and mechanistic dissection of the human microtubule branching machinery. Journal of Cell Biology, 221(7): e202109053, July 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/221/7/e202109053/213234/Reconstitution-and-
mechanistic-dissection-of-the.

Zhang:2021:MCE

[ZHW^{+}21] Jingjing Zhang, Ying Hu, Yanli Wang, Lin Fu, Xiumei Xu, Chunxia Li, Jie Xu, Chengbin Li, Linqiang Zhang, Rendan Yang, Xue Jiang, Yingjie Wu, Pingsheng Liu, Xiaoju Zou, and Bin Liang. mmBCFA C17iso ensures endoplasmic reticulum integrity for lipid droplet growth. Journal of Cell Biology, 220(11):e202102122, November 1, 2021. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/11/
e202102122/212690/mmBCFA-C17iso-ensures-endoplasmic-
reticulum.

Zhou:2022:MFM

[ZJDR22] Jialin Zhou, Martin Jung, Kai S. Dimmer, and Doron Rapaport. The multi-factor modulated biogenesis of the mitochondrial multi-span protein Om14. Journal of Cell Biology, 221(4): e202112030, April 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/221/4/e202112030/213056/The-multi-factor-
modulated-biogenesis-of-the.

Zhu:2022:PPC

[ZJH22] Qian Zhu, Zhaodi Jiang, and Xiangwei He. Pcp1/pericentrin controls the SPB number in fission yeast meiosis and ploidy homeostasis. Journal of Cell Biology, 221(1):e202104099, January 3, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/221/1/e202104099/212751/Pcp1-pericentrin-
controls-the-SPB-number-in.

Zhu:2022:TIP

[ZLJ^{+}22] Yueyao Zhu, Shuixing Li, Alexa Jaume, Riddhi Atul Jani, Cédric Delevoye, Graça Raposo, and Michael S. Marks. Type II phosphatidylinositol 4-kinases function sequentially in cargo delivery from early endosomes to melanosomes. Journal of Cell Biology, 221(11):e202110114, November 7, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/221/

11/e202110114/213509/Type-II-phosphatidylinositol-4-
kinases-function.

Zhang:2023:LCT

[ZLJ^{+}23] Qianqian Zhang, Yuan Li, Youli Jian, Meijiao Li, and Xiaochen Wang. Lysosomal chloride transporter CLH-6 protects lysosome membrane integrity via cathepsin activation. Journal of Cell Biology, 222(6):e202210063, June 5, 2023. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/222/6/
e202210063/214033/Lysosomal-chloride-transporter-CLH-
6-protects.

Zaman:2021:EME

[ZLS^{+}21] Riasat Zaman, Andrew Lombardo, Cécile Sauvanet, Raghuvir Viswanatha, Valerie Awad, Locke Ezra-Ros Bonomo, David Mc-Dermitt, and Anthony Bretscher. Effector-mediated ERM activation locally inhibits RhoA activity to shape the apical cell domain. Journal of Cell Biology, 220(6):e202007146, June 7, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/
6/e202007146/211973/Effector-mediated-ERM-activation-
locally-inhibits.

Zhao:2023:PPR

[ZLW23] Ning Zhao, Ning Li, and Tao Wang. PERK prevents rhodopsin degradation during retinitis pigmentosa by inhibiting IRE1-induced autophagy. Journal of Cell Biology, 222 (5):e202208147, May 1, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://

rupress.org/jcb/article/222/5/e202208147/214004/PERK-
prevents-rhodopsin-degradation-during.

Zagryazhskaya-Masson:2020:ITF

[ZMMM^{+}20] Anna Zagryazhskaya-Masson, Pedro Monteiro, Anne-Sophie Macé, Alessia Castagnino, Robin Ferrari, Elvira Infante, Aléria Duperray-Susini, Florent Dingli, Arpad Lanyi, Damarys Loew, Elisabeth Génot, and Philippe Chavrier. Intersection of TKS5 and FGD1/CDC42 signaling cascades directs the formation of invadopodia. Journal of Cell Biology, 219(9):e201910132, September 7, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/219/9/e201910132/151952/Intersection-of-
TKS5-and-FGD1-CDC42-signaling.

Zewe:2020:PSD

[ZMS^{+}20] James P. Zewe, April M. Miller, Sahana Sangappa, Rachel C. Wills, Brady D. Goulden, and Gerald R. V. Hammond. Probing the subcellular distribution of phosphatidylinositol reveals a surprising lack at the plasma membrane. Journal of Cell Biology, 219(3):e201906127, March 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

Zappa:2022:SIS

[ZMW^{+}22] Francesca Zappa, Nerea L. Muniozguren, Maxwell Z. Wilson, Michael S. Costello, Jose Carlos Ponce-Rojas, and Diego Acosta-Alvear. Signaling by the integrated stress response kinase PKR is fine-tuned by dynamic clustering. Journal of Cell Biology, 221 (7):e202111100, July 4, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/221/7/e202111100/213199/Signaling-by-
the-integrated-stress-response-kinase.

Zheng:2023:RRS

[ZPG^{+}23] Hui Zheng, Kangfu Peng, Xiaomeng Gou, Chen Ju, and Hong Zhang. RNA recruitment switches the fate of protein condensates from autophagic degradation to accumulation. Journal of Cell Biology, 222(6):e202210104, June 5, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https:/

/rupress.org/jcb/article/222/6/e202210104/213995/RNA-
recruitment-switches-the-fate-of-protein.

Zeng:2021:PPP

[ZPŠS21] Longhui Zeng, Ivan Palaia, Anela Šarić, and Xiaolei Su. PLCγ1 promotes phase separation of T cell signaling components. Journal of Cell Biology, 220(6):e202009154, June 7, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/220/6/
e202009154/212040/PLC-1-promotes-phase-separation-of-
T-cell.

Zhou:2023:TPA

[ZRO^{+}23] Jianwen Zhou, Nikoline Lander Rasmussen, Hallvard Lauritz Olsvik, Vyacheslav Akimov, Zehan Hu, Gry Evjen, Stéphanie Kaeser-Pebernard, Devanarayanan Siva Sankar, Carole Roubaty, Pauline Verlhac, Nicole van de Beek, Fulvio Reggiori, Yakubu Princely Abudu, Blagoy Blagoev, Trond Lamark, Terje Johansen, and Jörn Dengjel. TBK1 phosphorylation activates LIR-dependent degradation of the inflammation repressor TNIP1. Journal of Cell Biology, 222(2):e202108144, February 6, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/2/
e202108144/213785/TBK1-phosphorylation-activates-LIR-
dependent.

Zhang:2021:RDG

[ZS21] Yijun Zhang and Joachim Seemann. Rapid degradation of GRASP55 and GRASP65 reveals their immediate impact on the Golgi structure. Journal of Cell Biology, 220(1):e202007052, January 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/1/e202007052/211583/Rapid-degradation-of-GRASP55-
and-GRASP65-reveals.

Zhu:2020:CDR

[ZSJE20] Lu Zhu, Richa Sardana, Daniel K. Jin, and Scott D. Emr. Calcineurin-dependent regulation of endocytosis by a plasma membrane ubiquitin ligase adaptor, Rcr1. Journal of Cell Biology, 219(8):e201909158, August 3, 2020. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/219/8/
e201909158/151785/Calcineurin-dependent-regulation-of-
endocytosis-by.

Zhu:2023:PRG

[ZTL^{+}23] Lingxin Zhu, Yi Tang, Xiao-Yan Li, Samuel A. Kerk, Costas A. Lyssiotis, Xiaoyue Sun, Zijun Wang, Jung-Sun Cho, Jun Ma, and Stephen J. Weiss. Proteolytic regulation of a galectin-3/Lrp1 axis controls osteoclast-mediated bone resorption. Journal of Cell Biology, 222(4):e202206121, April 3, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
222/4/e202206121/213915/Proteolytic-regulation-of-a-
galectin-3-Lrp1-axis.

Zhou:2021:HSC

[ZVC^{+}21] Xin Zhou, Camille Vachon, Mélissa Cizeron, Océane Romatif, Hannes E. Bülow, Maëlle Jospin, and Jean-Louis Bessereau.

The HSPG syndecan is a core organizer of cholinergic synapses. Journal of Cell Biology, 220(9):e202011144, September 6, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL https://rupress.org/jcb/article/
220/9/e202011144/212450/The-HSPG-syndecan-is-a-core-
organizer-of.

Zucca:2023:ACC

[ZVL^{+}23] Federico Zucca, Clara Visintin, Jiaming Li, Steven P. Gygi, and Rosella Visintin. APC/C Cdc20-mediated degradation of Clb4 prompts astral microtubule stabilization at anaphase onset. Journal of Cell Biology, 222(1):e202203089, January 2, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/1/
e202203089/213563/APC-CCdc20-mediated-degradation-of-
Clb4-prompts.

Zhang:2020:PRD

[ZVM^{+}20] Xuemei Zhang, Michael Vigers, James McCarty, Jennifer N. Rauch, Glenn H. Fredrickson, Maxwell Z. Wilson, Joan-Emma Shea, Songi Han, and Kenneth S. Kosik. The proline-rich domain promotes Tau liquid–liquid phase separation in cells. Journal of Cell Biology, 219(11):e202006054, November 2, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/219/11/e202006054/152134/The-proline-
rich-domain-promotes-Tau-liquid-liquid.

Zhu:2022:MBA

[ZWJ22] Jiajun Zhu, Hua Wang, and Xuejun Jiang. mTORC1 beyond anabolic metabolism: Regulation of cell death. Journal of Cell Biology, 221(12):e202208103, December 5, 2022. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/221/12/
e202208103/213609/mTORC1-beyond-anabolic-metabolism-
Regulation-of.

Zhang:2020:MWI

[ZXW^{+}20] Zhenguo Zhang, Songbo Xie, Ruoxi Wang, Shuqun Guo, Qiuchen Zhao, Hui Nie, Yuanyuan Liu, Fengguo Zhang, Miao Chen, Libo Liu, Xiaoqian Meng, Min Liu, Li Zhao, Monica P. Colaiácovo, Jun Zhou, and Jinmin Gao. Multivalent weak interactions between assembly units drive synaptonemal complex formation. Journal of Cell Biology, 219(5): e201910086, May 4, 2020. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.org/

jcb/article/219/5/e201910086/151585/Multivalent-weak-
interactions-between-assembly.

Zhou:2023:CSE

[ZXY^{+}23] Lingjian Zhou, Xutong Xue, Ke Yang, Zhi Feng, Min Liu, and José C. Pastor-Pareja. Convergence of secretory, endosomal, and autophagic routes in trans-Golgi–associated lysosomes. Journal of Cell Biology, 222(1):e202203045, January 2, 2023. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL https://rupress.org/jcb/article/222/1/
e202203045/213547/Convergence-of-secretory-endosomal-
and-autophagic.

Zhang:2021:CTH

[ZY21] Qi Zhang and Yihong Ye. Chaperoning transmembrane helices in the lipid bilayer. Journal of Cell Biology, 220(1): e202012041, January 4, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL https://rupress.

org/jcb/article/220/1/e202012041/211631/Chaperoning-
transmembrane-helices-in-the-lipid.